A Complete Bibliography of Publications in IEEE Spectrum (2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

18 January 2020 Version 1.00

Title word cross-reference

$1 [Ste08]. $10 [Cas05b]. $12.95 [Cas01e]. $13.00 [Che01b]. $150 [Swe00a]. $22.96 [Cas02d]. $23 [Gol01c]. $24 [Gep01b]. $25 [Bre01a, Gol01b, Nah01]. $26 [Che01a]. $26.95 [Kum01a]. $27.50 [Cas01d]. $27.95 [Cas01f]. $280 [Gui07a]. 3 [Bal04, Hel02b]. $30 [Zor01a]. $35 [Cas01l]. $36 [Zor01b]. $84 [Sni00]. $99.95 [Swe00a]. _{2} [ES04, Fai09b, Moo01d].

-D [Hel02b].

0 [Cas01l, Cas05b]. 0-13-578782-3 [Sni00]. 0-395-95563-7 [Nah01]. 0-596-00287-4 [Cas02d]. 0-7382-0532-X [Cas01l].

0-9738-2030-6 [Cas05b]. ’02 [Mas02].

100 [Hir04]. 100-GHz [BTR00]. 100-meter [Per05c]. 101/Save [Cas06m]. 150-Kilometer [Das08a].

2.0 [Swe09a, Hol06]. 2000 [Ano00a]. 2004 [Ano04i, Ano04j, Ano04l, Ano04s, Ano04t, Ano04q, Ano04r, Ano04-31, Ano04x, Ano04y]. 2005 [Ano05q, Ano05r, Ano05v]. 2006 [Ano06p, Ano06s, Ano06u, Ano06t, Ano06w, Ano06-31, Ano06v, Ano06-32, Ano06-33, Ano06z, Ano06-27]. 2007 [Ano07y, Ano07-44, Ano07-43, Ano07-46, Ano07-45, Ano07-48, Ano07-47, Ano07a, Ano07s, Ano07t, GKJ^{+}07, Obe08]. 2008 [Ano08-58, Ano08-57, Mur09]. 21st [Bor01, Har00, Has09a]. 21st-Century [Has09a]. 7.50 [Gep01a].

3 [Che09e]. 3-Deep [Sul05]. 3.0 [Dur05]. 30 [Kar09c]. 360 [Che09e]. 3G [BBS^{+}01, Bla02b, LL03c, Oli00].

5 [Fos01d].

6 [Cas05b].

7 [Ano04-82].

8-hour [Ros06b].

911 [Kus06f].

A-Twitter [McF07b]. ABM [Sho05a]. abolition [Kum02d]. Above [Kar08j]. absurdity [Fra02d]. academic [App01b]. Academy [Jon08c]. access [Bre00b, Gep02b, IS01, Roy00]. Accident [Ano06c, Cas04a]. accounting [Tei01]. accuracy [KP00]. Accurate [Ano06-58]. accused [Hel05c]. achievements [Luc09c]. acquisitions [Cha08]. Acronym [San06]. Across [Che04g, Ups08a]. Act [Ano04-90, Kum04a, Per08d, Che05f, Fra02a, Gep02a, HNPC05, Moc06, Sop03]. Acting [Ano05a]. activities [McF04a]. activity [Ban02, MG00b]. actor [Kar08o]. actress [Kar08b]. acts [Chi05]. Ad [And08b]. Adaptive [Pfl00]. add [BOT06]. Addiction [San06]. addicts [LL06]. Adding [Gra00]. Administration [Swe04b]. adopts [Ano05-40]. Adrenaline [Swe06a]. Advance [Pat09a]. advancement [Gol06b]. advances [Ano06-73]. advertisement [Ano06-48]. advice [Peh01b]. Aerospace [Pat09b, Sco00]. Affair [Kar09d, Ano08a, PS07]. Affairs [Bre01a]. affordable [Der09a]. Afghan [Jon08c]. Afghanistan [KK02]. AFM [Ano04f]. Africa [App02a, Gol04d, MM06]. African [Ano07e, MM06, Zac05]. After [Ano01b, Li01, Kar08l]. Again [Ano05i, Gol05c, Che03e, Fos01c, HBW02]. against [Hoo05a, Jon03b, Tei00, Tro04]. Age [Ano09r, Ges04, Has08c, Kar09c, Bre01a, Cas05d, HD03, Koo03, McF03b, Pos00, Rio02, Sch03b, Zor01a]. Agency [Mas02, Swe05b]. ages [Luc01e]. aging [Ano04-87, GG04]. Ago [Jon00b, Jon00e, Jon00f, Jon00g, Jon00h, Jon00i, Jon00j, Jon00k, Jon00l, Jon00m, Jon00n, Jon00o]. Ahead [Cas04c, Bre02b]. AI [Cas02e]. aid [Ano02r]. aids [Jon04e]. ailing [JL04]. ails [Coo02]. Air [Wei02a, Ano05-40, Bre01l, Bre01n, Che03f, Fai09a, Rie01e, Ros02, Ros04d, Sin02, Zbi05, Ort09]. airborne [Gui04c]. Airbus [Jon05a]. aircraft [FH01, Gui04d, Kar01c]. airline [Bre01a]. airliners [Kar03a]. Airlines [Per08g]. airport [Ano02l]. Airspeed [SMAS06]. Airwaves [SR09]. Akimbo [Per05n]. Al [Ano09-87, Ano04-75, Dur05]. Alamos [Ade08e]. alarm [FPP00]. alarming [Voe06b]. alarms [Obe03c]. album [Obe02b]. alchemists [Sav03c]. alert [Ros09c]. Algae [Ups08d]. Algebra [Kar08k]. Algorithm [Ano05-47]. alive [DiG03]. All-Electric [Voe08d]. all-electronic [Rie02a]. all-optical [Mul01a]. all-seeing [Ano04a, PP09a]. all-terrain [Gui05c]. Allen [Dig08a]. alliance [Ham02, Swe03c]. allocation [SW04]. Almost [Voe08d, Sul05, Wei02b]. Alone [Ups08a]. along [Cas03c]. Also [Lin08]. Alteration [Kar09k]. alternative [Fai05b, Gol01c]. Always [Ano01c, Agr01, Has09e]. always-on [Agr01, Has09e]. am [Bro08]. amateurs [Luc07b]. Amazing [Ano09a, Cas01e, Gep02a]. amazon.com [Kus02a]. America [Obe07a, Che01b, Mur09]. American [Bro02b, LL05a]. Americans [Kar09c]. Amiga [Wal01a]. amongst [Che06g]. Amped [Jon04a]. Amy [Cas01e]. analog [Che09d, Mar01]. Analysis [Ano01n, BSK^{+}01, DHGR01, BBS^{+}01, Bre00f, Bre00g, Cas04f, Com00a, Com00c, Den01, Gep00b, Gep00c, GRG^{+}01, HCJ^{+}01, JCF^{+}01, Kap00, Kor00, KLS^{+}01, Mir00a, MMG^{+}01, MMH^{+}01, Per00a, Per00b, Rie00b, Roo00, Sco00, SSB^{+}01, Sni00, Swe00f, ZMR^{+}01, Sni00]. Analyze [LW00]. Anatomy [Cas01b, Gui07b, Sel04a]. Andes [Ano04d, Fai04b]. Andy [Gep00a]. Anger [Che05a]. Angier [Ups07f]. Anglo [Bow03]. angry [Pos00]. animal [Gol02h]. Ann [Six05e]. Announcement [Ano00o]. annual [Hir05]. Answer [Swe04e, Das05, Swe00b]. ante [Avi02b]. Antennas [SR09, Hel05c]. anthrax [Ano02h, Dig06a]. anti [Sia01]. anti-terrorist [Sia01]. Anticipating [App01c]. anticlimactic [Par02a]. Antipiracy [Cas06a]. antitrust [Swe01b]. ants [Gol01b]. anxiety [McF03b]. Any [SMAS06, Kap02, Sel02b]. Anyone [Kar09c]. anywhere [Che09f]. apart [GG04]. Ape [Ano06-60]. Apogee [Cas05b]. Apollo [And08c, Cas03b]. App [Ros09a]. appeal [Per09g]. Appeals [BSK^{+}01]. appearance [Luc03b]. Apple [Cas01c]. applications [Bre02d]. appraisals [Gol01a]. Appraising [Gol01a]. Approach [Hug09]. April [Ano04-75, Ano04i, Ano06p, Ano07s, Ano07t]. Arabia [Ups09c]. Arachnoid [Ano08a]. Arbor [Six05e]. archeologist [Cas02c]. Archibald [Ano06-52]. Architecture [Ano08-66, MO00]. archive [Gol04b]. argument [Swe08a]. Arieta [Ups09a]. Arms [Kun09, DG00]. Array [Dig08a]. arrival [Bre02c]. Arrive [Li01]. Arrives [Dig07b]. Arriving [Ano00b]. Art [Kar09a, Cas01g, Cas06g, Gol02e, Per02d, Sel05a]. artful [Dix05]. artful- [Dix05]. Article [Lin08]. Artificial [Ano01-78, Pre05g]. Ash [Kum08a]. Asia [KS04]. ASIC [Gra00]. asked [App03b]. aspects [Swe00c]. assets [Swe01c]. Assigned [Mue02]. assignment [Ano09-58]. assistance [JA05].

Associations [Swe00a]. assuming [Has08e]. asteroids [Ing01b]. Astronaut [Min09, Cas04g]. AT&T [Ano05-52, Kum07a, Rio05a]. Athens [PS07]. Atlantic [Ano05-70]. atmosphere [Dig06b]. atmospheric [Bro03b]. atom [Gui06e, McC05]. Atomic [Ade08e, Swe05b, Gep05a, Gui05a]. atoms [Ano03f]. attached [Jon06f]. attack [Ade09a, Ast01]. Attic [Kum07a]. attraction [Lub00]. attractions [Ros04a]. auction [Cas05d]. Audacity [Has08a]. audio [Day02, Kus03a, Per09a, Put03]. auditors [Jon09a]. Aug [Ano06q]. augurs [Hel00]. Aural [Kar03e]. Ausra [Per08j]. Aussies [FS04]. authentication [JP06]. author [Ano01-47]. Authority [And08a]. Automate [Sch09a]. Automated [Jon09a, Pet09]. Automation [Mar01, Per05o, Gep00c, Kum05b]. Automotive [KMT00, Ber02]. autos [Ano02g]. Awards [Ano00a, Ano00-43, Ano00-43, Ano07a]. aware [Luc07a]. Away [Che09b, Ter07, Voe08b, Gui04c, Moo05c]. Ayanna [Cas05a].

B [And08a, Obe08, Rio07a]. Babbage [Col06]. Baby [Ano06-50, Ano08-94, Ano09-73, Che03c, Gui09a, Hel01a]. Back [Ano04b, Ano04c, Ano04d, Ano05a, Ano05c, Ano05d, Ano05g, Ano05b, Ano05e, Ano05f, Ano05h, Ano05j, Ano05i, Ano05k, Ano05l, Ano06f, Ano06g, Ano06h, Ano06c, Ano06d, Ano06i, Ano06k, Ano06j, Ano06l, Ano06e, Ano06a, Ano06b, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07b, Ano07c, Ano07d, Ano07x, Ano08f, Ano08g, Ano08e, Ano08h, Ano08b, Ano08c, Ano08d, Ano08s, Ano09b, Cas05b, Moo05a, Ups08b, Ano08j, Ano08-89, Ano09n, Ano09-62, Bla03b, Che03g, Mul04b, Sin03b]. backlash [Che05o]. Backstory [Ano08i]. backups [Bre04a]. backward [Ano00b].

Bad [And07a, Ano01a, McF06f, Min06c, Wal08c, Ano02o, Cas01c, Gui04c, Kum04d, Moo08b, Sin04b]. baggage [Bre02c]. Bailouts [Moo09a]. bait [Dum09b]. Balaji [Sin04a]. Balancing [Fra02a, HNPC05, Kum04a]. Ball [Ano04-70, Gui05b]. Balloon [GP06b]. ballot [Mer02]. balls [Gui05b]. Baltimore [Zor01b]. ban [Fai05a]. Band [Kus09a]. Bandwidth [Ano04-77, Che02b, Che04b]. Banerjee [Zor09b]. bang [Mas02]. Bar [Tes07b]. Barbados [Moo06a]. Barbot [Pet09]. Bare [Cas07c]. barred [Boy04d]. Barrow [Gep01a]. bars [Per05j]. bartender [Pet09]. Based [BBS^{+}01, CT05, DGKM05, Dix05, HNPC05, Kus06d, Sav04]. Basic [Cas01d]. Basics [Ups07f]. Batteries [Ano06m, Voe07b, HBW02, Rie01c]. battering [Kum01a]. Battery [PP08a, Voe08a, Voe09c, Rie01b]. Battery-Capacitor [PP08a]. Battle [Gep03a, Che05b, Gar03, Jon05a, San01]. battles [Cal00a]. Bauerlein [Kar09c]. Be [And09e, Cas04d, DiG07c, Kar08e, Tan09, Ade08e, Ade09f, Ano04-87, Boy07, Che09l, Coe04, Cor00, Cos03, Gol02g, Haw07, Hut01, KT08, KA01, Per05h, Rio05a, Sin04b, Zor02d]. Beam [BBS^{+}01, Ade09h]. Bear [Das01]. bearer [QP05]. Beast [Ano07-68]. Beat [Ano04e, Gep00g]. Beating [Boi06, Kus06a]. Beatles [Kus09a]. Beattie [Obe07a]. Beautiful [Ups07f, Ano04f, San03]. Beauty [Ano07-68]. beckons [Pre05a]. become [Wal04]. becomes [App02a, Ros06e]. Becoming [Sel07a]. bed [Wei01]. bedfellows [Hel05d]. Beer [Ano08j]. Beetles [Mul07]. Before [And08a, GP06b, BR02]. Beginning [Ano08-78]. behemoths [Che02h]. Behind [Ano03a, Per05j]. beholder [Ano04-65]. Being [Has08d]. Belgian [Hel05a]. believing [Far09]. Bell [Ano05-52, Kum05b, Rio05a, Rio06b]. bells [Che03c]. Ben [Mul06a]. Bendable [Ano09c, CT05]. Bender [Coh09]. Beneath [Rie01a, App01d]. Beranek [Zim08]. best [Ano04-87, Gol01d, Kar02j]. Bet [Che07a]. bets [Gep03c]. Better [Ano06-54, Cas06h, Fos05b, Moo01c, Per05m, Ano08-92, Dig06a, Gep03b, Kar02b, Kra01a, Mer02, Mul04a, Rie00a]. Betting [Sav02c]. between [Day02, Kle05b, Peh01a, Sch03b]. Beware [Che04f, Che04a, Coe04]. Beyond [Kar08j, Pre07a, Ye08]. biased [Wei03]. bicycles [Fai05a]. Big [Ano03b, Ano06m, Ano06n, Ano06o, Ano06h, Ano07o, Ano07n, Ano07p, Ano07r, Ano07q, Ano07m, Ano07-59, Ano08m, Ano08l, CT05, Che09g, Der08, Hun06, Kus04a, McN07, Moo08a, Mor06, Per09a, Ric07, Sim07, Sin06a, Swe07d, Ter07, Win06, Yeu06, Ano08a, Ano08n, Ano08k, Ano08p, Ano08o, Ano08-36, Ano08-38, Ano09d, Ano09p, Gep05e, Kor06, Obe00a, Per09c, Swe01a, BBS^{+}01, Das01]. Biggest [PP08b, Sin05]. bill [Swe05c, Per05a]. billion [Luc07b]. billionaire [KA01]. bind [Ano08-89]. Bingaman [BBS^{+}01]. biodegradable [Ano02e]. Bioethics [FWC03]. biofeedback [Moo06b]. biography [Kar03e, Per00c, Sav01]. Biological [Ast01, Kum03a]. Biology [Car01, Sar06]. biometric [Gui06a, JP06, Ros05b]. Biometrics [Ano04g]. Bionic [Ano07o, Sar06]. bioterrorism [Ano02h]. bipolar [Gep03b]. Bird [Col07b, Kum04c]. birds [Kum04c]. Birthday [Mur07]. Biscuit [Col05]. bite [Ano04-48]. Bits [Win06, Luc05c, Smi05]. Biz [Voe09c]. Black [Ges04, Jon04b, Per06b]. blacken [Per05j]. Blackout [Gol07c, Jon03a, Ano03b]. Blackout-Proof [Gol07c]. blackouts [BR02]. blade [Wri05]. Blades [Wri05]. Blah [McF03f]. Blast [Ano07q, Fos07a]. Bless [Kar06b, Kar06c]. blessing [Moo04a]. blind [Jon04e]. blink [Cas03g]. blinking [Luc08a]. Blix [Swe08a]. block [Cas05d, Gep01c, Hei09]. blockbuster [Jon00b]. blocking [Amb01c]. Blog [McF03f, Che03a]. Blogging [Ano07-53]. Bloggy [Ano07-56]. Blood [Jon06a]. Bloom [Ups08d, Ano09p]. blower [Kum04f]. Blowing [Swe04e, Das05]. blows [Fai03c]. Blu [Che08a]. Blu-ray [Che08a]. Blue [Pre07a, Ano03a, Mul04a, Per02b, Zor02b]. Blues [Gol07a, Luc07e, Moo05c]. Bluetooth [Sch00, Che08f, McF05a]. Blundell [Ade09c]. board [Bal00, EK00, Mak01]. Bob [Kar08a]. Bodies [Swe00a]. Body [Ros01a, GG05, Jon06i, Web05]. Boeing [Jon04c, Jon05a]. Bogus [PT06]. Bomb [Abr03, Kar07b, Ram00, Zor08a, Bro02a]. bombs [Ano08j]. Bons [Ano05-57]. Book [Ano03f, Bre01a, Bre01d, Bre01m, Cas01d, Cas01e, Cas01f, Cas01i, Cas02c, Cas02d, Cas02g, Cas03b, Cas03g, Cas03e, Cas03h, Cas03k, Cas03m, Cas04a, Cas04b, Cas06m, Che01a, Che01g, Che01b, Che02f, Che03h, Dum08a, Dur03, DR00b, Fea09, Fos01a, Fos02a, Fos02b, Gep01b, Ges04, Gol01b, Gol01e, Gol01d, Gol01c, Ham02, Hoo02, Kra01b, Kum01a, Kus02a, Lew01, Mak02, Mar00, Moo01a, O’N03, Obe02a, Obe02b, Obe03a, Par03a, Rie01i, Rio02, Ros01a, Ros01b, Rot01a, Sch01, Swe03c, Swe04a, Ull03, Zor01a, Zor01b, Dur05, Gep01a, Hoo05b, Luc09d, Ros00, Swe00a]. Book-Scanning [Dum08a]. Books [Ano01-50, Ano04h, Cas01d, Cas01e, Cas01l, Cas04c, Cas05b, Cas05c, Cas06b, Cas06c, Cas07a, Cas07b, Cas07c, Cer00, Che00a, Che05c, Che08b, Eas00, Ens00, Fos00b, Fos08a, Gep01a, Gep01b, Gli00, Har00, Has08c, Kar08k, Lev08, Lou00, Mor01a, Mur09, Nah01, Obe08, Pfl00, Pos00, Raj08, Ram00, Rio06a, Rio07a, Ros00, Roy00, Sch08c, Sni00, Son08b, Swe00c, Swe00d, Swe00a, Zim00, Zim08, ZR07, And08a, Ano00-56, Ano00-57, Ano00-58, Ano00-59, Ano00-60, Ano00-61, Ano00-62, Ano00-63, Ano00-64, Ano00-65, Ano00-66, Ano00-67, Che01c, Fos08b, Swe08a]. Boom [Ano04g, Kat09, NH05]. Boost [Ano03l, And08g]. Boosting [Kar03f]. borders [Ban02]. Borrow [Mul07]. borrowing [Sar06]. Boss [Luk09, Sel07a, Kar08k]. Boston [Bre00a]. bot [Ano08-63]. Bots [SBB08, GG05]. botsourcing [Ano04-61]. bottom [Gol05d]. boundary [Kle05b]. Bowl [Ano08p]. box [Mer02, SR02]. boxes [Jon04b]. BP [Fai03a]. bracadabra [Ano09-38]. brace [Rie01f]. Brain [Ano04-71, Haw07, Kar09k, Pat09c, Pec08, Sar06, Ade08f, Ano03o, Coh09, FWC03, Jon05c, Jon09f, MG00b, Moo02a, Moo06h, Pre04, Ros04f]. Brain-Machine [Pat09c, Pec08]. brains [Gol01b]. Brake [Man08]. Brake-by-Wire [Man08]. Brave [Mcf09a]. Brazil [Gui06e, Rie02a, Rie03b]. Breaches [PP08d]. Break [And09b, Ups08b]. Breakdown [Bar04a]. Breaking [Ano09m, Das08a, Zor02b]. breaks [Ano04-82]. Breakthrough [McN07]. breakthroughs [CG04]. Breast [Ano08-91, Moo01c, She02]. Breathe [Ano09n]. breed [Ack09b]. breeding [Moo06a]. Brew [Gui09a, Kar09g]. brews [LL05b]. Bridge [Kol04, Cas01d]. bridges [Gep01d, Pat09e]. Bridging [Peh01a]. Brief [Ano08-69, Ano08-67, Ano08-60, Ano08-65, Ano08-68, Ano08-71, Ano08-70, Ano08-74]. briefs [Ano06-51]. Bright [Ups09a, Cra00, Moo09d, Wal09b]. brightens [Cos05b]. brighter [Mul05c]. Brilliance [Cas03d]. Brilliant [Rio07a]. Bringing [And08a, Ano04-85, Bro02c]. brink [Moo05a]. Britain [Mac01, Ano09o, Gui06a, Mul01a]. Broadband [Che03c, Che03b, Che05d, Che08g, Ros06c, Che05b, Che05l, Che06a, Che06e, Coo02, Jon04a, Jon05b, JB05, Kum06c, Rie03e, Roy00, Sch09e, SR02]. broadcast [Ano02k]. broadcasting [Bla05b]. Broader [Che05d]. broil [Mul05b]. broken [San03]. Bronx [Jon05b].

Brooklyn [Cas03e]. Brooks [Bro08]. brother [Kus04a]. Brothers [Ano05c]. Brown [Kar08j]. browser [Bre01c]. browsers [Ack09b]. Bruno [Zor08b]. Bruzelius [Per09h]. BTUs [GS07]. Bubble [Gui06b, LTN05]. Bubbles [GP06b]. Buckyballs [And08g]. Budget [Bar04a, Bar04b, Ros08b, Wal09a, Ano01n]. buffet [Ano05-53]. Bug [Ull03]. Bugged [Gui05b]. bugs [Ros05a]. Build [Ano03l, Che07c, Fai03b, Jon07b, Wal07a, Sch09d]. Building [Ell01, Fos05b, Jon02a, Kat09, Pas01, Per04a, RM03, Zak05a]. buildings [Sno03]. builds [Bre00a]. Bulb [Per05m]. bullet [Hol03]. Buran [Obe08]. Burlington [Cas05b]. Bursting [Ano01m, GP06b]. Bury [Gol09b, Fai09b]. Busby [Kar08b]. buses [Jon00d]. Bush [Ano01n, Lan03b, Sia01, Sop03, Swe04b, Swe06d]. Business [Che09g, Dur05, Fra09, Mus09, BC02, Cas03h, Che02d, Hoo02, Kar02j, LL05a, Luc00b, Per01e, Sin04b]. Business-Method [Fra09]. bust [Ano03h]. busy [Che02b]. Butterfly [Mul06b]. Buttons [Che07b]. Buy [Cas06m, Sin06a]. By-wire [Bre01b]. bye [Ano09-40]. bypass [Che06e].

C [Dur05, Sch08c]. C. [Fos08b]. CA [Cas02d]. Cabinet [Fea09]. Cable [Bre00b, Cic01, IS01, Jon03a, Jon03b, Che02h, Jon06e]. cables [Man00]. cacophony [Fra05a]. calculations [CGC04]. calculator [Kus09b]. Calendar [Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano00y, Ano00z, Ano00-27]. Calibration [KP00]. California [ISV01, Fai07a, Gol09a]. Call [GS07, McF05b, Kum04b, Kus03f, Swe00b]. Calling [Ano04-95, MM06]. calls [Gui05b]. Calm [Moo06b]. Cambridge [Cas01l]. Camera [Ano07-59, Rag09, Ano08-92, Gei03b]. Cameras [Gwy00]. Camilleri [Ups08f]. Can [Ano05-47, Ano06-60, Ano08-59, Ber02, Cas04d, Fai04a, Fai08a, Fra02b, KT08, Per07b, Ano02g, Che05h, Gol02c, Kri06, Sel04b, Sul05]. Canada [Cas05b, Chi02a]. canaries [Ast01]. Canceled [Ano08p]. Cancer [Ano08-91, Hel05c, Jon08b, Moo01c, She02, Ups08i]. Candy [Ano08m]. Canned [Zor01c]. Canon [Ups07f, Boy04a]. Can’t [Cas03c, Haw07, Gep05d, Mue02]. Cap [Cha09a]. Capacitor [PP08a]. capacitors [CM03]. capital [Mul05b, SW05]. capitalist [KB01]. capitalists [TS01]. Caps [Bla07]. capture [Boy03b, Kar08o]. Capturing [Per02a]. Car [Ano06g, Ano07r, Bla08, Voe08d, Ano02f, Ano04-68, Ber02, Bla02a, Bre01c, Hoo05b, Jon04b, Jon05f]. Carbon [Ano04-32, Ano09p, Bla07, Cha07a, Gui06c, PP08f, Rom07, Ros04c, Avo04]. Carbon-crunching [Ros04c]. Carbon-Nanotube [PP08f]. card [Ano00b, Boy05]. cards [Gui06a]. care [Aar05b, Ros04d]. Career [Ano06c, PP08b, Che06c, Kar05a, Kum03e, Pre05c, Sel06d]. Careers [App01b, App01c, Cas07d, Cos06, Cos07, Kar01a, Kar03f, Kar04a, Kar06a, Kar06b, Kar06c, Kar07b, Kar07a, Kar08b, Kar08c, Kar08d, Kar08e, Kar08g, Kar08f, Kum01b, Kum01e, Kum01h, KA01, Kum04a, Kus03b, Obe06, PP06a, PP07a, PP08b, PP08c, Sel06a, Sel06b, Sel06c, Sel07a, Sel08b, Sel08a, Wu08a, Kar02j, Kar08h, Kar08i, PP09a, Sel08f]. Carlyle [Chi07]. carriers [Ano04-64]. carry [PP06b]. Cars [Bla07, Voe06a, Voe07a, Voe07c, Voe08e, Bre01b, Gep05b, Jon01c, Jon02a, Jon05d, Ros03a, Voe04, Voe05, Voe06c, Voe09b]. Case [Mul06c, Gol05e, Par03c]. Cash [Sin07b, LL04a]. Catalunya [Six05f]. Catastrophe [Fos08b]. catastrophists [Ano04-35]. Catastropic [Swe07a]. catch [Oli02]. Catches [Jon08e]. Catching [Ano06h]. caters [Tes03b]. causing [Hel05c]. Cautionary [SC06a]. Caveat [Sti03]. Caven [Per05h]. cavities [McF05a, Moo01d]. CD [Day02, Swe00a]. CDs [Gei03a]. celestial [Pre05g]. Cell [Fai08c, Swe00b, Bar01, Bla02b, Cur01, Rie03f, Swe08e]. Cellphone [Wes09, BJ02, Boy06c, Gui05f, Ros06e]. Cellphones [Ano07n, Kas02, Moo08c, PP09b, Pre04, Che09d, Gol06a, Jon06h, Sch09i, Zor06a]. Cells [Dum06a, Ups08b, Boy08a, Der09a, Fai09f, Jon09b, PP06b, Rie01c, Rie01e]. Cellular [Ano09q, Rie00a, Sin03a, Boy03b, Che05h, Oli00, Pas01]. Cellulars [BBS^{+}01]. Censorship [And07a, Rie03c]. Center [Obe05]. centers [Kus05a]. Centrino [Gep03c]. Century [Bor01, Neb00, Ede00, Har00, Has09a]. CERN [Ano04-33, Wei07]. Certification [Kar06a]. Ch [Wan02]. chair [Par03a]. chairs [Ano03f]. Challenge [JN09, Cas03i, Swe09c]. challenges [Rei02]. chameleon [Ver03]. chance [Ano07-56, Kar03c, Lan02]. Chances [Ano08-96]. Change [Mor07, Swe07a, Ano08-34, Ano09-69, Gep03e, Per02a, Ups08d]. Changes [BBS^{+}01]. Changing [Boy06a, McF06a, Sim01]. Chaos [NB07]. Chaotic [Mul06d]. Charge [Sch07, Yeu06, Fai05a]. charged [Zor05]. Charpak [Zor01a]. chart [Hel02a]. Chase [Wei09b, Jon09a, Rie03d]. Chasing [Ano07f]. Cheap [Bla06, Das09b, Moo06c, Per08j, Bro05, Che02i, FS04, Jon04d, Mul05a, Per09a]. Cheaper [Pre05a, Sav08b, Che02i, Gep05a, Mul05c]. cheats [Jon09a]. check [Gol01c]. Checking [Gol02a]. checks [App02b]. cheese [Ano09-75]. Cheney [BBS^{+}01]. Chernobyl [Rie03a]. chess [Cas01d, Ros03c, Ros04e]. Cheswick [Che05i]. Cheyenne [Obe05]. chiefs [McF02c]. Chile [Ano06e]. Chill [Bal09]. Chilling [Kum03f]. China [Ano02a, Ano04-47, Ano05-29, Ano05-40, Ano05-41, Ano05-44, Ano05-51, Ano05-69, Che05j, Cos05c, Dig07a, Fai05a, Fai08d, FZ03, Gep05d, Hoo02, Hoo05c, KSC^{+}05, LL02b, LL03a, LL03c, LL05b, LL06, NH05, Obe09b, Per05c, Swe01a]. Chinese [Obe01a, Obe01b, Ano04-66, LL02c, LL03b, LL04a, LL05a]. Chip [Ade08a, Ade09a, Ano04-34, Kum00a, Ros08a, ST05, Web04, Gep05e, Gep05d, Gol02e, Moo01d, Moo06f, Sav09b, Sin06a]. Chipmaking [Sav03a]. Chips [Bal04, Che09e, Rom07, San07, And09d, Ano04-46, GP00, Gep01c, LL02a, LL02c, Moo01f, Moo06c, Moo09b]. Choice [PP08b]. choose [Kar05b]. Choosing [Day02, Kum01b]. Chris [Ano07l]. Christensen [Dur05]. Chronic [Ano08-90]. Church [Mul08]. Ciencia [Six05f]. cinema [Lub00]. circuit [Bal00, Fai08b]. circuitry [Gra00]. Circuits [Fai03b, Jon08d, CT05, Gep00b]. Cisco [Che06b, LL03a, Tei01]. cities [Gol01b]. City [Cas07e, Che07c, Fai07c, Gol07c, Moo07d, Ups07e, Jon00d, Mul06e]. civil [Sin02]. Claim [Moo09c]. Clair [Ano05-56]. Clapp [Kar08l]. clarity [Den01]. Clashing [Ano04-35]. class [BM05, JA05]. classroom [Fos02a]. Claude [Gui05c]. Clean [Swe08f, Ano05-40, Kel01]. cleans [Sop03]. Clear [Bre02b, Kle05b]. clearing [Pat09b]. Cliche [Mcf08b]. click [Svo06a]. Clicks [McF06f]. Climate [Bar06d, Gai07, McF06a, Mor07, Swe06c, Swe07a, Ups08d, Ano04-35, Per02a]. Climate-change [Ups08d]. clock [Gep05a, Kum05b]. clocks [Per00f, Zor02e]. Close [Ade08h, Pav07, Bro02a, Gol04e]. Close-Ups [Pav07]. Closed [Ano04-75]. Closed-Source [Ano04-75]. Closer [Kim05]. Closing [Fai08b, Gui04a]. Cloud [Che09b, Luc09a, Ros09a, WOS09, Bro03b, Mcf08f]. clouds [Ros08c]. Club [Wu08a, Sch08b]. CO [ES04, Fai09b, Moo01d]. Coal [Jon07e, Swe08f, Fai09e]. coalition [Sin04b].

Code [Ano07-55, Has08b, Gui04a, Kle05a, Mul02a, Ros00]. Codebuster [Ano04-36]. Codemaker [Per07a]. codes [Gui04a, HBB04, Min03]. coffin [Rie03a]. Cogeneration [PP09c]. Cold [Kum03b, Moo07a, Mul04b, Ups07h]. Collaboration [Gol03a]. collapse [Mor01a]. collar [Fos01a]. collector [PP05b]. Collide [Kar09m]. Colon [Jon08b]. color [Gui04b]. colors [Ano09-71]. Columbia [Cas04a, Obe03b]. com [McF02b]. Combat [Ano08-90, Kum02a]. Come [Gol09b, HBW02, Sch09d]. comeback [Amb01a]. Comedy [Wu08a]. Comes [Man08, Sch08c, Boy05, Fai09b, Mar01]. comfort [Gol04e]. Coming [Ano09r, Das03, Gui04b, Che05h, Lub00, Per09f]. commander [Obe05]. Commandments [Ros03b]. Commentary [Che07b, Per07b, Ano08-61]. Commerce [Cha07a]. Commercial [Obe09a, Obe04]. Commercializing [Moo07b]. Committees [Sel06b]. common [CGC04]. commons [Luc06e]. communication [Eas00, Kar01c, Mul06d]. Communications [Luc01b, Obe09a, Rie00b, Ano02s, BOT06, Gol06f, Jac05, Sav05c]. community [Aar06a, Ano03b]. companies [Gol01d, Rio02, Ros04b, Tes03c]. Company [Sin03a, Che08b]. Competition [Ano07-59]. Competitiveness [Bar06a]. complex [Moo01b]. composites [Gui06c]. comprehensive [Zim00]. computation [Fog00, Moo01b]. compute [Kle05b]. Computer [Ano07-50, Ano08s, Cos02a, Haw07, Luc07c, Luc09b, Moo08a, Ros09b, WOS09, Cas01b, Cas02e, Che05i, Che09l, Der09c, Ens00, Kar08l, Kum03c, Kum06d, Kus03a, Kus03d, Mcf08f, Mor00, Per01b, Per01a, Pre05d, RWZ05, Ros03c, Ros04e, Ros08c, Wal01a]. computers [Fos02a, Heg09a, Pes02, Wal02b]. Computing [Ano08-37, BBS^{+}01, Che09b, Com00a, HM04, Ros09a, And09d, Cas05d, GG06a, Hel04b, Luc09a, Min06b, Mul01b, Sar06, Sin06d, Wol02a]. con [GG02]. concentrate [PP05b]. concern [Jon05b]. concerns [Mil01]. Concorde [Hoo00]. Concussion [Jon08f]. condenser [Swe06a]. conductors [Bre04b]. confessions [Obe02a]. conflict [Sel05b]. confront [PT06]. Congress [Peh01b, Sia01]. Congressional [Sam03]. connect [Ano04-66]. connected [Gol01b, Pat09k]. Connecticut [Swe08e]. connection [DY01, Kum05c, Kum06c, MM06]. connectivity [Che06a]. Connects [Sin06b]. Conquest [Cas05b]. conscious [KT08]. consciousness [Hor08]. conservation [Che05i, Swe06b]. Considered [Cas04b]. Construction [Jon08g]. Consumer [Jon01a, Per00a, Per08f, Gep05d, Gol05d, Pre05f]. consumers [Boy03b, Eng01, PT06]. consumption [Per01b]. Contact [Gep04c, Gol02c]. Contain [Dig07d]. Content [GR09, Svo06a]. Contents [Ano04x, Ano05r, Ano05w, Ano06r, Ano06t, Ano06w, Ano07y, Ano07u, Ano08t, Ano09s, Ano09t, Ano09u, Ano00-68, Ano00-69, Ano00-70, Ano00-71, Ano00-72, Ano00-73, Ano00-74, Ano00-75, Ano00-76, Ano00-77, Ano00-78, Ano00-79, Ano01-52, Ano01-53, Ano01-54, Ano01-55, Ano01-56, Ano01-57, Ano01-58, Ano01-59, Ano01-60, Ano01-61, Ano01-62, Ano04j, Ano04k, Ano04t, Ano04r, Ano04w, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-93, Ano05z, Ano05-27, Ano05-28, Ano05-67, Ano05-68, Ano05-66, Ano05-62, Ano05-63, Ano05-64, Ano05-65, Ano06-30, Ano06-63, Ano06-65, Ano06-66, Ano06-67, Ano06-68, Ano06-69, Ano06-70, Ano06-64, Ano07v, Ano07w, Ano07-60, Ano07-61, Ano07-64, Ano07-65, Ano07-66, Ano07-67, Ano07-62, Ano07-63, Ano08q, Ano08r, Ano08-80, Ano08-81, Ano08-82, Ano08-83, Ano08-84, Ano08-85, Ano08-86, Ano08-87]. contents [Ano08-88, Ano09-77, Ano09-78, Ano09-79, Ano09-80, Ano09-81, Ano09-82, Ano09-83, Ano09-84, Ano09-85]. Continue [Fai08a, Sin04b]. Contractors [Gui08a]. contraptors [Gui08e]. Contributors [Ano01o, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano03c, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano04-44, Ano04-45, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano08-28, Ano08-29, Ano08-30, Ano08-31, Ano08-32, Ano09v, Ano09w, Ano09x, Ano09y, Ano09z, Ano09-27, Ano09-28, Ano09-29, Ano09-30, Ano09-31, Ano09-32, Ano09-33]. Control [Ano05h, Wei02a, Boy06d, Bre01l, Bre01n, Bro05, DG00, Gai07, Luc06d]. Controlled [NB07]. Controller [Ano07-59]. Controlling [Fos01b]. controls [Gei03b, Gei03a, SK00]. controversy [Jon03b]. conundrum [Hor08]. converge [Ano04-84, Dig04, Sia01]. convergence [Per04g]. conversions [Sel01c]. Cooking [DiG07c]. cool [Boy03a, Cas01e, Ros03a]. coolest [Der09c]. Copernicus [Gep01b]. Copia [Swe07b]. cops [Kum02c]. copy [Avi02b, Gei03a, Per05e]. Copying [Obe08, Bla03a, Gei03b, Sam03]. copyright [Ano01-63, Che02c, Kus06a, Kus06e, Per03b, Ree03, TP06]. copyright-infringement [TP06]. copyrighted [Kus03d]. Core [Amb01a, Ric07]. Corn [Swe07b]. Corn-O-Copia [Swe07b]. corner [Bre01b]. Corporate [Che09a]. Corporation [Mue02]. corral [Sam03]. correcting [Gra00]. cosmos [HM04]. cost [Che06g, DR00a]. costlier [Kar02b]. costly [Swe02]. Could [And09e, Dum08b, Hel05a, Min03, Min06a, Mul05a, Mul07, Obe09b, Ano04-79, Ano09-73, Cos03, Jon09b, Mul05c, PP06b, Swe02]. Count [Che07d]. Countdown [Cla05]. counteraction [Ano02h]. counterfeit [PT06]. Counterfeiting [Ano06-53].

Countering [Zor08c]. counterparts [Ano04-66]. counterterrorist [Vat01]. Counting [Boy03a]. country [NH05, Sop03]. Course [Sni00, Kra02]. Courseware [BBS^{+}01]. Court [BSK^{+}01]. cover [Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01-28, Ano01-29, Ano01-30, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano04-62, Ano04m, Ano04n, Ano04o, Ano04p, Ano05-39, Ano05s, Ano05t, Ano05u, Ano07-44, Ano07-43, Ano07-38, Ano07-46, Ano07-39, Ano07-45, Ano07-40, Ano07-41, Ano07-48, Ano07-47, Ano07-42, Ano08-58, Ano08-57, Ano08-49, Ano08-50, Ano08-51, Ano08-52, Ano08-53, Ano08-59, Ano08-54, Ano08-55, Ano08-56, Ano08-48, Ano09-53, Ano09-54, Ano09-55, Ano09-56, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k]. cows [Ano09-66]. CPU [And09e, BD09, Ros08d]. crack [Min03]. cracker [Cas03e]. Cracking [Hsu07]. cracks [Ano02b]. Crafting [Kro07]. Craig [Kum05a]. cranks [Boy04c]. Crash [Gui07b, Sho05a, Ano02s, Bla02a, Jon04b]. Crash-Test [Gui07b]. Crashes [Dum08b]. crashing [Jon01c]. Crashproof [HBB04]. Crazy [Bar06a, TJ00]. Create [Sin03a]. created [Cas03k]. Creates [Ade07b, Hel03]. Creating [Che09e]. Creation [And08a]. Creative [Bla02a, Sel04b]. credit [Boy05]. Crewless [Sel01b]. Crime [SC07]. Crimeware [And08f]. Crippling [Obe07a]. Crisis [Ano04c, Gui06a, Sta01, Swe03a, Wal04]. Critical [Rei02, Kum01a, Swe04f]. Critics [Ade08i]. critters [Ros04c]. cross [Cas06g, Zor02g]. cross-pollinate [Cas06g]. crossed [DGKM05, Ham02]. Crosshairs [Swe07e]. Crossroads [LL02a, Swe04a, Cho00]. crown [Per02e]. CRT [O’D06]. Cruiser [Ano05-42]. Crunch [Ano01-32, Boi06]. crunching [Ros04c].

Crusade [Per07c, Cas02d]. Cryptographers [Heg09a]. Cryptography [Das08a, Ros00]. Crystal [Boh06, Hel02c]. Crystals [BBS^{+}01]. Cuddling [Fer01]. Cult [Min09]. Culture [Ano06g, Cas03k]. curb [Ano05-40, Tes03c]. Cure [Gol07a, JL04]. Curiosities [Fea09]. current [Par03a, Sta01]. curtail [Che05j]. Curve [Kar09l]. Cyber [Sho05b, Kum01i]. Cybercrime [Ano06-34]. cybersecurity [Far03]. cyberspace [Cas03e]. cyberterrorism [Che05m]. Cyc [Wal08a]. cycling [Mcf08a]. cyclists [Fai05a]. Czar [Voe08a].

D [And08c, Fos08a, Gep01a, Hir04, Kar08k, Obe07a, Bal04, Hel02b]. D.C. [BSK^{+}01]. dad [Che05i]. Daintiest [LB04]. Daisey [Kar08p]. Dale [Per05b]. Dam [And09a]. Damage [Jon08f]. damn [Che06b]. dams [Swe01a]. Dance [Ano08-33, Ano09-35]. Dancer [Kim05]. dancing [Gep04g]. Danes [Moo03a]. danger [FPP00]. Dangerous [Raj08, Swe03c]. Danish [Swe04c]. Dares [Sel06c]. dark [Luc01e, Ste09, Svo06b]. darkness [Luc03a]. DARPATech [Ano08g]. dash [Per05c]. Data [Ade08b, BK09, Cha06b, PP08d, Ups07a, Ade08c, Ade09a, Ano04-86, Bre04a, Che08c, Che09l, Dum09b, Gep04f, Jon04b, Kus05a, Moo01e, Moo08c, PP09c, Pat09d, Sav03c, Swe08d, Tes03b, Ups08b]. Database [Har00, Moo09c]. Dave [And08c, Per04b]. David [Sni00, Che03j, Per08a]. Dawn [Abr03, FBT00, Hoo02, Ras05, Sch00]. day [Hoo00]. Days [Sim07]. DC [Cas01f]. Dead [Ano07-55, Che03i, Hei02, DiG03, Dig06c, Mul04b]. deadline [Per05c]. deaf [Sel01e]. deal [Fra04, Obe05]. Dealing [Kar02a, Sel07b, Sel08c]. Death [Fra09, vS06, Har00]. deaths [Ano02j]. Debacle [Che06d]. Dec. [Ano04j, Ano06s]. December [Ano05q, Ano05r]. decoys [Kus03c]. decree [Sel01d]. decryption [Che01e]. Deep [Sul05, Ano03a, Gui05e, Hoo05c, Wal00]. Deeply [Wal07b, Ano09n]. Defeat [Ano08-59]. defend [LL04b]. Defense [BBS^{+}01, Fos08a, Gui08a, Kum06d, Sho05c, Sin07a, Kar03a]. defibrillator [KKG08]. Defragmenting [Rob00]. degree [Kar04b]. Delayed [Bre02c]. delegating [Sel05a]. Delhi [Sin07a]. Delicate [Dum08a]. Deliverance [Hoo07]. Delivers [Ano05g]. Dell [Son08c]. Delusional [Ros07]. Delving [Ven02]. Demand [Cas07j, Amb01c, Kar03g]. demands [Kus05a, Tes03b]. dementia [Ano09-73]. Demilitarized [Kum07d]. demonstrates [Ano04-60]. Denee [Kar08b]. denial [Com01b]. Dennis [Per04c]. Department [Ade09d]. Deployment [Sho05a]. Deploys [Sho05c]. Depression [Ups08g, Moo05c]. Derby [Eyr01]. deregulation [Jan01, Lan02, SB02]. desalination [Ups09f]. Design [Ade08a, Ano08s, Cas07e, Fos08a, Joh07, Van07, Ade09a, Ead00, Gep00c, Gep01d, LW00, Sin06a, Wei01]. Designs [Lew01]. desktop [Gui06d, PC05]. Despite [Pat09e, Fai05a]. destruction [Bla02a, Che01b]. detection [Ast01, Bel06, Moo01c, She02, WMM03]. detectives [Kum00a]. Detectors [Fos05b, Wei07]. deter [Kus03c]. Deterrent [Ano09l]. Development [Ano06-61, Kar07a, Pfl00, Ens00, Fra05b, Zac05]. Develops [Ano08-73]. device [Hel03, Moo05c, Moo06b, Per09g]. Devices [Gep00b, Ups08g, Cas01h, Jon05e, Sav05c, TS03, Ver03]. Devil [Ano08-35]. devise [Hel01b]. diagram [Ano08-97]. Dial [Che09i, Kar01c]. diary [Zor08a]. diboride [Das02]. did [Kar02f]. Dielectrics [Ade08k]. diet [Che09h]. Difference [Ano04-89, ZR07, Oli00, Par02b]. differences [Ano05-70]. Different [Pol06, Kri06]. Difficult [Sel08c, Ups08g]. diffuse [Wal04]. Dig [Ano05-47]. Digit [Fai08a, Cos05c]. Digital [Ano01-27, Ano09-72, Che05e, Dum06b, Fox01, Gwy00, Has08b, Hoo07, Kar09c, Kus03c, Lay01, LL05b, Luc06a, Nev01, Per05c, Per08b, Put03, Swe04d, Wal02a, Ano01-63, Ano08-74, Ano09o, App01a, Che00a, Che01f, Cla05, Eyr01, Lub00, Mar00, Per01d, Per05k, Ras05, Ree03, Rio02, Sam03, Smi05, Spr02, Ven02, Wei02b, Zor02e, And08c]. Digitize [Dum08a]. Dilbert [Kar04a]. Dilemma [Dur05, Kum04f, Per08b]. Dimension [Cas07h]. diminish [Ano09-73]. Dinosaur [Ano09-34]. diode [BC05, Hel05a]. diodes [Mul04a, Mul06b, Per03c]. Dioxide [Rio07b]. director [Ade09d]. Directory [Swe00a]. Dirty [Kus07a]. Disappearing [Ano09l]. disarm [Kum02d]. Disarmament [Swe08a]. disarming [Swe08a]. Disaster [Ano05-59, Kum07b, Cas03f, Kan02, Kum01j, Obe03b, Ros01b]. Disasters [Fos08b]. Discarded [Ano07n]. Discerning [Wei07]. Disciplined [Kum01a]. discontents [MS09]. discovery [Shr09]. Discs [Boy06b]. discussion [Ano01-47]. Disease [Dig07d]. diseases [SNB06]. disk [Das03, Hug02, Ros06e]. disorder [Luc03e]. Display [Boy04b, Ano09-61, Per06b]. Displays [Pol06, Ano04-65, Boy04d, Cra00, Hel00, Hel02c, Mul05c, Sul05]. disposable [Rie01b]. dispute [Cas02f]. disputed [Sin05]. Distance [RH06]. distributed [Swe01c]. distribution [Che01f]. ditches [Fai03a]. Dithers [Dum06b]. diversity [Kum01g]. Divide [Ano01-27, Ano08-74, Che04g, Peh01a, Ven02]. divided [CGC04]. divisiveness [Spr02]. DIY [Ano07-59, Rag09]. DMCA [vS06]. DNA [Mcf09b, Pre05c]. Do [Ano04-86, Ano07-51, Fra03, GP06a, Loe06, And08c, Jon09g, Luc02c, Pat09i, Sel04b, Wal07c]. Do-It-Yourself [Ano07-51, Loe06]. Doc [RH06]. Doctoring [Hel06b]. doctrine [Fra02c]. DOD [Ano08-34]. Dodgeball [Ano05d]. dodger [Dix05]. Does [Ano06-54, Luc04a, Den01, JB05, Wal04]. dog [Gol04a, Kus02a]. Dogs [Ups08e]. D’oh [Fra03]. Doing [Ano06d, Per06c, And08c, Che08d, Kus02a, PP08e]. Dollar [Ano06-50, BBKV09, Kar02d]. Dollars [And08b, Kum06d]. domain [Che01l, Min06a]. domination [Jon06d]. Done [Kar09e]. Donovan [Kum05b]. Don’t [Ano01-49, Cor00, Kar09c, Kle05b, Sel03]. Don’ts [Sel08e]. Doodlemeister [Kar09b]. doom [Cas03k]. Doomed [Ups08d]. Doomsday [Ano08-68]. door [Cas06a, Hel02d]. doors [Kar08d]. Doppelganger [Ano09-35]. Dot [Van07, McF02b]. dot-com [McF02b]. Dot-To-Dot [Van07]. Double [Ano03k, Col07a, Fai08a, Cos05c, Gep02b, Mac08]. Double-Digit [Fai08a, Cos05c]. Double-Edged [Ano03k]. Doubles [Fai08d]. dough [Bro02c]. Down [Che09a, FH01, Gol05a, Kum08b, Luc03a, Ano01-51, Ano02b, Heg08a, Jon09a]. Downey [Per08a]. download [Boy04c]. Downloading [McD04]. dowry [LL03c]. drag [WU01]. dragons [Coe04]. DRAM [Moo04b, CM09]. draw [Ano02a]. drawerful [Fea09]. Dreaded [Luc09b]. dreadlocks [Ano09-64]. Dream [Ano07f, Ano09-36, Che06c, Per04d, Per05d, Rot01b, Ada08, Ade08h, Ano08-62, Ano08-72, K.08, Kar08j, Kum08a, Luc05a, P.08, Per08a, Sav03b, Ups08f, Zor08b]. Dreamin [Gol09a]. DreamJobs [GKJ^{+}07]. Dreams [Ano07a, Per04c]. Drilling [Zor02c, Swe05d]. drive [Gol02d, Hug02, Jon05d, Luc08f, Wol03]. Drives [Boy08b, Das03, Hug02, MO00, Rie01a]. Driving [Ano08-34, Fai09a, Ano02f, Dig06c, NH05]. DRM [Per07b]. Dropping [Ano09-37]. Dross [Ano07h]. drought [Sin04a]. Drowning [Kum06a]. drumsticks [Ano09-37]. DSL [Ano01-49, Boy04c].

Duane [Obe05]. dubious [Hel02a]. Dude [Mor06]. Duel [Boy06b]. duke [Com00b]. dullard [Che05e]. Dumb [Fos08a, Kar09c, Pat09e]. Dumber [Kar09c]. Dumbest [Kar09c]. Dummy [Gui07b]. Dumpling [Ano09-76]. Dust [Ano08-35]. Dust-Devil [Ano08-35]. Dustup [Gui07a]. DVD [Day02, Gei03b, Chi02b, Hoo05b, Per05e]. DVD-audio [Day02]. Dying [Cha06b]. Dynamo [Ano08-35]. dynamos [LB04].

E-Bomb [Abr03]. E-bracadabra [Ano09-38]. E-cycling [Mcf08a]. e-junkies [LL06]. e-learning [Ube00]. E-mail [Li01, Che01h, Luc04c, Ros06d, Smi03, VN03]. E-Newspapers [Hoo07]. e-trash [LL04a]. E-Vote [Che07d]. E-Voting [Ano04-88]. E-waste [Mcf08a, Per06e, Ano02b]. E.T. [Dig08a, Ros06a]. Ear [Che07g, Zim08, Sar06]. Early [Enr07, Cal00a, Ros05c]. ears [Gei03a, Jon04e, Sel01e]. Earth [Luc01f, Moo03b, Cas05b, Das01, Dig06b, Dig08b, Hoo00, Luc08d, Mul06c, Ups07h]. Earth-Sensing [Ups07h]. Earth-size [Dig08b]. Earthlings [She04]. Earthquake [BF05]. Earthquakes [Enr07]. easier [Fos01e]. East [KS04]. Easy [Has08d, Jon05g, Sch09b, Swe07d, Coe04, FS04, Jon08b]. eat [Bla02b]. eats [Fai09e]. eBay [Svo04]. eBuyer [Che04a, Che04f]. ecology [McF03d]. economic [NH05, Sin04b]. Economics [CKVY08, Han08]. Economy [Kar03f, Kap02, Sch09c]. Edge [San08, Ros01b, SC06b, Wri05]. Edged [Ano03k]. edging [Rou01]. Edholm [Che04b]. Edison [Par03a, And08a, Cas05b]. education [Amb01a, Cho00, Gui06f, Kar03b, Per06f]. educational [JA05, Neb03a]. Edward [Cas01l]. EE [Cos06, Avi02a, Cos03, Cos05b, Gui08b, Jon00c, Kar08d]. EECOM [Cas03b]. EEE [Ano01-28, Ano01-29, Ano01-30]. EEs [Ano00-30, Ano00-28, Ano00-29, Jon01a, Rie00c, Rie00d, RBJ00a, RBJ00b, Rie00e, Rie00f, Rie00g, RB00a, RB00b]. Effect [Kum03f, Ten09, Che05j, Mul06b]. Effective [Kar01a]. effects [Per04h]. Efficient [PP08g, Bow03, Voe06c]. efforts [Fai05a]. egg [Ano09-34]. Einstein [Nah01]. election [Rie02a]. Electric [Mak02, Moo04a, Neb00, Sch06, Voe06a, Voe08d, Ano04-68, App02a, Bas05, Bre00c, Fai05a, Jon03e, OW01, Par03a, Rie01c, Voe05, Wal07a]. Electrical [Ano04c, Ano04-46, Swe00c, Wal04, Kar02j]. Electricity [Ade08b, And08a, Bas05, ISV01, Jon07g, Mac01, Mur09, Cas03a, Eng01, Jan01, Swe00g, Swe06b]. electrochemical [Bro05]. electrodynamic [Joh00a]. Electromagnetic [Gli00]. Electronic [Ano08-64, Cas01g, CS08a, Gep00c, Cas03j, Cas06a, Che05e, Kum02c, PT06, Per05p, Per06e, Rie02a, Ros05b]. electronics [Ano02e, Ber02, BM05, FBT00, Gol01e, Hel01a, Hel02d, Kap00, KMT00, Luc09d, PT06, Per00a, Per06f, Pre05f, Roo00, Sav05c]. Electrons [Hel05e]. Elemental [Ye08]. Elements [Ano01-31]. Elevated [Ano05b]. elevators [Edw05]. elude [Che01h]. Elusive [Luc08b]. Emanuel [Mor07]. Embedded [San01]. embrace [Rie03c]. embraces [Jon03c]. emerge [KH00]. Emergence [Gol01b]. Emergency [Swe06c, Rie02b]. Emission [IS03]. emissions [Ano05-40]. emitting [Mul06b, Per03c]. emotion [Che05a]. Empire [Ade09b, Cas03k]. Empirical [Ano04-75]. employees [Sch03b]. employer [PP09a]. employment [Ano04-83]. Empty [Che08a]. enabled [Bar01, Voe06b]. Enabler [PP07a]. Enables [Ano08-64]. Encryption [Cal00a, Cal00b]. encyclopedia [Svo06a]. End [Ano03m, Min09, Cla05, Ise03, Rio05a, SW04]. endeavor [Per05j]. ends [Sel02a]. Energiya [Obe08]. Energiya-Buran [Obe08]. Energy [Ade09d, Ano01-32, Ano04-47, Ano08-36, Fai08a, PP08g, SB01, Swe03d, Swe04a, Swe04e, Swe05b, Ups06, Ups08b, Che09k, Fai03c, Fai05b, Fai09d, LTN05, Lan03a, Mai03, Pre05a, Sop03, Swe00f, Swe05c, BBS^{+}01]. Energy-Efficient [PP08g]. Engine [Gui07d, Ups07a, Cas04f, Zak02b]. Engineer [Bor01, GS08, Has09a, Kar08f, Kum05a, Obe06, Wu08a, Cas04g, Gep04b, Gui05c, Kum03e, Kus03a, Obe02a, Pre05b, Sel04b, Son09a, TS01]. Engineering [Ano04c, Ano04-75, Ano04-87, Ano05-56, Bro03b, Cas03a, Ede00, GS08, Has07, Has09c, Hel06a, Jon07i, Kar02h, Kar05a, Kus05a, Kus08a, Luc03b, Luc04b, Luc09c, Ups08d, Ade08f, Amb01a, Ano01n, Cav04, Che06c, GG04, Gui06f, Kar02j, Kum00a, Kus04b, Luc02a, Luc06f, PP05a, PP08c, Per06f, Ros04a, Sel03, Sin03b, Wal04, Zor06b, Swe00c]. engineering/technology [Ros04a]. Engineers [Ano01-77, Gui08f, Jon08c, Kar07c, Kar08g, PP08e, Pat09f, Ube00, Cas01d, Eas00, Kar02d, KB01, Kus03e, Luc08d, Zor06c]. Engines [Gol01d, Cas04e]. enhancement [HNPC05]. Enlightened [Voe06a]. enough [Ano03i, Gol02f]. Enron [Sin05, Tei02]. ensemble [Kum02a]. ensures [KP00]. enterprise [Ano02r]. entertain [Kus04b]. entertainment [Kar02h, Kus04b]. entrepreneurial [App01b]. Environment [Voe08e, Per00b]. environmental [Rie03b]. envision [Ede00]. Eolas [Che04h]. Epileptic [Hug09]. equal [Pre05b]. Equation [De 01, Gui08h]. equipment [Kum01c, Swe05d]. equivalents [Fra02c]. Erlendur [Gui09b]. error [Gra00, Tei01]. error-correcting [Gra00]. errors [Gep04f]. ESSAY [Mer04, Aiz05, Ano04-70, Ano04-71, Ano04-72, Ano04-73, Ano05-42, Ano05-43, Col05, Hel05e, Kim05, Kit05, Kol04, McN05, Nak04, Pie05, She04, Wee05]. essence [Eas00]. establishes [JB05]. Eternal [Smi05]. Ethernet [PP08g, Che05f, Kap01a, Rob03b]. ethical [Cos02b]. ethics [Cas03a, Lou00]. Euclid [Che01a]. Eurasian [Swe08d]. Euro [Mas02]. Europa [Kum07e]. Europe [Ano02b, Ano04-48, BBS^{+}01, Bla03a, Bla03b, Cha04, Dum06b, Fai09b, Fai09c, Rio05b, Wei02a, Zak05a]. European [Bla07, Bre03a, Gol06b, Jan01]. EUV [Das08b, Rou01]. EV [Rie00h]. Even [Ups08b, Rie01f]. Ever [Ano06-60]. EverQuest [Kus05a]. Every [Cha09b, Che07d, Che01c]. Everybody [Sch08c]. Everything [Kum08a]. everywhere [Moc04, Rio05a]. evil [Ano08-93]. evolution [Ros00, Tes03b]. evolutionary [Fog00]. evolves [Kar03b]. EVs [Fai07a]. Exascale [Ano08-37]. Executioner [Par03a]. executive [Ano03g]. exists [Kle05b]. Exit [Obe07b]. exoskeletons [GG05]. expands [Kar03b]. Expect [Pat09g]. expedites [Cos02a]. Experience [Ano01-45, Bre00c]. experiment [Cas01k, Gui06f]. experiments [Fos00b]. Expert [Kar08e, Swe06d, Che05i]. Expertise [Kum01h]. Experts [Pat09g, Ede00]. explains [GG04]. Exploring [Ano07-52, App01d, Cas07h, Sim01]. explosive [Rie01a]. Exposing [Gol01e]. express [Kri06]. expressiveness [Luc04b]. Expressway [GG06b]. Exquisite [Zim08]. extend [Hoo05c]. Extending [Amb01b, Moo02b, Rie03e]. exterminators [Ros05a]. extraordinary [PP05a]. Extreme [Ano07-52, Hut01, Moo03b, Ree03, SNB06]. Extremist [Kar08g]. Eye [Ano01-33, Ano08m, Ano08-69, Ano04a, Ano04-65, Cas03g, Dig08b, Par09, Sch09f]. Eye-Pods [Ano08-69]. Eyeglasses [Jon04d]. Eyes [Cas07i, App01a, Kor06, Tea09].

Fab [MC07, Ano09-39, Rou01]. Faberge [Per05f]. fabrication [Lip05]. Face [Ano07g, Ano08-38, Ano08-93, Jon09e, Rod03]. faces [Per05c]. Fact [Luc01c, Koo03]. Factor [Ano05-59]. Factory [Jon09c]. Facts [Ano08-38]. failings [Mue02]. fails [Cha05]. Failure [Ano05-58, Cas06b, Cha05]. Failures [Hoo00, CCK00, Rie00i]. Fairchild [Mur07, Moo07c]. Fairer [Bri08]. Fairy [Per09i]. fake [Kus03c]. Fall [Ano01b, Ano04-97, GG04, Pat09j]. Falling [Par03b, Per05k]. Fallows [Bre01a]. False [FPP00]. Fame [Far03]. Family [Ade08e, Obe02b]. Famous [Luc06b]. farm [Man02]. farms [Fai09c]. Fast [Ade07b, Bro03a, Nev01, BM05, Che06a]. Faster [Bri08, Hol03, Kar02b]. Fate [Fai07a]. Father [Ano07i, Raj08, Kar04b]. Faults [CCK00, Hoo00, Rie00i]. Favorite [Rob09b]. favors [Wei02a]. FBI [Ano02d, Kum03c]. FCC [Bre00b, Jon03c, Ano02c]. fears [Ray01]. Feb [Ano04k, Ano08-57]. Feb. [Ano07-43]. February [Ano04l, Ano06u, Ano06t]. Feds [Ano02d]. feeling [Che09j, Jon09h]. feels [BSK^{+}01]. feet [Pra01, Rie02c, Roy00]. fell [Hoo00]. Fellows [GP06b, Ano09-65]. fences [DG00]. Festo [Fra02c]. few [App02d, Tie04]. Fi [Ano02q, Che03f, Che03l, Che04n, Che05h, Che06g, LL04b]. Fiber [Che04i, Jon05c, WG01, Bas05, Che09h, DY01, Gui06c, Hoo05c, Tes03c]. fiber-optic [Hoo05c]. Fiction [Cas05b, Joh07, Koo03, Luc01c]. Field [Ano00-43, Dum08b, Mul06c, Fos04a, Jan01, JB05]. fifteen [Cas06g]. Fight [SC07, Avi02b]. Fighting [Ano05-53, Kum01c, Kum02a, SNB06, Sin04a]. Fights [Kra01b, Wal08c]. Fiji [Ano09r]. File [GH07, Ano00-44, Gol05e]. file-sharing [Ano00-44]. files [Kus03c, Smi05]. Filling [McF02a]. fillings [Moo01d]. Film [Ano04-76, Fai03a]. films [Bro03a, Cas06g]. final [Put03]. finally [Moo05a]. find [Zor06c]. Finding [Ano05e, Cas04e, LCS04, Svo04]. Finds [BBS^{+}01, Kar03a]. fine [Che08d, Ups08i]. fine-tunes [Ups08i]. Fingerprint [Moo06d]. fingertips [Mcf09c]. Finland [Ano09-40, Ups09b]. Fire [Mor07, Swe03d, Swe04b, Cur01, Pre05g]. firefighters [Gol02c]. Firefox [Kus06b]. firms [Ano02e, Ano04-66]. First [Ano08g, BBS^{+}01, Der09a, Hel01a, Obe09b, San03, Sni00, Ste08, Eng01, Luc09d, Roy00, Zim00]. firsts [Ens00]. fish [Luc02b]. Fission [Moo07a]. Fix [McF04e, Cas03f]. Fixed [Sin03a]. Fixed-Wireless [Sin03a]. fixing [Moo04b]. fizzle [Ano04-83]. Flames [Luc03c]. Flapper [Has08f]. Flash [Cas03d, And08g, Gol06e, Per04a]. Flat [Bro05, Boy04a, Luc08d, Mul05a]. flesh [Che02a]. Flexible [Jon08d, BM05, CT05, Hel00]. Flight [Bre01n, Dum07b, Bre01a, Gui04d, HBB04, Per00e]. flights [Kra01b]. Florida [Shr01]. flow [Ead00, Jon01b]. Flowers [Ano08-67]. FluidFocus [HK04]. Fly [Ano05-54, Col07b, Has08f, Woo08, Zbi05, Pas01]. Fly-Size [Has08f]. Flying [Ano07r, Gui04c]. Flywheel [HBW02]. foe [Kum01f]. foes [Rie00a]. fog [Luc03a]. foggy [Gol05b]. Folk [McF06b]. follow [Mac01]. food [Gol03c]. Force [Tan09]. Forecast [Ano07-56, Che09b, App03b, Bre00f, Bre00g, Com00a, Com00c, Gep00b, Gep00c, Gui04c, Kap00, Kor00, Pat09b, Per00a, Per00b, Rie00b, Roo00, Sco00, Swe00f]. Forecasting [Ano06-58, Jon01b, App04b, Per05g]. Foreign [Kar09d, Tes07a]. forensics [Kar09a]. Forge [Moo03a]. Forget [Ano01-49]. forgot [Gui05a]. Format [Ano08-78]. Formular [Gep05b]. Fort [Smi03]. Fortier [Gol01c]. fortress [Gui05a]. fortune [Kum03e]. Forum [Ano00-31, Ano00-32, Ano00-33, Ano00-34, Ano00-35, Ano00-36, Ano00-37, Ano00-38, Ano00-39, Ano00-40, Ano00-41, Ano00-42, Ano01-34, Ano01-35, Ano01-36, Ano01-37, Ano01-38, Ano01-39, Ano01-40, Ano01-41, Ano01-42, Ano01-43, Ano01-44, Ano03d, Ano04-49, Ano04-50, Ano04-51, Ano04-52, Ano04-53, Ano04-54, Ano04-55, Ano04-56, Ano04-57, Ano04-58, Ano04-59, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano05-38, Ano06-35, Ano06-36, Ano06-37, Ano06-38, Ano06-39, Ano06-40, Ano06-41, Ano06-42, Ano06-43, Ano06-44, Ano06-45, Ano06-46, Ano07z, Ano07-27, Ano07-28, Ano07-29, Ano07-30, Ano07-31, Ano07-32, Ano07-33, Ano07-34, Ano07-35, Ano07-36, Ano07-37, Ano08-39, Ano08-40, Ano08-41, Ano08-42, Ano08-43, Ano08-44, Ano08-45, Ano08-46, Ano08-47, Ano09-41, Ano09-42, Ano09-43, Ano09-44, Ano09-45, Ano09-46, Ano09-47, Ano09-48, Ano09-49, Ano09-50, Ano09-51, Ano09-52, Fin05, Gol04g, JS08]. Forum [Kre05, Lee05, SOR08, Spr02, Kur08]. Forums [Swe00a]. forward [Moo05c]. Fouling [Ree09]. Found [Wil08, Che02g, Ort09]. foundations [Cas05d]. foundries [LL02c]. Foundry [Chi09]. Fountain [Cas04f]. Four [Kor06]. Fourier [Sni00, Sni00]. fourth [Hir05]. fractious [Sin04b]. fragment [Min06a]. frame [Gui08c]. France [Ort09]. Francois [Gol01c]. Frankenstein [Ano09-59]. Franklin [Mul06a]. Fraud [Pat09g, Kar09a]. Fred [Ano07l]. Free [Ano05a, Bre01n, Per08i, Ups08b, Cas02d, Mcf09e, YDLC08, Bre01a, Cas02d]. freedom [Cas02d, Che05j]. Freight [Man08]. French [Ano04-60, SC06b]. frequencies [Mul02b]. Frequency [Ros08d, RSS00]. Fresh [Sch08a, Svo07]. friend [Kum01f]. Fritz [Che05g]. Front [Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01-28, Ano01-29, Ano01-30, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano04-62, Ano04m, Ano04n, Ano04o, Ano04p, Ano05-39, Ano05s, Ano05t, Ano05u, Ano07-44, Ano07-43, Ano07-38, Ano07-46, Ano07-39, Ano07-45, Ano07-40, Ano07-41, Ano07-48, Ano07-47, Ano07-42, Ano08-58, Ano08-57, Ano08-49, Ano08-50, Ano08-51, Ano08-52, Ano08-53, Ano08-59, Ano08-54, Ano08-55, Ano08-56, Ano08-48, Ano09-53, Ano09-54, Ano09-55, Ano09-56, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k]. Frontier [Ups06, Har02, Kus03a, Moo06c, Put03]. frontiers [Cas03e]. Frozen [Hel06a]. Fruehling [Per04b]. Fruitless [Per09b]. fu [Ano09q]. Fuel [Ano08-60, Dum07a, Jon09b, Rie01c, Der09a, Mil01, PP06b, Per08g, Pre05e, Rie01e, Rie03f, Sch09c, Swe08e, Voe06c]. fuel-economy [Sch09c]. Fueling [Per01b]. fuels [Boy08a]. Fujitsu [Boy04b]. full [Gep01e]. Fun [Ano06-59, Bre02e]. Funding [Bar04b]. further [Ano00-84, Ano01-65, Ano01-66, Ano01-67, Ano01-68, Ano01-69, Ano01-70, Ano01-71, Ano01-72, Ano01-73, Ano01-74, Ano01-75, Ano01-76, Ano04-96]. fuse [Ano04-46, FB00]. Fusion [Gui06b, Jon09c, Wal09a, Mul04b, Swe04g]. Fuss [BBS^{+}01]. Futility [Ast02]. Future [Ano06-56, Fos08a, Has08b, Joh07, Kar09c, Luc08b, Moo09b, MC07, App01c, BC02, Boy06c, Gol01d, Has09c, Luc02a, Obe02a, Per04h, ST05, Wal09b, Ree09]. futuristic [Bla04].

G [Kar09k, Che03l]. GaAsing [Gol06a]. Gab [Raz08]. Gadgeteer [Kus03b]. Gadgets [Raz08, Cas01e, Cas01e]. Gage [Per04f]. Gagnon [Gui05c]. gain [And09d]. gains [Gui04e]. Galileo [Swe08c]. gallium [Gep04a, Zor02d]. Gambit [Chi07, Raj08]. Gamble [Che08g, Gep04a]. Game [Ano06-61, Ano08s, Che09e, McF07c, Per03a, Raj08, Swe07c, Bla03b, Cas01e, Cas02e, Gui08c, Kus06a, Mak01, Raj08]. game-frame [Gui08c]. gameboy [Ano03g]. Games [Cas06m, Mac02, Bre02d, Bre02e, Cas02e, Kus02d, Kus03d, Kus06d, Rot01a]. Gaming [Ros09a, Kus05a]. GaN [EM02]. Gang [Ano09-76]. Gap [Gui08b, Fai09d, Jon04g, Sch03b]. gaps [McF02a]. Gardener [Per05b]. Garrett [Cas05b]. Garriotts [Ano09a]. Garwin [Zor01a]. Gas [Ano08-61, Bar06b, Swe08b, Swe08d]. gates [Tei00]. Gatt [Kar08o]. gauge [Sch09c]. gay [Pre05b]. Gaza [Wei09a]. gaze [Che01b]. GE [Fai05b]. gear [Der09a]. Gearhead [Ano08f]. Geek [Ano09-57, Per06d, Cas04g]. Geeks [Mcf08d]. Gem [And09e]. gems [Six05a]. Gen [San07, KyOK09]. Gender [Gui08b]. Gene [Ano08-64, Hod00, Pat09a]. General [Per04e]. Generation [Kar08h, Kar09c, PP09c, BTR00, Gep00g, Gep03c, Kum06b, Moo06a, Oli02, Rou01, Swe01c]. genes [Moo01f]. Genetic [Has08b]. genie [Pos00]. Genius [Ano06d, Cas05d]. Genome [Cla01, App02c, Moo00c]. genomics [Moo01e]. geometry [Che01a]. Geophysics [Gui08h]. Geopositioning [Swe08c]. George [Par03a]. Georges [Zor01a]. Geospatial [Ups07a]. German [Ano06-47, Bla05a, Hel03]. Germany [Bla02c, Fai09d]. Germs [Fai03b]. Get [Ano06-60, Rub07, SBB08, Swe06b, Voe07a, Cas03c, Gar03, Hec09, Jon04b, Kap02, LL03c, LL04a, Obe09b, Per09f, Tes03b]. get-rich [LL04a]. Get-rich-quick [Swe06b]. Gets [Ano07-58, PP08f, Che05j, Gep04d, Gep05a, Has09b, Kum06c, Moo04a, Mul06e, Per01c, Pra01, Zak02b]. Getting [Aar05b, Cas06i, Che02c, Kar02c, Kar09e, MN05, Swe01b, Swe05d, Wal03, Bro02a]. Getting-It-Done [Kar09e]. Ghosts [Ano07k, Wal02b]. GHz [BTR00, Raz08]. giant [Chi09, Moo05c, Obe04]. giants [Gol02g, Luc01f]. Gifts [Cas05e, Cas06k, Jon02b, Jon03d, Jon04f, KDG^{+}08, SRJ^{+}07, Jon00c, Jon01a]. Gilding [Sel06a]. Gilmore [Gep01b]. Give [And09b, Lan02, PP06b, Sin05]. gives [Zak03]. Glacier [Ano07p]. glaciers [Swe05d]. glance [Ano00b, Ano05-29, Ano06-34]. glass [Bas05, Gre02]. glasses [Ano04-74]. Gliders [Gui08a]. Global [Ram00, Ups07a, Ups08b, Ano02i, Mor01a, Sin02]. globalization [Ano04-63]. Globe [Ano08l]. gloom [Moo09d]. glory [Bla05a, Kar03c]. glucose [TLT02]. Gnomes [Ano05f]. Go [Ade07c, Cas04h, Hsu07, McF06f, Per09c, TS03, Wol09, Zub09, Ano02e, Bal04, Boy04a, Cas05d, Che03b, Gol01f, Jon04a, Ros06d, Cas01d]. Go-Go [Cas04h, Che03b]. goal [Ano03e]. Goes [Ano05-60, Fai04d, Ups08e, Ano02p, Che05m, Dit00, Fos06, Moo08d, Per01a, Rob03a, Wal09c]. Going [Swe00d, TJ00]. Gold [Ano07h, Che05g, Gep05d]. Golden [Sel06a]. Goliath [Che03j]. Gone [Ano08n, McF06c, TD06, Rio05a]. Gonelevu [Ups09a]. gonna [Kus03f]. Good [Ano05-55, Ano06d, McF06f, Per06c, Ano01a, Ano02o, Ano09-40, Cas01c, PP08e]. good-bye [Ano09-40]. Goodbye [Luc04c, O’D06, Gep04f]. Google [Ups07b, And08b, Jon06d, Kar09k, Luc04a, McF03c]. Googled [Ano07-58]. Gordon [Kar08l, Per08d]. Got [Ano08e, Gei04b, Jon06c, Mor06, Per08g]. Gott [Nah01]. government [Lan03b, Sin04b]. GPS [Bre00h, Dig08c, Kar08h, Voe06b]. GPS-enabled [Voe06b]. grad [Kar05b]. graduate [Jon08c, Kar05b]. graffiti [Gol02e]. grails [Ano04-99]. grand [Mon09]. Grande [Ano01-45]. Graphene [Sav09a, Sav08a]. Graphical [Fos06]. graphics [Per01a]. graphs [Hel02a]. Grapple [Jon03f]. grating [Per04i]. gravitational [Bel06]. gravity [Bel06]. Great [Lev08, Luc07d, Ano09-62, Che04g, Gep04a, Har02, Per05p]. Green [Che07c, Kit05, Voe09a, DK09, Fai09d, Jon04b, Per05j]. Greener [Jon07f, Has09b].

Greenhouse [Swe08b, Chi05]. Greening [Ups07b, Fai05b]. Greetings [She04]. Greey [Cas01e]. Grid [Ano04-91, BBS^{+}01, Sav09b, Ade09b, Boy06d, Fai04c, GG06a, MN05, Moo08d, OW01, Swe06a, Swe08e]. gridlock [Bla03a]. Gridlocks [Ano01-46]. grids [MFD00, Swe03b]. Grip [Ano09q, MN05]. Groceries [Kit05]. Grokster [TP06]. Grounded [Ups07c]. Group [Ano06-48, Gol03a, Chi07]. Grove [Gep00a]. grow [Cos05c]. growing [Amb01c, Kar03a]. Grows [Ano06e]. Growth [Fai08a, Rie01a, Sel06d]. GSM [Ano09m]. guarantees [Svo06a]. Guarding [Tro04]. guards [Obe05]. guide [Cas03h]. guided [Moo02a]. guild [Gui08c]. Gun [Mei07, Wal07a]. guns [Hec09]. Guys [Ano05-54, Wal08c, Cas03k].

H [Swe08a]. hack [Ack09b]. Hacker [Cas03e, Zac05]. Hackers [WU01, Cio01, Kus09b]. Hacking [Ano03f, Car02a, McF04a, Mcf09b, Wal08b, Aar05a, Che03e, Tur08]. Half [Cas04c, Amb01b]. Half-Speed [Cas04c]. Halfway [Kum06b]. Hall [Sni00]. Halt [Ade08i, Dum08b]. Ham [Ano06i, McF04c, Gol05a]. Hand [Ano07o, BBS^{+}01, Mir00a]. handbag [Hoo05b]. handed [Hel02b]. Handheld [Rie01d, Sav03b, TS03, Ver03]. Handhelds [Com00b]. Hands [Gra07, Lan02, Six05e, Tur08, Fos09c, Jon01d, Sch08b]. Hands-On [Six05e, Tur08]. handset [KP09]. handsets [Che05h]. Hanford [PA02]. hangover [Gol03b]. Hangs [Swe03d]. happen [Ano02o]. Happy [Mur07]. harassment [Voe06b]. Hard [GH09, Jon05d, Jon07a, Das03, Hug02]. harder [Ack09b]. Hardware [Cas06j, Gol03b, SR00]. Hardwires [Ade08a]. Harmonizing [Moo01e]. Harnessing [Lan03a]. Harvard [Dur05, GS08]. harvest [Ups08h]. Hate [Per08f]. haul [Rie01c]. haunting [Kar08i]. have-nots [Jon03b]. haven [Shr01]. haves [Jon03b]. haystack [Cas04e]. Hazard [Kum01e]. HDTV [Wol03]. head [Gol02b, LL06]. headaches [Wan02]. Headed [Ade08j]. headhunters [Kar02a]. heads [BC04]. heal [Ano04-46, BD09]. healing [Ade08g]. health [Aar05b, Kas02, Ros04d]. healthcare [Moo02b]. Heart [Ano07e, Kri06, Che01j]. heartland [Hoo05c]. Hearts [Ano01-78]. Heat [Rom07, Ano04e, Jon06e, Zor01c]. Heats [Cos07]. heavens [Edw05]. heavier [Rie01f]. heavyweight [Kus05a]. heir [Che02a]. Helix [Col07a]. hell [Bre01a, Che03d]. helm [App04a]. Helmets [Jon07a]. Help [Dum08b, Cla01, Pes02]. Helping [BBS^{+}01, Pes02]. helps [Kum05c]. Hendricks [Obe08]. Her [Ano07i, Ano08g, App02b]. Herbert [Per02b]. Here [Ano04-73, Boy05, Coe04, Sch08c, Sch08c]. Hero [Per07d, App02b, Cas01d]. Hertz [Kar02i]. Herwig [Sav01]. Hidden [Jon08f, McF04d, Six05a, Moo01a]. hide [Jon05e]. High [Ano05-61, BCGM07, Che09c, Chi06, Heg08a, Jon07b, Kar08i, Nev01, Per05f, Tes07b, Ano04-64, Ano09-58, Bre00a, Bre04b, Che05f, Gol02c, Hel05c, Jon05h, Kum01c, Kum01f, LW00, McF03b, Mul06e, RSS00, Sav02a, Shr01, Swe05d, Ups09c, Zor06c]. high-frequency [RSS00]. high-IQ [Bre00a]. High-k [BCGM07]. high-power [Hel05c]. high-pressure [Ano09-58]. high-rises [Gol02c]. high-security [Mul06e]. high-speed [Kum01f, LW00, Sav02a]. High-Stakes [Ano05-61]. High-Tech [Chi06, Per05f, Che09c, Kar08i, Ano04-64, Kum01c, Shr01, Ups09c]. high-wire [Che05f]. Higher [Fos05a]. highway [Bre00a]. Hill [P.08]. hire [Kar02e]. Hired [Tes08a]. Hiring [Cos07, Kar01b]. history [Rio04, Zim00]. hit [Per05g]. Hitachi [Mor01c]. Hits [Lai03, Gep01e]. Hitting [Che01c]. Hobbyist [McF07e]. hobgoblins [Sel01a]. Hodge [Ade08e]. Hoffman [Raj08]. hoist [Edw05]. Hold [Kim05]. holds [Rie02a]. holes [Zor02c]. Holiday [Cas05e, Cas06k, Swe04d, SRJ^{+}07]. Holidays [Jon03d, Jon04f, Jon00c, Jon01a, Jon02b]. Hollywood [Kus05b, Avi02b]. Home [Boy08a, Per05o, Pre07b, Sch09a, Cas02h, Che02h, Che04i, Per09c, Wal07c]. homeland [Cas02h]. Homemade [Lip05]. Homo [Mcf08c]. ’hood [Jon05b]. hook [RWZ05]. hooks [Jon05b]. Hope [Ade08i]. hopes [Sin08]. Hopkins [Zor01b]. hospital [Wei02b]. Hot [Ano01-49, Fos07b, Gui09c, Ups09d, Per03c, WU01]. hotspot [Ano02q]. Houghton [Nah01]. hour [Ros06b]. House [Kum00b, Mul06a, Per09a]. Housewives [Ade09c]. how-tos [Kar01b]. Howard [Cas05a]. Huawei [LL03a]. Hubble [Cas04d]. Hubbub [Chi06, Wal02a]. huge [Per05c]. hulls [Ade08g]. Human [And08c, Ano05-59, GG04, Moo00c, Tei01, Cla01]. humanity [Car02a]. humans [Fos00b]. humidity [Bre03b]. Humor [Sel08e]. hundred [Mon09]. hundred-grand [Mon09]. Hundreds [Gui08f]. Hunt [Ade08d]. Hunters [And07c]. Hurricane [GP06a, Per05g, Das05]. HY [Per04e]. HY-wire [Per04e]. Hybrid [Ano01-47, Fai04d, PP08a, Jon00d, Jon03e, Jon05f, LM01, Voe05, Zor04b]. Hybrids [Jon03e, PP08a]. Hydrogen [Jon06b, Pre05a]. hype [Che01d]. Hyperencryption [Che01d]. hyperlinking [Che01e]. Hypersonic [Gui04d]. hyperspace [Che01a]. hysteria [Zor01b].

Ian [Per05h]. IBM [Ano06-55, Gui05d, Hel01b]. iBrain [Kar09k]. IC [Bro03a, Bal04, Gep01d, MG00b, Rou01]. ICANN [Mue02]. Ice [Dum06a, Mor07, Gui05e, Swe03c]. iconoclasts [Cas01d]. ICs [BTR00, Gep03b, Wal00, Hel01b, RSS00]. Id [Kus02d, Gui06a, Moo06d]. Idea [Cas06d, Kar02d, Kum04d, Moo06a, Sav02c]. idea-and [Kar02d]. ideas [Che02k, Gep01a, Sel04b]. Identical [Hel02a]. Identity [Ano04c, Ano02p, Gui06a, Wal04]. Idiot [KKG08]. Idiot-proofing [KKG08]. Idyll [Sch06]. IE6 [Gol09b]. IEDS [Zor08c, Ano08-59]. IEEE [Ano00a, Ano04-75, Ano07a, Ano07-53, Ano00-43, Ano06-48, Ano09-65, Che05f, GP06b, Has09c, LL04b, Ano00o, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano03o, Ano04i, Ano04j, Ano04l, Ano04k, Ano04m, Ano04n, Ano04o, Ano04p, Ano04s, Ano04t, Ano04q, Ano04r, Ano04u, Ano04v, Ano04-31, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05m, Ano05n, Ano05o, Ano05p, Ano06p, Ano06q, Ano06r, Ano06s, Ano06u, Ano06t, Ano06w, Ano06-31, Ano06v, Ano06-32, Ano06x, Ano06y, Ano06-33]. IEEE [Ano06z, Ano06-28, Ano06-27, Ano06-29, Ano06-30, Ano07y, Ano07-44, Ano07-43, Ano07-38, Ano07-46, Ano07-39, Ano07-45, Ano07-40, Ano07-41, Ano07-48, Ano07-47, Ano07-42, Ano07s, Ano07t, Ano07x, Ano07u, Ano07v, Ano07w, Ano08-58, Ano08-57, Ano08-49, Ano08-50, Ano08-51, Ano08-52, Ano08-53, Ano08-59, Ano08-54, Ano08-55, Ano08-56, Ano08q, Ano08r, Ano09-53, Ano09-54, Ano09-55, Ano09-56, Ano09l, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09e, Ano09f, GR09, Has09b, Hir05]. If [Ade08e, Ano04-71, Sch09d, Per07b]. Iffy [Kar03f]. II [Mul06e]. illegal [Che01e]. ills [Che02a]. Illuminated [Kum08a]. illusion [Che04j]. IM [Ano02r, Che02d, Che08d]. image [Cav04, Che05p, Per05h]. images [Sul05]. Imagine [Ano07-54, Per07b]. Imagining [Ano06-56]. Impact [Car01, Ram00]. imperial [Sel01d]. Imperiled [San07]. impermanence [Luc04d]. implant [Ano04-92]. Implants [Ano08-90]. Imports [Ade08i]. Impossible [Luc06c, Kum03c]. impress [Jon05d]. improve [LW00]. improvements [Rie01a]. Improving [CR02]. In-vehicle [Cur01]. incident [Swe02]. Included [Ano06m]. Incorporated [Che07f]. increase [BOT06]. increases [Cos05c]. Incredible [Mul06c]. indelible [Gep03d]. Index [Ano08-77]. India [Sin03b, Ano09-60, BBS^{+}01, Cos05c, Har02, Sin02, Sin04b, Sin05, Sin06a, Sin06b, Sin06c, Sin06d, Ram00]. Indian [Ano04-66, Sin03a, Sin07b]. indigestion [Ano05-53]. Indium [RSS00]. Indonesia [Ros09c]. induced [Gep04f]. Industrial [Fos08b, Kap00, Gol02b, Hod00, Luc03d]. Industry [Car01, Mak02, Cas02a, Eng01, Fra02b, Ise03, Per05c, Swe00g, Zim08]. inexpensively [Sin06d]. infections [Ano09-63]. infinite [Ano03f, Gol04b]. Infoglut [Zel09]. Information [Ano05a, Bow03, Gol06b, Mcf09c, Gol01c, HD03, Ken00, Koo03, Moo01e, Ros05b, WU01]. infrared [Dig06c]. Infrastructure [BBS^{+}01, Gol04d, Obe00b, Rie01a]. infringement [TP06]. Ingenieurs [Ano05-57]. Ingres [Moo09c]. Injection [Ano08-64]. injuries [Swe00c]. Ink [Ano09-61]. Inner [Kun01]. Innovation [Ano08-79, De 01, Ano01n, Jon00b, SW05]. Innovations [Che08b, Gep00d, Gep00e, Gep00f, Hel00, MG00a, Moo00a, Moo00b, MG00b, Rio05a]. innovative [Swe06b]. Innovator [Dur05]. innovators [Kar04c, Sch08b]. Insecurity [Sho05b]. insights [Mul05c]. Insomniacs [Kus06c]. inspections [Swe02]. Instant [Der09b]. instincts [App01b]. Institute [Cas01f]. Instrumentation [Rie00j]. instruments [Bro03b, Kus09b]. insulation [Bro03a]. Insurance [Ano06-48]. integrated [Kum02a]. integration [US03]. Intel [Gep03c, Gol05c, Moo08a, Per09g]. intellectual [FZ03, Hoo05b, Kar02d]. intelligence [Zor02a]. Intelligent [Ano08s]. interconnect [Bro03a]. Interconnection [Bre00c, Bal04, Mor00]. interconnects [Sav02a]. Interface [Pat09c, Pec08, Woh05, Kum01f, Moo02a]. interfaces [Oli00, Per05i]. international [Kar04c, Zak03]. Internet [Ree09, McF04b, Ack09a, And07a, Ano00-44, Ano04-67, Ano05-53, Ban02, Bar01, Bri08, Che01j, Che02a, Che02j, Che03e, CGC04, Che05b, Che05i, Che05j, Com00b, Com00c, DY01, DR00a, DR00b, Ell01, Gol04d, Hoo05c, Jac05, LL06, Lou00, Luc02d, Luc08c, McF03d, Min06a, Mue02, NB07, Per01b, Per05a, Per09g, Rie01a, Rot01b, Roy00, SS05, Sin06b, Swe00b, SC06b, TJ00]. Internet-enabled [Bar01]. Internet-style [DR00a]. internship [Kar02f]. Interplanetary [Ack09a, Jac05]. interview [Kar02e]. intrabody [Gol06f]. intranet [Ead00]. Introduction [Ano08-62, Ade09c]. intruders [Cas06a]. intrusions [Ken00]. Invent [Tes08a]. Invented [Mur09]. Invention [Ano05c, Bra06, Cas06d, Kar03g, Sav05a, Sti03, Tes06, Tes07a, Tes08d, Tes08c, Tes08b, Ano04f, Jon05h, Par03a, Tes08a]. Inventions [Sav05b, FS04]. Inventor [Cas03h, Sti03, Kar02i, Tes09c]. Inventors [Ges04, Kar04c]. Investigation [Cas04a, Hel02d]. investors [LL02c, Sin04b]. invisible [BM05]. Inviting [Ros01b]. iodide [Amb01c]. ion [Per08e, BBS^{+}01]. ions [Min03]. IP [Che05k, Swe07c]. iphone [Per09d, Che07b]. iPod [Cas04h, McF05c]. IQ [Ano03l, Bre00a]. Iran [Swe04f]. Iraq [Ade08b, Bow03, Bro02a, Zor06b, Zor06c].

Irradiation [Gol03c]. irrelevance [Rod03]. irresistible [Gol03d]. ISBN [Cas01l, Cas02d, Cas05b, Nah01, Sni00]. Ishmael [McF05b]. Island [Ano07j, Jon03a]. Islands [Ups09a]. isms [McF07d]. ISScapades [Obe07a]. issue [Bre00b]. issues [Che04m]. Itanium [BBS^{+}01]. I’ve [Jon06c].

J [Nah01]. J. [Ree09]. Jack [Ano05-56]. Jackie [Kar08m]. jam [Cic01]. James [Ano04-75, Bre01a, Kar08j]. jammer [Gui05f]. Jamming [And07b]. Jan [Ano08-58]. Jan. [Ano07-44]. January [Ano06-32]. Japan [Boy03b, Boy04c, Boy08a]. Japanese [Ano02e, Boy04d]. Java [Bre02d]. jaw [San03]. Jeannine [Kar08n]. Jeff [Kum01a]. Jeffrey [Che01b]. Jeopardizes [Kar09c]. Jeopardy [Col07a]. Jet [Kar08f]. Jim [Kum04b]. Jini [Mor00]. jitter [LW00]. job [Cos05b, Kar02d, McC05, Pat09b, Per04b]. Jobs [Cas07d, Per04d, Per05d, Ada08, Ade08h, Ano08-62, Ano08-72, Ano09-36, Cas01h, Che06c, K.08, Kar02g, Kar03d, Kar08j, Kum08a, P.08, Per08a, Sel06d, Ups08f, Zor08b]. Jockey [Kar08f, Ros06e]. John [Gep01a, Per04f]. Johns [Zor01b]. Johnson [Gol01b, Kar02i]. Johnstone [Rio07a]. join [Zak05a]. Joins [Ano09-60]. Joke [Kar08m]. Jon [Per00c]. Jones [Kar03e]. Jordan [Ups09f, Ups09e]. Joseph [Kar08o]. Joules [GS07]. journalism [Che03a]. journals [Rie03c]. journey [Cas03e]. Judy [Kar08q]. Juice [PP05b, Jon05h]. Julie [Cas04g]. July [Ano04q, Ano04r, Ano05v, Ano06v, Ano07-45]. Jump [Moo08a, Aar05b]. juncture [Swe04f]. June [Ano04s, Ano04t, Ano06w, Ano06-31, Ano06-33]. Jungfrau [Fos07b]. junkies [LL06]. Jurassic [Son08a]. Just [And08c, Ano05d, Ano07-54, Moo02c, Yeu06, Bre01c, Cas03c, Per02b, Wal01b].

Kailath [Per07f]. Kaleidoscope [Boh06]. Kammler [Sni00]. Karaoke [Ano08-63]. Karaoke-bot [Ano08-63]. Karl [Gep04b]. Katrina [Jon07i]. Keep [Gui07c]. Keeping [Com01a, Jon01c, Swe07c, TLT02]. Keeps [BBS^{+}01]. Keith [Kum05b]. Kelly [Cas01f]. Kenyon [Per09e]. kept [Che02j]. Key [Rie03c, Bow03, Bro02b, Mul01a]. Keya [Zor09b]. Keys [Moo06d, Moo07b]. Kick [Voe09c]. Kick-start [Voe09c]. kid [Wal01b, Kus06b]. Kids [Wal07c, Kum05c]. Kierstein [Zor04a]. Kilby [Ano05-56]. Kill [Ade08d]. killed [Gol05e]. Killer [Ros09a, Ros05c]. Kilometer [Das08a]. Kind [Wol09, Boy07]. Kindlers [Per09f]. Kindling [Son08b]. King [Fai09e, Raj08]. Kings [Raj08, Zor02g]. kiosk [Cos02a]. Kirch [Per00d]. Kiss [Kar08k]. Kisser [Kar08k]. Klein [Mur09]. Kluge [Per09e]. Knocks [Jon07a]. Knopf [Che01b, Zor01a]. Knotty [Ano05-48]. Know [GP06a, Mor07, Ano04-86, Kar05c, WMM03]. know-how [Kar05c]. knowing [Tes08c]. knowledge [Cas04f, Luc04d]. Kofoed [Zor04a]. Kogelnik [Sav01]. Koike [Kum09a]. Korea [Che05l, Kum07d, LL02b, Swe08e]. Korean [Swe03a]. Kroemer [Per02b]. Kung [Ano09q]. Kung-fu [Ano09q]. Kunio [Kum09a]. Kuroki [Gep04g]. Kyoto [Chi02a, Mor01a, Swe09a].

L [Zim08, Zor01a]. Lab [Kar06b, Kar06c, Luc05b, Bre01e, Sav02b]. Laboratory [Ano09-59]. labs [Gol01d, Rio02, Rio06b]. Lady [Per08e]. Laissez [Ano05-57]. Lake [Gol05a, Kum08b]. lamps [Fai04b]. Land [Fai07b, Zor04a]. lander [Cas01f]. landmark [Mul06a]. Landov [Mor06]. Lands [BBS^{+}01]. Landscape [Boy06a]. Lane [Ade07b]. Language [Ano04-77, McF06a, Wes09, Sel01e]. Lanhan [Kum01a]. Laptop [Ano08k, Per07c, Ros06b, Gep03c, Luc05d, Tea08]. Large [Cas01h]. largest [Che02e, Gep05d, Rie03b]. Laser [Das09b, Moo06e, Chi02b, HF06, Mul02b, Mul04a]. Lasers [Ano09-62, EK00, Bru02, McC05, Sav04]. Last [Ano06i, Che03m, Gre02, Ing01a, Joh00b, Kar02f, Kur08, Swe05c]. Later [Ano04-91, GH07]. Latest [Jon07i, Gep01a]. Latin [Ros07]. Latte [Ano01-45]. Lau [Zor04a]. Launch [Ano06-54, Ano04-64, Cas02a]. launched [Pre05c]. Lauper [Wei09b]. Laureate [Kar03e]. law [Cas03h, Che04b, Fra02c, Kus06e, Ade08c, Ano06-57, BOT06, Tum06]. laws [Ban02, Kus06e, Ros03b]. lawsuits [Jon09g, Ree03]. Layman [Ups07f]. LCD [Per07e]. lead [Gui05d, Hel05a, Mul05c]. leading [Ano05-69]. League [Ano09-60, PP05a]. leagues [Man00]. Leaking [CM03]. Lean [Has08d]. leap [Gep04e, Kum09b, Min06b, Obe04]. learn [Sel03]. Learning [Ano05-58, Cas06b, Cio01, Sel06d, Ube00]. lease [Tes03b]. Leave [Ano02f]. Leaves [Gep04c]. Leaving [Kum01d]. LED [Cof05, Fai04b, Zor02d, Ste09]. LEDs [Sav02c, Che05g, Kit05, Mul07]. Left [Hel02b, Ups07h]. Left-handed [Hel02b]. Legacy [Tes03b]. legal [Kle05a, Swe00c]. Legislation [Swe03d, Bla03a, Per01d, Sia01]. Lego [Zor04a, Col06]. lengthen [FB00]. lens [HK04, Jon04d]. lenses [Mul05a]. Leonard [Che01a]. Less [Che04e, Jon06e, PP06b, Tei02, Ups08c]. Lessons [Wou01, Car02b, Ros01b, TP06]. Let [Boy07, GS07, Kus04b, Per08h, Zor02d]. lets [Ano04-46]. level [Jan01, JB05]. Levitating [Ano03f]. Li [Per08e, Ano05k]. Li-ion [Per08e]. liability [Tro04]. liberated [GR09]. licking [Che02j]. Lie [Fos05b]. liens [Ros06c]. lies [App01d, Che00a]. Life [Ano07-49, App04a, Car02b, Kum04c, Luc05c, Zim08, Ade09h, Amb01b, Ano09-39, Che05n, Kus03e, Sin05, Tes03b, Zor02e]. lifespan [Pre05d]. lifestlye [Mcf09e]. lift [Woh05, Zak03]. liftoff [Obe01a]. Lifts [Has08f]. Light [Ano04-70, Ano08-33, Das08b, Hel06a, Per05m, Sav04, Tan09, Ano08o, Boy07, Chi09, Hel02b, Hel04a, Jon04b, Luc08a, Mul05c, Mul06b, PP06b, Per03c, Per04i, Sav02a, Sch03a, Zor02d]. lighter [Sel01c]. Lighting [Ano04d, Fai04b, Rio07a]. lightning [Pat09f, Pre05g]. Lights [Gei03b, Ups09a]. Ligth [Cof05]. Like [Haw07, Luc04a, Ano02q, Ano04-94, Col07b, Gol04f, Hel03, McC05, Moo05c, Sel08d, TJ00, Zbi05]. Limbo [Per05n]. Limelight [Kus02b]. Limit [Das08a]. Line [BBS^{+}01, Ano02a, Che01k]. Lines [Ano01b, Ano01a, Ano01-51, Ano01-77, Ano02m, Ano03k, Ano03l, Ano03m, Ano03n, Ano04-85, Ano04-86, Ano04-87, Ano04-89, Ano04-90, Ano04-88, Ano04-91, Ano04-92, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano05-55, Ano05-56, Ano05-58, Ano05-57, Ano05-59, Ano05-60, Ano05-61, Ano06-52, Ano06-53, Ano06-54, Ano06-56, Ano06-55, Ano06-59, Ano06-58, Ano06-57, Ano06-60, Ano06-61, Ano06-62, Ano07-50, Ano07-51, Ano07-52, Ano07-53, Ano07-54, Ano07-57, Ano07-55, Ano07-56, Ano07-58, Ano07-59, Ano08-78, Ano08-79, Fos05b, Has01, Has08b, Has08c, Has08f, Joh07, Jon06h, Mur07, Nev01, Zor06a, Ano08-34, Ano09-66, Ash00, Che01a, Cor00, Has08e, Has08d, Has09b, Jon04a, Per08b, Per08g, Sch09e, Voe08d]. lingo [McF03e]. Lining [AS09]. Linked [Che02f, Pre04]. Linking [Sav02a]. Linux [Cas03i]. lip [Cos05a]. lip-syncing [Cos05a]. lips [Cos05a]. Liquid [BBS^{+}01, Boh06, Hel02c]. Liquid-Crystal [Boh06, Hel02c]. List [Ano08e]. Listening [Cas03j]. lists [Luc09c]. lite [FS04]. Lithium [Ano04-68, Voe07b]. Lithography [Das08b, Das09b, Gep04d, Hut01, Sch03a]. Little [Ano02g, Ano08e, Per09g, Ups09a, Cas01h, Che01d, Gol06e, Sch03a].

Littlefield [Kum01a]. Live [Ano06n, Ano07k, Ano07-55, BM07, Che03i, Son08a, Che09d]. lives [Gol01b, Kum01a]. Living [Ano01c, Ano01-48, Kus02b]. localism [Jon03c]. Location [Kus06d, Dix05, WMM03]. location-based [Dix05, Kus06d]. locator [KH00]. Logic [Svo07, Ye08, TS03]. Lohr [Cas01d]. London [Gol01c, Cha06a, Mul05b, Mul06a, Mul06e]. Long [Che03i, Che09d, Jon03a, Wee05, Ano09-40, Rie01c, Wol03, Ano07-55]. longer [Com01b]. Lonnie [Kar02i]. Look [Ano05d, Jon01d, Kum01a, Per01c, Mir00a]. Lookin [McN05]. Looking [Bla02b]. looks [Ano04-94, McC05]. Loop [Ano08-36]. Loser [Per08f, Ros08a, Swe08c, Tea09, Ups08d, Ups08e]. Losers [Ano05l, Ano06-62, Ano07-68, Zor08d, Ano04-99, HD03, Ros09d]. Losing [Ano05-61]. Lost [Che02g, Gep04c, Obe07a, Bla05a, Rio04]. Love [McN05, Mei07, Son08b, Aar05a]. Low [Jon07c, Moo09b, Sar06]. low-power [Moo09b]. lower [Ano02a]. lowers [Che06g]. Lunar [Ano09-60]. lunch [Bla02b]. Lure [Sin07b, Gep05d]. lurk [Rie00a].

M [And08a, Che09i, Dur05, Kar09c, Mur09]. ma [Jon01d, Rio05a]. Mac [Kar08a]. Mach [Ano04-82]. Machine [And08c, Che09e, Pat09c, Pec08, Ver03, Ano01-31, Fos09c, Kri06, Kum02b, Sin08, Ups08c]. Machines [Gui08d, KT08, Rob00]. Machinima [Kus08b]. Macroproblems [KM07]. Mad [Mar00, Ade09d]. Made [Jon05g, Jon07d, Rio06a, She00, ZR07, Che05h, Fos01e, Fos09c, Moo05a, SA09]. made-for-TV [Moo05a]. Made-to-Measure [Jon07d]. Made-to-order [She00]. madness [Cas06f, Mar00]. Magazine [Ano04-31, Ano06-32, Ano06-33]. magic [Ano04-85, GP00, Sch03a]. Magical [Zor09b]. Maglev [Ano06-47, LJ03, Ros02]. magnesium [Das02]. Magnet [Jon08g]. Magnetic [Com00d, Dum08b, Ing01a]. Magnetoelectronic [Joh00b]. Mail [Mor06, Che01h, Li01, Luc04c, Ros06d, Smi03, VN03]. Maines [Zor01b]. majors [Cas01d]. Make [Gol07b, Kus08c, Amb01a, Cas01e, Cha09b, Mul05a, Oli00, Par02b, SB02, Cas01e]. makeover [Che05j]. Maker [Swe07e, Fra04, Per02c]. Makers [Moo09a, Mur09, Bla03b, Gep05d, Peh01a]. Makes [Moo08a, Per08j, BM05, Moo01d, Sar06, Sav09a, Tes09c]. Making [Ano04-89, Ano07k, Che01f, Che07d, Moo01f, Mul02a, Sch01, Sel06b, Zor02a, Ano04-34, Kus09a, Mul04a, ST05]. malaria [Aar05b]. malice [Cas01b]. malicious [Aar05a]. Maligned [Swe03d]. mammography [She02]. Man [CS08b, Kar08m, Kar08p, Pie05, Cas06d, Per09d, Pos00, Tes08d, Ano05j, Per08a]. Management [LL05a, Che05a, Gol05e, Kar04b, Kum01b, Per01e, Tei02]. Managers [Che08b]. Managing [Luk09, Ros04d, Coe04]. Manhattan [Vat01]. Maniac [Ups08f]. manned [Obe01a]. manufacture [Kum01f]. Manufacturing [CKVY08, Hel07, PT06]. many [Ens00, Kus03g, McF02c]. map [Pat09f]. Maple [Fos01c, Fos06]. mapping [Mor01c]. marathon [Zor02g, Per08a]. March [Ano04u, Ano06x, Ano09-67, GKK09, Ano07y]. Marches [P.08]. Marching [Cas05c]. Marco [Gui09d]. mariner [Per05b]. mark [Par02c, Per05g, K.08]. Market [Ano08-95, Ano05-40, Boy04d, Cos05b, Moo04b]. Marketing [Che01g]. Marks [Fos07b, Bre00h]. Markt [Bla02c]. Marriage [Swe05a]. mars [Ano04-84, Ano03h, Cas01i, Cas01a, De 09, DiG03, DiG07c, Dig04, GH09, Has09d, Heg09b, JN09, Kum06b, Kum06c, Obe09b, Rob09b, Rom09a, Wol09, Zor02c, Zub09, Cas05b]. Martian [Zak09]. Martling [Kar08m].

Mary [Ros00]. Mash [And08d]. Mass [Cas01l, Jon07d, Per09g]. masses [Sni00]. Master [Gui09b, Kum07b, Per04g, O’K02]. Masters [Cas03k, Moo07e]. mat [App04c]. match [Che05h, Tum06]. Material [Joh07, Hel02b, Kus03d]. materials [Mul05a]. Math [Fos03b, Fos05a, Kar08k, Luc07e, Moo08a, Pat09i, Cas01d, Fos03a, Fos09b, Kle05b]. Mathcad [Pra01]. Mathematica [Fos01b]. Mathematical [Fea09]. mathML [Fos01d]. Mathtype [Fos01d]. Matlab [Fos01e]. matrix [Hel00]. matter [Ano03f, RM03]. matters [Hel04b, Swe08a]. May [Ano04v, Ano05w, Ano06y, Li01, Tan09, Ano02a, Mul01a, Rio05a, Sin05, Ano07-46]. maybe [Bre00h]. Mayday [Hel07]. MBA [Kar04b]. McKellar [Kar08k]. MD [Kum01a, Zor01b]. Me [Gei04b, Kim05, Luc04a, McF05b, Per09f]. mean [Cas03e, Mer01]. means [Che02d, NH05]. Measure [Jon07d, Moo07d, Das01]. Measurement [Sel01d, Bre00f]. Mechanical [Ngu09, Bro05]. Medalists [Ano00a]. Media [And08d, Ano06-49, Cas06e, Kus05b, Per06d, Sel07b, Swe06c, Bla03a, Che02g, Che02e, Che04m, Kus05c, Sam03]. Medical [Roo00, Swe00c]. Medicine [Bor09, Moo01g]. Medium [Ano08-79, Com00d]. Meet [Col06]. meeting [Eas00]. Meets [Jon07g, Tum06]. Megacities [MG07]. Megacity [Ano07-49, Gol07b, Has07]. megaproject [Mor01c]. megatons [Zor01a]. Megawatts [Ram07, Zor01a]. Melon [Ano07q]. members [BOT06]. Memoirs [Cas01l]. Memories [Ric07, Ade09g, Gep03d, Gol06e, Joh00b]. Memory [Boy08b, Moo07e, Moo09a, Ros08a, She01b, And08g, Gra00, Gui04e, Hel05b, Kum01f, Moo04b]. Memristor [Wil08]. memristors [Ano09c]. MEMS [Cas01j, Das08c]. Men [Mur09, Rio06a, Pat09i, Per05i]. Mending [DG00]. Mental [Hei09]. Mentor [Sel08a]. Mercedes [Voe08c]. Mercury [Dig07b]. merger [Ano02c]. mesh [Cas05g, Che06g]. Message [Ano08-79, Kar02c, Wal03]. messaging [Che02i]. Messenger [Dig07b]. Metabolism [Moo07d]. Metal [Rie01e]. Metcalfe [Ano06-57, BOT06]. meter [Per05c]. Method [Fra09, Cas06f]. metro [Ano02i, Jon07f]. mettle [Ros03c]. Mexico [Moo04a]. MGM [TP06]. micro [PP06b, Zbi05]. micro-air [Zbi05]. microbatteries [Ano04-32, LB04]. Microchip [Ano08-64, O’N03, Rio06a, Rio06b]. microchips [San09]. microelectromechanical [Cas01h]. Microfluidics [Moo01b, Jon09d]. micromachine [Bor09]. microprocessor [BC02, GP00, Gep05e]. Microprocessors [Gep00g]. Microscope [Min06d]. Microscopic [Ano08-65]. Microsoft [Lai03, Mul08, Ano04-48, Bre02a, Jon06d, LL03b, Par02a, Par02c, Ray01, Ros06d]. Microwave [Ano08-65]. middle [Ise03]. Mifflin [Nah01]. might [Gol02g, Pat09e]. Mightier [Fos03a, Fos03b]. Mighty [Rie03f]. Migliari [Gui09d]. Mike [Kar08p, Gol04c]. Mile [Jon07b, Che03m, Gre02]. Mile-High [Jon07b]. Military [Jon08c, Kum01c, Kum02a, Sco00]. Million [Ano06-50, Gui07c, Gui07a, BBKV09, Kar02d, Kor06, Swe09c]. millions [Fai05a]. Mind [Cas02e, Sch03b, Kar09k]. Mindell [And08c]. Minding [MBP00]. minds [Ano08i, Kum01a]. Mindstorms [Wal01b]. minefield [Fra04]. Mini [Kar08l, Kar08n, Kar08m, Kar08o, Kar08p, Kar08a]. Mini-Profile [Kar08m, Kar08p, Kar08l, Kar08n, Kar08o, Kar08a]. Mining [App02c, Ing01b]. minorities [Kum01g]. Mir [Obe00c]. mirages [Ano03f]. mishap [Obe03c]. Miss [Bre04c, Par02c]. Missed [Rio06b, Rio05b]. Missile [Kar03a, Sho05c].

Missiles [Ram07]. Missing [Che07b, Wil08]. Mission [Ano05b, Ano05h, Cas01i, Kum03c, Rom09b]. Mistakes [Obe08, Sel05c]. mites [Rie03f]. Mixed [Swe05a, LL02a, Mul05a]. mixed-signal [LL02a]. Mixing [Boy08b]. Mlodinow [Che01a]. Mobile [Tes04, Ano02a, Ano09-63, Cas01k, Che02i, Che05l, Der09a, Dix05, FM00, Fos09a, Kus05c, Per09g]. mobile-dodger [Dix05]. Mobs [Mcf03a, Che03h]. mod [Kus03d]. Modders [Mcf08d]. model [Bre00c, Mac01]. Modeling [Ano06-58, Bal00, Gol06c, Wal00]. models [MG00b]. modem [Cic01]. Modern [Kar09k, Mur09, Per09i, Per05b]. modes [Per05i]. modification [Kus03d]. Modular [YZD02]. Moguls [Kus08b]. molding [Jon04d]. molecular [Hel01a]. molecules [Gep05c]. Mom [Kar07b]. moment [Sch09e]. momentum [Gui04e]. Moms [Per08f]. Money [Ano03m, Cas06c, Kar09f, PA02, Tes07a, Aar05a, Coe04, JP06, Pat09e, Sin08]. monitoring [Gui05e, Jon06i, Rie03b, Ros04d]. Monkey [Luc00b]. Monolith [Ano05-56]. Monoliths [Ano05h]. monster [BK09, Moo06f]. Monsters [And08d]. month [Che03d]. months [Par02b]. moods [Kri06]. Moon [Ano07-57, Obe09a, Rom09b, Zak09, Das01, Sin06c, Cas01f]. Moonstruck [Swe09b]. Moore [Ade08c, Per08d, Tum06]. Moraesg [Mor06]. Moraesg/Reuters/Landov [Mor06]. Morale [Kar03f]. Morgan [Che05g]. Morpheus [Par03b]. Morristown [Swe00a]. Morse [Ano07-55]. MOS [Rio02]. Mosely [Kar08n]. Moses [Kum08b]. Most [Jon08g, Raj08, Gep01e, Pat09e]. Mother [Hun06, She02]. motherboards [CM03]. motion [Jon05h, Kar08o]. motion-capture [Kar08o]. Motor [Ups08f]. Motorcycle [SS08]. Motorhead [Ano08f]. Motorola [Ros06e]. Motors [Per04e, Bre04b, Ros03d].

Motto [Ano06-55]. Mountain [Obe05, Kum05d, GG02]. Move [Gei04a, Gep05a, Cha09b, Che03g, She01b]. Movie [Kus08b, Cas06g, Luc01a]. Movies [Ano05h, Gei04a, Che03g, Per04h]. Moving [Gui07c]. MP3 [Kus03c]. Mr [Ano05k]. Mr. [Col06]. MRI [Sav08b]. Much [Che01d, Swe03d, BOT06, Che05j, Gol06f]. much-needed [Che05j]. muck [CM03]. mulls [Ano09o]. Multicore [Gol07a, Moo08b, SA09]. multifaceted [Swe00e]. Multimedia [Moo06f, Kus05c, Str03]. Murphy [Sch09g]. Museu [Six05f]. Museum [Six05e]. Museums [Six05b, Six05e, Six05f, Six05c, Six05d, Six07a, Six07b, Six05a]. mushrooms [Ano02q]. Music [Che09i, Moc04, Zor08b, Che01f, Che04d, Kus03c, Per01d, Per09d, Ros06e]. Musical [Jon09d]. must [Ade08e, Oli02]. My [Ano07r, Che05i, Kar08k, Luc05d, Rob09b, Voe09a, Cos05a]. Mystery [Ano07j]. myths [Che05k].

N [Ade08e, Ups07f]. N800

[Wal08c, Wal08b]. Nagoya [Aiz05]. Nakamura [Rio07a, Rio07a]. name [Che01l, Min06a, Sel02b]. Named [Kar08q, Gol04a]. Names [Mue02]. Nance [Kum05a]. Nano [Ano08-71, Ano08-67, Ano09-64, KM07]. Nano-dreadlocks [Ano09-64]. Nano-Flowers [Ano08-67]. Nano-pine [Ano08-71]. nanochromics [Per06b]. nanodevice [Gol05d]. nanoelectronics [Gep00h]. Nanoholes [Hel04a]. Nanolaser [Das09a]. Nanolasers [BBS^{+}01]. nanometer [Gep05c]. nanometer-size [Gep05c]. Nanopore [Pat09a]. nanotech [Jon08a, Moo05b]. Nanotechnology [Mer01, Ros08a]. Nanotools [Moo03a]. Nanotube [PP08f, Ros08a, Ama03, Hel01b]. Nanotubes [Rom07, Avo04]. NASA [Kra02, Ano04-82, Ano06-54, Cas01i, Cas03f, Cas05c, Kum08b, O’K02, Obe00a, Obe08]. NASDAQ [Cha06a]. Nation [Gol03c, Gol03a, Har00]. National [SB01, CGC04, Rie02a]. Nations [Ano09-60]. Natural [Ano05-59, Fos08b]. Navigation [Swe08c, Bre00h, Bre02b, Bre03a, Sin02]. Need [Ano06-54]. needed [Che05j, Swe03b]. needle [Cas04e]. needs [Peh01b]. Neglect [O’C01]. Nehro [Kum08a]. Neither [Ano02h]. Neon [Aiz05]. Neptune [Fai05c]. Nerd [PP07b]. Nerdus [Mcf08c]. Net [Che07e, Ree09, Che05j, Com01a, Luc03c]. nets [Cas01m, Rie01f]. Network [Bre01e, Dig07d, Sin03a, Cas05g, Che04k, Che05h, Che05j, Com01b, Gep01c, Gep01d, Kar04c, Kum03c, Kum04c, Kum05c, Mor00, Per04f, Rie01d, RWZ05]. Networked [Ano01c, Ano01-48, Ano01-50, Has08c, Che00a, Jac05, Per01f]. Networking [Kar05c, Swe01c, Ups08e, Dix05, Moo01g]. Networks [And08b, Ano05-50, Cas01n, Ano02i, Ano02q, BOT06, Che02b, Che02f, Che02h, Che06g, DR00a, Kap01b, Mor01b, Pas01, Tes03b]. Neuer [Bla02c]. Neural [Cav04, Ano04-92]. neuro [Mcf09a]. Neuroscience [Fos05b]. neutral [Per01d]. Never [Bre03c, Bre04c, Li01]. New-breed [Ack09b]. Newman [Kum04b]. NEWS [Boy07, Dig07b, Dig07d, Dum07b, Gui07a, Jon07g, Per07b, Rom07, Swe07a, Ade07b, Ade07c, And07a, Ano01n, Ano06-51, Ano08-68, Ano08-69, Ano08-71, Ano08-67, Ano08-70, BSK^{+}01, BBS^{+}01, Che07b, DHGR01, Fai07a, GRG^{+}01, HCJ^{+}01, JCF^{+}01, Jon07a, Jon07f, KLS^{+}01, MMG^{+}01, Moo07c, MMH^{+}01, Per07d, SSB^{+}01, Swe07e, Voe07a, ZMR^{+}01, Ano01a, Ano08-60, Ano08-65, Ano08-74, Moo08b]. Newslog [Ano00-45, Ano00-46, Ano00-47, Ano00-48, Ano00-49, Ano00-50, Ano00-51, Ano00-52, Ano00-53, Ano00-54, Ano00-55]. Newspapers [Hoo07]. Next [Ano04-90, Ano06-71, Che06d, Fos08b, ISV01, KyOK09, Per08d, San07, Swe03a, Swe07d, Che03h, Che09l, Ede00, Kar02d, Kum06b, Moc06, Moo06c, Per02e, Ros04b, Ros05c, Rou01, Zak05a]. Next-Gen [San07, KyOK09]. next-generation [Rou01]. Ney [Ano08-72]. Niagara [Gep05e]. Nifty [Boy03b]. night [Dig06c]. Nightmare [McF05d]. Nights [K.08]. Nine [SC06a, Bro03b]. Nissan [Cas02f]. nitride [Gep04a, Zor02d]. nixes [Ano02c]. Nixies [Zor02e]. NJ [Sni00, Swe00a]. No [Ano04j, Ano04r, Ano04w, Ano04z, Ano05w, Ano05-28, Ano05-67, Ano05-68, Ano05-66, Ano06c, Ano06u, Ano06w, Ano06-32, Ano06z, Ano06-27, Ano06-65, Ano06-66, Ano06-67, Ano06-68, Ano06-69, Ano06-70, Ano06-64, Ano07t, Ano07v, Ano07-61, Ano07-64, Ano07-65, Ano07-66, Ano07-67, Ano07-62, Ano07-63, Com01b, Jon05e, Jon06f, McF04b, Obe07b, Per07b, Swe08c, Far03, Jon01d, Kle05b, Mir00a, Son09a, Svo06a]. Nobel [Ano09-65, MS09]. nodes [Che06g]. noise [Ano04f, HNPC05]. Nokia [Bla02b, Wal08b]. Noncombatant [Ano02j]. nonpartisan [Peh01b]. nonvolatile [Hel05b]. nor [Ano02h]. Norman [Fos08a]. North [LL02b]. Noses [CS08a]. Note [Lai03]. Notebook [Ano05e, Ano07i, Che09l]. Notebooks [Der08]. Nothing [Ade08k, Ano03m, Cas06c, Che07e, Gep01a, Kar09f, PA02]. notion [Gol05b]. nots [Jon03b]. Nov [Ano04w, Ano05x]. Nov. [Ano07-47]. Novels [Rob09b]. November [Ano06z]. nowhere [Ros05b]. Noyce [Ano05e, BC05]. NT [Rob00]. NTT [Ano08-73]. Nuclear [Ade08i, Ano01-77, Ano05-60, Fai07b, Per01c, Ram00, Swe02, Swe08a, Amb01a, Amb01b, Ano04-63, DG00, FB00, FPP00, GG02, GH03, Gui06e, Kum03a, LTN05, LB04, LL02b, Mil01, MBP00, PA02, Pos00, Pre05e, Svo06b, Swe00d, SK00, Swe03a, Swe03c, Swe04f, Ups09b, Wei01, Zor01c, Zor01a, ZM01, Ade08e]. Nuclear-Waste [Ade08i]. Nukes [ST09]. Num [Ano04-31, Ano05r]. Number [Ano04t, Ano04k, Ano04-28, Ano04-29, Ano04-30, Ano07w, Koo03]. Numbers [Ade08b, Che05c, MG07, Mue02]. Nut [Kar08a].

O [Swe07b, Ano08-73]. oasis [Ups09c]. oats [BSK^{+}01]. objects [Gol04a]. oblivious [Che08b]. observatory [Fai05c, McD04]. Obvious [Tes06]. Occupational [Kum01e]. Ocean [Jon08e]. oceans [ES04]. Oct [Ano04x]. Oct. [Ano04y, Ano07-48]. October [Ano04-31, Ano06-27]. odyssey [Per05a, Cas01a]. Off [GS07, Has08f, LJ03, Rom07, Ade09b, Che02l, Eas03, Gep00g, Kar01c, Lan02, Moo08d]. off-beat [Gep00g]. offers [Sta01]. Office [Kar02b, Sel08b, Mcf09e, Sel04a, Sel05b]. office-free [Mcf09e]. officials [LL04b]. Offshore [LL02c, Fra05b, Man02]. Oikarinen [Gep05b]. Oil [Gui08h, Heg08b]. oiled [Ano01-31, Ups08c]. Ojanpera [Che09i]. Old [Dum07a, GG06a, And09c, Fai09c]. Olin [Gui06f]. Olive [Kus08c]. OLPC [Tur08]. ominous [Aar05a]. Omnivorous [Gui07d]. on-line [Che01k]. One [Ano04-91, Ano08-78, Bar03, Gui08f, Kar08p, Svo06a, Swe07d, Che01g, Gep05b, Moo02c, Pos00, Swe00a]. One-click [Svo06a]. One-Man [Kar08p]. one-on-one [Che01g]. One-party [Bar03]. Online [Ano02d, Ano06-31, Ano06-28, Ano07-53, Ano07-56, Kar03b, Che04d, Che05m, GR09, Kar05c, Kus05a, Ros04f, Svo06a]. Only [Ano04-71, Cas07i, Pat09h, Cos05c, Gei03a, Par09]. Ont [Cas05b]. Ontario [Jon07e]. Open [And08e, Ano04-75, Ano08-94, Car01, Cha07b, Kun09, Kus05c, Cio01, IS01, Ray01, Wol02a]. Open-Source [Ano04-75, Ano08-94, Cha07b, And08e, Car01, Kus05c, Cio01, Wol02a]. Opening [Mai03]. opens [Cas06a, Dig08b, Kar08d, Mas02, Mul06a]. Opera [K.08]. Operating [Per01d, San01]. Operations [Obe05]. operator [Bre00c]. opportunities [Pre05b]. optic [Bas05, Hoo05c]. Optical [Rie01f, Man00, Mul01a, Sav02a, Sch03a]. Optics [Ano01-49, WG01]. Options [Ups08b, Roy00]. Oracle [And09e, Mor01c]. Orbiting [Ano09a]. orbits [Ham02]. Orchestra [Ano08k]. Order [Luc03e, She00]. Orders [Cas05c, Lan03b]. O’Reilly [Cas02d]. Organic [DK09, Gui04e, FBT00, Moo02c, Mul05c, Sav02c]. organics [Fai04a]. organization [Cos02b, Luc04e]. Organizations [Sch08c]. Organizing [Sch08c]. orgasm [Zor01b]. origins [Gep01a]. Oscar [Kar08c, Per01a]. Oscar-winning [Kar08c]. oscilloscope [App02b]. Oslo [Jon07f]. Ossi [Gep05b]. Other [Ano07r, Ano08-74, LTN05, Sel02b]. Our [Ano07-59, Fos08b, Kar09c, McF03e, Ree09]. Out-Cheney [BBS^{+}01]. Outback [Ups08a, Zor02g]. Outer [Swe07a]. outfit [Kum02a]. outlet [Bla04]. Outlook [Has01]. outsourcing [Ano04-61]. Oven [Ano08-65]. Overclockers [Mcf08d]. Overload [Ano09-68]. Overlord [Ano09-68]. Own [Gep04c, Kus08c, Ree09, Sel07a, Boy04a, Kum01d, Kus06a, Rie02c]. owns [Che04m]. oxymoron [Luc03d].

P [Cas05b, Raj08, Zor01b]. Pacemaker [Sch01, Moo05c]. pacemaker-like [Moo05c]. Pack [De 09]. package [Tum06]. Packages [Gol01f]. packets [Che05h]. page [Cra00]. pages [Cas05b]. Pain [Ano08-90]. pains [Ano08-61]. Paintable [Hel02c]. paladins [Com00e]. palm [Moo06b, Per08c]. Pan [Cas03d]. panda [Kum05c]. panels [Boy04a]. Pantheon [Gep01a]. Paper [Ano04-69, Ano08-75, BBS^{+}01, Kar08n, Ano01-47]. Paperboy [Obe02a]. Paradise [Gui08g, Gep04b]. Parallel [Aar06a, Che01a]. paramecium [Luc02b]. Parched [Ade08j]. Pardon [Gei04b]. Parking [Jon06g]. part [Jon09g]. participation [Ano07-56]. Particles [KM07]. partners [Mor01c]. partnership [San03]. party [Bar03]. passage [Sin03b]. Passes [Tes04]. passives [US03]. Passport [Ros05b]. Past [Fos07a, Wou01, Wal02b]. Patent [Fra02d, Fra03, Fra05a, Gol06d, JL04, Kar02b, Sav05a, Sav05b, Sav05c, Tes07a, Tes07b, Tes09a, Bro02b, Che04l, Che04m, Fra02b, Fra02c, FZ03, Gar03, Moo05b, Tes08d, Tes08c, Che08b]. Patenting [Tes09b]. Patently [Tes06, Che08b]. Patents [FS04, Fra09, Tes08b, Cha04, Hei02, Hol06, Kle05a, Kle05b, Kus03g, Loe06, Par03c, Tes03a, Hol06]. path [Per00e]. Patriot [Kar03c]. Paulson [Ano04-75]. Paving [Gre02]. Pay [GH07, Che01f]. Payette [Cas04g]. Payne [Mul01a]. Payoff [Swe08c]. Pays [And08f, Ano01-78, Per06e]. PC [Dit00, Pre05d]. PCB [US03]. PCs [Wol02b]. Peace [Swe05b, Sel05b]. Pebble [Wei01]. Pebble-bed [Wei01]. peek [Boy06c]. Pen [BBS^{+}01, Wol02b]. Pencier [Cas01e]. pencil [Sav08a]. pending [Kus03g]. Penguins [P.08]. People [Gui07c, Jon07c, Lin08, McF05c, Sel08c, Wol09, And08a, Luc06b, SS05]. Perfect [Cas07e, Kro07, Gui04a, Sav03b]. performance [Gol01a, Rie01d]. Performer [Ano07m]. peril [O’C01]. Perils [Pre05e, Che04c]. perks [Moo06g]. permission [IS03]. permit [Hel04a]. Perrow [Fos08b]. Perseus [Cas01l]. Persist [Kum03d]. Person [Son08a]. Personal [KH00, Ano02p, Der09c, Gol04a, Kar01c, Pre05d, Wal09c]. Perspective [Tes09a]. Pesky [Che02h]. Pet [Per08i]. Phase [Ano09-69, Hel05b]. Philippe [Wei09b]. Phillips [Ano09-87]. Philosophy [Ano01-50]. phish [Ros06d, Sch08a]. Phishin [McF06c]. Phone [Ano08-73, Kar09g, Tes04, Ano02a, Boy03b, Boy05, Tes03b, Tes03c]. Phone-y [Kar09g]. Phones [Dur01, Rie01b, Bla02b, Bre02d, Cur01, FM00, Swe00b]. phosphide [RSS00]. PHOTO [Mer04, Aiz05, Ano04-70, Ano04-71, Ano04-72, Ano04-73, Ano05-42, Ano05-43, Col05, Hel05e, Kim05, Kit05, Kol04, McN05, Nak04, Pie05, She04, Wee05]. Photograph [McN05, Mor06]. Photographers [Wu08b]. Photographic [Boy06a]. photographs [Luc06a]. Photon [Ups08a]. photonics [Gep04e, PK05, Ros06f]. photovoltaic [Gol05b, PP05b]. photovoltaics [Fai03a]. physical [Ano01n]. physicists [DHGR01]. Physics [Kar09l, Kar09m]. Picks [Ano06-49, Cas06e, Cas07f, Cas07g]. Picture [Ano06m, Ano06n, Ano06o, Ano07o, Ano07n, Ano07p, Ano07r, Ano07q, Ano07m, Ano08m, Ano08l, Hun06, McN07, Mor06, Ric07, Sim07, Ter07, Yeu06, Ano08a, Ano08n, Ano08k, Ano08p, Ano08o, Ano08-36, Ano08-38, Ano09d, Ano09p, Hel02a, Kor06]. Picture- [Ano07m]. pictures [Per05h]. Pieces [Ano08n]. pill [Ano08-92]. pillars [Ano04-32]. Pilot [Per08c]. Pilots [Gep04c]. Pinchuk [Cas01e]. pine [Ano08-71]. pioneer [FS04]. Pipe [Rot01b]. Pirates [HOD^{+}09, Jon03f]. pirating [Kus03c]. pitfalls [Tes08c]. pits [Jon03b]. PJM [Bre00c]. Place [Ano09-70, Bre03b, Fos02b, Jon05e, LL02c, US03]. Places [Ade08j]. Plan [Bar06d, CM09, Che06b, Kar02b, O’K02, Ups09f]. planarization [Bro05]. Plane [Jon06h, Wei04, Zor06a]. Planet [And07c, Ups09g, Dig07a]. planetary [Kum06b]. Plans [San07, Chi05]. Plant [Pre07b, Amb01b, Gui05a, Per02d, Sin05]. Planting [Per01e]. Plants [Ano09-86, Swe02]. Plaque [McF04d]. Plasma [Boy04b, Boy04d]. Plastic [Fai09f, Jon09e, CT05, Moo06g]. plastic-based [CT05]. plastics [Moo02c]. Platform [BBS^{+}01, Wol02a]. play [Ano09-70, Bre02d, Gol02a, Mac02]. Player [Kus07b, Per02e]. players [Cas01d, Che02e, Ros04e, Sin06a]. playground [Che02e]. Playing [Kus07a, Wal08a, Jan01, JB05]. PlayStation [Che09e]. playtime [Zor04a]. Please [Ano04-74]. pledges [Sin04b]. Pleo [Per08i]. Plug [SS08, Voe09a, Gol02a, Jon05f, Kar01c]. plug-and-play [Gol02a]. Plug-In [Voe09a, SS08, Jon05f]. Plugging [Voe08b]. plunge [Kum06d]. plus [Swe00a]. plutonium [Gui05a, Pre05e]. Podcast [Ano06-49, Cas06e, Cas07f, Cas07g]. Podfather [Per08c]. Pods [Ano08-69]. point [Kum02a, Zor01a]. Poky [Moo06g]. Polaroid [Ade09e]. police [Kum02c]. policies [Ano05-70]. Policy [SB01, BR02, JL04, Peh01a]. policy-makers [Peh01a]. Political [Swe07e]. politically [Wei03]. politicians [Ano03b]. pollinate [Cas06g]. polling [Che04c]. pollute [Gol02h]. pollution [Ano05-40]. Polymer [Hel00, Moo06g]. polynomial [Fos09a]. ponder [Ano04-78]. Poop [Per08i]. Poop-Free [Per08i]. Poor [CS08b, Tes08d]. pop [Cas03k]. poppy [McF02b]. pornographic [Che05p]. portability [Wan02]. portable [Cra00]. portends [Ano04-78]. Portraits [Ano05-44]. Poseidon [Shr09]. positioning [Sin02]. possibilities [Arn09]. PostDoc [PP06a]. posthuman [McF03e]. potassium [Amb01c]. Potential [Ups08b]. Power [Ada08, Ade07a, Ano07p, BR02, BBS^{+}01, Cho00, Dur01, Gep05c, Gol06d, Gui08g, Jon06c, Jon07c, Jon08e, Mak02, Pre07b, Rie01b, Sch08c, Swe00f, Swe06e, Zac07, Zub02, Amb01b, And08a, Ano03b, Ano03j, Ano04-32, Ano09-66, Ash00, Bas05, Boy06d, Cor00, Fai03c, Fai04c, Gol01c, Gol02h, Gol05b, Gui06e, Hel05c, Jon03b, Jon04a, Jon06e, Jon09b, KMT00, LTN05, MFD00, Moo06a, Moo09b, Mur09, Per01b, Per01c, Rie03f, Ros06c, Sar06, Sch09e, She00, Sin05, Sta01, Swe02, Swe03c, Swe03b, Swe09c, Ups08h, Zor06b, And08a]. powered [Zor02g]. Powerful [Jon08g, Gep01e]. Powerless [Wei09a]. pp [Cas01l, Cas02d, Nah01, Sni00, Swe00a]. pp. [Ano04-75]. Practical [And08a]. practices [Per05j]. Pragmatic [Mil01]. praise [Fai09b]. Pre [Kar08k, McF04e]. Pre-Algebra [Kar08k]. precious [Luc01d]. Precision [Bre03a]. Predicting [Hug09]. Prediction [Ano06k, Das05]. predominate [Pat09i]. Prentice [Sni00]. Prentice-Hall [Sni00]. prepare [Obe01a]. prepares [Mul05b]. Preparing [Fos08b]. prescription [JL04]. presentation [Eas00]. preservation [Smi05]. preserving [CR02]. President [Lan03b]. Press [Cas01f, Dur05]. pressure [Ano09-58]. prevention [Gei03b, Sam03, Smi03]. Prey [Dur03]. Price [Heg08b, Moo04b, Ano03j, Pre05f]. prices [Che09k, Moo04b]. Primary [Ano09-71]. prime [Moo05a]. primes [Ano06-72]. printable [Hel05a]. printed [Bal00]. printed-circuit [Bal00]. priorities [Kum00b]. prison [Per05j]. Prius [Voe08b]. privacy [Ano01-51, CR02, Che01b, Che04j, Har00, Luc08e]. Private [Che06f, VN03]. privatizes [Bre01l]. Prize [Swe05b, BBKV09, MS09]. prizes [Ano09-65]. pro [GG02]. probably [Che09l]. Probe [Ano01-66, Ano01-67, Ano01-69, Ano01-70, Ano01-71, Ano01-72, Ano01-73, Ano01-74, Ano01-75, Ano01-76, Ano00-84, Ano01-65, Ano01-68, Ano04-96, Moo01f]. probes [Gui04c]. Problem [Ano05-48, Cav04, Tes03a, Ups09e]. Problems [Kum03d, Luc09f, Sch03b]. processing [Per05h]. processor [Aar06a, Eyr01]. processors [Gep01c]. production [EK00, Gol02h, Gui05a, Zor06b]. Products [Ano04-75]. profession [PP05a, Wal04]. professional [Cas06f]. professionals [Kum01a, Ros04a]. Professor [Fea09]. Profile [Gep00a, Kar08m, Kar08p, Per00d, Cas04g, Cas06f, Kar08a, Kar08l, Kar08n, Kar08o]. profiling [Dig06b]. profiteers [Che04l]. Program [Bar06c, Swe06d, Swe04f, Ano06-48, Obe07a]. programmable [Ano03f, TS03]. programme [Obe01a]. programmers [Cas01d]. Programming [Lai03, BBKV09, Cio01]. programs [Kar04b]. Progress [Ano04-77, Ano04-75, Ano04-76, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-45, Ano05-46, Kan02]. project [Sin05, Tur09a, Cla01, Vat01]. projector [Tur09a]. Proliferation [Ram00]. promises [Moo06b]. promising [Gep04d]. promote [Che05j]. prompt [Swe02]. Proof [Gol07c]. proofing [KKG08]. proofs [Boy06d]. property [FZ03, Gol02b, Hoo05b, Kar02d, LL03a]. propose [Zak05b]. propulsion [Bre04b, Gui04d, Joh00a]. prospects [Kum03e]. prosthetics [Hoo05a]. prosthetized [Ade09f]. Protect [Che08b, FS04]. Protecting [Swe07d]. protection [Avi02b, FZ03, Kus06a, MFD00, Per05e]. protein [Mor01c]. protein-mapping [Mor01c]. protocol [Che05b, Chi02a, Mor01a]. proved [Tei01]. Proven [Bel00]. provisionals [Tes03a]. Psychiatry [Moo06h]. Psyching [Ros04e]. Public [Bre01a]. publications [Luc08c]. Publishing [Kum03f]. Pulse [Ano06k]. pump [Ano06-72]. punched [Ano00b]. punched-card [Ano00b]. Pundits [Ano04-78]. purpose [Ano03e, Zub02]. push [Obe00a]. pushes [Che05l]. Putin [Chi05]. Puts [BBS^{+}01, LL02b]. Putt [Ano06-52]. Putting [Bre03b, Che03g, Eng01, Fos02b, Jon07g, US03, Ups08h, Wol02b]. Putzeys [Zor08b]. puzzle [Obe03b]. Puzzles [Fea09]. PV [Fai04a].

Qatar [Kar08d]. Qrio [Ano04-79]. Quads [GS07]. quake [Boy06d]. quake-proofs [Boy06d]. quality [DR00a]. quants [Pat09j]. Quantum [Ade07b, Das08a, Gep00h, Kar09h, Kum09b, Min06b, Moo07b, Ano03f, Gep04e, Heg09a, Hel04b, Mul01b, Ros00]. quarks [Gep01b]. quartz [Gep05a]. Queen [Ros00]. quest [Ste08, Zor01e]. quick [Swe06b]. Quiet [Cas07j, Fos02b, Rie00j]. quit [Gol02c]. Quits [Jon07e]. quiz [Pat09i].

R [Che05i, Hir04, Cas02a, Mcf03a]. R&D [BBS^{+}01, Hir03, Hir04, Hir05, HR07, Sav02b]. R-D [Hir04]. Race [Che09e, Boy03b, Gep05b, Gol05d, Jon05d]. races [Gep05b]. Rachel [Zor01b]. Radar [Zim00, Jon05e]. radars [Kas02]. Radiance [Cas02g]. Radiation [Ups08i, Ano02a, Gep04f]. radical [JL04, Rob09a]. Radio [Cas06j, Dum06b, Gep04c, Gol02c, Jon03f, Kum07c, Laz09, Oli00, Sch09f, Ano09o, Che02l, Che06e, Dig08b, Hel05c, Kum05d, Lay01, Luc01d, Ngu09, Pas01, Rie01g, Rie02b]. radioactive [Ups09e]. Radios [Rub07, Ade09h]. rails [Eas03]. rain [Ano02h]. rainbow [Zor02f]. rainmaking [Moo04a]. Raise [BBS^{+}01]. Rallies [Ano04-72]. RAM [Gep04f, YDLC08]. Rambus [Kum01f, Par03c]. Random [Gep02b]. Ranger [Ada08]. ranking [CGC04]. Rap [Kar09m]. rapture [Ano08-98, Jon08a]. rat [Luc02b]. ratify [Chi02a]. rats [Moo02a]. Ray [Hec09, Che08a]. rays [FGBM07]. Re [Ano07-54, Son08b, Zor06b]. Re-engineering [Zor06b]. Re-Imagine [Ano07-54]. Re-Kindling [Son08b]. reach [Moo02b, Rie03e, Sul05]. Reaches [Dig07a, Swe04f]. reactor [Wei01]. Reactorland [Mak01]. Reactors [Ano05f, Gui06e]. reacts [Chi05, Hel02b]. Read [Cos05a, Lin08]. readable [Per06b]. Reading [Ano04h, Ano04-80]. Ready [Gol07d, MMPJ03, Pav07, Dit00, Hut01, Jon04a, Moo05a, Wol03]. ready-to-wear [Dit00]. Real [Ano06-60, Kum09a, PP08f, SC07, WOS09, Ano04-97, Ban02, Hec09, Jon00p]. Realism [Zor09b, Cas01g]. Reality [Jon08f, Per04c, SM00]. Really [Ano06h, Rob00]. Reap [Man02]. rebirth [Rie01g]. reborn [Ros09b]. recall [Che05n]. Recession [AS09]. recharger [Dur01]. Recipe [Kar07a]. reclaim [Bla05a]. reclaims [Gui05d]. recognition [Che05a, Jon09e]. reconfigure [TS03]. Record [Ano04-81]. recorders [Jon04b]. recording [Che05n]. Records [And09b]. recover [Che09k]. Recovering [Sel05c]. Recycling [Kum04d, Per05j, Gol02h, Per02d, Per06e]. Red [Kit05, Per03c, Ups09g, Chi02b, Dig07a, Ups09f]. red-laser [Chi02b]. Redefining [Kum01g]. reduce [Moo06b]. Reducing [Fos08b]. reduction [Chi05]. reeling [Ano02s]. refillable [Dur01]. reflectance [Das01]. Reflections [Luc00a, Luc00b, Luc00c, Luc00d, Luc00e, Luc00f, Luc01b, Luc01c, Luc01d, Luc01e, Luc01a, Luc01f, Luc02a, Luc02b, Luc02c, Luc02d, Luc03a, Luc03b, Luc03c, Luc03d, Luc03e, Luc03f, Luc04a, Luc04b, Luc04c, Luc04d, Luc04e, Luc04f, Luc05a, Luc05c, Luc05b, Luc05d, Luc05e, Luc05f, Luc06a, Luc06b, Luc06c, Luc06d, Luc06e, Luc06f, Luc07b, Luc07c, Luc07d, Luc07e, Luc07f, Luc07a, Luc08a, Luc08b, Luc08c, Luc08d, Luc08f, Luc08e, Luc09a, Luc09b, Luc09c, Luc09d, Luc09e, Luc09f]. reform [BR02, Fra05a]. reform-blackouts [BR02]. reforms [Sin04b]. refrigerator [Ing01a]. regional [Swe03b]. registration [Che01l, Cos02a]. regulation [Swe03b]. regulations [SC06b]. Regulators [Jon03f]. regulatory [Bla03a, Laz09, SC06b]. rehab [LL06]. Reinventing [Ano08-76]. Reis [Ano08-72]. relations [Gol01c]. reliability [Ber02, GG04, Svo06a]. relicensing [Amb01b]. Relics [Ano09-72]. Relief [Ano09-72]. Reloaded [Fos03b]. Remailers [Che01h]. remarkable [Hel04a]. Remote [Luc06d, Ros04d]. remotely [Moo02a]. remove [Ano04-74]. Renaissance [Ano05j, McF07e]. render [Sul05]. rent [Sch09h]. replace [Fai04a]. replaces [Fai09c]. replicate [Mul06e]. reply [Ano01-47]. Report [KSC^{+}05, Swe00a, Ano09-36, Cas04a]. Reports [Cas01i]. representation [Kum01g]. Republished [Cas05b]. Rescue [Sch09g, Jon03e, Mor00]. Research [Gui06b, Ano06-72, Gol01d, Hel02d, Kum06d, Luc03d, Obe01b, Rie03c, Rio02]. Researchers [Sav08a, Hel01b, Kum05c]. reservation [PA02]. resist [Gep05d]. Resources [And07b, Bre01g, Bre01h, Bre01i, Bre01j, Bre01k, BJC01, Cas06m, Che07g, Cos04, Cos05b, Cos05c, Cos06, Cos07, Dur05, Fos01e, Fos07a, Kus07b, PP05a, PP07b, Rie01h, She01a, SRJ^{+}07, Tes07a, Tes07b, Wal07a, ZR07]. respect [LL03a]. Rest [Per05o, Rom09a, Sch09c]. Restoring [Swe08f]. Restructuring [Swe00g, Mac01]. Résumé [Pat09g]. resumes [Kar01a]. resuscitate [Sin05]. Retaking [Fos04a]. retreat [Kus03a]. Retrospective [Luc04e]. Return [Zak09]. returns [Ano06-52, Bar03, Wei01]. Reuters [Mor06]. Reveals [Jon08f]. revelation [LL02b]. Reversal [Kum03e]. Reverse [Ade08f, Kum00a]. Review [Ano03f, Bre01a, Bre01d, Bre01m, Cas01d, Cas01e, Cas01f, Cas01i, Cas02c, Cas02d, Cas02g, Cas03b, Cas03g, Cas03e, Cas03h, Cas03k, Cas03m, Che01a, Che01g, Che01b, Che02f, Che03h, Dur01, Dur03, DR00b, Fos01a, Fos02a, Fos02b, Gep01a, Gep01b, Gol01b, Gol01e, Gol01d, Gol01c, Ham02, Hoo02, Kar09c, Kar09m, Kra01b, Kum01a, Kus02a, Lew01, Mak02, Mar00, Moo01a, Mor07, O’N03, Obe02a, Obe02b, Obe03a, Par03a, Rie01b, Rie01i, Rio02, Ros01a, Ros01b, Rot01a, Sch01, Swe03c, Zor01a, Zor01b, Ade08e, Ade09c, And08a, And08c, Cas04a, Cas04b, Che08b, Dur05, Fea09, Fos08a, Fos08b, Ges04, Kar08k, Kar09k, Lan03b, McF04a, Mur09, Obe07a, Obe08, Raj08, Ree09, Rio07a, Sch08c, Swe04a, Swe08a, Ull03, Ups07f, Zim08]. Reviewed [Cas06c]. Reviews [Eic00, Fos00a, Fos03b, Lai03, Mir00a, Mir00b, Pus00, DR00c]. revival [Jon09b]. revnu [Che08c]. Revolution [KSC^{+}05, Rio07a, Ade09f, Ano05-29, Ano05-41, Che03h, Fos02b, Hod00, Hoo02, Moo01a, NH05, Rie00j, Cas01d]. Revolutionaries [Kum07c]. Revolutionary [Sav03a]. reward [Zor06c]. RF [Gep01d]. RFID [Ano07-51, And09d, FJ07]. RFIDs [Hel05a]. Rhythms [Per06d]. rich [Che09h, LL04a, Swe06b]. Richard [Cas02d, Zor01a, Swe00e]. Richart [Nah01]. riches [Far03]. Ricochet [Che02m]. ride [Zak02a]. rider [Sch09b]. Riding [Ros02, Zim08]. Right [Ano05d, Zub09, Che02c, Che03k, Che09f, Ano06-57]. rights [LL03a]. Rile [Mer04]. Ring [Mul06e]. Rip [Per08h]. RISC [Per02c]. Rise [Gui08d, Pat09g, Avi02a, GG05, PP08c, Pat09j, Rio02, Voe06b]. rises [Gol02c]. rising [Fai05c, KS04, Mul05b, PT06]. Risk [Moo08c, Fra05b, Zor06c]. risks [Fos00b]. Risky [Mus09]. rivals [Rou01]. River [Sni00]. Road [Ano05i, Jon07g, Voe07b, Wee05, Ano04-68, Jon05d, Lay01]. roadblocks [Laz09]. Robert [Gep01b, BC05, Luc05e]. Robin [Sch09g]. Robinson [Ano08-72]. Robocopters [Dum09a]. Robogolf [Kar09i]. Robosoccer [Cas01k]. Robot [Gui07a, Gui08a, Woo08, Ano04-79, Bro08, Cas05a, Kri06]. Robotic [Ano09-73, Kum07d, Wu08b, Ano02g, GG05, Jon05d, Kum02c]. Roboticist [Sch09g, Wal05]. robotics [Cas01k]. Robots [Ano06-60, Che07f, Dum08a, Gui08f, Rie02c, Ano09-74, Fer01, Gep04g, TD06, YZD02].

Rock

[Kar08q, Kus09a, Mer04, Kar09h, And09e]. Rocket [Pat09h, Pie05, Obe01b, Obe02b, Zak02b]. Rockets [Cas02a, Ups09g]. rocks [Per04b, Ups09d]. Rodney [Bro08]. rods [WU01]. Roger [P.08]. roll [Fai09f]. rollout [Fox01]. Room [Sch09h, Cas01e]. roost [Swe08d]. rope [Obe00c]. Rosen [Che01b]. rot [Luc07c]. rough [Jon05d]. Rouler [Ano05-57]. Roundabout [Dig06b]. router [Rob09a]. routers [Mul01a]. rover [Kum06b]. Rowman [Kum01a]. royal [San01]. Rubber [Jon07g]. Rubinstein [Per00c]. Rule [Fai07a, Bar03, Che02e, Coe04]. rules [Ros03b, Swe08d, Luc05e]. run [Ano04-81, Bre02d]. Running [Ano09-74, Eas03, Hoo05a]. runs [Jon04c, Luc02d]. Runway [Dum08b]. Rupturing [Jon08a]. rush [Gep05d]. Russia [Cas02a, Chi02a, MBP00, Obe00b, Swe08d, Zak03, Zak05a]. Russian [Ano05j, Ham02, Zak02b, Zak09]. Russians [Zak05b].

S [Ade08e, Ade09c]. Saddle [Sni00]. Safe [Cas06h, Cor00, FM00, Kum03a]. Safer [Dig06c, Jon02a]. safety [Gol03c]. sag [Jon06e]. Sages [Kar09j]. Saget [Kar08a]. SAIC [Swe00a]. Sake [Per04j]. salaried [Kum01a]. salaries [Cos03, Cos06, PP08c]. sale [Luc05b]. sales [Che02g]. salesman [Sel08d]. Salinee [Ada08]. Salvi [Ano08-72]. Sam [Cas02d]. Sampler [Ano07-53]. Samsung [Boy04d]. Sanctions [Kum03f]. Sand [Col05, Heg08b, Kum04e]. sandbots [GKK09]. sanity [Sko03]. Santa [Sch09d]. Satellite [Bar06c, Per08f, Ano02c, Voe06b]. Satellites [DiG06d, Obe09a, Ups07h]. satisfaction [Zor01b]. Saturn [Cas06m]. Saucer [Ano07r]. Saudi [Ups09c]. Save [Cas06m]. Saved [Cas04d]. Saving [Obe00c, OZC03, VN03]. Savvy [Che08b].

Say [Ano09-75, Mcf09d]. says [Gep04f]. Scanning [Dum08a]. Scarcity [Moo08c, SW04]. Scheme [Ups08d, LL04a, Swe06b]. Schiffer [And08a]. schmesto [Fra02c]. Schmidt [Kum01a]. Scholarly [Kum03f]. Schön [Hel02d]. School [Ano06f, Dur05, Kar02j, Kar05b, Kum05c, Sel03]. schooler [Jon05h]. schools [Bel00]. Schubin [K.08]. Schuster [Che01a]. Sci [Cas06g, Six05e, Six05f, Six05c, Six05d, Six07a, Six07b]. Sci-tech [Cas06g]. Sci/Tech [Six05e, Six05f, Six05c, Six05d, Six07a, Six07b]. Science [Aar06b, Ano07-50, Cas05b, Joh07, PP09b, Pat09h, Swe04b, Ups07f, Zim08, Ade09d, Cas06g, Che02f, Che05e, Kum06d, Peh01b, Six05a, Ups07f]. sciences [Ano01n]. Scientific [And08a, Cas01l, Kar09j, Wol02a]. scientist [App04a]. scientists [Cas01d, Das01]. SciTech [Six05b]. Score [Swe07c]. Scores [Ano07-59, Fos01c]. Scots [Ros00, Gol02d]. Scott [Kar03e]. scramjet [Gui04d, Ano04-82]. Scraping [Bar01]. Scratchy [Ano04-76]. Screen [Win06, Wol02b]. screener [Dig06a]. Screening [Ano08-91, Bre02c, Jon08b]. screens [Jon09h, PP09d]. Scriber [Gol01b]. Scrutiny [Gui06b]. Sculptor [Kar08n]. sea [Gep03e, Kum04b, Man00, Swe09c, Ups09f]. seagoing [TD06]. seal [Ano09-73]. seams [Ano01m]. Search [Dig08a, Ros06a, Ups07a, Cas04e, Mor01b]. Searches [And08b, Tes08c]. Searching [Rom08]. seas [Bre04b, ES04]. Season [GP06a]. Sebastopol [Cas02d]. Second [Ano02k, BTR00, Chi02b, Kap02, Kar03c, Per01c, Rie03a]. secrecy [Ros00]. Secret [Ano03i, Fos00b, Gol02e, Ste09]. Secrets [Che00a, Ros01a]. sector [BR02]. Securing [Tea08]. Security [Ade08a, NB07, Sko03, Swe06d, Ano01-64, Ano02l, CR02, Cas02h, Che00a, Che02j, Che05i, Com01b, Mul06e, Rie00a, RWZ05, Ros06d]. See [Ano08-70, Gwy00, Ups07e, Cos05c]. seeds [Per01e]. Seeing [App01a, Far09, Mac08, Ano04a, PP09a]. seek [Bla05a]. sees [Chi09]. Segregation [Ges04]. Seismic [Kum07b]. Seizures [Hug09]. Self [Ade08g, Swe04g, Fos09c, Rie03c]. self-censorship [Rie03c]. Self-healing [Ade08g]. self-made [Fos09c]. Self-sustaining [Swe04g]. sell [Boy08a]. Selling [Che04d, Fai05a]. Semcad [Fos07b]. Semiconductor [CKVY08, Bla03b, Gol02d, Hel05d, Moo06g, Sav02b]. Semiconductors [Ade08k, BM05, Gep03e, Gep05d, Moo02c]. Senate [BBS^{+}01]. Sense [Jon07a]. sensibility [KC04]. sensible [Das02]. Sensing [Ups07h]. sensor [Che05p, Kum04c, Ups08i]. Sensors [Bet06, Dum08b, KC04, Ups08h]. sensory [HNPC05]. Sentry [Kum07d]. Separate [Che04n, Koo03]. Separates [Ano06-62]. Sept [Ano04z, Ano05y, Ano06-29]. September [Swe00a]. Sequencing [Pat09a, Hod00]. serendipity [Vin04]. Sergio [Mor06]. Serpent [Hun06]. servers [Wri05]. Service [Per01f, Aar06a]. services [KH00, Oli02]. Serviss [Cas05b]. set [Ano02i, Mul06e, Tes03c]. sets [Moo01e]. Setting [Dum09b, Moo01e]. settlement [Bre02a]. Seven [Che05k, ES04]. sexual [Zor01b]. Shake [BBS^{+}01]. Shaky [LJ03]. Shale [Heg08b]. Sham [Mak02]. Shaman [Kar06b, Kar06c]. Shame [Mak02]. Shannon [Per09h]. shapers [Che02b]. shapes [Kum01a]. shaping [Per04h]. Share [BBS^{+}01, Jon03g]. sharing [Ano00-44]. sharply [Ano02a, HK04, PP08c]. Shedding [Mul05c]. Sheen [Swe08f]. Shelden [Per04c]. Shielding [MFD00]. shifting [Cal00b]. ship [Bre04b]. Shirky [Sch08c]. Shock [Kro07]. shocked [Cor00]. Shocking [Ano03j, Moo06h]. shook [San09]. Shoots [Sin06c]. Shop [Sch09a]. shops [Web05]. Short [Ade09c]. shortage [Zor06b]. shortages [Sin05]. Shot [Bre04c, Svo06b]. shots [Web04]. should [Kar02e]. Show [Kar08p, GP00, Smi03]. Showdown [PC05, Ups06]. Showing [JA05, Kar08k]. shown [Hel02d]. shows [Ros03c]. Shrink [Per06a]. Shrinking [Arn09, Mul06c]. Shrinks [Das09b]. Shuji [Rio07a]. shut [Tea09]. shuttle [Zak05a, Zak05b, Obe03b, Obe08]. side [Sel01c, Tur09b]. Sidling [Cer00]. Sight [Jon04e]. Sights [Ano00-44, Che00b, CC00b, Che00c, Che00d, Che00e, Che00f, CC00a, Che01i, Com00f, Com00g, Jon00a]. Sign [Ano04-77, Wes09, Che02l]. signal [Eyr01, LL02a, Mir00a]. Signals [Dig08c]. Signs [Dig08c, Vin08, Sel01a]. Sigrid [Ade08h]. SIivers [Jon08d]. silence [Gui05f, Sch09i]. Silicon [Ano04-83, Gep05d, Jon08d, MG00b, Moo06e, Rio07b, Ros03c, Ros06f, Ye08, Cas02c, Cof05, Fai04a, PK05, Sav04, Sin07b]. silicon-based [Sav04]. SiLK [Gol03e]. Silver [AS09]. Simon [Che01a]. Simple [Per08j, Mul05a, SA09]. simplicity [Nor08]. Simulating [Kap01b]. simulation [Bal00]. Since [App03b]. single [Moo06f]. Singular [Nor08, ST05]. Singularity [Ano08-77, Ano08-99, Han08, Has08e, Kur08, Vin08]. sinking [Che09j]. sinks [Bla02c]. Sins [Hel05c]. Site [Ano06-54, Ano02k]. Sites [Cas01n, Che01d, Che01k]. situation [Ano04-83]. Size [Has08f, Dig08b, Gep05c]. Sizing [Per07e, Tre04]. sizzle [Ano04-83]. skies [Bre02b, Bre03a, Obe03a, Obe05, Pat09b]. skills [And09d, Bel00, Eas00]. skin [Lev02]. Skull [GG06b]. sky [McD04, Per02b, Per05k, Rom08, Sch09f]. Skyscraper [Jon07b]. slammed [Che03e]. Slams [Swe06d]. Slashdot [Kus07c]. slate [Kel01]. Slicing [McF04c]. slightly [Cos05c]. slip [Ano04-97]. Slipping [Ano01-51]. Slips [Gol03e]. slope [Ano01-51]. slouches [IS01]. Slow [Ano02l, Mor01a, Sch00]. slump [Bla02c]. Small [Cas04b, Der08, Kar09k, Ano09-62, Hoo05c]. small-town [Hoo05c]. smaller [Gep05a]. Smart [Bet06, Che03h, Kol04, Rub07, SBB08, Sno03, Voe07a, Gol04a, Ups08h, Zor04b]. smarter [Zor02a]. Smashing [Min06c]. smears [Ray01]. Smell [Ano08-73]. Smell-o-Phone [Ano08-73]. Smiles [Ano06j]. Smithsonian [Cas01f]. smoke [Cur01]. smooth [Bre00c]. Snap [Gui08a]. Sneak [Boy06c, Ade09a]. Sniff [CS08a]. Snipers [Che08e]. Snow [Ano08l, Ano02h]. Snowclone [Mcf08b]. Snowsuit [Ano06o]. Soar [Ano07a]. Soars [Bar04b]. Soccer [Per08f]. Social [And08b, McF07d, Ups08e, Che03h, Dix05, Kus03e]. Social-isms [McF07d]. society [Gol01c]. soft [Gep04f]. Software [Ano04-75, Ano05-58, Ano06-61, Ano08-95, Cas05f, Cas06j, DR00c, Eic00, Fos00a, Fos01b, Fos01e, Fos03b, Fos04b, Fos05a, Fos06, Kar08c, Kle05b, Kus08c, Lai03, Mir00a, Mir00b, Pfl00, Pra01, Pus00, Tes08b, Cas01d, Cas06a, Cha05, Cha04, Che01e, Che05n, Com00e, Fra05b, Gol03a, Gol05e, HBB04, Hol06, Hoo05b, Kle05a, Kor00, Kus05c, Ort09, Ros05a, Zac05, Cas02d, Gol01b]. Solar [Ano07-58, Fai08c, Moo03b, Per08j, Ste08, Ups08b, Wei04, Fai09f, Gol05b, MFD00, PP05b, Zor02g, Fai03a]. Solar-Cell [Fai08c]. soldiers [Mac02, PP06b]. Solid [Boy08b, Gol06e, Mul02b, Zor02f]. Solid-State [Boy08b, Gol06e, Mul02b, Zor02f]. Solution [BCGM07, Per04j, Rio07b, Joh00a, PK05, Ups09b]. solutions [Sta01]. solves [Wan02]. Solving [Ano05-48, De 01, Gui08h]. Some [Moo09d, Jon05a, JA05]. Something [Ano08h]. Son [Raj08]. song [Che04d, Kus03c]. Sonnenfeldt [Swe00e]. Sony [Kus06a]. soon [Che05h, Das03, Gol06e, Gui04b, Kap02]. sop [ES04]. sore [Jon04e]. Sorry [Cas06h, Che05c, Koo03, Obe00b]. Soul [Voe08c, Kum01a]. soul-battering [Kum01a]. Sound [Che08e, Jon09f, Kum06a, Zim08, Zor08b, Den01, Kra01a, Per05l]. sounds [Obe03c]. Source [Ano04-75, Ano08-94, Cha07b, Das08b, Fra05b, And08e, Car01, Cio01, Kus05c, Ray01, Wol02a]. sources [LTN05]. South [Che05l, Swe08e]. Soviet [Obe08, Swe00d]. Soyuz [Obe03c]. Space [Che09g, Kum05d, Lew01, Mas02, Obe07a, Obe08, Sim01, Swe07a, Zak03, Bre01m, Cas01j, Cas04g, DGKM05, Edw05, Ham02, Jac05, Joh00a, Obe01a, Obe01b, Obe04, PP08h, Sel01b, Zak02a, Zak05a, Zak05b, Zub02, Obe00a, Obe03b]. space-based [DGKM05]. Spacecraft [Ano04-84, Dig04]. Spaceflight [And08c]. Spacesuit [Ano06o]. spacious [Obe03a]. Spam [McF04c, Che06b]. spammers [Ros06d]. spark [Sel04b]. Spaser [Das09a]. spawn [Gep03c]. speakers [Kra01a]. Speaking [McF03f, McF04d, McF04b, McF04e, McF04c, McF05b, McF05c, McF05d, McF05e, McF06a, McF06b, McF06c, McF06e, McF06f, McF06d, McF07a, McF07b, McF07c, McF07e, McF07d, Mcf08d, Sel00b, Sel00d, Sel01c, Sel01e, McF07f, Mcf08a, Mcf08b, Mcf08c, Mcf08f, Sel00a, Sel00c, Sel00e, Sel00f]. Speakout [Cor00, IN00, KB01, MHLS00, Peh01b, SC00, TS01]. Special [KSC^{+}05, Swe00a, Ano09-36, Per04h]. species [SR00]. Spectral [Ano01b, Ano01a, Ano01-51, Ano01-77, Ano02m, Ano03k, Ano03l, Ano03m, Ano03n, Ano04-85, Ano04-86, Ano04-87, Ano04-89, Ano04-90, Ano04-88, Ano04-91, Ano04-92, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano05-55, Ano05-56, Ano05-58, Ano05-57, Ano05-59, Ano05-60, Ano05-61, Ano06-52, Ano06-53, Ano06-54, Ano06-56, Ano06-55, Ano06-59, Ano06-58, Ano06-57, Ano06-60, Ano06-61, Ano06-62, Ano07-50, Ano07-51, Ano07-52, Ano07-53, Ano07-54, Ano07-57, Ano07-55, Ano07-56, Ano07-58, Ano07-59, Ano08-34, Ano08-78, Ano08-79, Fos05b, Has01, Has08b, Has08e, Has08c, Has08d, Has08f, Joh07, Jon06h, Mur07, Per08g, Voe08d, Zor06a, Has09b, Per08b]. Spectrum [Ano01-28, Ano01-29, Ano01-30, Ano07-53, Ano09-76, Bar01, Lan03b, Luc01d, RSS00, SW04, Hir04, Ano00o, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano03o, Ano04i, Ano04j, Ano04l, Ano04k, Ano04m, Ano04n, Ano04o, Ano04p, Ano04s, Ano04t, Ano04q, Ano04r, Ano04u, Ano04v, Ano04-31, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05m, Ano05n, Ano05o, Ano05p, Ano06p, Ano06q, Ano06r, Ano06s, Ano06u, Ano06t, Ano06w, Ano06-31, Ano06v, Ano06-32, Ano06x, Ano06y, Ano06-33]. Spectrum [Ano06z, Ano06-28, Ano06-27, Ano06-29, Ano06-30, Ano07y, Ano07-44, Ano07-43, Ano07-38, Ano07-46, Ano07-39, Ano07-45, Ano07-40, Ano07-41, Ano07-48, Ano07-47, Ano07-42, Ano07s, Ano07t, Ano07x, Ano07u, Ano07v, Ano07w, Ano08-58, Ano08-57, Ano08-49, Ano08-50, Ano08-51, Ano08-52, Ano08-53, Ano08-59, Ano08-54, Ano08-55, Ano08-56, Ano08q, Ano08r, Ano08-79, Ano09-53, Ano09-54, Ano09-55, Ano09-56, Ano09l, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09e, Ano09f, GR09, Has09b, Hir05]. speech [Per05l]. Speed [Boy08b, Cas04c, EK00, Gep02b, Heg08a, Kum01f, LW00, Sav02a]. speeding [Hol03]. speeds [Boy04c, Ead00, Rob00]. Speedup [Moo09c]. Speedy [Bre04a]. spenders [Hir03]. Spending [Sin07a, Hir04]. spent [Pre05e]. SPIN [Zor01e, Ano08-66, BBS^{+}01, Hel06b, Svo07, And09c]. splash [Gep05e]. spontaneous [DR00b]. Spore [Kus08a]. sports [Ano04-68, Hoo05a]. Spot [Che08b, Dig08c, Fos07b, Bre00b, Bre00h, Gol04a]. spotlight [LL02b]. Spots [Ano08-95, Moo09d]. Spotting [Che08e]. spray [Ano08o]. Spree [Sin07a]. spring [Ano00-44]. Sprint [Che08g]. sprout [Ano02q]. spun [Gui06e]. Spurn [Cha06a]. spurts [Rie01c]. spying [Cas03j]. Spyware [McF05d]. Squabble [Fai08c]. squad [Zor08a]. squeezed [Ano01n, Gar03]. squirts [Kar02i]. SR [Swe00a]. SR-3776 [Swe00a]. St [Ano05-56]. stage [Tes03c]. Stakes [Ano05-61]. Stakkope [Gep04b]. Stalked [Voe06b]. Stalled [Bar06b, Ros08d]. Stallman [Cas02d]. Stand [Ano06i, Hel05c, Rie02c]. standard [Ano02i, Che05f, Che06g, Fra02b, LL04b, QP05]. standard-bearer [QP05]. Standardizing [Pec08]. Standards [Bar06b, Swe00a, BJ02, Moo01e, MO00]. Star [Ade08h, Ano05-42, DGKM05, Ham02]. Star-crossed [DGKM05, Ham02]. Stargazing [Jon05g]. starlight [LH03]. starring [Kar08l]. Stars [Voe06a, App01a, Kum05a]. Start [BBS^{+}01, Cha06a, LJ03, PP08b, Ros03d, Sin07b, Coe04, Per01e, Rio02, Voe09c]. Start-Up [BBS^{+}01, Coe04, Per01e, Rio02]. Start-ups [Cha06a, Sin07b]. Starts [Dig08a, Jon09c, Bre01e, Ros09c]. State [Boy08b, Per02d, Fos00b, Gol06e, Mul02b, SK00, Zor02f, Sav02b]. State-of-the-art [Per02d]. States [Avi02a, Bar03, Cos05c, JB05, Kar03a, Min06c]. static [Gep04f]. Station [Cas07a, OZC03, Obe00a, Zak03]. stats [Jon05a]. stay [Pat09e]. Steady [Fai03c]. Steal [Hoo05b]. Stealing [Pre05g]. Steam [Mur09]. steampunk [Gui08e]. steel [Mul06e]. Stellar [Ano07m, Obe06, Sch09h]. Step [Ter07, Moo05c]. steps [Hel01a]. Steve [Cas01d, Per04h, Per00d]. Steven [Gol01b, Ups08f]. Stewart [Fea09, Fea09]. Still [Ros08a, Rie00a, Tie04]. stimulation [Moo06h]. stimulus [Pat09e]. Stinkers [Ano08-95]. STM [Ano04f]. Stock [Ano08-95]. Stock-Market [Ano08-95]. Stop [Man09, Ree09]. Stopgaps [Ano01-46]. storage [Com00d, Fai03c, GG02, Pre05e, Zor01c]. store [MBP00]. Storm [Ano06-58, GP06a, Moo03b]. storms [MFD00]. Story [Ano04b, Ano04c, Ano04d, Ano05a, Ano05c, Ano05d, Ano05g, Ano05b, Ano05e, Ano05f, Ano05h, Ano05j, Ano05i, Ano05k, Ano05l, Ano06f, Ano06g, Ano06h, Ano06c, Ano06d, Ano06i, Ano06k, Ano06j, Ano06l, Ano06e, Ano06a, Ano06b, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07b, Ano07c, Ano07d, Ano07x, Ano08e, Ano08s, Per09h, Ano08j, Ano08f, Ano08g, Ano08h, Ano08b, Ano08c, Ano08d, Ano08-89, Ano09n, Ano09b, Ano09-62, Cas01d, Che01a, Hoo02]. Stowaway [Rom09b]. strand [Pre05c]. Strange [Cas07b, Hel05d]. Strategic [Tes09b, Kar02b]. strategies [Gol01c]. Strategy [Sim01]. Streaming [Che01i, Che04m]. Streamlining [DY01]. Street [Rag09, Voe07a]. Street-View [Rag09]. streetcar [Jon09b]. streets [Cas03e, Jon00d, Kum01c]. stress [Moo06b]. stretched [Swe00g]. Stretches [Sin03a]. Stricter [Bar06b]. stride [Gep01e]. Strikes [Cas05b, Moo03b, Kum01j]. Striking [Che01j]. strings [Jon06f]. strip [WU01]. Stripe [Pol06]. Struck [Ade08h]. struggle [Mor01a]. Struggles [And08a, Cha04]. stuck [Gep04d]. students [LL05a, Ros04a]. studies [Amb01a, Kar05b]. Studio [Gui09d]. Study [Ano04-75, Kum04c]. Stuff [Ano03n, Sel03, Bre01d]. Stupid [Voe08e, Zor09a]. stupidity [Zor09a]. Stupifies [Kar09c]. style [DR00a, Ead00, LL05a, Swe00d]. Submersible [Gui08a]. submicron [Wal00]. submillimeter [Kum05d]. Success [CS08a, Cer00, Che01i]. successfully [LL04b]. Sudoku [Aar06b]. suffers [LL03b]. Suffix [Mcf09d]. sugarcane [Moo06a]. Sullivan [Per04h]. Sultan [Per05l]. Summer [Ano09-57, Boy03a, Kar02f, Sin05]. Sun [Zor02g, Ano08h, Gol02f, And09e, Gep05e, Wei09b]. Sunshine [Ano08-60]. Super [Bre04b, Cas01m, Zor05, Day02, Ano08p]. superatoms [McC05]. Supercheap [Cas03l]. Supercomputer [Ano03l, Cas03l, Gep01e, Gui05d, Gui06d, Moo06f, Per04a, Sin08, Wal09c]. supercomputers [Cas01m, Moo08b]. superconducting [Swe06a]. superconductivity [Ade09c, Ade09c]. Superconductor [BTR00, Swe07e, Bre04b, Bro02b, DHGR01, Das02]. Superficial [Wal07b]. Supermarket [Bla04]. Supertubes [Avo04]. supplies [Rie03f]. Supply [Moo08c, Eng01, Swe00g]. support [Kar03a, Kus03f, Mil01]. Supremacy [Kus07c]. Surf [Gol04d]. surgeon [SM00]. surgery [Jon05c, Jon09e]. surprise [Ano01n]. surprises [Ano03b]. surround [Kra01a]. Surveillance [DiG06d, Swe06d, Ano04-97, Che01h, Kus04a]. survey [App03b, Hir05, Ros04a, GP06b]. Surviving [Kar09k]. Survivor [Sel08b]. suspicion [Han00]. sustaining [Swe04g]. Suttle [Obe03b]. Swarm [Nak04]. sweatshop [Fos01a]. Swedish [BBS^{+}01]. Sweepstakes [Swe04d]. sweet [Moo06a]. swimmin [TD06]. Swimming [Kum07e]. Switch [Ade08d, Gep05c, Kar05a]. switching [Ano04-67]. Sword [Ano03k]. Syn [Fai07c]. synchronous [Swe06a]. syncing [Cos05a]. syndrome [McF02b]. Synthetic [Lev02, Vin04]. System [LJ03, Sho05c, Tes04, Ano00b, Ano02d, Ano04-64, Boy06d, Gol05b, Kum01a, Min06a, MFD00, MO00, Rie03b, Ros05c, Ros09c, Sin02, Tum06, Swe08c, Wei02a]. system-on-package [Tum06]. Systems [BBS^{+}01, Ano02p, BF05, Boy03b, Bre02b, Cas01h, Dig06c, Gui05b, JP06, Ken00, Oli00, Rie02b, San01].

T [FGBM07]. T-rays [FGBM07]. Table [Ano00-68, Ano00-69, Ano00-70, Ano00-71, Ano00-72, Ano00-73, Ano00-74, Ano00-75, Ano00-76, Ano00-77, Ano00-78, Ano00-79, Ano01-52, Ano01-53, Ano01-54, Ano01-55, Ano01-56, Ano01-57, Ano01-58, Ano01-59, Ano01-60, Ano01-61, Ano01-62, Ano04j, Ano04k, Ano04t, Ano04r, Ano04w, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-93, Ano05z, Ano05-27, Ano05-28, Ano05-67, Ano05-68, Ano05-66, Ano05-62, Ano05-63, Ano05-64, Ano05-65, Ano06-30, Ano06-63, Ano06-65, Ano06-66, Ano06-67, Ano06-68, Ano06-69, Ano06-70, Ano06-64, Ano07u, Ano07v, Ano07w, Ano07-60, Ano07-61, Ano07-64, Ano07-65, Ano07-66, Ano07-67, Ano07-62, Ano07-63, Ano08q, Ano08r, Ano08-80, Ano08-81, Ano08-82, Ano08-83, Ano08-84, Ano08-85, Ano08-86, Ano08-87, Ano08-88, Ano09-77, Ano09-78, Ano09-79, Ano09-80, Ano09-81, Ano09-82, Ano09-83, Ano09-84, Ano09-85]. Tablet [Wol02b]. Tabletop [Min06d, Jon04d]. tackle [Wol02a]. tactics [Cal00b, KK02]. tag [Per05p]. Taikonauts [Obe01a]. Taiwan [Chi05, Chi06, Chi07, CM09, Gep05d]. Taiwanese [Ano08-95]. Take [Jon05f, Kar01c, McF06e, Rom07, Voe07b, Fai05a, Heg09a, Per05e, Swe04c]. taken [Che04k]. takeoff [Ano02l, Gui06c]. Takes [Dum07b, Jon07f, Aar05a, Ano04-48, Ano04-68, App03a, App04a, Che03l, Che08f, GFB03, Lan03a, Lay01, Moo05c, Per01f]. Taking [Ano02n, Jon06i, LH03, Nak04, SS05, Zak02a]. Tale [Per09i, Jon05b, Tei02]. talent [Sin03b]. Talented [Ano05k]. Tales [Rio07a, SC06a]. Talk [Bla06, Che02i, Kum02b, Bla08, Che06g, Jon06h, Sel08d, Zor06a]. Talking [Ano01-63, Ano09-86]. Tall [McF02b]. tangle [Ano04-47]. Tangled [Hel05e]. Tank [Jon07h]. Tapping [Kum01h]. Tar [Heg08b]. Targets [Per06a]. Taser [Ups07d]. Tata [Sin08]. Tavaran [Ada08]. tax [Cha09a]. taxes [Cas03h]. teaches [LL05a]. Teaching [Kel01]. team [Ano04-60, Hel03]. Tech [Ano05-51, Bla02c, Cas06l, Cha06a, Chi06, Dum07a, GP06b, Heg08b, Kat09, KSC^{+}05, Lev08, Per05f, Six05e, Six05f, Six05c, Six05d, Six07a, Six07b, Voe07c, Voe08e, Ano04-64, Ano05-70, Cas06g, Che09c, Kar08i, Kum01c, McF03b, McF05f, NH05, Ros04b, Shr01, Ups09c, Voe04, Voe05, Voe06c, Voe09b]. Technica [Six05f]. Technical [Ano00-43, Kor00, Luc08c, Par03c, Sel01e, Eas00, Kar08b, Kus03f]. Technically [McF03f, McF04d, McF04b, McF04e, McF04c, McF05b, McF05c, McF05d, McF05e, McF06a, McF06b, McF06c, McF06e, McF06f, McF06d, McF07a, McF07b, McF07c, McF07e, McF07d, McF07f, Mcf08d, Mcf08c, Mcf08f, Sel00a, Sel00b, Sel00c, Sel00d, Sel00e, Sel00f, Sel01c, Sel01e, Mcf08a, Mcf08b]. Technique [Sav03a, SSM07]. techniques [Kar02e, Moo06h, Per05h]. Techno [Kum02c, Ros03a]. techno-cool [Ros03a]. technocrat [Ano06-52]. Technological [Kar09k, Ano05-41, CG04]. Technologies [Swe00a, Ano02o]. Technologies/SAIC [Swe00a]. technologists [Ano05-69, DHGR01, Peh01a]. Technology [Ade08j, Ano00-80, Ano00-81, Ano00-82, Ano00-83, Ano05-61, Ano06-56, Ano08-59, Ano08-79, Ano08-91, AK04, Bar06d, Bre00f, Bre00g, Bre00d, Bre00e, Cas00, Com00a, Com00c, Gep00b, Gep00c, GS00, Has08a, HOD^{+}09, Jon00d, Kap00, Kor00, Kum00b, Lar07, Luc01a, Per00a, Per00b, Rie00b, Rie00k, Rie00l, Roo00, San08, Sco00, Sin03a, Sta01, Swe00f, Swe00h, Ano05-29, App03b, App04b, Boy04c, Cas01g, Che06g, Chi02b, Coe04, DR00b, Ede00, Gol04a, Gol06b, HNPC05, Hel05d, Hoo02, KK02, Kar04b, Kum01f, Kum02c, Kus04b, Lip05, Moo04a, Moo05a, Mor00, NH05, Peh01b, Per01e, Per05l, Rio02, Ros01b, Ros04a, Ros03b, Sin04a, Six05a, Zor01b, Rio07a]. tee [Gol02d]. Telcordia [Swe00a]. Telecom [Ano04-90, BBS^{+}01, Ber03, Che03k, Che09h, GFB03, Lan02, Luc01e, Luc04f, Tes03b]. Telecommunications [Che08g, Swe00a, Gol04d]. Telecommuting [Che09a]. telemedicine [Har02]. Telephone [Bla05b, Ise03]. Telephony [Moc06, Ano04-67, Bar01, TJ00]. Teleporting [Hel04b]. Telescope [Dig08a, Dum07b, Dig08b, Cas04d]. telescopes [Kum05d]. Television [Swe04d, Che05b]. Tell [Ano06-62, Ano06-73, Ano08-100]. Teller [Cas01l]. temperature [Jon06i]. temperatures [Mul05b]. Ten [Ano05-69]. tentacles [Hoo05c]. Terabit [Das03]. terahertz [Ano04-60, Mul02b]. terminology [McF04a]. Termites [Jon07h]. terms [Kum01d, McF05f]. Tero [Che09i]. terrain [Gui05c]. terrestrial [Fox01]. Terror [Ano06-71, Che05m, Zor01d]. terrorism [Ano02n, Kum03a, Tro04]. Terrorist [Fos08b, Sia01]. Terrorists [Gol06c, FGBM07]. Teska [Che08b]. Tesla [CS08b]. Test [Bre00f, Gui07b, Gui07e, Jon06a, Jon08f, Tes04, Ano04-82, Bre01e, MO00]. tested [Ack09a]. Testing [Ano05-60, Che01k, Fos01e, Ber02, Bla02a, O’C01, Svo06b]. tests [LL03a, Rie03b, Rie01d]. tether [Joh00a]. Texas [Kus09b, Tes04]. Text [Lar07, Che02i]. textbooks [Neb03a]. Texts [Dum08a, Neb03a, Neb03b]. Thanks [Ade09g]. theaters [Per05k]. Their [Gep04c, Boy04a, Gol01d, Kum01a, LL02c, Pav07, Per05g, US03]. them [Fai05a]. themselves [Ano04-46, Che06g, Pes02]. Theories [Cas07c]. theory [GG04]. There [DiG07c, Kum03b, Per07b, Ano02p, Boy07, Rob03a, Zor02d]. Thermal [Per08j]. thieves [Dum09b]. thin [Fai03a, Swe00g]. thin-film [Fai03a]. thin-stretched [Swe00g]. Thing [GS07]. Things [Ano06l, Cas04b, Ano02o, Fos08a]. Think [Ano06-55, And08c]. Thinking [Ano01-64, Luc03f]. Third [Cas07h, Oli02]. Third-generation [Oli02]. Thirst [Zac07]. Thirty [Jon00b, Jon00e, Jon00f, Jon00g, Jon00h, Jon00i, Jon00j, Jon00k, Jon00l, Jon00m, Jon00n, Jon00o]. Thomas [Par03a, Per07f]. Thorsteinsson [Gui09b]. those [Kus02c]. Thought [And07a]. thoughts [Luc07d]. Thread [Cas07c]. Thread-Bare [Cas07c]. threat [Ano09-63]. threats [Kum03a]. Three [GFB03, Gui08f, Ano09-65]. Throwing [Kar09l]. thumb [Ros03b]. thumbs [Cas02b]. thwarts [Ade09a, SW05]. thx [Che08c]. Thyroid [Amb01c]. Thyroid-blocking [Amb01c]. thyself [BD09]. TIA [Che03i]. ticket [Rie03d]. ticking [Che02j]. tidal [Lan03a]. tide [PT06]. tides [Mul05b]. ties [Ano08-89]. Tiger [Cas05f]. Tighter [Swe03b]. Til [Jon09g]. Time [Ano07g, Gei04b, Kum09a, Min06c, Ros08b, Son09a, SC07, Tur09b, Ups08g, Che02l, Gui05a, Hut01, Kap02, Kus02a, Moo05a, Nah01]. times [Kar01a]. Tinny [Kim05]. tintinnabulator [Kum05b]. Tiny [Ano08-90, Gui04c, Min03]. Tips [Sel08f]. Tired [McF07f]. tires [Bre01c]. Titan [Ano04-95]. Titans [Kat09]. today [Ede00]. Tokyo [Boy03a, Boy06d, Nak04]. Tolls [Ano05-52]. Tom [Cas01f]. Tomorrow [Per04i, Ede00, Gol01d]. Tongue [Ups07g]. Too [Der08, Gol04e, Gol06e, McF02c, Gol06f, Kum05c]. Took [Che02j]. tool [Rou01]. Tools [Bre01f, Bre01g, Bre01h, Bre01i, Bre01j, Bre01k, BJC01, Bre03c, Bre04c, Cas04h, Cas06h, Cas06i, Cas06j, Cas06k, Cas06l, Cas07h, Cas07j, Cas07i, Che04e, Che07g, Gei04a, Gei04b, Gei05, Gol04e, Jon04f, Jon05g, LCS04, Per04j, Per05n, Per05m, Per05o, Per08h, Per08i, Rie01h, Ros08b, She01a, Son08b, Son08c, Swe05a, Tea08, Wal05, Wal07a, Wal08b, Wal08c, Wal08a, Wu08b, Ano00-30, Ano00-28, Ano00-29, Ano08-63, Che08d, Kum01b, Moo06h, Rie00c, Rie00d, RBJ00a, RBJ00b, Rie00e, Rie00f, Rie00g, RB00a, RB00b, SS08, Shr09]. toothless [Ast02]. Top [LL04b, Ros03a, Voe04, Voe05, Voe06c, Voe07c, Voe08e, Voe09b, App04b, Hir03, Sav02b]. topsy [Mul01b]. topsy-turvy [Mul01b]. Toronto [Cas01e]. tos [Kar01b]. Toshiba [Boy04a]. Total [Che05n]. Touch [Bre03c, Jon09h, Gol06f, JP06, PP09d, Sul05]. Touches [Kum08b]. tough [Cos04, Gui05b, Swe01b]. toughest [EM02, Min03]. Tour [Ups07f]. tourists [Zak02a]. tower [Kum05b]. towers [Ano02k]. town [Hoo05c, LL04a]. Toy [Per09h, And09c, Ano00-28, Ano00-29, Cas01e, Wal01b]. Toys [Bre01f, Bre01g, Bre01h, Bre01i, Bre01j, Bre01k, BJC01, Bre03c, Bre04c, Cas04h, Cas06h, Cas06i, Cas06j, Cas06k, Cas06l, Cas07h, Cas07j, Cas07i, Che04e, Che07g, Gei04a, Gei04b, Gei05, Gol04e, Jon04f, Jon05g, LCS04, Per04j, Per05n, Per05m, Per05o, Per08h, Rie01h, Ros08b, She01a, Son08b, Son08c, Swe05a, Tea08, Wal05, Wal07a, Wal08a, Wu08b, Ano00-30, Ano08-63, Che08d, Per08i, Rie00c, Rie00d, RBJ00a, RBJ00b, Rie00e, Rie00f, Rie00g, RB00a, RB00b, SS08, Shr09, Wal08b, Wal08c]. Track [Jon06b, Jon07f, Dum09b]. Tracking [Per00e, Per05p, Gui05b]. Trade [Ano04-72, Cha09a]. Trademark [Kar02b, Cas02f]. tradeoffs [Luc06c]. Trading [Gui07e, Ano05-40, Mai03]. Traffic [Wei02a, Bre01l, Bre01n, Cic01, Jon01b, Rie01f]. Trafficking [Mcf08e]. Tragedy [Ano06-47, Luc06e]. Trail [Ano06j, Ano07-59, Ken00]. Train [Man09, Hol03, Ros02]. trainer [SM00]. Trains [Man08]. Transactions [Ano04-75]. transfer [Bas05]. transformed [Cas03k]. transforms [KK02]. Transistor [Ano08-75, Ano04-60, EM02, Gep02a, Gep03b, Gol03d, Hel03, HF06, Rio04, Rio05b, Zor01e]. transistor-like [Hel03]. Transistors [Ade07c, Ano05-49, Das09b, PP08h, Ano04-69, BM05, EM02, Gep00h, Mul05c, Sav08a, Sav09a]. Transit [Jon07d]. transition [Ano09o]. Transmeta [GP00]. transmission [Hel04a, Hel05c, Smi03]. transmitter [Sel02b]. Transparent [BM05, Mor01b]. Transportation [Bre00g]. Trap [SSM07, Tes07a, Kum04e, Moo05b]. Trapped [San08]. trapping [Min03]. trash [LL04a]. Travel [Che09a, Sel08f, Nah01, Obe04, Bre01a]. Traveler [Pat09k]. Traveling [PP06b]. Travels [Ade08e]. Treasured [Neb03b, Neb03a]. treat [Moo05c]. treaties [Ban02]. treatment [LL06]. treatments [Ups08i]. treaty [Ast02]. trees [Ano08-71]. trenches [Cos04]. Trends [Swe08b, AK04]. triage [Luc01b]. trial [LL05b]. Trials [Cas05f]. trick [Obe00c]. Tries [Gol05c]. trillion [Gui04b]. trillion-color [Gui04b]. triple [LL03b]. Tritium [Swe03c]. trouble [Che02h, PP09d]. Troubled [Bar06c, CM09]. Troubles [ISV01, GFB03]. troubling [SK00]. trumpeted [Tei02]. Trust [Kar09c, Ber02]. truth [Coh09]. Try [Cas02h, Cas06g, Wal07c]. Tsunami [DiG06d, Ros09c, Ros05c]. Tu [Ano04-88]. Tubular [Ano05-49]. Tuesday [Ade08e]. Tulane [Jon07i]. tumors [Pre04]. Tunable [Bru02, Sav09a]. tunes [Ups08i]. tunnel [BC05]. Tunneling [Ade07b]. turbines [Swe04c]. turbo [Gui04a]. Turbocharged [Ano04-90]. turbulence [Jon04c]. turn [Aar05a, Bre01b, Che03l, Jon03a, Lan03a, Swe04c, Ube00]. turn-on [Jon03a]. Turned [Kar08f]. Turning [BC04, Per04c, Zor01a]. Turns [Moo07c, Ano04-33, Jon05h, Kar08l]. turvy [Mul01b]. tutor [Che05e]. TV [Ano01-27, Ano02c, Bla05b, Cas06l, Che09d, Cla05, Fox01, Gui04b, Jon05b, LL05b, Moo05a, Per04i, Per05c, Per07e, Per08f, Ras05, Smi03]. twinkle [LH03]. Twitter [McF07b, Luc09e]. Two [Das09b, Kra01a, McF06e, Ano08i, Cas03k, Luc09c, Rie02c]. Two-Laser [Das09b]. Type [Boy04b, Hel03]. Typhoons [Dig08c].

U.S.

[Ade08i, Ade09d, Ano02c, Bar04b, Bar06b, Bar06d, Ham02, Jon03f, Kap02, Kum03d, Kum03f, Kum06d, LL04b, PP08c, Pat09e, Sho05c, Swe03d, Swe05c, Ups07h, Voe09c]. U.S.Military [Fai04d]. UCITA [Kan02]. ugly [Cas01c]. UK [Bre01l, Mul05b]. Ultimate [Ade08k]. Ultracapacitors [Sch07, Zor05]. ultrasound [KyOK09]. Ultraviolet [Ade09h, Hut01]. Ultrawide [Jon04g]. Ultrawideband [Bra06, Jon04g]. Un-assuming [Has08e]. unable [RM03]. unbreakable [Mul02a]. uncertain [Kar01a]. Uncertainty [Kar06a]. unclog [Che02b]. Uncommon [Kus06e]. Unconventional [ZM01]. undersea [Fai05c, Man00]. Understanding [Moo00c]. Undue [Fos00b]. unemployment [Avi02a, Che09c]. Unequal [Che04n]. unfortunate [Swe04c]. Unhooking [Moo01g]. unit [Rie01d, Sel01c]. Unite [Dum09a, Rod03, Kar04c]. United [Avi02a, Bar03, Cos05c, JB05, Kar03a, Min06c, Wei02a]. universal [KP09]. Universe [Gui09b, Rie01i, Gep01a, Nah01]. universities [Bla05a]. university [LL05a, Kar08d]. Unix [Cas03i]. unknown [Ano06-52]. unleash [RSS00]. Unleashed [Swe06e]. unlit [Fai09f]. Unplugged [Pat09c, McF04a, Str03]. unreadable [Ano04-80]. unrecognizably [Wal04]. Unruly [Ano04-91, Fai04c]. Unsafe [SMAS06, Ash00]. Unseen [Ups07e]. Unsqueezed [Son09b]. Unsticking [Das08c]. Unsuspected [DHGR01]. unswitched [Ano04-67]. Unsystematic [Luc06f]. Untangling [Ano08-91]. unthinkable [GH03]. unusable [Mul02b]. Unveiled [Rie01i]. Unveils [Boy04b]. unwanted [Che01b]. Update [Ade08j, Ade08i, And08g, And08f, Ano08-75, Ano08-92, Ano08-93, Ano08-94, Ano08-95, Ano08-96, Ano08-97, Boy08b, Che08f, Das08a, Dig08a, Dig08b, Dig08c, Dum08b, Fai08d, Gui08a, Jon08e, Jon08f, Jon08g, Kum08b, PP08f, PP08g, PP08h, Pec08, Per08j, Ros08c, Sav08a, Sin08, Swe08e, Ups08g, Ups08h, Ups08i, Ade08g, And08e, Bla08, Boy08a, Das08c, Fai09c, Heg08a, Heg09a, Jon08b, Jon09f, Moo08d, PP09d, Ros09c]. Upgrade [Luc09b, Ano02d]. upgrades [Jon06e]. Upper [Sni00]. Ups [Pav07, Avi02b, Cha06a, Sin07b]. Upset [Bra06]. USB [Jon06f]. use [Che05h, Das01, Sin02]. user [Moo02a, Per05i, Swe00a, Woh05]. uses [Ano00-44]. Using [Mul02b, Bre04b, Gol01d]. utility [Bre00c]. utilization [SW04]. utopia [Che06a]. UV [BBS^{+}01, Min06d].

v [Cas02f, TP06]. Vacation [Ade08e]. vacuums [Gep01a]. Valley [Ano04-83, Cas02c, Sin07b]. Valuable [Che08b]. value [BOT06]. valve [Per04i]. vanishing [Gep02a]. vast [Wal04]. Vatican [Hel05c]. Vault [Ano08-68]. VectorWise [Moo09c]. Vegas [Cas05g, Kus06f]. vegetables [Fai09e]. Vehicle [Ter07, Cur01, Zbi05]. Vehicles [Fai04d, Ano01-47, Jon03e, Jon05f, LM01, Rie01c, Voe05, Zor02g]. Venkataraman [Sin04a]. venture [KB01, SW05, TS01]. verdict [Par02c]. Verso [Gol01c]. Versus [Che03j]. Vertical [Ade07c, Bal04, Gol01f]. Very [Ade09c, Kus08c]. Vexed [Cas06i]. VI [Ros09d]. Vibrations [Ano05-55]. vibrator [Zor01b]. Victim [Jon07i]. Victor [Swe05d]. Victory [Che08a, Has08f, Bow03, Par02a]. Video [Cas06m, Kar09m, Ano04-97, Kus02d, Per02e, Rot01a]. VideoSave [Ano04-97]. vie [Jon06d]. View [Rag09, App04b, KB01, Sch09h, TS01]. village [App03a]. villages [Fai09f]. Villas [Gol04c]. Vinci [Kar04a]. Vintage [Rot01a]. Virginia [LJ03, Ros02]. Virtual [And07b, Ano05-60, Ban02, Jon08f, Kus08c, Gol05e, Jon06d, McD04, SM00]. Virtual-Reality [Jon08f]. Virtuality [Gol01c]. Virtually [Che06f]. Virtuoso [Gui09d]. viruses [Cas01b, Che02a]. Vis [Obe08]. Visas [Kum03d]. Vision [Ups07g, Dig06c, Luc04f]. visual [Jon04e]. Visualize [Fos04b]. Visualizing [Cas01n, OW01]. vital [Wal04]. Viva [Cas05g]. Vivendi [Ano02s]. voice [Che05k, Den01, Kar08l]. voids [Gep01a]. VoIP [Che05o]. Vol [Ano04j, Ano04t, Ano04r, Ano04-31, Ano04w, Ano04z, Ano05r, Ano05w, Ano05-28, Ano05-67, Ano05-68, Ano05-66, Ano06u, Ano06w, Ano06-32, Ano06z, Ano06-27, Ano06-65, Ano06-66, Ano06-67, Ano06-68, Ano06-69, Ano06-70, Ano06-64, Ano07t, Ano07-61, Ano07-64, Ano07-65, Ano07-66, Ano07-67, Ano07-62, Ano07-63]. volcanic [Pat09f]. Voltage [Son08c, SNB06]. volume [Ano04k, Ano04-28, Ano04-29, Ano04-30, Ano07v, Ano07w]. Vorgan [Kar09k]. Vote [Che07d]. voter [Cos02a]. Voting [Ano04-88, Che06d, Rot01b, And08e, Che01k]. vs [Ano04-75, Che04h, FGBM07, HOD^{+}09, Jon05a, Kar07b, Kus09b, McF07f]. Vulnerabilities [Fos08b]. vulnerable [Tei01]. Vying [Per02e].

W [Ano04-75, Sni00, Swe00e]. wages [Cos05c]. wait [Oli02]. Waiting [Ano08-98, Bel06, Ros05c, Ros08a]. waits [Son09a]. walk [Pos00]. walked [Luc01f]. walking [Hei02]. Walks [Wu08a]. wallet [Boy05]. Wandering [Kus03b]. Want [Kar08e, Che05p]. Wanted [And07c, Ano01-77, Cla01]. wants [Jon00d, KA01]. WAP [Kun01]. War [Ano08-78, Che03a, Kar07c, Sel05b, Ups08c, Zor06c]. ware [MMPJ03]. Warehouse [Gui08f]. Warfare [Cha07b, Ast01, KK02, Kum01i]. Warming [Ups08b, Mor01a]. Warms [Hel06a]. Warning [Enr07, BF05, Ros05c]. Warrior [Cas01l]. wars [Cal00a, Cal00b, Per03b]. Was [And07a, Ano04-97, Per04j, Bow03]. Washington [Bre00b, Cas01f, Ada00a, Ada00b, Ada00c, Ada00d, Ada00e, Ada00f, Ano00-85, KA00, SA00a, SA00b, SA00c, Voe09c]. Waste [Ade08i, Fai07b, Gol02h, Ano02b, GG02, Mcf08a, PA02, Per06e, Ups09b, Zor01c]. Watch [Ada00f, Bre00b, Che05p, Jon00d, Sav05a, Sav05b, SA00a, SA00b, Ada00a, Ada00b, Ada00c, Ada00d, Ada00e, Ano00-85, Ano05-69, Bre00d, Bre00e, Cas00, Gol04a, Gol04f, KA00, Kus02c, Rie00h, Rie00k, Rie00l, Sav05c, Swe00h, SA00c, TLT02]. watchers [Per05g]. Watching [Ama03, Cas07a, Kus02c, Per00f]. watchman [Obe05]. Watchwords [McF05e]. Water [Ade08j, Moo08d, Jon05h, Ups09e]. Watt [Ste08]. Watts [Fai08d, Rob03b, Ups08b, Jon05h, Per06f]. Wave [Ano06h, Jon08e, Bel06, Ros05c]. Waves [Gli00, Zim08, Jon05h, Jon09f]. way [Boy04a, Che01c, Che09c, Dig06b, FS04, Jon08b, Mul04a, Rob03b, Svo04]. Ways [Cas07b, Kap01a, LL05a]. weaker [Sav08b]. weapon [Kum01i, Svo06b]. Weaponry [Bar04b, Ade08e]. Weapons [Bre01m, Ano04-63, Cha08, DGKM05, GH03, Kum02d, Swe03c, ZM01]. Wear [Gol07d, Heg09b, Dit00]. Weather [Bar06c, Per00e]. Weave [Hel05e]. Weaving [Che02k]. Web [Ano00-44, Bre01c, Cas01n, Che01d, Che01i, Che01k, Che02k, Che03a, Che04h, Hel05e, Kum01i, McF06e, Ano09-87, Che04j, Ano04-98, BBS^{+}01, Che00b, CC00b, Che00c, Che00d, Che00e, Che00f, CC00a, Che02l, Com00f, Com00g, Ead00, Jon00a]. Web-Based [BBS^{+}01]. webbed [Pra01]. Websights [And07c, Boi06, Che01l, Che03j, Che04f, Che06f]. Weed [Ano06e]. Weight [Ano05-43, PP06b]. Weinberger [Ade08e]. weirdness [Ano03f]. Welcome [Agr01, BBS^{+}01, Jon00p, Wei02b, App04c]. Well [Pat09k, Per06c, Ano01-31, PP08e, Sel02a, Ups08c]. Well-connected [Pat09k]. well-oiled [Ano01-31, Ups08c]. wellstone [Cas03m]. went [Che02m]. Were [Ano04-73, Cas01d, Luc03f]. Western [LL05a]. Westinghouse [Par03a]. wet [Ano04-34]. whammy [LL03b]. Whatever [Ano08-79]. Wheel [Ano08-76]. Where [Ano06l, Ano07a, Ano09-87, Cas07d, GP06a, Jon07g, Kar02g, Kar03d, Kar09m, Luc04f, Ano04-86, Cur01, Sel06d, WMM03]. which [Che01c]. while [Cos05c]. Whirligig [Ups07f]. whistle [Kum04f]. whistle-blower [Kum04f]. White [Fos01a, Kum00b, Per05p, Per06b]. White-collar [Fos01a]. Who [Ano01-78, Ano02p, Ano08-99, Che04m, Gol05e, ISV01, Kar08k, KA01, Kus03f, LL03c, Lin08, Luc02d, Luc07f, Mur09, Per06e, Rio06a, Rob03a, Sel06c, Wei03, Wol09, Cas01d, Kus02c]. Whole [GS07]. Wi [Ano02q, Che03f, Che03l, Che04n, Che05h, Che06g, LL04b]. Wi-Fi [Ano02q, Che03f, Che03l, Che04n, Che05h, Che06g, LL04b]. Wicked [Luc09f]. wide [Ano09-87, Tea09]. Wideband [Str03]. Widening [Gui08b]. wider [App04c]. width [Ano05-70]. WiFi [Che08f]. Wiki [McF06d]. Wild [Ano06l, Ano08-72, Man02]. Will [Ano02r, Ano06-60, BBS^{+}01, DiG07c, Kap02, Kum03f, Par02c, Per08f, Voe09c, Wol09, Ade09f, Che05j, Che09k, Che09l, Hut01, Jan01, LL03c, Mer01, Sch09d, Sin04b]. William [Che05i]. Williams [Cas02d]. WiMax [Che04n]. win [Gol01d, Par03c]. Wind [Fai08a, Fai08d, Gui08g, Swe04e, Chi02b, Das05, Fai03c, Fai09c, Hoo05a, Kap02, Man02, McF05f, Swe04c]. Winding [Wee05]. window [Che01a]. Windows [Rob00]. Wing [Has08f, Ade09d]. Winged [Has08f]. Winging [Bre01n]. Winner [Ade08k, Che08g, Gui08h, Kus08c, Shr09, Swe08f]. Winners [Ano04-99, Ano05l, Ano06-62, Ano07-68, HD03, Ros09d]. Winning [Ano05-61, Ups08b, Kap01a, Kar08c]. Wins [Sel06c, Swe04d, Bro02b, Sav02b]. Wire [Man08, Bre01b, Che05f, FH01, Per04e]. Wired [McF07f]. Wireless [Ade07b, Ano01-45, Ano01-49, Moo06d, SR02, Sin03a, Sin06b, Boy05, Cas05g, Che03l, Che03m, Che06e, Che06g, Ell01, Gep03b, Jon06f, Kum04c, Kum05c, MM06, Moo01g, Moo06c, Oli02, Ups08h]. wireless-G [Che03l]. Wiretapping [Ups07i]. Wiring [Ano08-97, Hoo05c, PP08f, FH01, Rie01d, Wal00]. Wisdom [McF06b]. Wise [Hug02]. Wish [Ano04-73]. within [Ano04-98]. Without [KM07, Sch08c, WG01]. Witness [Kar08e]. Wizard [Per06f, Gep01b]. wizardry [Kus02d]. wizards [Cas01d]. Woes [Ups07i, SB01]. Woman [Ano07-50]. women [App02d, Kar08d, PP05a, Tie04, Zor01b]. Wonder [Per07b]. Wonk [PP07b]. won’t [Bre02a]. Woodcock [Per05a]. Word [Che07g, Che09f, Kur08, Moo02c]. Words [McF05f, Mcf08e]. Wordsmithing [Luc05f]. Work [Ano06-59, Ano07-50, Sel06b, Bre02a, Bre02e, Fos02b, Kum01b, Kus04a, Luc02c, Svo04, SB02, Ups08h, Zak02b]. Workaholism [Sel07c]. Working [Has09e, Zor06c]. workplace [Kum01g, Kus04a]. Works [Ups07d, Bre01d]. World [Ade08e, Ano01c, Ano01-50, Ano05-43, Der09c, Gep01e, Jon08g, Kum01h, Kus08c, McF06d, Mur09, Raj08, Sim07, Wei04, Agr01, Ano04-34, Ano09-87, Che00a, CGC04, GG06a, Gep05d, Gol01d, Gol04c, Has09e, Jon00p, Jon06d, Kar04c, Kus03d, Mcf09a, Mul01b, Per01f, Rie03b, San09, Zor01d, NH05]. Worldcom [Ano02s]. worms [RWZ05]. worry [Sin04b]. worse [Cos03]. Worsens [Ano01-32]. worst [Fos08b]. worth [Ano01-47, LM01]. worthwhile [Fos02a]. Would [Ano04-92]. wouldn’t [Com00d, Wal01a]. wows [DHGR01]. wrangler [Cas05a]. wreckage [Ort09]. Wright [Ano03n]. Wringing [Jon05h]. Wrist [Gei05]. Write [CG04]. Writers [Ano04h]. Writing [Cas05c, Eas00]. Wrong [Che05c, Heg08b, Ber03, BOT06, Cha08, Che02m, CG04, Koo03, Pre05f, Ano06-57]. WWW [Fos01d].

X [Bre00h, Fos07b, Cas01c, Cas01l]. XBox [Che09e]. Xs [Gep03a].

ya [Kus03f]. Yahoo [Ano06-72, Che06b]. Yahooligan [Ano06-72]. Yangtze [Swe06e]. yardstick [Bel00]. Year [Ano04-91, FZ03, Has01, Obe01b, Cas02c, Sin06d]. Years [Ade08c, Jon00b, Sel06a, Jon00e, Jon00f, Jon00g, Jon00h, Jon00i, Jon00j, Jon00k, Jon00l, Jon00m, Jon00n, Jon00o, Kus02a, Ros04b]. York [Bre01a, Cas01d, Che01a, Che01b, Gep01a, Gep01b, Gol01b, Mul06e, Nah01, Sav02b, Zor01a, Ano06n, Jon00d]. Yoshihiro [Gep04g]. Young [Kar04c, Kar09c]. you’re [And08c]. Yourself [Ano07-51, LCS04, Loe06, Wal07a]. YouTube [Kar09m]. You’ve [Mor06]. Yucca [GG02].

Zagorodnov [Swe05d]. Zap [Ano09-88, SNB06]. Zapping [Moo05c]. Zero [Luc08e, Per07d, Per09e]. Zink [Per09i]. Zipf [Luc08f]. Zittrain [Ree09]. Zn [Rie01e]. Zn-air [Rie01e]. Zone [Gui05f, Mul06e, Zor06c, Kum07d]. Zooming [Ano05-50]. Zoos [Dig07d].

References

Aaronson:2005:LMM

[Aar05a] L. Aaronson. For love of money — malicious hacking takes an ominous turn. IEEE Spectrum, 42(11):17–19, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aaronson:2005:GJM

[Aar05b] L. Aaronson. Getting a jump on malaria [health care]. IEEE Spectrum, 42(9):16, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aaronson:2006:PPC

[Aar06a] L. Aaronson. Parallel processor [community service]. IEEE Spectrum, 43(5):61–64, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aaronson:2006:SS

[Aar06b] L. Aaronson. Sudoku science. IEEE Spectrum, 43(2): 16–17, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Abrams:2003:DB

[Abr03] M. Abrams. Dawn of the ebomb. IEEE Spectrum, 40(11): 24–30, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2009:IIT

[Ack09a] E. Ackerman. Interplanetary Internet tested. IEEE Spectrum, 46(7):13–14, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ackerman:2009:NBB

[Ack09b] E. Ackerman. New-breed browsers are harder to hack. IEEE Spectrum, 46(9):19, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:2000:WWa

[Ada00a] J. Adam. Washington watch. IEEE Spectrum, 37(1):24I2– 24I3, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:2000:WWb

[Ada00b] J. Adam. Washington watch. IEEE Spectrum, 37(3):24I4, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:2000:WWc

[Ada00c] J. Adam. Washington watch. IEEE Spectrum, 37(4):24I6, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:2000:WWd

[Ada00d] J. Adam. Washington watch. IEEE Spectrum, 37(5):24I4–24I6, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:2000:WWe

[Ada00e] J. Adam. Washington watch.

IEEE Spectrum, 37(6):32I4–32I5, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:2000:WWf

[Ada00f] J. Adam. Washington watch.

IEEE Spectrum, 37(12):28I2, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adams:2008:STP

[Ada08] D. Adams. Salinee Tavaran — power ranger [dream jobs 2008]. IEEE Spectrum, 45(2): 42–43, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2007:P

[Ade07a] S. Adee. Power up. IEEE Spectrum, 44(9):18–20, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2007:QTC

[Ade07b] S. Adee. Quantum tunneling creates fast lane for wireless [news]. IEEE Spectrum, 44 (10):18–20, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2007:TGV

[Ade07c] S. Adee. Transistors go vertical [news]. IEEE Spectrum, 44 (11):14–16, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:CDH

[Ade08a] S. Adee. Chip design hardwires security. IEEE Spectrum, 45(1):16, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:DIE

[Ade08b] S. Adee. The data — Iraq electricity, by the numbers. IEEE Spectrum, 45(7):56, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:DYM

[Ade08c] S. Adee. The data: 37 years of Moore’s Law. IEEE Spectrum, 45(5):56, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:HKS

[Ade08d] S. Adee. The hunt for the kill switch. IEEE Spectrum, 45(5):34–39, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:ITM

[Ade08e] S. Adee. If it’s Tuesday, this must be Los Alamos [review of a Nuclear Family Vacation: Travels in the World of Atomic Weaponry (S. Weinberger and N. Hodge; 2008)]. IEEE Spectrum, 45(6):28, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:REB

[Ade08f] S. Adee. Reverse engineering the brain. IEEE Spectrum, 45 (6):51–53, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:SHH

[Ade08g] S. Adee. Self-healing hulls [update]. IEEE Spectrum, 45 (11):16, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:SCS

[Ade08h] S. Adee. Sigrid close — star struck [dream jobs 2008]. IEEE Spectrum, 45(2):30–31, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:UUC

[Ade08i] S. Adee. Update — U.S. critics hope to halt nuclear-waste imports. IEEE Spectrum, 45(7):11– 12, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:UNW

[Ade08j] S. Adee. Update: New water technology headed for parched places. IEEE Spectrum, 45 (4):11–12, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2008:WSU

[Ade08k] S. Adee. Winner: Semiconductors the ultimate dielectrics is … nothing. IEEE Spectrum, 45 (1):39–42, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:CDT

[Ade09a] S. Adee. Chip design thwarts sneak attack on data. IEEE Spectrum, 46(11):16, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:EG

[Ade09b] S. Adee. Empire off the grid. IEEE Spectrum, 46(8): 32–39, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:HSR

[Ade09c] S. Adee. Housewives of superconductivity [review of Superconductivity: A Very Short Introduction (Blundell, S.; 2009)]. IEEE Spectrum, 46(12):21, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:NDU

[Ade09d] S. Adee. New director for U.S. Energy Department’s mad science wing. IEEE Spectrum, 46 (11):15, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:P

[Ade09e] S. Adee. Polaroid 2.0. IEEE Spectrum, 46(2):21, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:RWP

[Ade09f] S. Adee. The revolution will be prosthetized. IEEE Spectrum, 46 (1):44–48, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:TM

[Ade09g] S. Adee. Thanks for the memories. IEEE Spectrum, 46 (5):48–51, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adee:2009:URB

[Ade09h] S. Adee. Ultraviolet radios beam to life. IEEE Spectrum, 46(5):18, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Agre:2001:WAW

[Agr01] P. E. Agre. Welcome to the always-on world. IEEE Spectrum, 38(1):10–13, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aizawa:2005:PEN

[Aiz05] T. Aizawa. Photo essay: Neon in nagoya. IEEE Spectrum, 42(5):22–23, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2004:TT

[AK04] A. Applewhite and J. Kumagai.

Technology trends 2004. IEEE Spectrum, 41(1):14–19, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Amaratunga:2003:WN

[Ama03] G. Amaratunga. Watching the nanotube. IEEE Spectrum, 40 (9):28–32, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Amber:2001:CSM

[Amb01a] D. P. Amber. Core studies make comeback [nuclear engineering education]. IEEE Spectrum, 38 (11):52–53, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Amber:2001:ELH

[Amb01b] D. P. Amber. Extending life by half [nuclear power plant relicensing]. IEEE Spectrum, 38 (11):48–51, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Amber:2001:TBP

[Amb01c] D. P. Amber. Thyroid-blocking potassium iodide in growing demand. IEEE Spectrum, 38(11): 24–27, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2007:ICB

[And07a] M. Anderson. Internet censorship: As bad as you thought it was [news]. IEEE Spectrum, 44 (11):20–22, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2007:RVJ

[And07b] M. Anderson. Resources: Virtual jamming. IEEE Spectrum, 44(7):53–56, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2007:WPH

[And07c] M. Anderson. Websights: Planet hunters wanted. IEEE Spectrum, 44(3):49–52, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2008:BPP

[And08a] M. Anderson. Bringing power to the people (review of Power Struggles: Scientific Authority and the Creation of Practical Electricity Before Edison by M. B. Schiffer; 2008) [books]. IEEE Spectrum, 45(10):26, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2008:GSA

[And08b] M. Anderson. Google searches for ad dollars in social networks. IEEE Spectrum, 45(12):16, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2008:JWD

[And08c] M. Anderson. Just what do you think you’re doing, Dave? [review of Digital Apollo: Human and Machine in Spaceflight (D. A. Mindell; 2008)]. IEEE Spectrum, 45(6):26–28, June 2008.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2008:MMM

[And08d] M. Anderson. Media: The mash monsters. IEEE Spectrum, 45(5):22–23, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2008:OSV

[And08e] M. Anderson. Open-source voting — [update]. IEEE Spectrum, 45(10):13–14, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2008:UCP

[And08f] M. Anderson. Update — crimeware pays. IEEE Spectrum, 45(7):13, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2008:UBB

[And08g] M. Anderson. Update: Buckyballs to boost flash memory. IEEE Spectrum, 45(6):20, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2009:DI

[And09a] M. Anderson. Dam it! IEEE Spectrum, 46(11):21, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2009:GYR

[And09b] M. Anderson. Give your records a break. IEEE Spectrum, 46 (3):24, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2009:NSO

[And09c] M. Anderson. A new spin on an old toy. IEEE Spectrum, 46 (11):18–19, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2009:RCG

[And09d] M. Anderson. RFID chips gain computing skills. IEEE Spectrum, 46(5):15, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2009:SRC

[And09e] M. Anderson. Sun’s Rock CPU could be a gem for Oracle. IEEE Spectrum, 46(6):12–13, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:IMA

[Ano00a] Anonymous. The 2000 IEEE Medalists [Awards]. IEEE Spectrum, 37(6):62–67, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:BGA

[Ano00b] Anonymous. A backward glance.

Arriving at the punched-card system. IEEE Spectrum, 37 (2):58–61, February 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFa

[Ano00c] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 37(1):c1, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFb

[Ano00d] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 37(2):c1, February 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFc

[Ano00e] Anonymous. IEEE Spectrum — [front cover]. IEEE Spectrum, 37(3):c1, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFd

[Ano00f] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37(4):c1, April 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFe

[Ano00g] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37(5):c1, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFf

[Ano00h] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37(6):c1, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFg

[Ano00i] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37(7):c1, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFh

[Ano00j] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37(8):c1, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFi

[Ano00k] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37 (9):c1, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFj

[Ano00l] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37(10):c1, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFk

[Ano00m] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37 (11):c1, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISFl

[Ano00n] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 37 (12):c1, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ISA

[Ano00o] Anonymous. IEEE Spectrum announcement. IEEE Spectrum, 37(12):1, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ca

[Ano00p] Anonymous. Calendar. IEEE Spectrum, 37(1):16–120R18, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Cb

[Ano00q] Anonymous. Calendar. IEEE Spectrum, 37(2):18–23, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Cc

[Ano00r] Anonymous. Calendar. IEEE Spectrum, 37(3):18–80R19, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Cd

[Ano00s] Anonymous. Calendar. IEEE Spectrum, 37(4):16–64R15, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ce

[Ano00t] Anonymous. Calendar. IEEE Spectrum, 37(5):16–24I12, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Cf

[Ano00u] Anonymous. Calendar. IEEE Spectrum, 37(6):16–80R12, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Cg

[Ano00v] Anonymous. Calendar. IEEE Spectrum, 37(7):14–68R10, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ch

[Ano00w] Anonymous. Calendar. IEEE Spectrum, 37(8):16–76R14, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ci

[Ano00x] Anonymous. Calendar. IEEE Spectrum, 37(9):30–80R24, September 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Cj

[Ano00y] Anonymous. Calendar. IEEE Spectrum, 37(10):30–80R31, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ck

[Ano00z] Anonymous. Calendar. IEEE Spectrum, 37(11):28–72R22, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Cl

[Ano00-27] Anonymous. Calendar. IEEE Spectrum, 37(12):24–28I8, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ETTb

[Ano00-28] Anonymous. EEs’ tools toy. IEEE Spectrum, 37(2):78–80, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ETTc

[Ano00-29] Anonymous. EEs’ tools toy. IEEE Spectrum, 37(3):89–90, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ETTa

[Ano00-30] Anonymous. EEs’ tools toys. IEEE Spectrum, 37(1):117– 120, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fa

[Ano00-31] Anonymous. Forum. IEEE Spectrum, 37(1):6–8, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fb

[Ano00-32] Anonymous. Forum. IEEE Spectrum, 37(2):6–8, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fc

[Ano00-33] Anonymous. Forum. IEEE Spectrum, 37(3):8–13, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fd

[Ano00-34] Anonymous. Forum. IEEE Spectrum, 37(4):6–8, April 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fe

[Ano00-35] Anonymous. Forum. IEEE Spectrum, 37(5):6–8, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ff

[Ano00-36] Anonymous. Forum. IEEE Spectrum, 37(6):6–8, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fg

[Ano00-37] Anonymous. Forum. IEEE Spectrum, 37(7):6–9, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fh

[Ano00-38] Anonymous. Forum. IEEE Spectrum, 37(8):6–10, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fi

[Ano00-39] Anonymous. Forum. IEEE Spectrum, 37(9):6–8, September 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fj

[Ano00-40] Anonymous. Forum. IEEE Spectrum, 37(10):6–8, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fk

[Ano00-41] Anonymous. Forum. IEEE Spectrum, 37(11):8–12, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Fl

[Ano00-42] Anonymous. Forum. IEEE Spectrum, 37(12):8–10, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:ITF

[Ano00-43] Anonymous. The IEEE Technical Field Awards [awards]. IEEE Spectrum, 37(8):58–70, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:MUS

[Ano00-44] Anonymous. More uses spring up for Internet file-sharing [web sights]. IEEE Spectrum, 37 (9):80, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Na

[Ano00-45] Anonymous. Newslog. IEEE Spectrum, 37(1):1, January 2000.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Nb

[Ano00-46] Anonymous. Newslog. IEEE Spectrum, 37(2):1, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Nc

[Ano00-47] Anonymous. Newslog. IEEE Spectrum, 37(3):1, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Nd

[Ano00-48] Anonymous. Newslog. IEEE Spectrum, 37(4):1, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ne

[Ano00-49] Anonymous. Newslog. IEEE Spectrum, 37(5):1, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Nf

[Ano00-50] Anonymous. Newslog. IEEE Spectrum, 37(6):1, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ng

[Ano00-51] Anonymous. Newslog. IEEE Spectrum, 37(7):1, July 2000.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Nh

[Ano00-52] Anonymous. Newslog. IEEE Spectrum, 37(8):1, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ni

[Ano00-53] Anonymous. Newslog. IEEE Spectrum, 37(9):1, September 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Nj

[Ano00-54] Anonymous. Newslog. IEEE Spectrum, 37(10):1, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Nk

[Ano00-55] Anonymous. Newslog. IEEE Spectrum, 37(11):1, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBa

[Ano00-56] Anonymous. Recent books. IEEE Spectrum, 37(1):24I4– 24I7, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBb

[Ano00-57] Anonymous. Recent books. IEEE Spectrum, 37(2):24I4, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBc

[Ano00-58] Anonymous. Recent books.

IEEE Spectrum, 37(3):24I7–24I8, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBd

[Ano00-59] Anonymous. Recent books.

IEEE Spectrum, 37(4):64R4, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBe

[Ano00-60] Anonymous. Recent books.

IEEE Spectrum, 37(5):24I8– 64R10, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBf

[Ano00-61] Anonymous. Recent books.

IEEE Spectrum, 37(6):80R16– 80R19, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBg

[Ano00-62] Anonymous. Recent books.

IEEE Spectrum, 37(7):20I6–20I8, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBh

[Ano00-63] Anonymous. Recent books.

IEEE Spectrum, 37(8):76R17– 76R19, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBi

[Ano00-64] Anonymous. Recent books. IEEE Spectrum, 37(9):80R6– 80R10, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBj

[Ano00-65] Anonymous. Recent books. IEEE Spectrum, 37(10):80R10– 80R12, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBk

[Ano00-66] Anonymous. Recent books. IEEE Spectrum, 37(11):72R4– 72R8, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:RBl

[Ano00-67] Anonymous. Recent books. IEEE Spectrum, 37(12):64R18– 64R20, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCa

[Ano00-68] Anonymous. Table of contents. IEEE Spectrum, 37(1):2–4, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCb

[Ano00-69] Anonymous. Table of contents. IEEE Spectrum, 37 (2):2–4, February 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCc

[Ano00-70] Anonymous. Table of contents. IEEE Spectrum, 37(3):2–4, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCd

[Ano00-71] Anonymous. Table of contents. IEEE Spectrum, 37(4):2–4, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCe

[Ano00-72] Anonymous. Table of contents. IEEE Spectrum, 37(5):2– 4, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCf

[Ano00-73] Anonymous. Table of contents. IEEE Spectrum, 37(6):2– 4, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCg

[Ano00-74] Anonymous. Table of contents. IEEE Spectrum, 37(7):2– 4, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCh

[Ano00-75] Anonymous. Table of contents.

IEEE Spectrum, 37(8):2–4, August 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCi

[Ano00-76] Anonymous. Table of contents. IEEE Spectrum, 37(9): 2–4, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCj

[Ano00-77] Anonymous. Table of contents. IEEE Spectrum, 37(10):2–4, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCk

[Ano00-78] Anonymous. Table of contents. IEEE Spectrum, 37(11): 2–4, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:TCl

[Ano00-79] Anonymous. Table of contents. IEEE Spectrum, 37(12):4–6, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Ta

[Ano00-80] Anonymous. Technology 2000 — 1. IEEE Spectrum, 37(1):32, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Tb

[Ano00-81] Anonymous. Technology 2000 — 2. IEEE Spectrum, 37(1):57, January 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Td

[Ano00-82] Anonymous. Technology 2000 — 3. IEEE Spectrum, 37(1):80, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:Te

[Ano00-83] Anonymous. Technology 2000 — 4. IEEE Spectrum, 37 (1):103, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:PF

[Ano00-84] Anonymous. To probe further. IEEE Spectrum, 37(1): 115–116, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2000:WW

[Ano00-85] Anonymous. Washington watch.

IEEE Spectrum, 37(2):24I2, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:BNG

[Ano01a] Anonymous. 2001: Bad news, and good [spectral lines]. IEEE Spectrum, 38(12):9, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:AFS

[Ano01b] Anonymous. After the fall [spectral lines]. IEEE Spectrum, 38(10):9, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ALN

[Ano01c] Anonymous. Always on: Living in a networked world. IEEE Spectrum, 38(1):41, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFa

[Ano01d] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 38(1):c1, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFb

[Ano01e] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 38(5):c1, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFc

[Ano01f] Anonymous. IEEE Spectrum front cover. IEEE Spectrum, 38(6):c1, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFd

[Ano01g] Anonymous. IEEE Spectrum front cover. IEEE Spectrum, 38(7):c1, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFe

[Ano01h] Anonymous. IEEE Spectrum front cover. IEEE Spectrum, 38(8):c1, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFf

[Ano01i] Anonymous. IEEE Spectrum front cover. IEEE Spectrum, 38 (9):c1, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFg

[Ano01j] Anonymous. IEEE Spectrum front cover. IEEE Spectrum, 38(10):c1, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFh

[Ano01k] Anonymous. IEEE Spectrum front cover. IEEE Spectrum, 38 (11):c1, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ISFi

[Ano01l] Anonymous. IEEE Spectrum front cover. IEEE Spectrum, 38 (12):c1, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:BS

[Ano01m] Anonymous. Bursting at the seams. IEEE Spectrum, 38 (1):43, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:BBS

[Ano01n] Anonymous. Bush budget surprise: physical sciences, engineering innovation squeezed [news analysis]. IEEE Spectrum, 38(4):29–30, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Ca

[Ano01o] Anonymous. Contributors. IEEE Spectrum, 38(1):6, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Cb

[Ano01p] Anonymous. Contributors. IEEE Spectrum, 38(2):7, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Cc

[Ano01q] Anonymous. Contributors. IEEE Spectrum, 38(3):7, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Cd

[Ano01r] Anonymous. Contributors. IEEE Spectrum, 38(4):7, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Ce

[Ano01s] Anonymous. Contributors. IEEE Spectrum, 38(5):7, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Cf

[Ano01t] Anonymous. Contributors. IEEE Spectrum, 38(6):7, June 2001.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Cg

[Ano01u] Anonymous. Contributors. IEEE Spectrum, 38(7):7, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Ch

[Ano01v] Anonymous. Contributors. IEEE Spectrum, 38(8):7, August 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Ci

[Ano01w] Anonymous. Contributors. IEEE Spectrum, 38(9):7, September 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Cj

[Ano01x] Anonymous. Contributors. IEEE Spectrum, 38(10):7, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Ck

[Ano01y] Anonymous. Contributors. IEEE Spectrum, 38(11):7, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Cl

[Ano01z] Anonymous. Contributors. IEEE Spectrum, 38(12):7, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:DTD

[Ano01-27] Anonymous. Digital TV divide. IEEE Spectrum, 38(5):10–12, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ESFa

[Ano01-28] Anonymous. Eee spectrum — [front cover]. IEEE Spectrum, 38(2):c1, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ESFb

[Ano01-29] Anonymous. Eee spectrum — [front cover]. IEEE Spectrum, 38(3):c1, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ESFc

[Ano01-30] Anonymous. Eee spectrum — [front cover]. IEEE Spectrum, 38(4):c1, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:EWO

[Ano01-31] Anonymous. Elements of a welloiled machine. IEEE Spectrum, 38(1):65, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ECW

[Ano01-32] Anonymous. An energy crunch worsens. IEEE Spectrum, 38 (1):79, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:EN

[Ano01-33] Anonymous. An eye for the new.

IEEE Spectrum, 38(5):85–86, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fa

[Ano01-34] Anonymous. Forum. IEEE Spectrum, 38(1):10, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fb

[Ano01-35] Anonymous. Forum. IEEE Spectrum, 38(2):8, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fc

[Ano01-36] Anonymous. Forum. IEEE Spectrum, 38(3):10–12, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fd

[Ano01-37] Anonymous. Forum. IEEE Spectrum, 38(4):10–12, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fe

[Ano01-38] Anonymous. Forum. IEEE Spectrum, 38(6):10–12, June 2001.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Ff

[Ano01-39] Anonymous. Forum. IEEE Spectrum, 38(7):10–12, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fg

[Ano01-40] Anonymous. Forum. IEEE Spectrum, 38(8):10–12, August 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fh

[Ano01-41] Anonymous. Forum. IEEE Spectrum, 38(9):10, September 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fi

[Ano01-42] Anonymous. Forum. IEEE Spectrum, 38(10):10–12, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fj

[Ano01-43] Anonymous. Forum. IEEE Spectrum, 38(11):10–12, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:Fk

[Ano01-44] Anonymous. Forum. IEEE Spectrum, 38(12):10–12, December 2001. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:GLW

[Ano01-45] Anonymous. The grande latte wireless experience. IEEE Spectrum, 38(5):9, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:GS

[Ano01-46] Anonymous. Gridlocks and stopgaps. IEEE Spectrum, 38 (1):89, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:HVW

[Ano01-47] Anonymous. Hybrid vehicles are worth it! [paper discussion and authors’ reply]. IEEE Spectrum, 38(5):79–84, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:NL

[Ano01-48] Anonymous. Networked living.

IEEE Spectrum, 38(1):101, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:ODH

[Ano01-49] Anonymous. Optics and DSL are hot, but don’t forget wireless. IEEE Spectrum, 38(1):58, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PNW

[Ano01-50] Anonymous. Philosophy for a networked world [books]. IEEE Spectrum, 38(1):127–128, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:SPS

[Ano01-51] Anonymous. Slipping down the privacy slope [spectral lines]. IEEE Spectrum, 38 (9):9, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCa

[Ano01-52] Anonymous. Table of contents. IEEE Spectrum, 38(1):1–4, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCb

[Ano01-53] Anonymous. Table of contents. IEEE Spectrum, 38 (2):1–3, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCc

[Ano01-54] Anonymous. Table of contents. IEEE Spectrum, 38(3):3–5, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCd

[Ano01-55] Anonymous. Table of contents. IEEE Spectrum, 38(4):3–5, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCe

[Ano01-56] Anonymous. Table of contents. IEEE Spectrum, 38(5):3– 5, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCf

[Ano01-57] Anonymous. Table of contents. IEEE Spectrum, 38(6):3– 5, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCg

[Ano01-58] Anonymous. Table of contents. IEEE Spectrum, 38(7):3– 5, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCh

[Ano01-59] Anonymous. Table of contents.

IEEE Spectrum, 38(8):3–5, August 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCi

[Ano01-60] Anonymous. Table of contents.

IEEE Spectrum, 38(10):3–5, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCj

[Ano01-61] Anonymous. Table of contents. IEEE Spectrum, 38(11): 3–5, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TCk

[Ano01-62] Anonymous. Table of contents. IEEE Spectrum, 38(12):3–5, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TAD

[Ano01-63] Anonymous. Talking about digital copyright. IEEE Spectrum, 38(6):9, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:TAS

[Ano01-64] Anonymous. Thinking about security. IEEE Spectrum, 38 (11):9, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFa

[Ano01-65] Anonymous. To probe further. IEEE Spectrum, 38(1): 157–159, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFb

[Ano01-66] Anonymous. To probe further. IEEE Spectrum, 38(2): 103, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFc

[Ano01-67] Anonymous. To probe further. IEEE Spectrum, 38(3):101, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFd

[Ano01-68] Anonymous. To probe further. IEEE Spectrum, 38(4):103, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFe

[Ano01-69] Anonymous. To probe further. IEEE Spectrum, 38(5):111, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFf

[Ano01-70] Anonymous. To probe further. IEEE Spectrum, 38(6):87, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFg

[Ano01-71] Anonymous. To probe further. IEEE Spectrum, 38(7):86, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFh

[Ano01-72] Anonymous. To probe further.

IEEE Spectrum, 38(8):71, August 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFi

[Ano01-73] Anonymous. To probe further. IEEE Spectrum, 38(9): 88, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFj

[Ano01-74] Anonymous. To probe further. IEEE Spectrum, 38(10):87, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFk

[Ano01-75] Anonymous. To probe further. IEEE Spectrum, 38(11): 95, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:PFl

[Ano01-76] Anonymous. To probe further. IEEE Spectrum, 38(12):95, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:WNE

[Ano01-77] Anonymous. Wanted: Nuclear engineers [spectral lines]. IEEE Spectrum, 38(7):9, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:WPA

[Ano01-78] Anonymous. Who pays for artificial hearts? IEEE Spectrum, 38(8):9–11, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:CMD

[Ano02a] Anonymous. China may draw a sharply lower line on mobile phone radiation. IEEE Spectrum, 39(9):33, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:ECW

[Ano02b] Anonymous. Europe cracks down on e-waste. IEEE Spectrum, 39(5):46–51, May 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:FNU

[Ano02c] Anonymous. FCC nixes U.S.

satellite TV merger. IEEE Spectrum, 39(12):16, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:FOF

[Ano02d] Anonymous. Feds online [fbi it system upgrade]. IEEE Spectrum, 39(9):36–37, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:JFG

[Ano02e] Anonymous. Japanese firms go for biodegradable electronics. IEEE Spectrum, 39(10): 18–20, October 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:LDY

[Ano02f] Anonymous. Leave the driving to your car? IEEE Spectrum, 39(4):9, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:LRA

[Ano02g] Anonymous. Little robotic autos that can. IEEE Spectrum, 39(7):24–25, July 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:NSN

[Ano02h] Anonymous. Neither snow, nor rain, nor anthrax [bioterrorism counteraction]. IEEE Spectrum, 39(9):38–39, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:NGS

[Ano02i] Anonymous. New global standard set for metro networks. IEEE Spectrum, 39(8):21, August 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:ND

[Ano02j]  Anonymous.  Noncombatant

deaths. IEEE Spectrum, 39 (4):32–33, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:SSB

[Ano02k] Anonymous. Second site [broadcast towers]. IEEE Spectrum, 39 (9):40–41, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:STA

[Ano02l] Anonymous. Slow takeoff [airport security]. IEEE Spectrum, 39(9):37–38, September 2002.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:SL

[Ano02m] Anonymous. Spectral lines.

IEEE Spectrum, 39(10):9–12, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:TT

[Ano02n] Anonymous. Taking on terrorism. IEEE Spectrum, 39(9): 35–36, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:WBT

[Ano02o] Anonymous. When bad things happen to good technologies. IEEE Spectrum, 39(12):9, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:WGT

[Ano02p] Anonymous. Who goes there?

[personal identity systems]. IEEE Spectrum, 39(9):39–40, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:WFH

[Ano02q] Anonymous. Wi-Fi hotspot networks sprout like mushrooms. IEEE Spectrum, 39(9): 18–20, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:WIA

[Ano02r] Anonymous. Will im aid the enterprise? IEEE Spectrum, 39(11):9, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2002:WVR

[Ano02s] Anonymous. Worldcom and vivendi reeling from communications crash. IEEE Spectrum, 39 (8):18–19, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:BDB

[Ano03a] Anonymous. Behind deep blue. IEEE Spectrum, 40(6):40, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:BBS

[Ano03b] Anonymous. Big blackout surprises politicians, but not the power community. IEEE Spectrum, 40(9):9, September 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:C

[Ano03c] Anonymous. Contributors. IEEE Spectrum, 40(12):7, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:F

[Ano03d] Anonymous. Forum. IEEE Spectrum, 40(12):8–10, December 2003. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:GP

[Ano03e] Anonymous. A goal, not a purpose. IEEE Spectrum, 40(3):11, March 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:HML

[Ano03f] Anonymous. Hacking matter:

Levitating chairs, quantum mirages and the infinite weirdness of programmable atoms [book review]. IEEE Spectrum, 40 (4):50–51, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:EG

[Ano03g] Anonymous. Is this the executive’s gameboy? IEEE Spectrum, 40(2):54, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:MB

[Ano03h] Anonymous. Mars or bust.

IEEE Spectrum, 40(10):9, October 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:SEY

[Ano03i] Anonymous. Secret enough for you? IEEE Spectrum, 40(4):9, April 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:SPP

[Ano03j] Anonymous. Shocking price power. IEEE Spectrum, 40 (1):89, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:SLA

[Ano03k] Anonymous. Spectral lines: Another double-edged sword. IEEE Spectrum, 40(11):11, November 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:SLB

[Ano03l] Anonymous. Spectral lines boost your iq, build a supercomputer. IEEE Spectrum, 40(7):7, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:SLE

[Ano03m] Anonymous. Spectral lines the end of money for nothing? IEEE Spectrum, 40(7):7, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:SLW

[Ano03n] Anonymous. Spectral lines: The Wright stuff. IEEE Spectrum, 40(12):11, December 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2003:YBI

[Ano03o] Anonymous. This is your brain on IEEE Spectrum. IEEE Spectrum, 40(6):9, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ASE

[Ano04a] Anonymous. The all-seeing eye. IEEE Spectrum, 41(7):8, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:BS

[Ano04b] Anonymous. The back story.

IEEE Spectrum, 41(10):7, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:BSE

[Ano04c] Anonymous. The back story:

Electrical engineering’s identity crisis. IEEE Spectrum, 41 (11):9, November 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:BSL

[Ano04d] Anonymous. The back story:

Lighting up the andes. IEEE Spectrum, 41(12):5, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:BH

[Ano04e] Anonymous. Beat the heat.

IEEE Spectrum, 41(5):38–43, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:BNI

[Ano04f] Anonymous. A beautiful noise [invention of STM and AFM].

IEEE Spectrum, 41(5):50–52, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:BB

[Ano04g] Anonymous. Biometrics boom. IEEE Spectrum, 41(3):13, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:BWW

[Ano04h] Anonymous. Books: What the writers are reading. IEEE Spectrum, 41(8):54–55, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISA

[Ano04i] Anonymous. IEEE Spectrum — April 2004. IEEE Spectrum, 41(4):01, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISD

[Ano04j] Anonymous. IEEE Spectrum — Dec. 2004 — vol 41, no 12 — table of contents. IEEE Spectrum, 41(12):1–3, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISFb

[Ano04k] Anonymous. IEEE Spectrum — feb. 2004 — volume, 41 number 2 — table of contents. IEEE Spectrum, 41(2):1–3, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISFa

[Ano04l] Anonymous. IEEE Spectrum — February 2004. IEEE Spectrum, 41(2):01, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISFc

[Ano04m] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 41(8):01, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISFd

[Ano04n] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 41 (9):01, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISFe

[Ano04o] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 41 (11):01, November 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISFf

[Ano04p] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 41 (12):01, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISJc

[Ano04q] Anonymous. IEEE Spectrum — July 2004. IEEE Spectrum, 41(7):01, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISJd

[Ano04r] Anonymous. IEEE Spectrum — July 2004 — vol 41, no 7 — table of contents. IEEE Spectrum, 41(7):1–3, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISJa

[Ano04s] Anonymous. IEEE Spectrum — June 2004. IEEE Spectrum, 41(6):01, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISJb

[Ano04t] Anonymous. IEEE Spectrum — June 2004 — vol 41 number 6 — table of contents. IEEE Spectrum, 41(6):3–5, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISMa

[Ano04u] Anonymous. IEEE Spectrum — march 2004. IEEE Spectrum, 41(3):01, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISMb

[Ano04v] Anonymous. IEEE Spectrum — may 2004. IEEE Spectrum, 41(5):01, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISN

[Ano04w] Anonymous. IEEE Spectrum — nov. 2004 — vol 41, no 11 — table of contents. IEEE Spectrum, 41(11):5–7, November 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISOa

[Ano04x] Anonymous. IEEE Spectrum — Oct 2004 — contents. IEEE Spectrum, 41(10):3, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISOb

[Ano04y] Anonymous. IEEE Spectrum — Oct. 2004 — table of contents. IEEE Spectrum, 41(10):5, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISS

[Ano04z] Anonymous. IEEE Spectrum — sept. 2004 — vol 41, no 9 — table of contents. IEEE Spectrum, 41(9):3–5, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:IST

[Ano04-27] Anonymous. IEEE Spectrum — table of contents. IEEE Spectrum, 41(8):1–3, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISVa

[Ano04-28] Anonymous. IEEE Spectrum — volume 41, number 3 — table of contents. IEEE Spectrum, 41(3):3–5, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISVb

[Ano04-29] Anonymous. IEEE Spectrum — volume 41, number 4 — table of contents. IEEE Spectrum, 41(4):2–3, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISVc

[Ano04-30] Anonymous. IEEE Spectrum — volume 41, number 5 — table of contents. IEEE Spectrum, 41(5):1–3, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ISMc

[Ano04-31] Anonymous. IEEE Spectrum magazine October 2004 vol: 41 num. 10. IEEE Spectrum, 41 (10):01, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:CPP

[Ano04-32] Anonymous. Carbon pillars power microbatteries. IEEE Spectrum, 41(10):20–22, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:CT

[Ano04-33] Anonymous. CERN turns 50. IEEE Spectrum, 41(10):18, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:CMW

[Ano04-34] Anonymous. Chip making’s wet new world. IEEE Spectrum, 41(5):29–33, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:CCC

[Ano04-35] Anonymous. Clashing climate catastrophists. IEEE Spectrum, 41(5):20–22, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Cf

[Ano04-36] Anonymous. Codebuster. IEEE Spectrum, 41(5):59–60, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Ca

[Ano04-37] Anonymous. Contributors. IEEE Spectrum, 41(1):6, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Cb

[Ano04-38] Anonymous. Contributors. IEEE Spectrum, 41(2):5, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Cc

[Ano04-39] Anonymous. Contributors. IEEE Spectrum, 41(3):7, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Cd

[Ano04-40] Anonymous. Contributors. IEEE Spectrum, 41(4):5, April 2004.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Ce

[Ano04-41] Anonymous. Contributors. IEEE Spectrum, 41(5):5, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Cg

[Ano04-42] Anonymous. Contributors. IEEE Spectrum, 41(6):7, June 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Ch

[Ano04-43] Anonymous. Contributors. IEEE Spectrum, 41(7):5, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Ci

[Ano04-44] Anonymous. Contributors. IEEE Spectrum, 41(8):5, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Cj

[Ano04-45] Anonymous. Contributors. IEEE Spectrum, 41(9):7, September 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:EFL

[Ano04-46] Anonymous. Electrical fuse lets chips heal themselves. IEEE Spectrum, 41(10):16–20, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ETC

[Ano04-47] Anonymous. Energy tangle in China. IEEE Spectrum, 41 (10):12–14, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ETB

[Ano04-48] Anonymous. Europe takes bite out of Microsoft. IEEE Spectrum, 41(5):10–12, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fa

[Ano04-49] Anonymous. Forum. IEEE Spectrum, 41(2):8, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fb

[Ano04-50] Anonymous. Forum. IEEE Spectrum, 41(3):10, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fd

[Ano04-51] Anonymous. Forum. IEEE Spectrum, 41(4):6, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fe

[Ano04-52] Anonymous. Forum. IEEE Spectrum, 41(5):6, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Ff

[Ano04-53] Anonymous. Forum. IEEE Spectrum, 41(6):8, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fg

[Ano04-54] Anonymous. Forum. IEEE Spectrum, 41(7):6, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fh

[Ano04-55] Anonymous. Forum. IEEE Spectrum, 41(8):6, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fi

[Ano04-56] Anonymous. Forum. IEEE Spectrum, 41(9):8, September 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fj

[Ano04-57] Anonymous. Forum. IEEE Spectrum, 41(10):8, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fk

[Ano04-58] Anonymous. Forum. IEEE Spectrum, 41(11):12, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:Fl

[Ano04-59] Anonymous. Forum. IEEE Spectrum, 41(12):6, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:FTD

[Ano04-60] Anonymous. French team demonstrates terahertz transistor. IEEE Spectrum, 41(5):16, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:OB

[Ano04-61] Anonymous. From outsourcing to botsourcing? IEEE Spectrum, 41(5):8, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:FC

[Ano04-62] Anonymous. [front cover]. IEEE Spectrum, 41(1):1, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:GNW

[Ano04-63] Anonymous. The globalization of nuclear weapons. IEEE Spectrum, 41(6):10, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:HTL

[Ano04-64] Anonymous. A high-tech launch system for carriers. IEEE Spectrum, 41(1):24–25, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:EBD

[Ano04-65] Anonymous. In the eye of the beholder [displays]. IEEE Spectrum, 41(5):24–28, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:IIF

[Ano04-66] Anonymous. Indian IT firms connect with Chinese counterparts. IEEE Spectrum, 41(1): 26–27, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:ITS

[Ano04-67] Anonymous. Internet telephony: switching to unswitched. IEEE Spectrum, 41(1):25, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:LTR

[Ano04-68] Anonymous. Lithium takes to the road [all electric sports car]. IEEE Spectrum, 41(12):11, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PT

[Ano04-69] Anonymous. Paper transistors. IEEE Spectrum, 41(8):48, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PEB

[Ano04-70] Anonymous. Photo essay: Ball of light. IEEE Spectrum, 41 (3):26–27, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PEI

[Ano04-71] Anonymous. Photo essay: If I only had a brain . . .. IEEE Spectrum, 41(12):24, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PET

[Ano04-72] Anonymous. Photo essay: Trade rallies. IEEE Spectrum, 41 (10):26–27, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PEW

[Ano04-73] Anonymous. Photo essay: Wish you were here. IEEE Spectrum, 41(2):26–27, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PRY

[Ano04-74] Anonymous. Please remove your glasses. IEEE Spectrum, 41(7): 60, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:POV

[Ano04-75] Anonymous. Progress: Open vs. closed [An Empirical Study of Open-Source and Closed-Source Software Products, by James W.

Paulson et al., IEEE Transactions on Software Engineering, April 2004, pp.]. IEEE Spectrum, 41(8):50, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PSF

[Ano04-76] Anonymous. Progress: Scratchy film. IEEE Spectrum, 41(7):56, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PBS

[Ano04-77] Anonymous. Progress: The bandwidth of sign language. IEEE Spectrum, 41(7):52, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PPW

[Ano04-78] Anonymous. Pundits ponder what 2004 portends. IEEE Spectrum, 41(1):23, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:QRC

[Ano04-79] Anonymous. Qrio, the robot that could. IEEE Spectrum, 41(5):34–37, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:RU

[Ano04-80] Anonymous. Reading the unreadable. IEEE Spectrum, 41 (5):54–56, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:RR

[Ano04-81] Anonymous. Record run. IEEE Spectrum, 41(1):24–25, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SBM

[Ano04-82] Anonymous. Scramjet breaks Mach 7 in NASA test. IEEE Spectrum, 41(5):17, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SVS

[Ano04-83] Anonymous. Silicon valley: sizzle or fizzle? [employment situation]. IEEE Spectrum, 41 (9):15, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SCM

[Ano04-84] Anonymous. Spacecraft converge on mars. IEEE Spectrum, 41 (1):22–23, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLB

[Ano04-85] Anonymous. Spectral lines:

Bringing you the magic. IEEE Spectrum, 41(11):14, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLD

[Ano04-86] Anonymous. Spectral lines: Do you know where your data are?

IEEE Spectrum, 41(12):8, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLE

[Ano04-87] Anonymous. Spectral lines: Engineering and aging: The best is yet to be. IEEE Spectrum, 41(9):10, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLTb

[Ano04-88] Anonymous. Spectral lines: Et tu, e-voting? IEEE Spectrum, 41(10):10, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLM

[Ano04-89] Anonymous. Spectral lines: Making a difference. IEEE Spectrum, 41(2):7, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLTa

[Ano04-90] Anonymous. Spectral lines: Telecom’s turbocharged next act. IEEE Spectrum, 41(3):9, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLU

[Ano04-91] Anonymous. Spectral lines: The unruly grid, one year later. IEEE Spectrum, 41(8):8, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:SLW

[Ano04-92] Anonymous. Spectral lines:

Would you have a neural implant? IEEE Spectrum, 41(4):8, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:TCa

[Ano04-93] Anonymous. Table of contents.

IEEE Spectrum, 41(1):3–5, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:WLL

[Ano04-94] Anonymous. This is what 40 looks like. IEEE Spectrum, 41(1):9, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:TCb

[Ano04-95] Anonymous. Titan calling. IEEE Spectrum, 41(10):28–33, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:PF

[Ano04-96] Anonymous. To probe further. IEEE Spectrum, 41(1): 96–97, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:WSF

[Ano04-97] Anonymous. Was that slip and fall for real? [video surveillance and VideoSave]. IEEE Spectrum, 41(5):18, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:WWW

[Ano04-98] Anonymous. The Web within the Web. IEEE Spectrum, 41 (2):42–46, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2004:WLG

[Ano04-99] Anonymous. Winners losers or grails. IEEE Spectrum, 41 (1):35, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSA

[Ano05a] Anonymous. The back story: Acting to free information. IEEE Spectrum, 42(9):5, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSE

[Ano05b] Anonymous. The back story: An elevated mission. IEEE Spectrum, 42(8):7, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSB

[Ano05c] Anonymous. The back story: Brothers of invention. IEEE Spectrum, 42(11):7, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSDa

[Ano05d] Anonymous. The back story: Dodgeball: Just the right look.

IEEE Spectrum, 42(3):5, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSF

[Ano05e] Anonymous. The back story:

Finding Noyce’s notebook. IEEE Spectrum, 42(5):7, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSG

[Ano05f] Anonymous. The back story:

Gnomes and reactors. IEEE Spectrum, 42(4):7, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSDb

[Ano05g] Anonymous. The back story:

He delivers. IEEE Spectrum, 42(12):7, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSM

[Ano05h] Anonymous. The back story:

Movies, monoliths, and mission control. IEEE Spectrum, 42(7): 5, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSRb

[Ano05i] Anonymous. The back story: On the road again. IEEE Spectrum, 42(10):7, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSRa

[Ano05j] Anonymous. The back story: Russian Renaissance Man. IEEE Spectrum, 42(2):5, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BST

[Ano05k] Anonymous. The back story: The talented Mr Li. IEEE Spectrum, 42(6):5, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:BSW

[Ano05l] Anonymous. The back story: Winners & losers 2005. IEEE Spectrum, 42(1):5, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISa

[Ano05m] Anonymous. IEEE Spectrum. IEEE Spectrum, 42(1):01, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISb

[Ano05n] Anonymous. IEEE Spectrum. IEEE Spectrum, 42(5):7, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISc

[Ano05o] Anonymous. IEEE Spectrum. IEEE Spectrum, 42(6):01, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISd

[Ano05p] Anonymous. IEEE Spectrum.

IEEE Spectrum, 42(8):01, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISDa

[Ano05q] Anonymous. IEEE Spectrum — December 2005. IEEE Spectrum, 42(12):01, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISDb

[Ano05r] Anonymous. IEEE Spectrum — December 2005 — vol 42 num. 12 — contents. IEEE Spectrum, 42(12):1–5, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISFa

[Ano05s] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 42(4):01, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISFb

[Ano05t] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 42(5):1, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISFc

[Ano05u] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 42(10):01, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISJ

[Ano05v] Anonymous. IEEE Spectrum — July 2005. IEEE Spectrum, 42(7):01, July 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISM

[Ano05w] Anonymous. IEEE Spectrum — may 2005 — vol 42, no 5 — contents. IEEE Spectrum, 42(5):2– 5, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISN

[Ano05x] Anonymous. IEEE Spectrum — nov. 2005. IEEE Spectrum, 42 (11):01, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISS

[Ano05y] Anonymous. IEEE Spectrum — sept. 2005. IEEE Spectrum, 42 (9):01, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISTa

[Ano05z] Anonymous. IEEE Spectrum — table of contents. IEEE Spectrum, 42(2):1–2, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISTb

[Ano05-27] Anonymous. IEEE Spectrum — table of contents. IEEE Spectrum, 42(6):1, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:ISV

[Ano05-28] Anonymous. IEEE Spectrum — vol 42, no 4 — table of contents. IEEE Spectrum, 42(4):3–5, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:CGT

[Ano05-29] Anonymous. China at a glance [technology revolution]. IEEE Spectrum, 42(6):26–27, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fa

[Ano05-30] Anonymous. Forum. IEEE Spectrum, 42(2):6, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fb

[Ano05-31] Anonymous. Forum. IEEE Spectrum, 42(3):6, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fd

[Ano05-32] Anonymous. Forum. IEEE Spectrum, 42(4):8, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fe

[Ano05-33] Anonymous. Forum. IEEE Spectrum, 42(5):8, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Ff

[Ano05-34] Anonymous. Forum. IEEE Spectrum, 42(7):6, July 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fg

[Ano05-35] Anonymous. Forum. IEEE Spectrum, 42(9):6, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fh

[Ano05-36] Anonymous. Forum. IEEE Spectrum, 42(10):8, October 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fi

[Ano05-37] Anonymous. Forum. IEEE Spectrum, 42(11):8, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Fj

[Ano05-38] Anonymous. Forum. IEEE Spectrum, 42(12):8, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:FC

[Ano05-39] Anonymous. [front cover]. IEEE Spectrum, 42(2):c1, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:MCA

[Ano05-40] Anonymous. A market for clean air: China adopts emissions trading to curb pollution. IEEE Spectrum, 42(6):60, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:MCT

[Ano05-41] Anonymous. More on China [technological revolution]. IEEE Spectrum, 42(6):67, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:PES

[Ano05-42] Anonymous. Photo essay: Star cruiser. IEEE Spectrum, 42 (1):22–23, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:PEW

[Ano05-43] Anonymous. Photo essay: The weight of the world. IEEE Spectrum, 42(10):24–25, October 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:PC

[Ano05-44] Anonymous. Portraits of China.

IEEE Spectrum, 42(6):20–25, June 2005. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Pa

[Ano05-45] Anonymous. Progress. IEEE Spectrum, 42(2):52, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:Pb

[Ano05-46] Anonymous. Progress. IEEE Spectrum, 42(2):52, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:PAY

[Ano05-47] Anonymous. Progress: An algorithm you can dig. IEEE Spectrum, 42(10):66, October 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:PSK

[Ano05-48] Anonymous. Progress: Solving a knotty problem. IEEE Spectrum, 42(7):58, July 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:PTT

[Ano05-49] Anonymous. Progress: Tubular transistors. IEEE Spectrum, 42(10):65, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:PZN

[Ano05-50] Anonymous. Progress: Zooming in on networks. IEEE Spectrum, 42(10):66, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLA

[Ano05-51] Anonymous. Spectral lines: All the tech in China. IEEE Spectrum, 42(6):8, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLB

[Ano05-52] Anonymous. Spectral lines: Bell tolls for AT&T. IEEE Spectrum, 42(3):8, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLFa

[Ano05-53] Anonymous. Spectral lines:

Fighting indigestion at the Internet buffet. IEEE Spectrum, 42(5):10, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLFb

[Ano05-54] Anonymous. Spectral lines:

Fly guys. IEEE Spectrum, 42 (11):10, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLG

[Ano05-55] Anonymous. Spectral lines:

Good vibrations. IEEE Spectrum, 42(4):10, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLJ

[Ano05-56] Anonymous. Spectral lines: Jack St. Clair Kilby (1923–2005): Engineering monolith. IEEE Spectrum, 42(8):10, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLLb

[Ano05-57] Anonymous. Spectral lines: Laissez les bons ingenieurs rouler. IEEE Spectrum, 42(10):10, October 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLLa

[Ano05-58] Anonymous. Spectral lines: Learning from software failure. IEEE Spectrum, 42(9): 8, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLNa

[Ano05-59] Anonymous. Spectral lines: Natural disaster and the human factor. IEEE Spectrum, 42(2):8, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLNb

[Ano05-60] Anonymous. Spectral lines: Nuclear testing goes virtual. IEEE Spectrum, 42(12):10, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLW

[Ano05-61] Anonymous. Spectral lines: Winning and losing in high-stakes technology. IEEE Spectrum, 42(1):8, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TCa

[Ano05-62] Anonymous. Table of contents.

IEEE Spectrum, 42(1):1–3, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TCb

[Ano05-63] Anonymous. Table of contents. IEEE Spectrum, 42(6):3, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TCc

[Ano05-64] Anonymous. Table of contents.

IEEE Spectrum, 42(8):3–5, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TCd

[Ano05-65] Anonymous. Table of contents.

IEEE Spectrum, 42(10):1–3, October 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TCVc

[Ano05-66] Anonymous. Table of contents — vol 42 no 11. IEEE Spectrum, 42 (11):1–3, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TCVa

[Ano05-67] Anonymous. Table of contents — vol 42, no 7. IEEE Spectrum, 42(7):1–3, July 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TCVb

[Ano05-68] Anonymous. Table of contents — vol 42, no 9. IEEE Spectrum, 42(9):1–3, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:TWC

[Ano05-69] Anonymous. Ten to watch [China’s leading technologists]. IEEE Spectrum, 42(6):37, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:WAD

[Ano05-70] Anonymous. The width of the Atlantic [differences in tech policies]. IEEE Spectrum, 42(5):13, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSa

[Ano06a] Anonymous. The back story. IEEE Spectrum, 43(1):6, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSd

[Ano06b] Anonymous. The back story. IEEE Spectrum, 43(8):7, August 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSCc

[Ano06c] Anonymous. The back story: A career that’s no accident. IEEE Spectrum, 43(9):7, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSG

[Ano06d] Anonymous. The back story: A genius for doing good. IEEE Spectrum, 43(12):7, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSWb

[Ano06e] Anonymous. The back story: A weed grows in Chile. IEEE Spectrum, 43(3):9, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSB

[Ano06f] Anonymous. The back story:

Back to school. IEEE Spectrum, 43(5):7, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSCa

[Ano06g] Anonymous. The back story:

Car culture. IEEE Spectrum, 43(4):9, April 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSCb

[Ano06h] Anonymous. The back story:

Catching a really big wave. IEEE Spectrum, 43(7):7, July 2006.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSH

[Ano06i] Anonymous. The back story: Ham’s last stand. IEEE Spectrum, 43(10):7, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BST

[Ano06j] Anonymous. The back story: On a trail of smiles. IEEE Spectrum, 43(6):7, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSP

[Ano06k] Anonymous. The back story: Pulse prediction. IEEE Spectrum, 43(11):7, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BSWa

[Ano06l] Anonymous. The back story: Where the wild things are. IEEE Spectrum, 43(2):7, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BPB

[Ano06m] Anonymous. The big picture: Batteries included. IEEE Spectrum, 43(9):22–23, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BPL

[Ano06n] Anonymous. The big picture:

Live from New York! IEEE Spectrum, 43(5):22–23, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:BPS

[Ano06o] Anonymous. The big picture:

Spacesuit, snowsuit. IEEE Spectrum, 43(8):18–19, August 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISAa

[Ano06p] Anonymous. IEEE Spectrum — April 2006. IEEE Spectrum, 43(4):01, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISAb

[Ano06q] Anonymous. IEEE Spectrum — aug. 2006. IEEE Spectrum, 43(8):01, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISC

[Ano06r] Anonymous. IEEE Spectrum — contents. IEEE Spectrum, 43 (10):1–5, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISD

[Ano06s] Anonymous. IEEE Spectrum — Dec. 2006. IEEE Spectrum, 43 (12):c1, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISFb

[Ano06t] Anonymous. IEEE Spectrum — February 2006 — contents. IEEE Spectrum, 43(2):1–5, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISFa

[Ano06u] Anonymous. IEEE Spectrum — February 2006 vol. 43 no. 2. IEEE Spectrum, 43(2):01, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISJc

[Ano06v] Anonymous. IEEE Spectrum — July 2006. IEEE Spectrum, 43(7):01, July 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISJa

[Ano06w] Anonymous. IEEE Spectrum — June 2006, vol 43, no 6 — contents. IEEE Spectrum, 43(6):1– 2, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISMb

[Ano06x] Anonymous. IEEE Spectrum — march 2006. IEEE Spectrum, 43(3):01, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISMc

[Ano06y] Anonymous. IEEE Spectrum — may 2006. IEEE Spectrum, 43(5):01, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISN

[Ano06z] Anonymous. IEEE Spectrum — November 2006, vol. 43, no. 11. IEEE Spectrum, 43 (11):c1, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISOb

[Ano06-27] Anonymous. IEEE Spectrum — October 2006, vol. 43, no. 10. IEEE Spectrum, 43(10):c1, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISOa

[Ano06-28] Anonymous. IEEE Spectrum — online. IEEE Spectrum, 43(1):5, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISS

[Ano06-29] Anonymous. IEEE Spectrum — sept. 2006. IEEE Spectrum, 43 (9):01, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:IST

[Ano06-30] Anonymous. IEEE Spectrum — table of contents. IEEE Spectrum, 43(1):1–3, January 2006.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISJb

[Ano06-31] Anonymous. IEEE Spectrum June 2006 — online. IEEE Spectrum, 43(6):5, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISMa

[Ano06-32] Anonymous. IEEE Spectrum magazine — January 2006 — vol. 43 no. 1. IEEE Spectrum, 43(1):01, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:ISMd

[Ano06-33] Anonymous. IEEE Spectrum magazine — June 2006. IEEE Spectrum, 43(6):01, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:CG

[Ano06-34] Anonymous. Cybercrime at a glance. IEEE Spectrum, 43(4):17, April 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fa

[Ano06-35] Anonymous. Forum. IEEE Spectrum, 43(1):8, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fb

[Ano06-36] Anonymous. Forum. IEEE Spectrum, 43(2):8, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fc

[Ano06-37] Anonymous. Forum. IEEE Spectrum, 43(3):11, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fd

[Ano06-38] Anonymous. Forum. IEEE Spectrum, 43(4):10, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fe

[Ano06-39] Anonymous. Forum. IEEE Spectrum, 43(5):8, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Ff

[Ano06-40] Anonymous. Forum. IEEE Spectrum, 43(6):8, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fg

[Ano06-41] Anonymous. Forum. IEEE Spectrum, 43(7):8, July 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fh

[Ano06-42] Anonymous. Forum. IEEE Spectrum, 43(8):8, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fi

[Ano06-43] Anonymous. Forum. IEEE Spectrum, 43(9):8, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fj

[Ano06-44] Anonymous. Forum. IEEE Spectrum, 43(10):8, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fk

[Ano06-45] Anonymous. Forum. IEEE Spectrum, 43(11):8, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:Fl

[Ano06-46] Anonymous. Forum. IEEE Spectrum, 43(12):8, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:GMT

[Ano06-47] Anonymous. German Maglev tragedy. IEEE Spectrum, 43 (11):22, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:IGI

[Ano06-48] Anonymous. IEEE Group Insurance Program [advertisement]. IEEE Spectrum, 43 (11):51, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:MPP

[Ano06-49] Anonymous. Media: Podcast picks. IEEE Spectrum, 43 (10):58, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:MDB

[Ano06-50] Anonymous. Million dollar baby. IEEE Spectrum, 43 (3):22–23, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:NB

[Ano06-51] Anonymous. News briefs. IEEE Spectrum, 43(2):18, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLA

[Ano06-52] Anonymous. Spectral lines:

Archibald putt: The unknown technocrat returns. IEEE Spectrum, 43(4):12, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLC

[Ano06-53] Anonymous. Spectral lines:

Counterfeiting and you. IEEE Spectrum, 43(5):10, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLD

[Ano06-54] Anonymous. Spectral lines: Does NASA need a better launch site? IEEE Spectrum, 43(10):10, October 2006. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLIb

[Ano06-55] Anonymous. Spectral lines: IBM’s new motto: Think … about how others think. IEEE Spectrum, 43(12):10, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLIa

[Ano06-56] Anonymous. Spectral lines: Imagining the future of technology. IEEE Spectrum, 43 (9):10, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLMc

[Ano06-57] Anonymous. Spectral lines: Metcalfe’s Law: Right? Wrong? IEEE Spectrum, 43(11):10, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLMb

[Ano06-58] Anonymous. Spectral lines: Modeling toward accurate storm forecasting. IEEE Spectrum, 43(8):9, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLMa

[Ano06-59] Anonymous. Spectral lines: More fun at work. IEEE Spectrum, 43(2):10, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLR

[Ano06-60] Anonymous. Spectral lines:

Robots can ape us, but will they ever get real? IEEE Spectrum, 43(7):10, July 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLS

[Ano06-61] Anonymous. Spectral lines:

Software development: Game over? IEEE Spectrum, 43(6):10, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:SLW

[Ano06-62] Anonymous. Spectral lines:

What separates winners from losers? You tell us. IEEE Spectrum, 43(1):9, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TC

[Ano06-63] Anonymous. Table of contents. IEEE Spectrum, 43(11): 1–5, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TCVg

[Ano06-64] Anonymous. Table of contents — vol 43, no 12. IEEE Spectrum, 43(12):1–5, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TCVa

[Ano06-65] Anonymous. Table of contents — vol 43, no 3. IEEE Spectrum, 43(3):3–7, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TCVb

[Ano06-66] Anonymous. Table of contents — vol 43, no 4. IEEE Spectrum, 43(4):3–7, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TCVc

[Ano06-67] Anonymous. Table of contents — vol 43, no 5. IEEE Spectrum, 43(5):1–5, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TCVd

[Ano06-68] Anonymous. Table of contents — vol 43, no 7. IEEE Spectrum, 43(7):2–5, July 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TCVe

[Ano06-69] Anonymous. Table of contents — vol 43, no 8. IEEE Spectrum, 43(8):1–5, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TCVf

[Ano06-70] Anonymous. Table of contents — vol 43, no 9. IEEE Spectrum, 43(9):1–5, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:TWN

[Ano06-71] Anonymous. Terror: What’s next. IEEE Spectrum, 43(9):

24–25, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:YYP

[Ano06-72] Anonymous. Yahooligan: Yahoo primes its research pump. IEEE Spectrum, 43(3):61–62, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:YTU

[Ano06-73] Anonymous. You tell us [it advances]. IEEE Spectrum, 43 (1):50–54, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:IAW

[Ano07a] Anonymous. 2007 IEEE Awards [where dreams soar]. IEEE Spectrum, 44(7):c1–c3, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSb

[Ano07b] Anonymous. The back story.

IEEE Spectrum, 44(7):7, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSd

[Ano07c] Anonymous. The back story.

IEEE Spectrum, 44(11):5, November 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSe

[Ano07d] Anonymous. The back story. IEEE Spectrum, 44(12):5, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSA

[Ano07e] Anonymous. The back story: African at heart. IEEE Spectrum, 44(5):5, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSC

[Ano07f] Anonymous. The back story: Chasing the dream. IEEE Spectrum, 44(2):5, February 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSF

[Ano07g] Anonymous. The back story: Face time. IEEE Spectrum, 44(4):5, April 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSG

[Ano07h] Anonymous. The back story: Gold out of dross. IEEE Spectrum, 44(1):5, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSH

[Ano07i] Anonymous. The back story: Her father’s notebook. IEEE Spectrum, 44(3):5, March 2007.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSI

[Ano07j] Anonymous. The back story: Island of mystery. IEEE Spectrum, 44(6):5, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSM

[Ano07k] Anonymous. The back story:

Making the ghosts live. IEEE Spectrum, 44(8):7, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BSW

[Ano07l] Anonymous. The back story:

When chris met fred. IEEE Spectrum, 44(10):5, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BPS

[Ano07m] Anonymous. The big picturestellar performer. IEEE Spectrum, 44(8):22–23, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BPD

[Ano07n] Anonymous. The big picture: 426 000 discarded cellphones. IEEE Spectrum, 44(9): 24–25, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BPB

[Ano07o] Anonymous. The big picture: Bionic hand. IEEE Spectrum, 44(10):22, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BPG

[Ano07p] Anonymous. The big picture: Glacier power. IEEE Spectrum, 44(4):18–19, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BPMb

[Ano07q] Anonymous. The big picture melon blast. IEEE Spectrum, 44 (12):22–23, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:BPMa

[Ano07r] Anonymous. The big picture: My other car is a flying saucer. IEEE Spectrum, 44 (11):24–25, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:ISAa

[Ano07s] Anonymous. IEEE Spectrum — April 2007. IEEE Spectrum, 44(4):c1, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:ISAb

[Ano07t] Anonymous. IEEE Spectrum — April 2007 — vol. 44 — no. 4. IEEE Spectrum, 44(4):1, April 2007. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:IST

[Ano07u] Anonymous. IEEE Spectrum — table of contents. IEEE Spectrum, 44(4):2, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:ISVa

[Ano07v] Anonymous. IEEE Spectrum — volume 44 no 8 — table of contents. IEEE Spectrum, 44(8):3–5, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:ISVb

[Ano07w] Anonymous. IEEE Spectrum — volume 44, number 9 — table of contents. IEEE Spectrum, 44 (9):3–5, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:ISB

[Ano07x] Anonymous. IEEE Spectrum:

The back story. IEEE Spectrum, 44(9):7, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:CIS

[Ano07y] Anonymous. Contents — IEEE Spectrum March 2007. IEEE Spectrum, 44(3):1–3, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fa

[Ano07z] Anonymous. Forum. IEEE Spectrum, 44(1):6, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fb

[Ano07-27] Anonymous. Forum. IEEE Spectrum, 44(2):6, February 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fc

[Ano07-28] Anonymous. Forum. IEEE Spectrum, 44(3):6, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fd

[Ano07-29] Anonymous. Forum. IEEE Spectrum, 44(4):6, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fe

[Ano07-30] Anonymous. Forum. IEEE Spectrum, 44(5):6, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Ff

[Ano07-31] Anonymous. Forum. IEEE Spectrum, 44(6):6, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fg

[Ano07-32] Anonymous. Forum. IEEE Spectrum, 44(7):8, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fh

[Ano07-33] Anonymous. Forum. IEEE Spectrum, 44(8):8, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fi

[Ano07-34] Anonymous. Forum. IEEE Spectrum, 44(9):8, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fj

[Ano07-35] Anonymous. Forum. IEEE Spectrum, 44(10):6, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fk

[Ano07-36] Anonymous. Forum. IEEE Spectrum, 44(11):6, November 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:Fl

[Ano07-37] Anonymous. Forum. IEEE Spectrum, 44(12):6, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIc

[Ano07-38] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 44(3):c1, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIe

[Ano07-39] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 44(6):c1, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIg

[Ano07-40] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 44(8):c1, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIh

[Ano07-41] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 44(9):c1, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIk

[Ano07-42] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 44(12):c1, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIb

[Ano07-43] Anonymous. [front cover — IEEE Spectrum — Feb. 2007]. IEEE Spectrum, 44(2):c1, February 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIa

[Ano07-44] Anonymous. [front cover — IEEE Spectrum — Jan. 2007].

IEEE Spectrum, 44(1):c1, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIf

[Ano07-45] Anonymous. [front cover — IEEE Spectrum — July 2007]. IEEE Spectrum, 44(7):c1, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCId

[Ano07-46] Anonymous. [front cover — IEEE Spectrum — May 2007]. IEEE Spectrum, 44(5):c1, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIj

[Ano07-47] Anonymous. [front cover — IEEE Spectrum — Nov. 2007]. IEEE Spectrum, 44 (11):c1, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:FCIi

[Ano07-48] Anonymous. [front cover — IEEE Spectrum — Oct. 2007]. IEEE Spectrum, 44(10):c1, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:LM

[Ano07-49] Anonymous. Life in the megacity. IEEE Spectrum, 44(6):42–47, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLC

[Ano07-50] Anonymous. Spectral lines: Computer science, a woman’s work. IEEE Spectrum, 44(5):8, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLD

[Ano07-51] Anonymous. Spectral lines: Doit-yourself rfid. IEEE Spectrum, 44(3):8, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLE

[Ano07-52] Anonymous. Spectral lines: Exploring the extreme. IEEE Spectrum, 44(9):10, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLI

[Ano07-53] Anonymous. Spectral lines: Ieee spectrum online: A blogging sampler. IEEE Spectrum, 44(4): 8, April 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLJ

[Ano07-54] Anonymous. Spectral lines: Just re-imagine. IEEE Spectrum, 44(6):8, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLMb

[Ano07-55] Anonymous. Spectral lines: Morse code is dead. Long live Morse code. IEEE Spectrum, 44(2):8, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLO

[Ano07-56] Anonymous. Spectral lines: Online forecast: Bloggy with a 100% chance for participation. IEEE Spectrum, 44(7):10, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLMa

[Ano07-57] Anonymous. Spectral lines: Over the moon. IEEE Spectrum, 44(1):8, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLS

[Ano07-58] Anonymous. Spectral lines: Solar gets googled. IEEE Spectrum, 44(10):8, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:SLT

[Ano07-59] Anonymous. Spectral lines: Trail camera controller scores big in our DIY competition. IEEE Spectrum, 44(11):8, November 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TC

[Ano07-60] Anonymous. Table of contents.

IEEE Spectrum, 44(12):1–2, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TCVa

[Ano07-61] Anonymous. Table of contents — vol 44, no 1. IEEE Spectrum, 44(1):1–3, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TCVf

[Ano07-62] Anonymous. Table of contents — vol 44, no 10. IEEE Spectrum, 44(10):1–3, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TCVg

[Ano07-63] Anonymous. Table of contents — vol 44 no 11. IEEE Spectrum, 44 (11):1–3, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TCVb

[Ano07-64] Anonymous. Table of contents — vol 44, no 2. IEEE Spectrum, 44(2):1–3, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TCVc

[Ano07-65] Anonymous. Table of contents — vol 44 no 5. IEEE Spectrum, 44(5):1–3, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TCVd

[Ano07-66] Anonymous. Table of contents — vol 44 no 6. IEEE Spectrum, 44(6):1–3, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:TCVe

[Ano07-67] Anonymous. Table of contents — vol 44, no 7. IEEE Spectrum, 44(7):3–5, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2007:WLB

[Ano07-68] Anonymous. Winners losers 2007: Beauty and the beast. IEEE Spectrum, 44(1):20–21, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:AAB

[Ano08a] Anonymous. Arachnoid affair [the big picture]. IEEE Spectrum, 45(11):20–21, November 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BSa

[Ano08b] Anonymous. Back story. IEEE Spectrum, 45(1):6, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BSb

[Ano08c] Anonymous. Back story. IEEE Spectrum, 45(2):6, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BSc

[Ano08d] Anonymous. Back story. IEEE Spectrum, 45(3):4, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BSHb

[Ano08e] Anonymous. Back story — he’s got a little list. IEEE Spectrum, 45(7):4, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BSG

[Ano08f] Anonymous. Back story: A gearhead’s motorhead. IEEE Spectrum, 45(4):4, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BSHa

[Ano08g] Anonymous. Back story: Her first darpatech. IEEE Spectrum, 45(5):4, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BSS

[Ano08h] Anonymous. Back story: Something new under the sun. IEEE Spectrum, 45(8):4, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BTM

[Ano08i] Anonymous. Backstory: of two minds. IEEE Spectrum, 45(6):4, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BBB

[Ano08j] Anonymous. Beer, not bombs — [back story]. IEEE Spectrum, 45(10):4, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BPL

[Ano08k] Anonymous. The big picture — laptop orchestra. IEEE Spectrum, 45(1):22–23, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BPSb

[Ano08l] Anonymous. The big picture — snow globe. IEEE Spectrum, 45(7):18–19, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BPE

[Ano08m] Anonymous. The big picture [eye candy]. IEEE Spectrum, 45(3):18, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BPG

[Ano08n] Anonymous. The big picture:

Gone to pieces. IEEE Spectrum, 45(4):18–19, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BPSc

[Ano08o] Anonymous. The big picture:

spray of light. IEEE Spectrum, 45(8):18–19, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:BPSa

[Ano08p] Anonymous. The big picture: “Super Bowl” canceled?

IEEE Spectrum, 45(5):16–17, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:ISTa

[Ano08q] Anonymous. IEEE Spectrum — table of contents. IEEE Spectrum, 45(1):1–3, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:ISTb

[Ano08r] Anonymous. IEEE Spectrum — table of contents. IEEE Spectrum, 45(2):1–2, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:CGI

[Ano08s] Anonymous. A computer game’s intelligent design [back story]. IEEE Spectrum, 45 (9):4, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cl

[Ano08t] Anonymous. Contents. IEEE Spectrum, 45(12):1–3, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Ca

[Ano08u] Anonymous. Contributors. IEEE Spectrum, 45(1):4, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cb

[Ano08v] Anonymous. Contributors. IEEE Spectrum, 45(2):4, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cc

[Ano08w] Anonymous. Contributors. IEEE Spectrum, 45(3):6, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cd

[Ano08x] Anonymous. Contributors. IEEE Spectrum, 45(4):6, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Ce

[Ano08y] Anonymous. Contributors. IEEE Spectrum, 45(5):6, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cf

[Ano08z] Anonymous. Contributors. IEEE Spectrum, 45(6):6, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cg

[Ano08-27] Anonymous. Contributors. IEEE Spectrum, 45(7):6, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Ch

[Ano08-28] Anonymous. Contributors. IEEE Spectrum, 45(8):6, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Ci

[Ano08-29] Anonymous. Contributors. IEEE Spectrum, 45(9):6, September 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cj

[Ano08-30] Anonymous. Contributors. IEEE Spectrum, 45(10):6, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Ck

[Ano08-31] Anonymous. Contributors. IEEE Spectrum, 45(11):6, November 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Cm

[Ano08-32] Anonymous. Contributors. IEEE Spectrum, 45(12):6, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:DL

[Ano08-33] Anonymous. Dance into the light. IEEE Spectrum, 45 (12):4, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:DDT

[Ano08-34] Anonymous. Driving the DOD toward change [spectral lines]. IEEE Spectrum, 45 (11):9, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:DDD

[Ano08-35] Anonymous. Dust-devil dynamo. IEEE Spectrum, 45(3):12, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:ELB

[Ano08-36] Anonymous. Energy loop [the big picture]. IEEE Spectrum, 45 (9):18–19, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:EC

[Ano08-37] Anonymous. Exascale computing by 2015? IEEE Spectrum, 45(12):12, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FFB

[Ano08-38] Anonymous. Face facts [the big picture]. IEEE Spectrum, 45 (2):20–21, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fa

[Ano08-39] Anonymous. Forum. IEEE Spectrum, 45(1):10, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fb

[Ano08-40] Anonymous. Forum. IEEE Spectrum, 45(2):10, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fc

[Ano08-41] Anonymous. Forum. IEEE Spectrum, 45(4):10, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fd

[Ano08-42] Anonymous. Forum. IEEE Spectrum, 45(5):10, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fe

[Ano08-43] Anonymous. Forum. IEEE Spectrum, 45(6):10, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Ff

[Ano08-44] Anonymous. Forum. IEEE Spectrum, 45(8):10, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fg

[Ano08-45] Anonymous. Forum. IEEE Spectrum, 45(9):10, September 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fh

[Ano08-46] Anonymous. Forum. IEEE Spectrum, 45(11):10, November 2008.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:Fi

[Ano08-47] Anonymous. Forum. IEEE Spectrum, 45(12):10, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCa

[Ano08-48] Anonymous. [front cover]. IEEE Spectrum, 45(5):c1, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIc

[Ano08-49] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(3):c1, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCId

[Ano08-50] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(4):c1, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIe

[Ano08-51] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(6):C1, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIf

[Ano08-52] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(7):c1, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIg

[Ano08-53] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(8):c1–c2, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIi

[Ano08-54] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(10):c1, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIj

[Ano08-55] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(11):c1, November 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIk

[Ano08-56] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 45(12):c1, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIb

[Ano08-57] Anonymous. [front cover — IEEE Spectrum — Feb 2008]. IEEE Spectrum, 45 (2):C1, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIa

[Ano08-58] Anonymous. [front cover — IEEE Spectrum — Jan 2008]. IEEE Spectrum, 45(1):c1, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FCIh

[Ano08-59] Anonymous. [front cover — IEEE Spectrum: Can technology defeat IEDs?]. IEEE Spectrum, 45(9):fc, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:FSN

[Ano08-60] Anonymous. Fuel from sunshine [news brief]. IEEE Spectrum, 45(2):17, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:GPC

[Ano08-61] Anonymous. Gas pains [commentary]. IEEE Spectrum, 45 (11):19, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:IDJ

[Ano08-62] Anonymous. Introduction [dream jobs 2008]. IEEE Spectrum, 45(2):29, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:KBT

[Ano08-63] Anonymous. Karaoke-bot [tools toys]. IEEE Spectrum, 45 (2):25, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:MEE

[Ano08-64] Anonymous. Microchip enables electronic gene injection. IEEE Spectrum, 45(2):17, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:MMO

[Ano08-65] Anonymous. Microscopic microwave oven [news brief]. IEEE Spectrum, 45(1):20, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NSA

[Ano08-66] Anonymous. A new spin on architecture. IEEE Spectrum, 45 (12):20–21, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NBNb

[Ano08-67] Anonymous. News brief — nanoflowers. IEEE Spectrum, 45(7): 15, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NBD

[Ano08-68] Anonymous. News brief: Doomsday vault. IEEE Spectrum, 45(4):15, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NBE

[Ano08-69] Anonymous. News brief: [eyepods]. IEEE Spectrum, 45(3):15, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NBS

[Ano08-70] Anonymous. News brief: I see you. IEEE Spectrum, 45(5):13, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NBNa

[Ano08-71] Anonymous. News brief: Nanopine trees. IEEE Spectrum, 45(6):20, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NRS

[Ano08-72] Anonymous. Ney Robinson Salvi dos Reis — into the wild [dream jobs 2008]. IEEE Spectrum, 45 (2):39–40, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:NDS

[Ano08-73] Anonymous. Ntt develops a smell-o-phone. IEEE Spectrum, 45(1):14, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:ODD

[Ano08-74] Anonymous. The other digital divide [news brief]. IEEE Spectrum, 45(1):16, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:PTU

[Ano08-75] Anonymous. Paper transistor [update]. IEEE Spectrum, 45 (9):15, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:RW

[Ano08-76] Anonymous. Reinventing the wheel. IEEE Spectrum, 45 (2):15, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:SI

[Ano08-77] Anonymous. Singularity index. IEEE Spectrum, 45(6):92, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:SLO

[Ano08-78] Anonymous. Spectral lines — one format war is over. Is another beginning? IEEE Spectrum, 45(2):9, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:SLT

[Ano08-79] Anonymous. Spectral lines — technology innovation is Spectrum’s message, whatever the medium. IEEE Spectrum, 45 (1):9, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCa

[Ano08-80] Anonymous. Table of contents. IEEE Spectrum, 45(3):1–3, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCb

[Ano08-81] Anonymous. Table of contents. IEEE Spectrum, 45(4):1–3, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCc

[Ano08-82] Anonymous. Table of contents. IEEE Spectrum, 45(5):1– 3, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCd

[Ano08-83] Anonymous. Table of contents. IEEE Spectrum, 45(6):1– 3, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCe

[Ano08-84] Anonymous. Table of contents. IEEE Spectrum, 45(7):1– 3, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCf

[Ano08-85] Anonymous. Table of contents.

IEEE Spectrum, 45(8):1–3, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCg

[Ano08-86] Anonymous. Table of contents. IEEE Spectrum, 45(9): 1–3, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCh

[Ano08-87] Anonymous. Table of contents. IEEE Spectrum, 45(10):1–3, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TCi

[Ano08-88] Anonymous. Table of contents. IEEE Spectrum, 45(11): 1–3, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TBB

[Ano08-89] Anonymous. The ties that bind [back story]. IEEE Spectrum, 45(11):4, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:TIC

[Ano08-90] Anonymous. Tiny implants combat chronic pain. IEEE Spectrum, 45(12):18, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UNB

[Ano08-91] Anonymous. Untangling a new breast cancer screening technology. IEEE Spectrum, 45(2): 13–14, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UBC

[Ano08-92] Anonymous. Update: A better camera pill. IEEE Spectrum, 45(8):12, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UFE

[Ano08-93] Anonymous. Update: Face of evil. IEEE Spectrum, 45(8):15, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UOS

[Ano08-94] Anonymous. Update: Opensource baby. IEEE Spectrum, 45(6):14, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UTS

[Ano08-95] Anonymous. Update: Taiwanese software spots stockmarket stinkers. IEEE Spectrum, 45(5):12, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UWC

[Ano08-96] Anonymous. Update: What are the chances? IEEE Spectrum, 45(4):12, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UWD

[Ano08-97] Anonymous. Update: Wiring diagram. IEEE Spectrum, 45(8):14, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:WR

[Ano08-98] Anonymous. Waiting for the rapture. IEEE Spectrum, 45 (6):32–35, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:WWS

[Ano08-99] Anonymous. Who’s who in the singularity. IEEE Spectrum, 45(6):42–43, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:YTU

[Ano08-100] Anonymous. You tell us. IEEE Spectrum, 45(1):64, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:AOG

[Ano09a] Anonymous. The amazing orbiting garriotts. IEEE Spectrum, 46(6):70–72, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:BS

[Ano09b] Anonymous. Back story. IEEE Spectrum, 46(6):4, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:BM

[Ano09c] Anonymous. Bendable memristors. IEEE Spectrum, 46(7):15, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:BP

[Ano09d] Anonymous. The big picture. IEEE Spectrum, 46(12): 18–19, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISa

[Ano09e] Anonymous. IEEE Spectrum.

IEEE Spectrum, 46(6):c1, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISb

[Ano09f] Anonymous. IEEE Spectrum.

IEEE Spectrum, 46(10):c1, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISFa

[Ano09g] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 46(4):C1, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISFb

[Ano09h] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 46(5):c1, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISFc

[Ano09i] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 46(8):c1, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISFd

[Ano09j] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 46 (11):c1–c2, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISFe

[Ano09k] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 46 (12):c1, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ISD

[Ano09l] Anonymous. IEEE Spectrum: Disappearing deterrent. IEEE Spectrum, 46(3):fc, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:BG

[Ano09m] Anonymous. Breaking gsm. IEEE Spectrum, 46(12):13, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:BDB

[Ano09n] Anonymous. Breathe deeply [back story]. IEEE Spectrum, 46(1):4, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:BMD

[Ano09o] Anonymous. Britain mulls over digital radio transition. IEEE Spectrum, 46(9):16, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:CBB

[Ano09p] Anonymous. Carbon in bloom — [the big picture]. IEEE Spectrum, 46(1):20–21, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:CKF

[Ano09q] Anonymous. Cellular kungfu grip. IEEE Spectrum, 46 (2):14, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:CAF

[Ano09r] Anonymous. Coming of age in fiji. IEEE Spectrum, 46 (2):4, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cc

[Ano09s] Anonymous. Contents. IEEE Spectrum, 46(3):1–3, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cg

[Ano09t] Anonymous. Contents. IEEE Spectrum, 46(6):1–3, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cl

[Ano09u] Anonymous. Contents. IEEE Spectrum, 46(10):1–3, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Ca

[Ano09v] Anonymous. Contributors. IEEE Spectrum, 46(1):6, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cb

[Ano09w] Anonymous. Contributors. IEEE Spectrum, 46(2):6, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cd

[Ano09x] Anonymous. Contributors. IEEE Spectrum, 46(3):6, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Ce

[Ano09y] Anonymous. Contributors. IEEE Spectrum, 46(4):6, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cf

[Ano09z] Anonymous. Contributors. IEEE Spectrum, 46(5):6, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Ch

[Ano09-27] Anonymous. Contributors. IEEE Spectrum, 46(6):6, June 2009.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Ci

[Ano09-28] Anonymous. Contributors. IEEE Spectrum, 46(7):6, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cj

[Ano09-29] Anonymous. Contributors. IEEE Spectrum, 46(8):6, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Ck

[Ano09-30] Anonymous. Contributors. IEEE Spectrum, 46(9):6, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cm

[Ano09-31] Anonymous. Contributors. IEEE Spectrum, 46(10):6, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Cn

[Ano09-32] Anonymous. Contributors. IEEE Spectrum, 46(11):6, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Co

[Ano09-33] Anonymous. Contributors. IEEE Spectrum, 46(12):6, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:DE

[Ano09-34] Anonymous. Dinosaur egg. IEEE Spectrum, 46(8):16, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:DDa

[Ano09-35] Anonymous. Doppelganger dance. IEEE Spectrum, 46(7):20– 21, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:DJS

[Ano09-36] Anonymous. Dream jobs 09 special report. IEEE Spectrum, 46(2):23, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:DDb

[Ano09-37] Anonymous. Dropping the drumsticks. IEEE Spectrum, 46 (9):4, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:B

[Ano09-38] Anonymous. E-bracadabra. IEEE Spectrum, 46(9):25, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:FLb

[Ano09-39] Anonymous. A fab life. IEEE Spectrum, 46(5):20–21, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:FLG

[Ano09-40] Anonymous. Finland’s long good-bye. IEEE Spectrum, 46 (12):4, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fa

[Ano09-41] Anonymous. Forum. IEEE Spectrum, 46(1):10, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fb

[Ano09-42] Anonymous. Forum. IEEE Spectrum, 46(2):8, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fc

[Ano09-43] Anonymous. Forum. IEEE Spectrum, 46(3):10, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fd

[Ano09-44] Anonymous. Forum. IEEE Spectrum, 46(4):10, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fe

[Ano09-45] Anonymous. Forum. IEEE Spectrum, 46(5):10, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Ff

[Ano09-46] Anonymous. Forum. IEEE Spectrum, 46(6):8, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fg

[Ano09-47] Anonymous. Forum. IEEE Spectrum, 46(7):12, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fh

[Ano09-48] Anonymous. Forum. IEEE Spectrum, 46(8):10, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fi

[Ano09-49] Anonymous. Forum. IEEE Spectrum, 46(9):12, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fj

[Ano09-50] Anonymous. Forum. IEEE Spectrum, 46(10):12, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fk

[Ano09-51] Anonymous. Forum. IEEE Spectrum, 46(11):12, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Fm

[Ano09-52] Anonymous. Forum. IEEE Spectrum, 46(12):10, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:FCIa

[Ano09-53] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 46(1):c1, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:FCIb

[Ano09-54] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 46(2):C1, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:FCIc

[Ano09-55] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 46(7):C1, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:FCId

[Ano09-56] Anonymous. [front cover — IEEE Spectrum]. IEEE Spectrum, 46(9):C1, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:GS

[Ano09-57] Anonymous. Geek summer.

IEEE Spectrum, 46(3):4, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:HPA

[Ano09-58] Anonymous. A high-pressure assignment. IEEE Spectrum, 46 (11):4, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:FLa

[Ano09-59] Anonymous. In Frankenstein’s laboratory. IEEE Spectrum, 46 (3):18–19, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:IJL

[Ano09-60] Anonymous. India joins league of lunar nations. IEEE Spectrum, 46(6):62–63, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ID

[Ano09-61] Anonymous. Ink display. IEEE Spectrum, 46(6):13, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:LGS

[Ano09-62] Anonymous. Lasers, great and small [back story]. IEEE Spectrum, 46(7):4, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:MIT

[Ano09-63] Anonymous. The mobile infections threat. IEEE Spectrum, 46(5):12, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:ND

[Ano09-64] Anonymous. Nano-dreadlocks.

IEEE Spectrum, 46(2):11, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:NPT

[Ano09-65] Anonymous. Nobel prizes for three IEEE fellows. IEEE Spectrum, 46(11):15, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:CPL

[Ano09-66] Anonymous. Of cows and power lines. IEEE Spectrum, 46(7):14, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:M

[Ano09-67] Anonymous. On the march.

IEEE Spectrum, 46(10):18, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:OO

[Ano09-68] Anonymous. Overlord of overload. IEEE Spectrum, 46 (10):4, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:PCb

[Ano09-69] Anonymous. Phase change.

IEEE Spectrum, 46(11):14, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:PP

[Ano09-70] Anonymous. Place and play. IEEE Spectrum, 46(9):22–23, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:PCa

[Ano09-71] Anonymous. Primary colors. IEEE Spectrum, 46(4):18–19, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:RDR

[Ano09-72] Anonymous. Relief for digital relics. IEEE Spectrum, 46 (2):10, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:RBS

[Ano09-73] Anonymous. Robotic baby seal could diminish dementia. IEEE Spectrum, 46(5):14, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:RR

[Ano09-74] Anonymous. Running with robots. IEEE Spectrum, 46(4):4, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:SC

[Ano09-75] Anonymous. Say cheese. IEEE Spectrum, 46(9):19, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:SDG

[Ano09-76] Anonymous. The spectrum dumpling gang. IEEE Spectrum, 46(8):4, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCa

[Ano09-77] Anonymous. Table of contents.

IEEE Spectrum, 46(1):1–3, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCb

[Ano09-78] Anonymous. Table of contents. IEEE Spectrum, 46 (2):1–3, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCc

[Ano09-79] Anonymous. Table of contents. IEEE Spectrum, 46(4):1–3, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCd

[Ano09-80] Anonymous. Table of contents. IEEE Spectrum, 46(5):1– 3, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCe

[Ano09-81] Anonymous. Table of contents. IEEE Spectrum, 46(7):1– 3, July 2009. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCf

[Ano09-82] Anonymous. Table of contents. IEEE Spectrum, 46(8):1–3, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCg

[Ano09-83] Anonymous. Table of contents. IEEE Spectrum, 46(9): 1–3, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCh

[Ano09-84] Anonymous. Table of contents. IEEE Spectrum, 46(11): 1–3, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TCi

[Ano09-85] Anonymous. Table of contents. IEEE Spectrum, 46(12):1–3, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:TP

[Ano09-86] Anonymous. Talking plants. IEEE Spectrum, 46(6):16, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:WWW

[Ano09-87] Anonymous. Where in the world wide web is al phillips? IEEE Spectrum, 46(5):4, May 2009.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:Z

[Ano09-88] Anonymous. Zap! IEEE Spectrum, 46(7):17, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Appell:2001:SSD

[App01a] D. Appell. Seeing stars with digital eyes. IEEE Spectrum, 38(7):62–67, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2001:AEI

[App01b] A. Applewhite. An academic with entrepreneurial instincts [careers]. IEEE Spectrum, 38 (10):59–62, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2001:AFC

[App01c] A. Applewhite. Anticipating the future [careers]. IEEE Spectrum, 38(4):81–82, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2001:EWL

[App01d] A. Applewhite. Exploring what lies beneath. IEEE Spectrum, 38 (9):65–67, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2002:ABE

[App02a] A. Applewhite. Africa becomes electric. IEEE Spectrum, 39 (8):54–56, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2002:HCH

[App02b] A. Applewhite. The hero checks her oscilloscope. IEEE Spectrum, 39(2):66–69, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2002:MG

[App02c] A. Applewhite. Mining the genome. IEEE Spectrum, 39 (4):69–71, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2002:WFW

[App02d] A. Applewhite. Why so few women? IEEE Spectrum, 39 (5):65–66, May 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2003:ITV

[App03a] A. Applewhite. It takes a village. IEEE Spectrum, 40(9): 40–45, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2003:SYA

[App03b] A. Applewhite. Since you asked [technology forecast survey]. IEEE Spectrum, 40(1): 11–16, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2004:LST

[App04a] A. Applewhite. Life scientist takes helm at MIT. IEEE Spectrum, 41(11):18, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2004:VTT

[App04b] A. Applewhite. The view from the top [technology forecasting]. IEEE Spectrum, 41(11): 36–38, November 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Applewhite:2004:WWM

[App04c] A. Applewhite. A wider welcome mat. IEEE Spectrum, 41(6):53–57, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Arnold:2009:SP

[Arn09] B. Arnold. Shrinking possibilities. IEEE Spectrum, 46 (4):26–56, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Apte:2009:RSL

[AS09] P. Apte and G. Scalise. The recession’s silver lining. IEEE Spectrum, 46(10):46–51, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ashley:2000:PLU

[Ash00] J. R. Ashley. Are power lines unsafe? IEEE Spectrum, 37(7):21– 23, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aston:2001:BWC

[Ast01] C. Aston. Biological warfare canaries [biological attack detection]. IEEE Spectrum, 38 (10):35–40, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aston:2002:FTT

[Ast02] C. Aston. Futility of a toothless treaty. IEEE Spectrum, 39 (1):15–16, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Avioli:2002:EUR

[Avi02a] D. Avioli. EE unemployment on the rise in United States. IEEE Spectrum, 39(9):21–22, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Avioli:2002:HUA

[Avi02b]  D. Avioli.  Hollywood ups

ante in copy protection fight. IEEE Spectrum, 39(7):24–25, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Avouris:2004:SCN

[Avo04] P. Avouris. Supertubes [carbon nanotubes]. IEEE Spectrum, 41 (8):40–45, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Balaram:2000:MPC

[Bal00]  A. Balaram.  Modeling for

printed-circuit board simulation. IEEE Spectrum, 37(10):

70–74, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baliga:2004:CGV

[Bal04] J. Baliga. Chips go vertical [3D IC interconnection]. IEEE Spectrum, 41(3):43–47, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Balandin:2009:C

[Bal09] A. A. Balandin. Chill out. IEEE Spectrum, 46(10):34–39, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Banisar:2002:VBR

[Ban02] D. Banisar. Virtual borders, real laws [Internet activity and treaties]. IEEE Spectrum, 39 (10):51–52, October 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2001:SSI

[Bar01] S. Barlas. Scraping up spectrum for Internet-enabled cell telephony. IEEE Spectrum, 38 (6):34–37, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2003:OPR

[Bar03] S. Barlas. One-party rule returns to United States. IEEE Spectrum, 40(2):22–24, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2004:BB

[Bar04a] S. Barlas. Budget breakdown. IEEE Spectrum, 41(4):11, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2004:FSW

[Bar04b] S. Barlas. Funding soars for weaponry in U.S. budget. IEEE Spectrum, 41(4):10–12, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2006:CAC

[Bar06a] S. Barlas. Crazy about competitiveness. IEEE Spectrum, 43(4):18–20, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2006:SUG

[Bar06b] S. Barlas. Stricter U.S. gas standards stalled. IEEE Spectrum, 43 (9):12–15, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2006:TWS

[Bar06c] S. Barlas. Troubled weather satellite program. IEEE Spectrum, 43(6):14–16, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2006:UCT

[Bar06d] S. Barlas. U.S. climate technology plan. IEEE Spectrum, 43 (11):12–14, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Basanskaya:2005:EGF

[Bas05] A. Basanskaya. Electricity over glass [fiber optic to transfer electric power]. IEEE Spectrum, 42(10):18, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:2009:MDP

[BBKV09] R. M. Bell, J. Bennett, Y. Koren, and C. Volinsky. The million dollar programming prize. IEEE Spectrum, 46(5):28–33, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2001:NAE

[BBS^{+}01] J. Blau, S. Barlas, V. Singh, L. Geppert, H. Goldstein, Samuel K. Moore, M. J. Riezenman, and S. Miller. News analysis: (1) Europe’s cellulars to share infrastructure? (2) Why the fuss about 3G? (3) Senate changes will shake telecom, defense R&D, Energy. (4) In Energy, will Bingaman out-Cheney Cheney? (5) India’s power grid finds helping hand. (6) The Itanium platform lands its first computing systems. (7) Ion beam keeps liquid crystals in line. (8) How to raise UV nanolasers.

(9) Swedish start-up puts new spin on pen and paper. (10) Big welcome for MIT’s Web-based courseware. IEEE Spectrum, 38(7):20–37, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bass:2002:FMB

[BC02] M. J. Bass and C. M. Christensen. The future of the microprocessor business. IEEE Spectrum, 39(4):34–39, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bar-Cohen:2004:TH

[BC04] Y. Bar-Cohen. Turning heads. IEEE Spectrum, 41(6):28–33, June 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berlin:2005:RNT

[BC05] L. Berlin and C. Casey. Robert Noyce and the tunnel diode. IEEE Spectrum, 42(5):49–53, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bohr:2007:HKS

[BCGM07] M. T. Bohr, R. S. Chau, T. Ghani, and K. Mistry. The high-k solution. IEEE Spectrum, 44(10):29–35, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blaauw:2009:CHT

[BD09] D. Blaauw and S. Das. CPU, heal thyself. IEEE Spectrum, 46 (8):40–56, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:2000:PSN

[Bel00] T. E. Bell. Proven skills: the new yardstick for schools. IEEE Spectrum, 37(9):63–67, September 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:2006:WGG

[Bel06] T. E. Bell. Waiting for gravity [gravitational wave detection]. IEEE Spectrum, 43(7):40– 46, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berger:2002:CYT

[Ber02] I. Berger. Can you trust your car? [automotive electronics reliability testing]. IEEE Spectrum, 39(4):40–45, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernstein:2003:WWT

[Ber03] P. A. Bernstein. What’s wrong with telecom. IEEE Spectrum, 40(1):26–29, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Betts:2006:SS

[Bet06] B. Betts. Smart sensors. IEEE Spectrum, 43(4):50–53, April 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bleier:2005:EEW

[BF05] T. Bleier and F. Freund. Earthquake [earthquake warning systems]. IEEE Spectrum, 42(12):

22–27, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bing:2002:CAS

[BJ02] B. Bing and N. Jayant. A cellphone for all standards. IEEE Spectrum, 39(5):34–39, May 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:TTRf

[BJC01] E. A. Bretz, W. D. Jones, and Stephen Cass. Tools toys [resources]. IEEE Spectrum, 38 (10):69–71, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blas:2009:DM

[BK09] A. D. Blas and T. Kaldewey. Data monster. IEEE Spectrum, 46(9):46–51, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2002:CDC

[Bla02a] J. Blau. Creative destruction [car crash testing]. IEEE Spectrum, 39(11):55–56, November 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2002:LEN

[Bla02b] J. Blau. Looking to eat Nokia’s lunch [3G cell phones]. IEEE Spectrum, 39(4):20–21, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2002:TSS

[Bla02c] J. Blau. Tech slump sinks Germany’s Neuer Markt. IEEE Spectrum, 39(11):24–25, November 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2003:ERG

[Bla03a] J. Blau. Europe’s regulatory gridlock [media copying legislation]. IEEE Spectrum, 40 (5):34–35, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2003:ESM

[Bla03b] J. Blau. Europe’s semiconductor makers are back in the game. IEEE Spectrum, 40(2): 18–19, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2004:SFO

[Bla04] J. Blau. Supermarket’s futuristic outlet. IEEE Spectrum, 41(4):21–22, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2005:GUS

[Bla05a] J. Blau. German universities seek to reclaim lost glory. IEEE Spectrum, 42(1):10–12, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2005:TTT

[Bla05b] J. Blau. Telephone TV [TV broadcasting]. IEEE Spectrum, 42(6):16–17, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2006:TC

[Bla06] J. Blau. Talk is cheap. IEEE Spectrum, 43(10):14–15, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2007:CCE

[Bla07] J. Blau. Carbon caps for European cars. IEEE Spectrum, 44(6):16–18, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blau:2008:CTU

[Bla08] J. Blau. Car talk — [update]. IEEE Spectrum, 45(10):16, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2005:TTN

[BM05] J. Boyd and Samuel K. Moore. Transparent transistors: a new class of semiconductors makes for fast, flexible, and invisible electronics. IEEE Spectrum, 42 (3):18–19, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bolter:2007:ILI

[BM07] J. D. Bolter and B. Macintyre. Is it live or is it ar? IEEE Spectrum, 44(8):30–35, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bohley:2006:LCK

[Boh06] C. Bohley. Liquid-crystal kaleidoscope. IEEE Spectrum, 43 (1):18–19, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boillot:2006:WBC

[Boi06] N. Boillot. Websights: Beating the crunch. IEEE Spectrum, 43(6):55–57, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bordogna:2001:CE

[Bor01] J. Bordogna. The 21st century engineer. IEEE Spectrum, 38 (1):17–19, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Borenstein:2009:MM

[Bor09] J. T. Borenstein. Medicine by micromachine. IEEE Spectrum, 46 (11):36–41, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Briscoe:2006:MLW

[BOT06] Bob Briscoe, Andrew Odlyzko, and Benjamin Tilly. Metcalfe’s Law is wrong — communications networks increase in value as they add members — but by how much? IEEE Spectrum, 43(7):34–39, July 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bowden:2003:IWK

[Bow03] M. Bowden. Information was key to efficient US–Anglo victory in Iraq. IEEE Spectrum, 40(6):18–19, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2003:CCT

[Boy03a] J. Boyd. Counting on a cool Tokyo summer. IEEE Spectrum, 40(6):17, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2003:NNC

[Boy03b] J. Boyd. Nifty new cellular phone systems race to capture japan’s consumers. IEEE Spectrum, 40 (10):14–18, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2004:CTG

[Boy04a] J. Boyd. Canon and Toshiba go their own way in flat panels. IEEE Spectrum, 41(11): 24, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2004:FUN

[Boy04b] J. Boyd. Fujitsu unveils new type of plasma display. IEEE Spectrum, 41(3):18–21, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2004:JCD

[Boy04c] J. Boyd. Japan cranks up download speeds [DSL technology]. IEEE Spectrum, 41(12): 10–12, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2004:SPD

[Boy04d] J. Boyd. Samsung’s plasma displays barred from Japanese market. IEEE Spectrum, 41(6):20, June 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2005:HCW

[Boy05]  J. Boyd.  Here comes the

wallet phone [wireless credit card]. IEEE Spectrum, 42(11): 12–14, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2006:CPL

[Boy06a] J. Boyd. Changing photographic landscape. IEEE Spectrum, 43 (12):18, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2006:DD

[Boy06b] J. Boyd. Duel of the discs. IEEE Spectrum, 43(9):14–16, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2006:SPC

[Boy06c] J. Boyd. Sneak peek at cellphone future. IEEE Spectrum, 43(6):15, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2006:TPQ

[Boy06d] J. Boyd. Tokyo power quakeproofs its grid control system. IEEE Spectrum, 43(1):15, January 2006. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2007:LTN

[Boy07] J. Boyd. Let there be (a new kind of) light [news]. IEEE Spectrum, 44(7):12–14, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2008:HFC

[Boy08a] J. Boyd. Home fuels cells to sell in Japan — [update]. IEEE Spectrum, 45(10):20, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyd:2008:UMM

[Boy08b] J. Boyd. Update — mixing memory to speed solid-state drives. IEEE Spectrum, 45(7): 15, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baker:2002:PSR

[BR02] G. Baker and F. Ramirez. Power sector reform-blackouts before policy. IEEE Spectrum, 39 (5):14–15, May 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brackenridge:2006:IUU

[Bra06]  B. Brackenridge.  Invention:

Ultrawideband upset. IEEE Spectrum, 43(9):65–66, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:BBH

[Bre00a] E. A. Bretz. Boston builds a high-IQ highway. IEEE Spectrum, 37(8):47–52, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:CAI

[Bre00b] E. A. Bretz. Cable access issue has fcc on the spot [washington watch]. IEEE Spectrum, 37(9): 36I2–36I4, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:PIM

[Bre00c] E. A. Bretz. Pjm interconnection: model of a smooth operator [electric utility experience]. IEEE Spectrum, 37 (6):50–55, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:TWa

[Bre00d] E. A. Bretz. Technology watch.

IEEE Spectrum, 37(7):24, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:TWb

[Bre00e] E. A. Bretz. Technology watch.

IEEE Spectrum, 37(10):32, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:TMT

[Bre00f] E. A. Bretz. Test and measurement [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1):75–79, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:TTA

[Bre00g] E. A. Bretz. Transportation [technology 2000 analysis and forecast]. IEEE Spectrum, 37 (1):91–96, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2000:XMS

[Bre00h] E. A. Bretz. X marks the spot, maybe [gps navigation]. IEEE Spectrum, 37(4):26–36, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:FFA

[Bre01a] E. A. Bretz. Free flight: from airline hell to a new age of travel, James Fallows, Public Affairs, New York, $25 [book review]. IEEE Spectrum, 38 (11):66, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:WCT

[Bre01b] E. A. Bretz. By-wire cars turn the corner. IEEE Spectrum, 38(4):68–73, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:CJW

[Bre01c] E. A. Bretz. The car: just a web browser with tires. IEEE Spectrum, 38(1):92–94, January 2001.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:HSW

[Bre01d] E. A. Bretz. How stuff works [book review]. IEEE Spectrum, 38(11):65, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:NTS

[Bre01e] E. A. Bretz. Network test starts in the lab. IEEE Spectrum, 38 (1):77–78, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:TT

[Bre01f] E. A. Bretz. Tools toys. IEEE Spectrum, 38(11):70–72, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:TTRa

[Bre01g] E. A. Bretz. Tools toys [resources]. IEEE Spectrum, 38(1): 136–138, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:TTRb

[Bre01h] E. A. Bretz. Tools toys [resources]. IEEE Spectrum, 38 (3):82–84, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:TTRc

[Bre01i] E. A. Bretz. Tools toys [resources]. IEEE Spectrum, 38 (6):84–86, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:TTRd

[Bre01j] E. A. Bretz. Tools toys [resources]. IEEE Spectrum, 38 (7):84–85, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:TTRe

[Bre01k] E. A. Bretz. Tools toys [resources]. IEEE Spectrum, 38(9): 73–74, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:UPA

[Bre01l] E. A. Bretz. UK privatizes air traffic control. IEEE Spectrum, 38(6):37–39, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:WSB

[Bre01m] E. A. Bretz. Weapons in space [book review]. IEEE Spectrum, 38(11):65, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2001:WIF

[Bre01n] E. A. Bretz. Winging it [free flight, air traffic control]. IEEE Spectrum, 38(1):97–99, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bresnahan:2002:WMS

[Bre02a] T. F. Bresnahan. Why the Microsoft settlement won’t work.

IEEE Spectrum, 39(9):13–14, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2002:CSA

[Bre02b] E. A. Bretz. Clear skies ahead [navigation systems]. IEEE Spectrum, 39(1):78–81, January 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2002:DAU

[Bre02c]  E. A. Bretz.  Delayed ar-

rival for us baggage screening? IEEE Spectrum, 39 (5):16–19, May 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2002:NPP

[Bre02d] E. A. Bretz. New phones play games and run Java applications. IEEE Spectrum, 39(2): 62–63, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2002:WWF

[Bre02e] E. A. Bretz. When work is fun and games. IEEE Spectrum, 39 (12):50, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2003:PNE

[Bre03a] E. A. Bretz. Precision navigation in European skies. IEEE Spectrum, 40(9):16, September 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2003:PHP

[Bre03b] E. A. Bretz. Putting humidity in its place. IEEE Spectrum, 40(3):28–30, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2003:TTN

[Bre03c] E. A. Bretz. Tools toys: Never out of touch. IEEE Spectrum, 40(12):49, December 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2004:SDB

[Bre04a]  E. A. Bretz.  Speedy data

backups. IEEE Spectrum, 41 (1):94–96, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2004:SCH

[Bre04b] E. A. Bretz. Super conductors on the high seas [ship propulsion using superconductor motors]. IEEE Spectrum, 41(1): 60–67, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bretz:2004:TTN

[Bre04c] E. A. Bretz. Tools toys: Never miss another shot. IEEE Spectrum, 41(3):62, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Briscoe:2008:FFI

[Bri08] B. Briscoe. A fairer, faster Internet. IEEE Spectrum, 45(12): 42–47, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Broder:2002:HCI

[Bro02a] J. Broder. How close is Iraq to getting the bomb? IEEE Spectrum, 39(12):21–23, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:2002:ASW

[Bro02b] A. S. Brown. American superconductor wins key patent. IEEE Spectrum, 39(12):20–21, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:2002:BD

[Bro02c] A. S. Brown. Bringing in the dough. IEEE Spectrum, 39(6):66, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:2003:FFI

[Bro03a] A. S. Brown. Fast films [ic interconnect insulation]. IEEE Spectrum, 40(2):36–40, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:2003:ECN

[Bro03b] I. K. Brown. Engineering cloud nine [atmospheric instruments]. IEEE Spectrum, 40 (4):60–61, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:2005:FCU

[Bro05] A. S. Brown. Flat, cheap, and under control [electrochemical mechanical planarization]. IEEE Spectrum, 42(1):40–45, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brooks:2008:RBR

[Bro08] R. Brooks. I, Rodney Brooks, am a robot. IEEE Spectrum, 45(6):68–71, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bruce:2002:TL

[Bru02] E. Bruce. Tunable lasers. IEEE Spectrum, 39(2):35–39, February 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barlas:2001:DCA

[BSK^{+}01] S. Barlas, W. Sweet, J. Kumagai, J. F. Mason, and Y. Chiu. D.C. Appeals Court feels its oats [news analysis]. IEEE Spectrum, 38(4):30–41, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brock:2000:SIG

[BTR00] D. K. Brock, E. K. Track, and J. M. Rowell. Superconductor ics: the 100-ghz second generation. IEEE Spectrum, 37(12): 40–46, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caloyannides:2000:EWE

[Cal00a] M. A. Caloyannides. Encryption wars: early battles. IEEE Spectrum, 37(4):37–43, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caloyannides:2000:EWS

[Cal00b] M. A. Caloyannides. Encryption wars: shifting tactics. IEEE Spectrum, 37(5):46–51, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlson:2001:OSB

[Car01] R. Carlson. Open-source biology and its impact on industry. IEEE Spectrum, 38(5):15–17, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlson:2002:HH

[Car02a] R. Carlson. Hacking humanity. IEEE Spectrum, 39(12): 47–48, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlson:2002:LL

[Car02b] R. Carlson. Life lessons. IEEE Spectrum, 39(12):46, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2000:TW

[Cas00] Stephen Cass. Technology watch.

IEEE Spectrum, 37(12):28, December 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:MO

[Cas01a] Stephen Cass. 2001: a Mars Odyssey. IEEE Spectrum, 38 (4):58–65, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:AMC

[Cas01b] Stephen Cass. Anatomy of malice [computer viruses]. IEEE Spectrum, 38(11):56–60, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:AXG

[Cas01c] Stephen Cass. Apple OS X the good, the bad, and the ugly. IEEE Spectrum, 38(9): 68–70, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:GSM

[Cas01d] Stephen Cass. Go to: the story of the math majors, bridge players, engineers, chess wizards, scientists and iconoclasts who were the hero programmers of the software revolution, Steve Lohr, Basic Books, New York, $27.50 [book review]. IEEE Spectrum, 38(11):66, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:MCG

[Cas01e] Stephen Cass. Make cool gadgets for your room and make amazing toy and game gadgets, by Amy Pinchuk, Greey de Pencier Books, Toronto, $12.95 [book review]. IEEE Spectrum, 38 (11):68, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:MLT

[Cas01f] Stephen Cass. Moon lander, Tom Kelly, Smithsonian Institute Press, Washington, DC, $27.95 [book review]. IEEE Spectrum, 38(11):66, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:ERA

[Cas01g] Stephen Cass. Electronic realism [art and technology]. IEEE Spectrum, 38(3):68–73, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:LJL

[Cas01h] Stephen Cass. Large jobs for little devices [microelectromechanical systems]. IEEE Spectrum, 38(1):72–73, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:MNM

[Cas01i] Stephen Cass. Mars: The NASA mission reports [book review]. IEEE Spectrum, 38(4):85, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:MS

[Cas01j] Stephen Cass. MEMS in space. IEEE Spectrum, 38(7):56–61, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:RMR

[Cas01k] Stephen Cass. Robosoccer [mobile robotics experiment]. IEEE Spectrum, 38(5):75–77, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:SWE

[Cas01l] Stephen Cass. A scientific warrior [Edward Teller, Memoirs, Perseus, Cambridge, Mass., $35, 592 pp., ISBN: 0-7382-0532-X] [books]. IEEE Spectrum, 38 (10):65–66, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:SNS

[Cas01m] Stephen Cass. Super nets for supercomputers. IEEE Spectrum, 38(1):116–118, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2001:VNW

[Cas01n] Stephen Cass. Visualizing networks [web sites]. IEEE Spectrum, 38(1):122, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:RRU

[Cas02a] S. A. Cass. Rockets R us [Russia launch industry]. IEEE Spectrum, 39(2):52–58, February 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:AT

[Cas02b] Stephen Cass. All thumbs.

IEEE Spectrum, 39(2):63, February 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:AYS

[Cas02c] Stephen Cass. An archeologist’s year in silicon valley [book review]. IEEE Spectrum, 39 (1):86–88, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:FFR

[Cas02d] Stephen Cass. Free as in freedom:

Richard Stallman’s crusade for free software, by Sam Williams, O’Reilly, Sebastopol, CA, 2002, ISBN 0-596-00287-4, 240 pp., $22.96 [book review]. IEEE Spectrum, 39(6):56–57, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:MGC

[Cas02e] Stephen Cass. Mind games [computer game ai]. IEEE Spectrum, 39(12):40–44, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:NVN

[Cas02f] Stephen Cass. Nissan v. Nissan [trademark dispute]. IEEE Spectrum, 39(10):53–54, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:RBR

[Cas02g] Stephen Cass. Radiance [book review]. IEEE Spectrum, 39 (9):75, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:THH

[Cas02h] Stephen Cass. Try this at home [homeland security]. IEEE Spectrum, 39(9):42, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Casazza:2003:EEE

[Cas03a] J. Casazza. Engineering, ethics electricity. IEEE Spectrum, 40 (7):11–12, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:AEB

[Cas03b] Stephen Cass. Apollo EECOM [book review]. IEEE Spectrum, 40(6):41, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:CWA

[Cas03c] Stephen Cass. Can’t we all just get along? IEEE Spectrum, 40 (1):47–50, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:FBF

[Cas03d] Stephen Cass. Flash of brilliance or flash in the pan? IEEE Spectrum, 40(7):18, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:HCJ

[Cas03e] Stephen Cass. Hacker cracker:

A journey from the mean streets of Brooklyn to the frontiers of cyberspace [book review]. IEEE Spectrum, 40(1):83–85, January 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:HFN

[Cas03f] Stephen Cass. How to fix the NASA disaster. IEEE Spectrum, 40(10):10–12, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:BEB

[Cas03g] Stephen Cass. In the blink of an eye [book review]. IEEE Spectrum, 40(6):41, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:IGL

[Cas03h] Stephen Cass. Inventor’s guide to law, business and taxes [book review]. IEEE Spectrum, 40 (8):45–48, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:LCU

[Cas03i] Stephen Cass. Linux’s challenge to Unix. IEEE Spectrum, 40(6):15–16, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:LES

[Cas03j] Stephen Cass. Listening in [electronic spying]. IEEE Spectrum, 40(4):32–37, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:MDH

[Cas03k] Stephen Cass. Masters of doom: How two guys created an empire and transformed pop culture [book review]. IEEE Spectrum, 40(5):43, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:SS

[Cas03l] Stephen Cass. Supercheap supercomputer. IEEE Spectrum, 40(7):17, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:WBR

[Cas03m] Stephen Cass. The wellstone [book review]. IEEE Spectrum, 40(6):40, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:BRC

[Cas04a] Stephen Cass. Book review: The Columbia Accident Investigation Report. IEEE Spectrum, 41(3):58, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:BRS

[Cas04b] Stephen Cass. Book review:

Small things considered. IEEE Spectrum, 41(3):56–58, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:BHS

[Cas04c] Stephen Cass. Books: Halfspeed ahead. IEEE Spectrum, 41 (10):55–57, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:CHT

[Cas04d] Stephen Cass. Can the Hubble Telescope be saved? IEEE Spectrum, 41(3):14–15, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:FNH

[Cas04e] Stephen Cass. Finding the needle in the haystack [search engines]. IEEE Spectrum, 41(12): 12–14, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:FKA

[Cas04f] Stephen Cass. Fountain of knowledge [analysis engine]. IEEE Spectrum, 41(1):68–71, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:JPG

[Cas04g] Stephen Cass. Julie payette: geek in space [astronaut engineer profile]. IEEE Spectrum, 41 (2):38–39, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2004:TTI

[Cas04h] Stephen Cass. Tools toys: iPod à go-go. IEEE Spectrum, 41 (8):49–51, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2005:AHR

[Cas05a] Stephen Cass. Ayanna howard: robot wrangler. IEEE Spectrum, 42(2):21–22, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2005:BES

[Cas05b] Stephen Cass. Books: The Earth strikes back: [Edison’s Conquest of Mars, by Garrett P. Serviss. Apogee Books Science Fiction, Burlington, Ont., Canada. Republished 2005, 264 pages, US $10. ISBN 0-9738-2030-6]. IEEE Spectrum, 42(9):64–68, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2005:BWN

[Cas05c] Stephen Cass. Books: Writing NASA’s marching orders. IEEE Spectrum, 42(3):63, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2005:GBF

[Cas05d] Stephen Cass. Genius on the block: the foundations of the computing age go up for auction. IEEE Spectrum, 42(7):40– 45, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2005:HG

[Cas05e] Stephen Cass. Holiday gifts.

IEEE Spectrum, 42(11):59–68, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2005:STT

[Cas05f] Stephen Cass. Software: Tiger trials. IEEE Spectrum, 42(7):59– 60, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2005:VMV

[Cas05g] Stephen Cass. Viva mesh Vegas [mesh wireless network]. IEEE Spectrum, 42(1):48–53, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:ASO

[Cas06a] Stephen Cass. Antipiracy software opens door to electronic intruders. IEEE Spectrum, 43 (1):12–13, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:BLF

[Cas06b] Stephen Cass. Books: Learning from failure. IEEE Spectrum, 43(6):51–52, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:BMN

[Cas06c] Stephen Cass. Books: Money for nothing reviewed. IEEE Spectrum, 43(10):60–62, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:IIM

[Cas06d] Stephen Cass. Invention: The idea man. IEEE Spectrum, 43 (8):45–48, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:MPP

[Cas06e] Stephen Cass. Media: Podcast picks. IEEE Spectrum, 43(6):53–54, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:MMP

[Cas06f] Stephen Cass. A method out of madness [professional profile]. IEEE Spectrum, 43(1): 55–57, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:STM

[Cas06g] Stephen Cass. Sci-tech: the movie — fifteen films that try to cross-pollinate art and science. IEEE Spectrum, 43(7):57– 58, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:TTB

[Cas06h] Stephen Cass. Tools and toys:

Better safe than sorry. IEEE Spectrum, 43(12):54–55, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:TTG

[Cas06i] Stephen Cass. Tools toys: Getting vexed. IEEE Spectrum, 43(5):68–69, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:TTHa

[Cas06j]  Stephen Cass.  Tools toys:

Hardware for your software radio. IEEE Spectrum, 43(10): 51–54, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:TTHb

[Cas06k] Stephen Cass. Tools toys: Holiday gifts. IEEE Spectrum, 43 (11):49–58, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:TTT

[Cas06l] Stephen Cass. Tools toys: Tech TV. IEEE Spectrum, 43(7):54– 55, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2006:VGS

[Cas06m] Stephen Cass. Video games 101/save a saturn: Buy this book [resources]. IEEE Spectrum, 43(5):66–67, May 2006.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:BSWa

[Cas07a] Stephen Cass. Books: Station watching. IEEE Spectrum, 44(2):57, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:BSWb

[Cas07b] Stephen Cass. Books: Strange ways. IEEE Spectrum, 44 (4):54–55, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:BTB

[Cas07c] Stephen Cass. Books: Threadbare theories. IEEE Spectrum, 44(1):55–58, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:CWJ

[Cas07d] Stephen Cass. Careers: Where the jobs are. IEEE Spectrum, 44 (2):51–57, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:HDP

[Cas07e] Stephen Cass. How to design a perfect city. IEEE Spectrum, 44(6):59, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:PPa

[Cas07f] Stephen Cass. Podcast picks. IEEE Spectrum, 44(1):59, January 2007. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:PPb

[Cas07g] Stephen Cass. Podcast picks.

IEEE Spectrum, 44(4):53, April 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:TTE

[Cas07h] Stephen Cass. Tools toys:

Exploring the third dimension. IEEE Spectrum, 44(2):58, February 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:TTY

[Cas07i] Stephen Cass. Tools toys: For your eyes only. IEEE Spectrum, 44(3):52–54, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2007:TTQ

[Cas07j] Stephen Cass. Tools toys: Quiet in demand. IEEE Spectrum, 44(4):52, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cavuoto:2004:NEI

[Cav04] J. Cavuoto. Neural engineering’s image problem. IEEE Spectrum, 41(4):32–37, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:WSg

[CC00a] S. Cherry and Stephen Cass.

Web sights. IEEE Spectrum, 37 (12):64, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:WSb

[CC00b] S. Cherry and R. Comerford. Web sights. IEEE Spectrum, 37(3):80, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:FF

[CCK00] S. Cherry, R. Comerford, and J. Kumagai. Faults failures. IEEE Spectrum, 37(6):30–32, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ceruzzi:2000:SSB

[Cer00] P. Ceruzzi. Sidling to success [books]. IEEE Spectrum, 37(6):10–12, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:WWT

[CG04] S. Cherry and Erico Guizzo. Write and wrong [technological breakthroughs]. IEEE Spectrum, 41(11):74–80, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:WDC

[CGC04] S. M. Cherry, H. Goldstein, and Stephen Cass. A world divided by a common Internet [national ranking calculations]. IEEE Spectrum, 41(2):56–57, February 2004. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Chatterjee:2004:ESS

[Cha04] A. C. Chatterjee. Europe struggles over software patents. IEEE Spectrum, 41(9):61–63, September 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2005:WSF

[Cha05] R. N. Charette. Why software fails [software failure]. IEEE Spectrum, 42(9):42–49, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chang:2006:TSU

[Cha06a] S. A. Chang. Tech startups spurn NASDAQ for London. IEEE Spectrum, 43(12): 14–15, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2006:DD

[Cha06b] R. N. Charette. Dying for data. IEEE Spectrum, 43(10): 22–27, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chang:2007:CC

[Cha07a] S. A. Chang. Carbon commerce. IEEE Spectrum, 44 (4):12–15, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2007:OSW

[Cha07b] R. N. Charette. Open-source warfare. IEEE Spectrum, 44(11): 26–32, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2008:WWW

[Cha08]  R. N. Charette.  What’s

wrong with weapons acquisitions? IEEE Spectrum, 45(11): 33–39, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chang:2009:CTC

[Cha09a] S. A. Chang. Cap and trade or cap and tax? IEEE Spectrum, 46(4):8, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Charette:2009:EMY

[Cha09b] R. N. Charette. Every move you make. IEEE Spectrum, 46 (12):7, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:SLD

[Che00a] S. Cherry. Secrets and lies: digital security in a networked world [books]. IEEE Spectrum, 37 (10):15–16, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:WSa

[Che00b] S. Cherry. Web sights. IEEE Spectrum, 37(1):122, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:WSc

[Che00c] S. Cherry. Web sights. IEEE Spectrum, 37(4):63–64, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:WSd

[Che00d] S. Cherry. Web sights. IEEE Spectrum, 37(5):66, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:WSe

[Che00e] S. Cherry. Web sights. IEEE Spectrum, 37(8):73, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2000:WSf

[Che00f] S. Cherry. Web sights. IEEE Spectrum, 37(11):71–72, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:EWS

[Che01a] S. Cherry. Euclid’s window:

the story of geometry from parallel lines to hyperspace, by Leonard Mlodinow, Simon & Schuster, New York, $26 [book review]. IEEE Spectrum, 38 (11):67, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:UGD

[Che01b] S. Cherry. The unwanted gaze: the destruction of privacy in America, By Jeffrey Rosen, Knopf, New York $13.00 [book review]. IEEE Spectrum, 38 (11):67, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:HBE

[Che01c] S. Cherry. Hitting the books every which way. IEEE Spectrum, 38(3):81, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:HMH

[Che01d] S. M. Cherry. Hyperencryption: Much hype about little that is new [web sites]. IEEE Spectrum, 38(4):87, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:HDS

[Che01e] S. M. Cherry. Is hyperlinking to decryption software illegal? IEEE Spectrum, 38 (8):64–65, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:MMP

[Che01f] S. M. Cherry. Making music pay [digital music distribution]. IEEE Spectrum, 38(10): 41–46, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:MOO

[Che01g] S. M. Cherry. Marketing oneon-one [book review]. IEEE Spectrum, 38(12):61–62, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:REM

[Che01h] S. M. Cherry. Remailers elude e-mail surveillance. IEEE Spectrum, 38(11):69, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:STS

[Che01i] S. M. Cherry. Streaming toward success [web sights]. IEEE Spectrum, 38(10):68, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:SIH

[Che01j] S. M. Cherry. Striking at the Internet’s heart. IEEE Spectrum, 38(12):66–68, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:TLV

[Che01k] S. M. Cherry. Testing on-line voting [web sites]. IEEE Spectrum, 38(2):81, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2001:WDN

[Che01l] S. M. Cherry. Websights [domain name registration]. IEEE Spectrum, 38(6):83, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:AIF

[Che02a] S. M. Cherry. All the ills that flesh is heir to [Internet viruses]. IEEE Spectrum, 39 (11):52, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:BSU

[Che02b] S. M. Cherry. Bandwidth shapers unclog busy networks. IEEE Spectrum, 39(2):61, February 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:GCR

[Che02c] S. M. Cherry. Getting copyright right. IEEE Spectrum, 39 (2):47–51, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:IMB

[Che02d] S. M. Cherry. Im means business. IEEE Spectrum, 39(11): 28–32, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:LPR

[Che02e] S. M. Cherry. The largest players rule the media playground. IEEE Spectrum, 39 (7):32–33, July 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:LNS

[Che02f] S. M. Cherry. Linked: the new science of networks [book review]. IEEE Spectrum, 39 (6):57–58, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:LFM

[Che02g] S. M. Cherry. Lost and found:

media sales. IEEE Spectrum, 39(7):45, July 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:PHN

[Che02h] S. M. Cherry. Pesky home networks trouble cable behemoths. IEEE Spectrum, 39 (4):60–62, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:TCT

[Che02i] S. M. Cherry. Talk is cheap, text is cheaper [mobile messaging]. IEEE Spectrum, 39(5):55, May 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:TLK

[Che02j] S. M. Cherry. Took a licking, kept on ticking [Internet security]. IEEE Spectrum, 39 (12):49, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:WWI

[Che02k] S. M. Cherry. Weaving a web of ideas. IEEE Spectrum, 39 (9):65–69, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:WRT

[Che02l] S. M. Cherry. Web radio: time to sign off? IEEE Spectrum, 39(8):53, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2002:WWW

[Che02m] S. M. Cherry. What went wrong at ricochet? IEEE Spectrum, 39(3):60–61, March 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:BWW

[Che03a] S. M. Cherry. The blog of war [web journalism]. IEEE Spectrum, 40(6):48, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:BGG

[Che03b] S. M. Cherry. Broadband a gogo. IEEE Spectrum, 40(6):20–25, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:BBB

[Che03c] S. M. Cherry. Broadband and the baby bells. IEEE Spectrum, 40(4):57, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:HM

[Che03d] S. M. Cherry. Hell month. IEEE Spectrum, 40(10):48, October 2003. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:ISA

[Che03e] S. M. Cherry. Internet slammed again [hacking]. IEEE Spectrum, 40(3):59, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:MAW

[Che03f] S. M. Cherry. More air for Wi-Fi?

IEEE Spectrum, 40(2):51, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:PMB

[Che03g] S. M. Cherry. Putting the move back in movies. IEEE Spectrum, 40(10):49, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:SMN

[Che03h] S. M. Cherry. Smart mobs:

The next social revolution [book review]. IEEE Spectrum, 40 (3):57–58, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:TDL

[Che03i] S. M. Cherry. Tia is dead — long live tia. IEEE Spectrum, 40(11):22, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:WDG

[Che03j]  S. M. Cherry.  Websights:

David and Goliath versus another David. IEEE Spectrum, 40(12):47, December 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:WRT

[Che03k] S. M. Cherry. What’s right with telecom. IEEE Spectrum, 40 (1):30–34, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:WFT

[Che03l] S. M. Cherry. Wi-Fi takes new turn with “wireless-G”. IEEE Spectrum, 40(8):12–13, August 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2003:WLM

[Che03m] S. M. Cherry. The wireless last mile. IEEE Spectrum, 40(9): 18–22, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:EB

[Che04a] S. Cherry. eBuyer, beware. IEEE Spectrum, 41(10):60, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:ELB

[Che04b] S. Cherry. Edholm’s law of bandwidth. IEEE Spectrum, 41(7):58– 60, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:PPb

[Che04c] S. Cherry. The perils of polling. IEEE Spectrum, 41 (10):34–40, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:SMS

[Che04d] S. Cherry. Selling music for a song [online music]. IEEE Spectrum, 41(12):56, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:TTW

[Che04e] S. Cherry. Tools toys: When less is more. IEEE Spectrum, 41 (9):65–66, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:WEB

[Che04f] S. Cherry. Websights: eBuyer, beware. IEEE Spectrum, 41 (10):60, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:AGD

[Che04g] S. M. Cherry. Across the great divide. IEEE Spectrum, 41 (1):36–39, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:EVW

[Che04h] S. M. Cherry. Eolas vs web. IEEE Spectrum, 41(1): 85–86, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:FH

[Che04i] S. M. Cherry. Fiber to the home. IEEE Spectrum, 41(1):

42–43, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:IWP

[Che04j] S. M. Cherry. The illusion of web privacy. IEEE Spectrum, 41(4):56–59, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:NT

[Che04k] S. M. Cherry. The network not taken. IEEE Spectrum, 41(6):60–61, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:PPa

[Che04l] S. M. Cherry. The patent profiteers. IEEE Spectrum, 41 (6):38–43, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:WOS

[Che04m] S. M. Cherry. Who owns streaming media? [patent issues]. IEEE Spectrum, 41 (3):61–62, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2004:WWF

[Che04n] S. M. Cherry. WiMax and Wi-Fi: Separate and unequal. IEEE Spectrum, 41(3):16, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:AME

[Che05a] S. Cherry. Anger management [emotion recognition]. IEEE Spectrum, 42(4):16, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:BBI

[Che05b] S. Cherry. The battle for broadband [Internet protocol television]. IEEE Spectrum, 42(1): 24–29, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:BSW

[Che05c] S. Cherry. Books: Sorry, wrong numbers. IEEE Spectrum, 42 (5):60–63, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:BB

[Che05d] S. Cherry. Broader broadband. IEEE Spectrum, 42(1): 17–20, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:DDE

[Che05e] S. Cherry. Digital dullard [electronic science tutor]. IEEE Spectrum, 42(1):46–47, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:EHW

[Che05f] S. Cherry. Ethernet’s highwire act [new IEEE standard]. IEEE Spectrum, 42(4):63–65, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:FML

[Che05g] S. Cherry. Fritz Morgan: LEDs into gold. IEEE Spectrum, 42 (2):22–23, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:MMP

[Che05h] S. Cherry. A match made in packets — coming soon: cellular handsets that can use a Wi-Fi network. IEEE Spectrum, 42(7):14–16, July 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:MDC

[Che05i] S. Cherry. My dad’s computer: a conservation with Internet security expert William R. Cheswick. IEEE Spectrum, 42 (8):55–56, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:NEC

[Che05j] S. Cherry. The net effect: as China’s Internet gets a muchneeded makeover, will the new network promote freedom or curtail it? IEEE Spectrum, 42 (6):38–44, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:SMA

[Che05k] S. Cherry. Seven myths about voice over ip. IEEE Spectrum, 42(3):52–57, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:SKP

[Che05l] S. Cherry. South Korea pushes mobile broadband. IEEE Spectrum, 42(9):14–16, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:TGO

[Che05m] S. Cherry. Terror goes online [cyberterrorism]. IEEE Spectrum, 42(1):72–73, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:TRL

[Che05n] S. Cherry. Total recall [life recording software]. IEEE Spectrum, 42(11):24–30, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:VB

[Che05o] S. Cherry. The VoIP backlash. IEEE Spectrum, 42(10): 61–63, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2005:WWY

[Che05p] S. Cherry. Watch what you want [pornographic image sensor]. IEEE Spectrum, 42(12): 15–16, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2006:BUF

[Che06a] S. Cherry. A broadband utopia [fast broadband connectivity]. IEEE Spectrum, 43(5):48–54, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2006:CYP

[Che06b] S. Cherry. Cisco and Yahoo’s plan to damn spam. IEEE Spectrum, 43(1):51, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2006:DJE

[Che06c] S. Cherry. Dream jobs [engineering career]. IEEE Spectrum, 43 (2):36–48, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2006:NVD

[Che06d] S. Cherry. The next voting debacle? IEEE Spectrum, 43 (10):12–14, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2006:RBW

[Che06e] S. Cherry. A radio bypass [wireless broadband]. IEEE Spectrum, 43(1):32–35, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2006:WVP

[Che06f] S. Cherry. Websights: Virtually private. IEEE Spectrum, 43 (12):52–55, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2006:WFN

[Che06g] S. Cherry. Wi-Fi nodes to talk amongst themselves [mesh technology standard that lowers cost of wireless networks]. IEEE Spectrum, 43(7):55–56, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2007:BI

[Che07a] S. Cherry. Bet on it. IEEE Spectrum, 44(9):48–53, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2007:CIM

[Che07b] S. Cherry. Commentary: The iPhone’s missing buttons [news]. IEEE Spectrum, 44(10):14, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2007:HBG

[Che07c] S. Cherry. How to build a green city. IEEE Spectrum, 44(6):26–29, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2007:MEV

[Che07d] S. Cherry. Making every evote count. IEEE Spectrum, 44 (1):13–14, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2007:NN

[Che07e] S. Cherry. Nothing but net. IEEE Spectrum, 44(1): 22–26, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2007:RI

[Che07f] S. Cherry. Robots incorporated. IEEE Spectrum, 44 (8):24–29, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2007:WYE

[Che07g] S. Cherry. A word in your ear [resources: Tools toys]. IEEE Spectrum, 44(12):65–66, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:BRE

[Che08a] S. Cherry. Blu-ray’s empty victory. IEEE Spectrum, 45(3):16, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:BPO

[Che08b] S. Cherry. Books: patently oblivious [review of Patent Savvy for Managers: Spot Protect Valuable Innovations in Your Company (Teska, K.; 2007)]. IEEE Spectrum, 45(8):25, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:DTR

[Che08c] S. Cherry. The data: thx 4 the revnu — [data]. IEEE Spectrum, 45(10):76, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:IDF

[Che08d] S. Cherry. Im doing fine [tools toys]. IEEE Spectrum, 45(1): 27–28, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:SSS

[Che08e] S. Cherry. Spotting snipers with sound. IEEE Spectrum, 45 (12):14, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:UWT

[Che08f] S. Cherry. Update: WiFi takes on Bluetooth. IEEE Spectrum, 45(8):14, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:WTS

[Che08g] S. Cherry. Winner: Telecommunications — Sprint’s broadband gamble. IEEE Spectrum, 45(1):50–54, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:CTT

[Che09a] S. Cherry. Corporate travel down, telecommuting up. IEEE Spectrum, 46(2):68, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:FCC

[Che09b] S. Cherry. Forecast for cloud computing: Up, up, and away. IEEE Spectrum, 46(10):68, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:HTU

[Che09c] S. Cherry. High-tech unemployment is up, but not way up. IEEE Spectrum, 46(5):64, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:LLA

[Che09d] S. Cherry. Long live analog TV — on cellphones. IEEE Spectrum, 46(11):84, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:RNG

[Che09e] S. Cherry. The race for a new game machine: Creating the chips inside the XBox 360 and the PlayStation 3. IEEE Spectrum, 46(3):24, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:RWA

[Che09f] S. Cherry. The right word, anywhere. IEEE Spectrum, 46(7):24, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:SBB

[Che09g] S. Cherry. Space is big business. IEEE Spectrum, 46(6):92, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:TDR

[Che09h] S. Cherry. A telecom diet rich in fiber. IEEE Spectrum, 46(7): 56, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:TOD

[Che09i] S. Cherry. Tero ojanpera dial m for music. IEEE Spectrum, 46(2):31, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:SF

[Che09j] S. Cherry. That sinking feeling. IEEE Spectrum, 46(1): 49–50, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:WWE

[Che09k] S. Cherry. When will energy prices recover? IEEE Spectrum, 46(9):68, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2009:YNC

[Che09l] S. Cherry. Your next computer will probably be a notebook — [the data]. IEEE Spectrum, 46(1):84, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2002:RCR

[Chi02a]  Yu-Tzu Chiu.  Russia and

Canada to ratify Kyoto Protocol. IEEE Spectrum, 39(10): 22–23, October 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2002:SWR

[Chi02b] Yu-Tzu Chiu. A second wind for red-laser DVD technology? IEEE Spectrum, 39 (11):18, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2005:PAT

[Chi05] Yu-Tzu Chiu. Putin acts, Taiwan reacts [Taiwan’s greenhouse reduction plans]. IEEE Spectrum, 42(1):20–21, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2006:THT

[Chi06] Yu-Tzu Chiu. Taiwan’s hightech hubbub. IEEE Spectrum, 43(8):16, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2007:CGT

[Chi07] Y. Chiu. Carlyle Group’s Taiwan gambit. IEEE Spectrum, 44(2):18, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2009:FGS

[Chi09] Y. Chiu. Foundry giant sees the light. IEEE Spectrum, 46(8):19, August 2009. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Chowdhury:2000:PEC

[Cho00] B. H. Chowdhury. Power education at the crossroads. IEEE Spectrum, 37(10):64–69, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ciciora:2001:CMT

[Cic01] W. S. Ciciora. Cable modem traffic jam. IEEE Spectrum, 38(6):48–53, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cioffi:2001:LHO

[Cio01] D. F. Cioffi. Learning from hackers [open-source programming]. IEEE Spectrum, 38 (6):77–79, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christensen:2008:NES

[CKVY08] C. M. Christensen, S. King, M. Verlinden, and W. Yang. The new economics of semiconductor manufacturing. IEEE Spectrum, 45(5):24–29, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Claverie:2001:HGP

[Cla01] Jean-Michel Claverie. The Human Genome Project: help wanted. IEEE Spectrum, 38(3):9, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Claudy:2005:CED

[Cla05] L. Claudy. Countdown to the end [digital TV]. IEEE Spectrum, 42 (10):30–31, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2003:LCM

[CM03] Yu-Tzu Chiu and Samuel K. Moore. Leaking capacitors muck up motherboards. IEEE Spectrum, 40(2):16–17, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chiu:2009:TTD

[CM09] Y. Chiu and Samuel K. Moore. Taiwan’s troubled DRAM plan. IEEE Spectrum, 46(6):11–12, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Coetzee:2004:HDR

[Coe04] F. M. Coetzee. Here be dragons — rule 1 in managing a technology start-up: beware the easy money. IEEE Spectrum, 41(9): 64–65, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Coffa:2005:LSL

[Cof05] S. Coffa. Ligth from silicon [led]. IEEE Spectrum, 42(10): 44–49, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cohen:2009:TAB

[Coh09] D. X. Cohen. The truth about bender’s brain. IEEE Spectrum, 46(5):24–25, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Colot:2005:PES

[Col05] A. Colot. Photo essay: Sand biscuit. IEEE Spectrum, 42(7):20– 21, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Colclasure:2006:MBM

[Col06] K. Colclasure. Mr. Babbage, meet Mr. Lego. IEEE Spectrum, 43(4):26–27, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cole:2007:DHJ

[Col07a]  S. A. Cole.  Double helix

jeopardy. IEEE Spectrum, 44 (8):44–49, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Colozza:2007:FLB

[Col07b] A. Colozza. Fly like a bird.

IEEE Spectrum, 44(5):38–43, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2000:CTA

[Com00a] R. Comerford. Computing [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1): 45–50, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2000:HDI

[Com00b] R. Comerford. Handhelds duke it out for the Internet. IEEE Spectrum, 37(8):35–41, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2000:ITA

[Com00c] R. Comerford. The Internet [technology 2000 analysis and forecast]. IEEE Spectrum, 37 (1):40–44, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2000:MSM

[Com00d] R. Comerford. Magnetic storage: the medium that wouldn’t die. IEEE Spectrum, 37(12): 36–39, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2000:NSP

[Com00e] R. Comerford. The new software paladins. IEEE Spectrum, 37(6):56–61, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2000:WSa

[Com00f] R. Comerford. Web sights. IEEE Spectrum, 37(2):60–61, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2000:WSb

[Com00g] R. Comerford. Web sights. IEEE Spectrum, 37(10):79, October 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2001:KN

[Com01a] R. Comerford. Keeping the net up. IEEE Spectrum, 38 (1):47–50, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:2001:NLD

[Com01b] R. Comerford. No longer in denial [network security]. IEEE Spectrum, 38(1):59–61, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cooper:2002:WAB

[Coo02] M. Cooper. What ails broadband? IEEE Spectrum, 39(9): 15–16, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Coren:2000:DSP

[Cor00] R. L. Coren. Don’t be shocked:

power lines are safe! [speakout]. IEEE Spectrum, 37(9): 22, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2002:CKE

[Cos02a] T. Costlow. Computer kiosk expedites voter registration. IEEE Spectrum, 39(10):26, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2002:EO

[Cos02b] T. Costlow. The ethical organization. IEEE Spectrum, 39 (12):56–58, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2003:ESI

[Cos03] T. Costlow. EE salaries: it could be worse. IEEE Spectrum, 40(7):46–47, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2004:RTT

[Cos04] T. Costlow. Resources — tough in the trenches. IEEE Spectrum, 41(7):51–52, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2005:RML

[Cos05a] T. Costlow. Read my lips [lipsyncing]. IEEE Spectrum, 42(7): 54, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2005:REJ

[Cos05b] T. Costlow. Resources EE job market brightens. IEEE Spectrum, 42(1):63–64, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2005:RWU

[Cos05c] T. Costlow. Resources: wages in the United States grow only slightly, while China and India see double-digit increases. IEEE Spectrum, 42(7):53–54, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2006:RES

[Cos06] T. Costlow. Resources: Ee salaries up all over [careers]. IEEE Spectrum, 43(7):52–54, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costlow:2007:RCH

[Cos07] T. Costlow. Resources: Careers:

Hiring heats up. IEEE Spectrum, 44(7):57–60, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2002:ISP

[CR02] Stephen Cass and M. J. Riezenman. Improving security, preserving privacy. IEEE Spectrum, 39(1):44–49, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Crawford:2000:BNP

[Cra00] G. P. Crawford. A bright new page in portable displays. IEEE Spectrum, 37(10):40–46, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Change:2008:ENS

[CS08a] J. B. Change and V. Subramanian. Electronic noses sniff success. IEEE Spectrum, 45 (3):50–56, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cherry:2008:PMT

[CS08b] S. Cherry and D. Schneider. A poor man’s tesla. IEEE Spectrum, 45(12):27, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chalamala:2005:BBF

[CT05] B. R. Chalamala and D. Temple. Big and bendable [flexible plastic-based circuits]. IEEE Spectrum, 42(9):50–56, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Curry:2001:VCP

[Cur01] D. G. Curry. In-vehicle cell phones: smoke, but where’s the fire? IEEE Spectrum, 38 (8):16–18, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2001:BBS

[Das01] S. R. Das. Big Bear scientists use Moon to measure Earth’s reflectance. IEEE Spectrum, 38(6):27–28, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2002:SSM

[Das02] S. R. Das. The sensible superconductor [magnesium diboride]. IEEE Spectrum, 39(7):34–37, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2003:CST

[Das03] S. R. Das. Coming soon: Terabit hard disk drives. IEEE Spectrum, 40(2):19–20, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2005:ABW

[Das05] S. R. Das. The answer is blowing in the wind [hurricane prediction]. IEEE Spectrum, 42 (9):20, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2008:BQC

[Das08a] S. R. Das. Breaking quantum cryptography’s 150-kilometer limit [update]. IEEE Spectrum, 45(9):15, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2008:NLS

[Das08b] S. R. Das. A new light source for EUV lithography. IEEE Spectrum, 45(3):14, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2008:UMU

[Das08c] S. R. Das. Unsticking MEMS — [update]. IEEE Spectrum, 45(10):18, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2009:SN

[Das09a] S. R. Das. The spaser nanolaser. IEEE Spectrum, 46 (10):14–16, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2009:TLL

[Das09b] S. R. Das. Two-laser lithography shrinks transistors on the cheap. IEEE Spectrum, 46(6):16, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Day:2002:CBD

[Day02]  R. Day.  Choosing between

dvd-audio and super audio cd? IEEE Spectrum, 39(5):60–62, May 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeMarco:2001:SIE

[De 01] S. J. De Marco. Solving the innovation equation. IEEE Spectrum, 38(4):9–10, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeWeck:2009:WPM

[De 09] O. L. De Weck. What to pack for Mars. IEEE Spectrum, 46(6):39, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Denisowski:2001:HDI

[Den01] P. Denisowski. How does it sound? [voice clarity analysis]. IEEE Spectrum, 38(2): 60–64, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2008:SBN

[Der08] D. P. Dern. Small is big in notebooks — but not too small. IEEE Spectrum, 45(12):26–27, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2009:FAF

[Der09a] D. P. Dern. First affordable fuel cells for mobile gear. IEEE Spectrum, 46(4):16, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2009:I

[Der09b] D. P. Dern. Instant on. IEEE Spectrum, 46(9):28, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dern:2009:WCP

[Der09c] D. P. Dern. World’s coolest personal computer. IEEE Spectrum, 46(8):23, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Diakov:2000:MNF

[DG00] A. S. Diakov and J. E. Goodby.

Mending nuclear fences [arms control]. IEEE Spectrum, 37 (3):54–58, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Deblois:2005:SCS

[DGKM05] B. M. Deblois, R. L. Garwin, S. Kemp, and J. C. Marwell. Star-crossed [space-based weapons]. IEEE Spectrum, 42 (3):40–49, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Das:2001:USW

[DHGR01] S. R. Das, A. Hellemans, P. Gwynne, and M. J. Riezenman. Unsuspected superconductor wows physicists, technologists [news analysis]. IEEE Spectrum, 38(4):18–26, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DiGregorio:2003:MDA

[DiG03] B. E. DiGregorio. Mars: dead or alive? IEEE Spectrum, 40(5):36–41, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2004:SCM

[Dig04] B. E. Digregorio. Spacecraft converge on Mars. IEEE Spectrum, 41(10):22–23, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2006:BAS

[Dig06a] B. E. Digregorio. A better anthrax screener? IEEE Spectrum, 43(3):18–21, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2006:RWP

[Dig06b] B. E. Digregorio. Roundabout way of profiling Earth’s atmosphere. IEEE Spectrum, 43 (4):22–23, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2006:SDD

[Dig06c] B. E. Digregorio. Safer driving in the dead of night [infrared vision systems]. IEEE Spectrum, 43(3):20–21, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DiGregorio:2006:TSS

[DiG06d] B. E. DiGregorio. Tsunami surveillance by satellites. IEEE Spectrum, 43(10):21, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2007:CRR

[Dig07a] B. E. Digregorio. China reaches for the Red Planet. IEEE Spectrum, 44(8):15–16, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2007:MAM

[Dig07b] B. E. Digregorio. Messenger arrives at mercury [news]. IEEE Spectrum, 44(12):14–16, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DiGregorio:2007:TWC

[DiG07c] B. E. DiGregorio. There will be cooking on Mars. IEEE Spectrum, 44(5):16, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2007:ZNC

[Dig07d] B. E. Digregorio. Zoos network to contain disease [news].

IEEE Spectrum, 44(12):20, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2008:ATA

[Dig08a] B. E. Digregorio. Allen Telescope Array starts search for E.T. [update]. IEEE Spectrum, 45(9):16, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2008:UES

[Dig08b] B. E. Digregorio. Update: Earthsize radio telescope opens its eye. IEEE Spectrum, 45(8):15, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Digregorio:2008:UGS

[Dig08c] B. E. Digregorio. Update: Gps signals spot signs of typhoons. IEEE Spectrum, 45(4):15, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ditlea:2000:PGR

[Dit00] S. Ditlea. The PC goes readyto-wear. IEEE Spectrum, 37 (10):34–39, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dixit:2005:AMD

[Dix05] J. Dixit. The artful- and mobiledodger [location-based social networking]. IEEE Spectrum, 42(3):59–60, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Divan:2009:OG

[DK09] D. Divan and F. Kreikebaum.

Organic (but not green). IEEE Spectrum, 46(11):48–53, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:2000:CQI

[DR00a] A. Dutta-Roy. The cost of quality in Internet-style networks. IEEE Spectrum, 37(9): 57–62, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:2000:IST

[DR00b] A. Dutta-Roy. The Internet — a spontaneous technology [book review]. IEEE Spectrum, 37(9): 10–12, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dutta-Roy:2000:SR

[DR00c] A. Dutta-Roy. Software reviews. IEEE Spectrum, 37(10): 95–97, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2006:CI

[Dum06a] M. Dumiak. Cells on ice. IEEE Spectrum, 43(9):18, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2006:EDD

[Dum06b] M. Dumiak. Europe dithers over digital radio. IEEE Spectrum, 43(10):20, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2007:NTO

[Dum07a] M. Dumiak. New tech, old fuel. IEEE Spectrum, 44(2):13, February 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2007:TTF

[Dum07b] M. Dumiak. A telescope takes flight [news]. IEEE Spectrum, 44(7):22, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2008:BSR

[Dum08a] M. Dumiak. Book-scanning robots digitize delicate texts. IEEE Spectrum, 45(1):18, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2008:UMF

[Dum08b] M. Dumiak. Update: Magnetic field sensors could help halt runway crashes. IEEE Spectrum, 45(4):13, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2009:RU

[Dum09a] M. Dumiak. Robocopters unite! IEEE Spectrum, 46(2):12, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dumiak:2009:SBT

[Dum09b] M. Dumiak. Setting bait to track data thieves. IEEE Spectrum, 46(5):11–12, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Durniak:2001:RRR

[Dur01]  A. Durniak.  The refillable

recharger [review — power to the phones]. IEEE Spectrum, 38(8):69, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Durniak:2003:PBR

[Dur03] A. Durniak. Prey [book review].

IEEE Spectrum, 40(2):48, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Durniak:2005:RID

[Dur05] A. Durniak. Resources: The Innovator’s Dilemma: 3.0 (Christensen, C. M., et al. Harvard Business School Press, 2004) [book review]. IEEE Spectrum, 42(4):58–60, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dixit:2001:SIF

[DY01] S. Dixit and Yinghua Ye.

Streamlining the Internet–fiber connection. IEEE Spectrum, 38(4):52–57, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eads:2000:WSI

[Ead00] R. Eads. Web style intranet speeds design flow. IEEE Spectrum, 37(6):75–79, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eastman:2000:ETC

[Eas00] M. Eastman. The essence of technical communication for engineers: writing, presentation, and meeting skills [books]. IEEE Spectrum, 37(8):11–14, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eastham:2003:RR

[Eas03] T. R. Eastham. Running off the rails. IEEE Spectrum, 40 (1):66–69, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eden:2000:ETT

[Ede00] M. Eden. Engineering tomorrow: today’s technology experts envision the next century. IEEE Spectrum, 37(1): 10–15, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Edwards:2005:HHS

[Edw05] B. C. Edwards. A hoist to the heavens [space elevators]. IEEE Spectrum, 42(8):36–41, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eichenberger:2000:SR

[Eic00] K. A. Eichenberger. Software reviews. IEEE Spectrum, 37(12): 65–68, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ehsani:2000:LSB

[EK00] A. R. Ehsani and M. Kesler.

Lasers speed up board production. IEEE Spectrum, 37 (5):40–45, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Elliott:2001:BWI

[Ell01] C. Elliott. Building the wireless Internet. IEEE Spectrum, 38(1):14–16, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eastman:2002:TTY

[EM02] L. F. Eastman and U. K. Mishra.

The toughest transistor yet [GaN transistors]. IEEE Spectrum, 39(5):28–33, May 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

English:2001:PCF

[Eng01] G. English. Putting consumers first [electricity supply industry]. IEEE Spectrum, 38(6):19–20, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Enriquez:2007:EWE

[Enr07] A. Enriquez. Early warning for earthquakes. IEEE Spectrum, 44 (2):14–16, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ensmenger:2000:MFC

[Ens00]  N. Ensmenger.  The many

firsts of computer development [books]. IEEE Spectrum, 37(12): 12–15, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Evans:2004:SSS

[ES04] M. Evans and W. Sweet. Seven seas sop up CO_{2} [CO_{2} in the oceans]. IEEE Spectrum, 41 (9):20, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eyre:2001:DSP

[Eyr01] J. Eyre. The digital signal processor derby. IEEE Spectrum, 38(6):62–68, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2003:BSD

[Fai03a] P. Fairley. BP Solar ditches thinfilm photovoltaics. IEEE Spectrum, 40(1):18–19, January 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2003:GBC

[Fai03b] P. Fairley. Germs that build circuits. IEEE Spectrum, 40(11): 36–41, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2003:SSB

[Fai03c]  P. Fairley.  Steady as she

blows [wind power, energy storage]. IEEE Spectrum, 40(8): 35–39, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2004:COR

[Fai04a] P. Fairley. Can organics replace silicon in PV? IEEE Spectrum, 41(1):28–30, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2004:LAL

[Fai04b] P. Fairley. Lighting up the andes [led lamps]. IEEE Spectrum, 41 (12):44–49, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2004:UPG

[Fai04c] P. Fairley. The unruly power grid. IEEE Spectrum, 41(8): 22–27, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2004:UMG

[Fai04d] P. Fairley. U.s.military goes for hybrid vehicles. IEEE Spectrum, 41(3):22–25, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2005:CCT

[Fai05a] P. Fairley. China’s cyclists take charge: electric bicycles are selling by the millions despite efforts to ban them. IEEE Spectrum, 42(6):54–59, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2005:GGA

[Fai05b] P. Fairley. The greening of GE [alternative energy]. IEEE Spectrum, 42(7):28–33, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairly:2005:NRU

[Fai05c] P. Fairly. Neptune rising [undersea observatory]. IEEE Spectrum, 42(11):38–45, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2007:CRF

[Fai07a] P. Fairley. California to rule on fate of EVs [news]. IEEE Spectrum, 44(11):10–12, November 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2007:NWL

[Fai07b]  P. Fairley.  Nuclear waste

land. IEEE Spectrum, 44(2): 38–44, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2007:SC

[Fai07c] P. Fairley. Syn city. IEEE Spectrum, 44(6):20–21, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2008:CWE

[Fai08a] P. Fairley. Can wind energy continue double-digit growth? IEEE Spectrum, 45(2):16, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2008:CC

[Fai08b] P. Fairley. Closing the circuit. IEEE Spectrum, 45(11): 52–60, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2008:SCS

[Fai08c] P. Fairley. Solar-cell squabble.

IEEE Spectrum, 45(4):36–40, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2008:UCD

[Fai08d] P. Fairley. Update: China doubles wind watts. IEEE Spectrum, 45(5):11–12, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2009:DA

[Fai09a] P. Fairley. Driving on air. IEEE Spectrum, 46(11):30–35, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2009:ECB

[Fai09b] P. Fairley. Europe comes to bury CO_{2}, not to praise it. IEEE Spectrum, 46(8):64, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2009:ERO

[Fai09c] P. Fairley. Europe replaces old wind farms — [update]. IEEE Spectrum, 46(1):13, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2009:GGE

[Fai09d] P. Fairley. Germany’s green — energy gap. IEEE Spectrum, 46(7):40–48, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2009:KCE

[Fai09e] P. Fairley. King Coal eats its vegetables. IEEE Spectrum, 46(7): 16, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fairley:2009:PSC

[Fai09f] P. Fairley. Plastic solar cells roll into unlit villages. IEEE Spectrum, 46(11):13–14, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Farber:2003:FNR

[Far03] D. Farber. Fame, but no riches, for cybersecurity. IEEE Spectrum, 40(1):51–52, January 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Farid:2009:SB

[Far09] H. Farid. Seeing is not believing. IEEE Spectrum, 46 (8):44–51, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feiveson:2000:HLN

[FB00] H. A. Feiveson and B. G.

Blair. How to lengthen the nuclear fuse. IEEE Spectrum, 37 (3):40–43, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forrest:2000:DOE

[FBT00] S. Forrest, P. Burrows, and M. Thompson. The dawn of organic electronics. IEEE Spectrum, 37(8):29–34, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feather:2009:PDR

[Fea09] C. Feather. Puzzles by the drawerful (review of Professor Stewart’s Cabinet of Mathematical Curiosities by I. Stewart; 2008) [book review]. IEEE Spectrum, 46(7):25, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ferorelli:2001:CR

[Fer01] F. Ferorelli. Cuddling up to robots. IEEE Spectrum, 38 (8):66–67, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Federici:2007:RVT

[FGBM07] J. F. Federici, D. Gary, R. Barat, and Z. Michalopoulou. T-rays vs. terrorists. IEEE Spectrum, 44(7):47–52, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Furse:2001:WAW

[FH01] C. Furse and R. Haupt. Down to the wire [aircraft wiring]. IEEE Spectrum, 38(2):34–39, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Finne:2005:F

[Fin05] R. Finne. Forum. IEEE Spectrum, 42(6):6, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2007:RI

[FJ07] Kenneth R. Foster and J. Jaeger. RFID inside. IEEE Spectrum, 44(3):24–29, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2000:MPS

[FM00] Kenneth R. Foster and J. E. Moulder. Are mobile phones safe? IEEE Spectrum, 37 (8):23–28, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fogel:2000:WEC

[Fog00] D. B. Fogel. What is evolutionary computation? IEEE Spectrum, 37(2):26–32, February 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2000:SR

[Fos00a] Kenneth R. Foster. Software reviews. IEEE Spectrum, 37(6):86, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2000:URS

[Fos00b] Kenneth R. Foster. Undue risks:

secret state experiments on humans [books]. IEEE Spectrum, 37(2):14–16, February 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:WCS

[Fos01a] K. Foster. White-collar sweatshop [book review]. IEEE Spectrum, 38(5):89–92, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:CMS

[Fos01b] Kenneth R. Foster. Controlling Mathematica [software]. IEEE Spectrum, 38(5):94–95, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:MSA

[Fos01c] Kenneth R. Foster. Maple scores again. IEEE Spectrum, 38 (12):63, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:MMW

[Fos01d] Kenneth R. Foster. Mathtype 5 with mathML for the WWW. IEEE Spectrum, 38(12):64, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2001:MME

[Fos01e] Kenneth R. Foster. Matlab made easier and testing with Matlab [software resources]. IEEE Spectrum, 38(2):87–89, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2002:CCW

[Fos02a] Kenneth R. Foster. Are classroom computers worthwhile? [book review]. IEEE Spectrum, 39(2):60, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2002:PWP

[Fos02b] Kenneth R. Foster. Putting work in its place: a quiet revolution [book review]. IEEE Spectrum, 39(7):44, July 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2003:MM

[Fos03a] Kenneth R. Foster. Mightier math. IEEE Spectrum, 40(2): 58–59, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2003:SRM

[Fos03b] Kenneth R. Foster. Software reviews mightier math reloaded. IEEE Spectrum, 40(7):40–41, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2004:RF

[Fos04a] K. Foster. Retaking the field.

IEEE Spectrum, 41(7):54–55, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2004:SV

[Fos04b] Kenneth R. Foster. Software: Visualize this! IEEE Spectrum, 41(8):52–54, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2005:SHM

[Fos05a] Kenneth R. Foster. Software:

Higher math. IEEE Spectrum, 42(6):74, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2005:SLB

[Fos05b] Kenneth R. Foster. Spectral lines: Building better lie detectors with neuroscience? IEEE Spectrum, 42(7):8, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2006:SMG

[Fos06] Kenneth R. Foster. Software:

Maple goes graphical. IEEE Spectrum, 43(2):64, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2007:RBP

[Fos07a] Kenneth R. Foster. Resources: Blast from the past. IEEE Spectrum, 44(10):62–64, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2007:SXJ

[Fos07b] Kenneth R. Foster. Semcad x jungfrau marks the spot (the hot spot). IEEE Spectrum, 44 (6):74–75, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2008:BDD

[Fos08a] Kenneth R. Foster. Books: In defense of dumb [review of The Design of Future Things (D. A. Norman; 2007)]. IEEE Spectrum, 45(5):21, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2008:PWR

[Fos08b] Kenneth R. Foster. Preparing for the worst (review of The Next Catastrophe: Reducing Our Vulnerabilities to Natural, Industrial, and Terrorist Disasters by Perrow, C.; 2007) [books]. IEEE Spectrum, 45 (1):28–29, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2009:MP

[Fos09a] Kenneth R. Foster. The mobile polynomial. IEEE Spectrum, 46(1):26, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2009:NM

[Fos09b] Kenneth R. Foster. New math.

IEEE Spectrum, 46(4):21–22, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2009:SMM

[Fos09c] Kenneth R. Foster. A self-made machine — [hands on]. IEEE Spectrum, 46(1):24–26, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fox:2001:DTR

[Fox01] B. Fox. Digital TV rollout [US digital terrestrial TV]. IEEE Spectrum, 38(2):65–67, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forden:2000:FAN

[FPP00] G. Forden, P. Podvig, and T. A.

Postol. False alarm, nuclear danger. IEEE Spectrum, 37 (3):31–39, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2002:BA

[Fra02a] S. J. Frank. Balancing act.

IEEE Spectrum, 39(6):65–66, June 2002. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2002:CYP

[Fra02b] S. J. Frank. Can you patent an industry standard? IEEE Spectrum, 39(3):14–15, March 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2002:FSP

[Fra02c] S. J. Frank. Festo, schmesto! [patent law and the doctrine of equivalents]. IEEE Spectrum, 39(7):38–39, July 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2002:PA

[Fra02d] S. J. Frank. Patent absurdity. IEEE Spectrum, 39(8): 48–49, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2003:PO

[Fra03] S. J. Frank. Patent do’s and d’oh! IEEE Spectrum, 40 (3):64–67, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2004:DMM

[Fra04] S. J. Frank. The deal maker’s minefield. IEEE Spectrum, 41 (4):58–59, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2005:PRC

[Fra05a] S. J. Frank. Patent reform cacophony. IEEE Spectrum, 42 (12):58–59, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2005:SRO

[Fra05b] S. J. Frank. Source out, risk in [offshore software development]. IEEE Spectrum, 42 (4):60–62, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2009:DBM

[Fra09] S. J. Frank. The death of business-method patents. IEEE Spectrum, 46(3):32–35, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2004:PLA

[FS04] S. J. Frank and W. A. Slater.

Patents lite — Aussies pioneer a cheap and easy way to protect inventions. IEEE Spectrum, 41 (11):94–95, November 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:2003:BB

[FWC03] Kenneth R. Foster, P. Root Wolpe, and A. L. Caplan. Bioethics the brain. IEEE Spectrum, 40(6):34–39, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:2003:YPI

[FZ03] S. J. Frank and Y. P. Zhang. Year of the patent [intellectual property protection, China]. IEEE Spectrum, 40(1):86–88, January 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gail:2007:CC

[Gai07] W. B. Gail. Climate control. IEEE Spectrum, 44(5):20–25, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gardella:2003:HGS

[Gar03] G. H. Gardella. How not to get squeezed [patent battle]. IEEE Spectrum, 40(6):46, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geier:2003:YEO

[Gei03a] M. J. Geier. For your ears only [copy controls for cds]. IEEE Spectrum, 40(5):25–26, May 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geier:2003:LCC

[Gei03b] M. J. Geier. Lights, camera, controls! [dvd copying prevention]. IEEE Spectrum, 40(5):28–31, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geier:2004:TTM

[Gei04a] M. J. Geier. Tools toys: Movies on the move. IEEE Spectrum, 41(6):57–59, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geier:2004:TTP

[Gei04b] M. J. Geier. Tools toys: Pardon me, got the time? IEEE Spectrum, 41(10):62–63, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geier:2005:TTA

[Gei05] M. J. Geier. Tools toys: It’s all on the wrist. IEEE Spectrum, 42(8):57–58, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:AGP

[Gep00a] L. Geppert. Andy Grove [profile].

IEEE Spectrum, 37(6):35–38, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:DCT

[Gep00b] L. Geppert. Devices and circuits [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1):63–69, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:EDA

[Gep00c] L. Geppert. Electronic design automation [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1):70–74, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:Ia

[Gep00d] L. Geppert. Innovations. IEEE Spectrum, 37(2):62–64, February 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:Ib

[Gep00e] L. Geppert. Innovations. IEEE Spectrum, 37(4):24, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:Ic

[Gep00f] L. Geppert. Innovations. IEEE Spectrum, 37(8):72, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:MBG

[Gep00g] L. Geppert. Microprocessors: the off-beat generation. IEEE Spectrum, 37(7):44–49, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:QTT

[Gep00h] L. Geppert. Quantum transistors: toward nanoelectronics. IEEE Spectrum, 37(9): 46–51, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2001:BNV

[Gep01a] L. Geppert. The book of nothing: vacuums, voids, and the latest ideas about the origins of the universe, John D. Barrow, Pantheon Books, New York, 7.50 [book review]. IEEE Spectrum, 38(11): 66–67, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2001:WQR

[Gep01b] L. Geppert. The wizard of quarks, by Robert Gilmore, Copernicus Books, New York, $24 [book review]. IEEE Spectrum, 38(11):68, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2001:NCB

[Gep01c] L. Geppert. The new chips on the block [network processors]. IEEE Spectrum, 38(1): 66–68, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2001:RBN

[Gep01d] L. Geppert. RF bridges to the network [RF IC design]. IEEE Spectrum, 38(1):69–71, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2001:WMP

[Gep01e] L. Geppert. World’s most powerful supercomputer hits full stride. IEEE Spectrum, 38(11): 29–31, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2002:AVT

[Gep02a] L. Geppert. The amazing vanishing transistor act. IEEE Spectrum, 39(10):28–33, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2002:RAD

[Gep02b] N. Geppert. Random access at double the speed. IEEE Spectrum, 39(8):20–21, August 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2003:BX

[Gep03a] L. Geppert. Battle of the xs. IEEE Spectrum, 40(8):16–17, August 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2003:BBT

[Gep03b] L. Geppert. A better bipolar transistor for wireless ics. IEEE Spectrum, 40(12):20–21, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2003:IBC

[Gep03c] L. Geppert. Intel bets on Centrino to spawn new laptop generation. IEEE Spectrum, 40 (5):13–15, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2003:NIM

[Gep03d] L. Geppert. The new indelible memories. IEEE Spectrum, 40(3):48–54, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2003:SCS

[Gep03e] L. Geppert. A sea change for semiconductors. IEEE Spectrum, 40(1):76–82, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2004:GGN

[Gep04a] L. Geppert. The great gallium nitride gamble. IEEE Spectrum, 41(1):52–55, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2004:KSE

[Gep04b] L. Geppert. Karl Stakkope: engineer in paradise. IEEE Spectrum, 41(2):33–34, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2004:LRC

[Gep04c] L. Geppert. Lost radio contact leaves pilots on their own. IEEE Spectrum, 41(11):16–17, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2004:PLG

[Gep04d] L. Geppert. A promising lithography gets stuck. IEEE Spectrum, 41(1):56–57, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2004:QLP

[Gep04e] L. Geppert. A quantum leap for photonics. IEEE Spectrum, 41(7):16–17, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2004:SRS

[Gep04f] L. Geppert. A static RAM says goodbye to data errors [radiation induced soft errors]. IEEE Spectrum, 41(2):16–17, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2004:YKD

[Gep04g] L. Geppert. Yoshihiro kuroki: dancing with robots. IEEE Spectrum, 41(2):34–35, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2005:MQA

[Gep05a] L. Geppert. Move over, quartz: the atomic clock gets smaller and cheaper. IEEE Spectrum, 42(5):20–21, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2005:OOR

[Gep05b] L. Geppert. Ossi oikarinen: At the races [formular one race cars]. IEEE Spectrum, 42(2):33, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2005:PMN

[Gep05c] L. Geppert. Power to the molecules [nanometer-size switch]. IEEE Spectrum, 42(4):18–19, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2005:SGR

[Gep05d] L. Geppert. Silicon gold rush:

Taiwan’s chip makers can’t resist the lure of China, the world’s largest consumer of semiconductors. IEEE Spectrum, 42(6):62–66, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2005:SBS

[Gep05e] L. Geppert. Sun’s big splash [Niagara microprocessor chip]. IEEE Spectrum, 42(1):56–60, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geselowitz:2004:BRB

[Ges04] M. N. Geselowitz. Book review: Black inventors in the age of segregation. IEEE Spectrum, 41(6):61–62, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guerin:2003:TTT

[GFB03] R. Guerin, F. Ferrante, and J. A.

Bellisio. Three takes on telecom’s troubles. IEEE Spectrum, 40(1):35–36, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garrick:2002:YMP

[GG02] B. J. Garrick and V. Gilinsky.

Yucca Mountain pro and con [nuclear waste storage]. IEEE Spectrum, 39(10):41–44, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gavrilov:2004:WWF

[GG04] L. Gavrilov and N. Gavrilova. Why we fall apart: engineering’s reliability theory explains human aging. IEEE Spectrum, 41(9): 31–35, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2005:RBB

[GG05] Erico Guizzo and H. Goldstein. The rise of the body bots [robotic exoskeletons]. IEEE Spectrum, 42(10):50–56, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gagliardi:2006:OWN

[GG06a] F. Gagliardi and F. Grey. Old world, new grid [grid computing]. IEEE Spectrum, 43(7):28– 33, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:EYS

[GG06b] Erico Guizzo and H. Goldstein. Expressway to your skull. IEEE Spectrum, 43(8): 34–39, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gover:2003:UNW

[GH03] J. E. Gover and P. G. Huray. Not so unthinkable [nuclear weapons]. IEEE Spectrum, 40(3):15–16, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gunderman:2007:FNP

[GH07] R. D. Gunderman and J. M.

Hammond. File now, pay later. IEEE Spectrum, 44 (6):77–78, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guterl:2009:MH

[GH09] F. Guterl and M. Heger. Mars is hard. IEEE Spectrum, 46 (6):30–39, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2007:D

[GKJ^{+}07] H. Goldstein, J. Kumagai, W. D.

Jones, G. Zorpette, Sandra Upson, Tekla S. Perry, P. E. Ross, and M. Hood. DreamJobs 2007. IEEE Spectrum, 44(2): 22–36, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldman:2009:MS

[GKK09] D. Goldman, H. Komsuoglu, and D. Koditschek. March of the sandbots. IEEE Spectrum, 46 (4):30–35, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gliekos:2000:EWB

[Gli00] G. Gliekos. Electromagnetic waves [books]. IEEE Spectrum, 37(4):12–13, April 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2001:APP

[Gol01a] H. Goldstein. Appraising the performance of performance appraisals. IEEE Spectrum, 38(11): 61–63, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2001:ECL

[Gol01b] H. Goldstein. Emergence: the connected lives of ants, brains, cities, and software, Steven Johnson, Scriber, New York, $25 [book review]. IEEE Spectrum, 38(11):66, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2001:VCP

[Gol01c] H. Goldstein. Virtuality check: power relations and alternative strategies in the information society, François Fortier, Verso, London, $23 [book review]. IEEE Spectrum, 38(11):66, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2001:ETH

[Gol01d] H. Goldstein. Engines of tomorrow: how the world’s best companies are using their research labs to win the future [book review]. IEEE Spectrum, 38(11):65, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2001:EEB

[Gol01e] H. Goldstein. Exposing electronics [book review]. IEEE Spectrum, 38(11):65, November 2001.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2001:PGV

[Gol01f] H. Goldstein. Packages go vertical. IEEE Spectrum, 38(8):46–51, August 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:CPP

[Gol02a] H. Goldstein. Checking the play in plug-and-play. IEEE Spectrum, 39(6):50–55, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:AYH

[Gol02b] H. Goldstein. It’s all in your head [industrial property]. IEEE Spectrum, 39(3):65–68, March 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:RCH

[Gol02c] H. Goldstein. Radio contact in high-rises can quit on firefighters. IEEE Spectrum, 39(4):24–27, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:STS

[Gol02d] H. Goldstein. Scots tee up for semiconductor drive. IEEE Spectrum, 39(11):17, November 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:SAC

[Gol02e] H. Goldstein. The secret art of chip graffiti. IEEE Spectrum, 39(3):50–55, March 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:SE

[Gol02f] H. Goldstein. The sun is not enough. IEEE Spectrum, 39 (4):62–63, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:TMG

[Gol02g] H. Goldstein. They might be giants. IEEE Spectrum, 39(9): 43–48, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2002:WPA

[Gol02h] H. Goldstein. Waste not, pollute not [animal waste recycling for power production]. IEEE Spectrum, 39(1):72–77, January 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2003:CNG

[Gol03a] H. Goldstein. Collaboration nation [group software]. IEEE Spectrum, 40(6):49–51, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2003:HH

[Gol03b] H. Goldstein. Hardware hangover. IEEE Spectrum, 40(1):

40–43, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2003:INF

[Gol03c] H. Goldstein. Irradiation nation [food safety]. IEEE Spectrum, 40(8):24–29, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2003:IT

[Gol03d] H. Goldstein. The irresistible transistor. IEEE Spectrum, 40 (3):42–47, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2003:SS

[Gol03e] H. Goldstein. SiLK slips. IEEE Spectrum, 40(12):14–15, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2004:DNS

[Gol04a] H. Goldstein. A dog named spot [smart personal objects technology watch]. IEEE Spectrum, 41 (1):72–73, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2004:IA

[Gol04b]  H. Goldstein.  The infinite

archive. IEEE Spectrum, 41 (1):74, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2004:MVW

[Gol04c] H. Goldstein. Mike Villas’s world. IEEE Spectrum, 41(7):45– 48, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2004:SAT

[Gol04d]  H. Goldstein.  Surf Africa

[telecommunications and Internet infrastructure]. IEEE Spectrum, 41(2):48–54, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2004:TTT

[Gol04e] H. Goldstein. Tools toys: Too close for comfort. IEEE Spectrum, 41(7):56–57, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2004:WLW

[Gol04f]  H. Goldstein.  We like to

watch. IEEE Spectrum, 41 (7):30–34, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldwasser:2004:F

[Gol04g] S. M. Goldwasser. Forum. IEEE Spectrum, 41(1):11–12, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2005:HL

[Gol05a] H. Goldstein. Down and out in Ham Lake. IEEE Spectrum, 42(5):55–57, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2005:FNS

[Gol05b] H. Goldstein. A foggy notion [solar photovoltaic power system]. IEEE Spectrum, 42 (1):30–31, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2005:ITT

[Gol05c] H. Goldstein. Intel tries, tries again. IEEE Spectrum, 42(1): 14–16, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2005:RBC

[Gol05d] H. Goldstein. The race to the bottom [consumer nanodevice]. IEEE Spectrum, 42 (3):32–39, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2005:WKV

[Gol05e] H. Goldstein. Who killed the virtual case file? [case management software]. IEEE Spectrum, 42 (9):24–35, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2006:GC

[Gol06a] H. Goldstein. GaAsing up cellphones. IEEE Spectrum, 43 (5):15–16, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2006:ITE

[Gol06b] H. Goldstein. Information technology for European advancement. IEEE Spectrum, 43(1):

52–53, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2006:MT

[Gol06c] H. Goldstein. Modeling terrorists. IEEE Spectrum, 43(9): 26–34, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2006:PP

[Gol06d] H. Goldstein. Patent power. IEEE Spectrum, 43(11):44–47, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2006:TLT

[Gol06e] H. Goldstein. Too little, too soon [solid-state flash memories]. IEEE Spectrum, 43(1): 30–31, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2006:TTM

[Gol06f] H. Goldstein. A touch too much [intrabody communications]. IEEE Spectrum, 43(1): 24–25, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2007:CMB

[Gol07a] H. Goldstein. Cure for the multicore blues. IEEE Spectrum, 44 (1):40–43, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2007:HMM

[Gol07b] H. Goldstein. How not to make a megacity. IEEE Spectrum, 44(6):58, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2007:HBP

[Gol07c] H. Goldstein. How to blackoutproof a city. IEEE Spectrum, 44(6):38–41, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2007:RW

[Gol07d] H. Goldstein. Not ready to wear. IEEE Spectrum, 44(1): 38–39, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2009:CD

[Gol09a]  H. Goldstein.  California

dreamin’. IEEE Spectrum, 46 (3):7, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2009:WCB

[Gol09b] H. Goldstein. We come to bury ie6. IEEE Spectrum, 46 (10):10, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:TMS

[GP00] L. Geppert and Tekla S. Perry.

Transmeta’s magic show [microprocessor chips]. IEEE Spectrum, 37(5):26–33, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gall:2006:HSD

[GP06a] R. Gall and D. Parsons. It’s hurricane season: Do you know where your storm is? IEEE Spectrum, 43(8):27–32, August 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gorbis:2006:IFS

[GP06b] M. Gorbis and D. Pescovitz. IEEE Fellows Survey: Bursting tech bubbles before they balloon. IEEE Spectrum, 43(9): 50–55, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2009:ISO

[GR09] H. Goldstein and J. J. Romero. IEEE Spectrum online: Content is liberated! IEEE Spectrum, 46(7):10, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gray:2000:AEC

[Gra00] K. Gray. Adding error-correcting circuitry to ASIC memory. IEEE Spectrum, 37(4):55–60, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Graafstra:2007:H

[Gra07] A. Graafstra. Hands on. IEEE Spectrum, 44(3):18–23, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Green:2002:PLM

[Gre02] P. E. Green. Paving the last mile with glass. IEEE Spectrum, 39 (12):13–14, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gwynne:2001:NA

[GRG^{+}01] P. Gwynne, D. Rosenbaum, H. Goldstein, S. Cherry, S. Siang, Samuel K. Moore, and A. Hellemans. News analysis. IEEE Spectrum, 38(10):20–34, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:T

[GS00] L. Geppert and W. Sweet. Technology 2000. IEEE Spectrum, 37 (1):26–31, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldstein:2007:JBQ

[GS07] H. Goldstein and W. Sweet.

Joules, BTUs, Quads — let’s call the whole thing off. IEEE Spectrum, 44(1):16, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:EHE

[GS08] Erico Guizzo and W. Sweet.

Engineering the Harvard engineer. IEEE Spectrum, 45 (4):46–52, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2004:CPC

[Gui04a]  Erico Guizzo.  Closing in

on the perfect code [turbo codes]. IEEE Spectrum, 41 (3):36–42, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

URL https://spectrum.ieee.
org/computing/software/closing-
in-on-the-perfect-code.

Guizzo:2004:CST

[Gui04b] Erico Guizzo. Coming out soon: trillion-color TV. IEEE Spectrum, 41(8):17–18, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2004:FFB

[Gui04c] Erico Guizzo. Flying away — forecast bad for tiny airborne probes. IEEE Spectrum, 41 (1):40–41, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2004:HFS

[Gui04d] Erico Guizzo. Hypersonic flight [scramjet aircraft propulsion]. IEEE Spectrum, 41(1):66, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2004:OMG

[Gui04e] Erico Guizzo. Organic memory gains momentum. IEEE Spectrum, 41(4):17–18, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2005:AFT

[Gui05a]  Erico Guizzo.  The atomic

fortress that time forgot [plutonium production plant]. IEEE Spectrum, 42(4):42–49, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2005:BBT

[Gui05b] Erico Guizzo. Bugged balls for tough calls [ball tracking systems]. IEEE Spectrum, 42 (6):14–15, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2005:CGA

[Gui05c] Erico Guizzo. Claude Gagnon:

the all-terrain engineer. IEEE Spectrum, 42(2):28–29, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2005:IRS

[Gui05d] Erico Guizzo. IBM reclaims supercomputer lead. IEEE Spectrum, 42(2):15–16, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2005:DII

[Gui05e] Erico Guizzo. Into deep ice [ice monitoring]. IEEE Spectrum, 42 (12):28–35, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2005:ZSC

[Gui05f] Erico Guizzo. Zone of silence [cellphone jammer]. IEEE Spectrum, 42(5):17–18, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:BIC

[Gui06a] Erico Guizzo. Britain’s identity crisis [biometric ID cards]. IEEE Spectrum, 43(1):42–43, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:BFR

[Gui06b] Erico Guizzo. Bubble fusion research under scrutiny. IEEE Spectrum, 43(5):16–21, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:CTC

[Gui06c] Erico Guizzo. Carbon takeoff [carbon fiber composites]. IEEE Spectrum, 43(1):26–29, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:DS

[Gui06d] Erico Guizzo. A desktop supercomputer. IEEE Spectrum, 43(1):54, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:HBS

[Gui06e] Erico Guizzo. How Brazil spun the atom [nuclear power reactors]. IEEE Spectrum, 43 (3):50–56, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:OEE

[Gui06f] Erico Guizzo. The olin experiment [engineering education experiment]. IEEE Spectrum, 43(5):30–36, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2007:MRD

[Gui07a] Erico Guizzo. $280 million robot dustup [news]. IEEE Spectrum, 44(12):10–12, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2007:ACT

[Gui07b] Erico Guizzo. Anatomy of a crash-test dummy. IEEE Spectrum, 44(10):42–49, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2007:HKM

[Gui07c] Erico Guizzo. How to keep 18 million people moving. IEEE Spectrum, 44(6):32–35, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2007:OE

[Gui07d] Erico Guizzo. The omnivorous engine. IEEE Spectrum, 44 (1):34–37, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2007:TT

[Gui07e]  Erico Guizzo.  The trading

test. IEEE Spectrum, 44(8): 51–54, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:DCS

[Gui08a] Erico Guizzo. Defense contractors snap up submersible robot gliders [update]. IEEE Spectrum, 45(9):11–12, September 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:EGG

[Gui08b] Erico Guizzo. The EE gender gap is widening. IEEE Spectrum, 45(12):23, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:GFG

[Gui08c] Erico Guizzo. The game-frame guild. IEEE Spectrum, 45 (8):44–52, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:RM

[Gui08d] Erico Guizzo. The rise of the machines. IEEE Spectrum, 45 (12):88, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:SC

[Gui08e] Erico Guizzo. The steampunk contraptors. IEEE Spectrum, 45 (10):48–55, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:TEH

[Gui08f] Erico Guizzo. Three engineers, hundreds of robots, one warehouse. IEEE Spectrum, 45(7):26– 34, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:WPP

[Gui08g] Erico Guizzo. Wind power in paradise. IEEE Spectrum, 45 (3):38–45, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:WGS

[Gui08h] Erico Guizzo. Winner: Geophysics solving the oil equation. IEEE Spectrum, 45(1): 32–36, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2009:BBB

[Gui09a]  Erico Guizzo.  Brew, baby,

brew. IEEE Spectrum, 46(1): 63–64, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2009:ETM

[Gui09b] Erico Guizzo. Erlendur thorsteinsson master of the universe. IEEE Spectrum, 46(2):32–33, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2009:H

[Gui09c] Erico Guizzo. Hot or not? IEEE Spectrum, 46(1):36–38, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2009:MMS

[Gui09d] Erico Guizzo. Marco Migliari studio virtuoso. IEEE Spectrum, 46(2):24–25, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gwynne:2000:DCS

[Gwy00] P. Gwynne. Digital cameras that see it all. IEEE Spectrum, 37 (12):21–22, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hammond:2002:SCO

[Ham02] B. Hammond. Star-crossed orbits: inside the U.S.–Russian space alliance [book review]. IEEE Spectrum, 39(3):57–58, March 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Handler:2000:US

[Han00] J. Handler. Under suspicion. IEEE Spectrum, 37(3):51–53, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hanson:2008:ES

[Han08] R. Hanson. Economics of the singularity. IEEE Spectrum, 45 (6):45–50, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harris:2000:DND

[Har00] L. E. Harris. Database nation; the death of privacy in the 21st century [books]. IEEE Spectrum, 37(11):16–18, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harris:2002:ITG

[Har02] G. Harris. India: telemedicine’s great new frontier. IEEE Spectrum, 39(4):16–17, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2001:NYN

[Has01] S. Hassler. A new year, a new outlook [spectral lines]. IEEE Spectrum, 38(1):9, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2007:EM

[Has07] S. Hassler. Engineering the megacity. IEEE Spectrum, 44 (6):22–23, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2008:AT

[Has08a] S. Hassler. The audacity of technology. IEEE Spectrum, 45 (12):9, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2008:FCD

[Has08b] S. Hassler. The future of code, digital and genetic [spectral lines]. IEEE Spectrum, 45 (9):9, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2008:SLB

[Has08c] S. Hassler. Spectral lines — books in a networked age. IEEE Spectrum, 45(7):9, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2008:SLE

[Has08d] S. Hassler. Spectral lines: It’s not easy being lean. IEEE Spectrum, 45(5):9, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2008:SLA

[Has08e] S. Hassler. Spectral lines: Unassuming the singularity. IEEE Spectrum, 45(6):9, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2008:SLW

[Has08f] S. Hassler. Spectral lines [winged victory: Fly-size wing flapper lifts off]. IEEE Spectrum, 45(3):9, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2009:CE

[Has09a] S. Hassler. The 21st-Century engineer. IEEE Spectrum, 46 (2):7, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2009:ISG

[Has09b] S. Hassler. IEEE Spectrum gets greener — [spectral lines]. IEEE Spectrum, 46(1):9, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2009:EFI

[Has09c] S. Hassler. Engineering the future at IEEE. IEEE Spectrum, 46(5):8, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2009:WMW

[Has09d] S. Hassler. Why Mars? Why now? IEEE Spectrum, 46 (6):26–29, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hassler:2009:WAW

[Has09e] S. Hassler. Working in an alwayson world. IEEE Spectrum, 46 (9):8, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hawkins:2007:WCC

[Haw07] J. Hawkins. Why can’t a computer be more like a brain? IEEE Spectrum, 44(4):21–26, April 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hess:2004:CCF

[HBB04] R. K. Hess, D. I. Bass, and J. B. Baca. Crashproof codes [flight software]. IEEE Spectrum, 41 (9):48–53, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hebner:2002:FBC

[HBW02] R. Hebner, J. Beno, and A. Walls. Flywheel batteries come around again. IEEE Spectrum, 39(4):46–51, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2001:NA

[HCJ^{+}01] A. Hellemans, Stephen Cass, W. D. Jones, J. F. Mason, B. Klein, and R. Comerford. News analysis. IEEE Spectrum, 38(3):20–39, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Homer-Dixon:2003:WLI

[HD03] T. Homer-Dixon. Winners and losers in the information age. IEEE Spectrum, 40(5):11–12, May 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hecht:2009:RGG

[Hec09] J. Hecht. Ray guns get real. IEEE Spectrum, 46(7):28–33, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heger:2008:HSU

[Heg08a] M. Heger. High speed down under [update]. IEEE Spectrum, 45 (11):13–14, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heger:2008:PWO

[Heg08b] M. Heger. The price is wrong for oil shale and tar sand tech. IEEE Spectrum, 45(12):11–12, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heger:2009:CTQ

[Heg09a] M. Heger. Cryptographers take on quantum computers — [update]. IEEE Spectrum, 46 (1):14, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heger:2009:WWM

[Heg09b] M. Heger. What to wear on Mars. IEEE Spectrum, 46(6):34, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heinze:2002:DPW

[Hei02] W. F. Heinze. Dead patents walking. IEEE Spectrum, 39 (5):52–54, May 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heingartner:2009:MB

[Hei09] D. Heingartner. Mental block.

IEEE Spectrum, 46(1):42–43, January 2009. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2000:PMA

[Hel00] A. Hellemans. Polymer matrix augurs flexible displays [innovations]. IEEE Spectrum, 37(12): 18–21, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2001:FBS

[Hel01a]  A. Hellemans.  First baby

steps toward molecular electronics. IEEE Spectrum, 38(11): 18–19, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2001:IRD

[Hel01b] A. Hellemans. IBM researchers devise nanotube ICs. IEEE Spectrum, 38(6):26–27, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2002:IGC

[Hel02a] A. Hellemans. Identical graphs chart a dubious picture. IEEE Spectrum, 39(7):20–21, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2002:LHM

[Hel02b] A. Hellemans. Left-handed material reacts to 3-D light. IEEE Spectrum, 39(10):24–25, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2002:PLC

[Hel02c] A. Hellemans. Paintable liquidcrystal displays. IEEE Spectrum, 39(8):28–29, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2002:SSD

[Hel02d] A. Hellemans. Schön shown the door [electronics research investigation]. IEEE Spectrum, 39 (11):20–21, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2003:GTC

[Hel03] A. Hellemans. German team creates new type of transistorlike device. IEEE Spectrum, 40 (1):20–21, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2004:NPR

[Hel04a] A. Hellemans. Nanoholes permit remarkable light transmission. IEEE Spectrum, 41(7):18, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2004:TWM

[Hel04b]  A. Hellemans.  Teleporting

what matters [quantum computing]. IEEE Spectrum, 41 (9):18, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2005:CBD

[Hel05a] A. Hellemans. Could Belgian diode lead to printable RFIDs?

IEEE Spectrum, 42(12):14, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2005:NPN

[Hel05b] A. Hellemans. A new phase in nonvolatile memory? IEEE Spectrum, 42(6):18, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2005:STV

[Hel05c] A. Hellemans. Sins of transmission? Vatican radio’s high-power antennas stand accused of causing cancer. IEEE Spectrum, 42 (10):12–15, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2005:SBS

[Hel05d] A. Hellemans. Strange bedfellows [semiconductor technology]. IEEE Spectrum, 42(2): 14–15, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heller:2005:PET

[Hel05e] E. J. Heller. Photo essay: The tangled web electrons weave. IEEE Spectrum, 42(3):20–21, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2006:EWF

[Hel06a]  A. Hellemans.  Engineering

warms to frozen light. IEEE Spectrum, 43(8):14–15, August 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2006:SD

[Hel06b] A. Hellemans. Spin doctoring. IEEE Spectrum, 43(11): 20, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2007:MM

[Hel07] A. Hellemans. Manufacturing mayday. IEEE Spectrum, 44 (1):10–13, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Holonyak:2006:TL

[HF06] N. Holonyak and M. Feng. The transistor laser. IEEE Spectrum, 43(2):50–55, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hira:2003:TRS

[Hir03] R. Hira. The top 100 R&D spenders. IEEE Spectrum, 40 (10):33–36, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hira:2004:SRD

[Hir04] R. Hira. Spectrum R-D 100 [R&D spending]. IEEE Spectrum, 41(11):61–65, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hira:2005:RIS

[Hir05]  R. Hira.  R&D 100 [IEEE

Spectrum’s fourth annual survey]. IEEE Spectrum, 42(12):

36–41, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hendriks:2004:TLS

[HK04] B. Hendriks and S. Kuiper. Through a lens sharply [Fluid-Focus lens]. IEEE Spectrum, 41 (12):32–36, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellemans:2004:CC

[HM04] A. Hellemans and M. Mukerjee. Computing the cosmos. IEEE Spectrum, 41(8): 28–34, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harry:2005:BAN

[HNPC05] J. D. Harry, J. B. Niemi, A. A. Priplata, and J. J. Collins. Balancing act [noise based sensory enhancement technology]. IEEE Spectrum, 42(4):36–41, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hodgson:2000:GSI

[Hod00] J. Hodgson. Gene sequencing’s industrial revolution. IEEE Spectrum, 37(11):36–42, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heger:2009:TVP

[HOD^{+}09] M. Heger, J. Oberg, M. Dumiak, Samuel K. Moore, and P. Patel-Predd. Technology vs. pirates. IEEE Spectrum, 46 (2):9–10, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Holmer:2003:FTS

[Hol03] P. Holmer. Faster than a speeding bullet train. IEEE Spectrum, 40(8):30–34, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hollaar:2006:PSP

[Hol06] Lee A. Hollaar. Patents 2.0 [software patents]. IEEE Spectrum, 43(2):61–64, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hood:2000:DCF

[Hoo00] M. Hood. The day Concorde fell to Earth [faults failures]. IEEE Spectrum, 37(9):24–28, September 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hood:2002:CDS

[Hoo02] M. Hood. China dawn: the story of a technology and business revolution [book review]. IEEE Spectrum, 39(8):50, August 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hood:2005:RAW

[Hoo05a] M. Hood. Running against the wind [sports prosthetics]. IEEE Spectrum, 42(6):13–14, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hood:2005:SSD

[Hoo05b] M. Hood. Steal this software and this DVD, and this book, and this handbag, and this car … [intellectual property]. IEEE Spectrum, 42 (6):52–53, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hood:2005:WST

[Hoo05c] M. Hood. Wiring small-town China: the fiber-optic tentacles of China’s Internet extend deep into the heartland. IEEE Spectrum, 42(6):45, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hood:2007:NDD

[Hoo07] M. Hood. E-newspapers: Digital deliverance? IEEE Spectrum, 44(2):10–12, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horgan:2008:CC

[Hor08] J. Horgan. The consciousness conundrum. IEEE Spectrum, 45(6):36–41, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hira:2007:R

[HR07] R. Hira and P. E. Ross. The R&D 100. IEEE Spectrum, 44 (12):38–41, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hsu:2007:CG

[Hsu07] F. Hsu. Cracking go. IEEE Spectrum, 44(10):50–55, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hughes:2002:WDH

[Hug02] G. F. Hughes. Wise drives [hard disk drive]. IEEE Spectrum, 39 (8):37–41, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hughes:2009:NAP

[Hug09] V. Hughes. A new approach to predicting epileptic seizures. IEEE Spectrum, 46(6):18, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hunter:2006:BPS

[Hun06] J. Hunter. The big picture: Serpent mother. IEEE Spectrum, 43 (11):24–25, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hutcheson:2001:EUL

[Hut01] G. D. Hutcheson. Extreme ultraviolet lithography: will it be ready in time? IEEE Spectrum, 38(11):15–16, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Introna:2000:S

[IN00] L. Introna and H. Nissenbaum.

Speakout. IEEE Spectrum, 37 (6):26–27, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ingebretsen:2001:LMR

[Ing01a] M. Ingebretsen. At last, a magnetic refrigerator? IEEE Spectrum, 38(11):21–23, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ingebretsen:2001:MA

[Ing01b] M. Ingebretsen. Mining asteroids. IEEE Spectrum, 38 (8):34–39, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ingebretsen:2001:CST

[IS01] M. Ingebretsen and M. Siegel. Cable slouches toward open access. IEEE Spectrum, 38(4): 74–79, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ingebretsen:2003:EP

[IS03] M. Ingebretsen and W. Sweet. Emission permission. IEEE Spectrum, 40(1):59–62, January 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Isenberg:2003:EMT

[Ise03] D. S. Isenberg. The end of the middle [telephone industry]. IEEE Spectrum, 40(1): 37–38, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ilic:2001:ETC

[ISV01] M. Ilic, P. Skantz, and P. Visudhiphan. Electricity troubles in california: Who’s next? IEEE Spectrum, 38(2):11–13, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:SSC

[JA05] W. D. Jones and L. Aaronson.

Showing some class [educational assistance]. IEEE Spectrum, 42 (11):16, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jackson:2005:IIN

[Jac05] J. Jackson. The interplanetary Internet [networked space communications]. IEEE Spectrum, 42(8):30–35, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jancher:2001:EED

[Jan01] L. Jancher. European electricity deregulation will not level the playing field. IEEE Spectrum, 38(7):15–16, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:USE

[JB05] W. D. Jones and A. Basanskaya.

United States establishes “level playing field” in broadband — or does it? IEEE Spectrum, 42(10):13, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2001:NA

[JCF^{+}01] W. D. Jones, Y. Chiu, M. Furlaneto, G. Weiss, S. Miller, B. Klein, C. Morrison, and J. Mullins. News analysis. IEEE Spectrum, 38(8):20–32, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jaffe:2004:PPR

[JL04] A. B. Jaffe and J. Lerner. Patent prescription: a radical cure for the ailing [us patent policy]. IEEE Spectrum, 41(12): 38–43, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jamieson:2009:MC

[JN09] L. H. Jamieson and J. Norberg. The Mars challenge. IEEE Spectrum, 46(6):73, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:2000:TSS

[Joh00a] L. Johnson. The tether solution [space propulsion, electrodynamic tether]. IEEE Spectrum, 37(7):38–43, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:2000:MML

[Joh00b] M. Johnson. Magnetoelectronic memories last and last. IEEE Spectrum, 37(2):33–40, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:2007:SLM

[Joh07] D. Johnson. Spectral lines: Material by design: Future science or science fiction? IEEE Spectrum, 44(8):10, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:WS

[Jon00a] R. Jones. Web sights. IEEE Spectrum, 37(6):92–93, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:BIT

[Jon00b] W. D. Jones. A blockbuster innovation [thirty years ago]. IEEE Spectrum, 37(9):80R2– 80R4, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:EGH

[Jon00c] W. D. Jones. EE gifts for the holidays. IEEE Spectrum, 37 (12):31–35, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:NYW

[Jon00d]  W. D. Jones.  New York

wants more hybrid buses on the city’s streets [technology watch]. IEEE Spectrum, 37 (9):36, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAa

[Jon00e] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37 (1):120R4–120R7, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAb

[Jon00f] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(2): 64R2–64R9, February 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAc

[Jon00g] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(3): 80R2–80R5, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAd

[Jon00h] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(4):64R2, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAe

[Jon00i] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(5): 64R2–64R4, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAf

[Jon00j] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(6): 80R2–80R4, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAg

[Jon00k] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(7): 68R2–68R4, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAh

[Jon00l] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(8): 76R2–76R4, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAi

[Jon00m] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(10): 80R4–80R8, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAj

[Jon00n] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(11): 32I9–32I10, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:TYAk

[Jon00o] W. D. Jones. Thirty years ago . . .. IEEE Spectrum, 37(12): 64R2–64R4, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2000:WRW

[Jon00p] W. D. Jones. Welcome to the real world. IEEE Spectrum, 37(7):70– 72, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2001:EGH

[Jon01a] W. D. Jones. EEs’ gifts for the holidays [consumer]. IEEE Spectrum, 38(12):48–53, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2001:FTF

[Jon01b] W. D. Jones. Forecasting traffic flow. IEEE Spectrum, 38 (1):90–91, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2001:KCC

[Jon01c] W. D. Jones. Keeping cars from crashing. IEEE Spectrum, 38 (9):40–45, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2001:LMN

[Jon01d] W. D. Jones. Look, ma, no hands. IEEE Spectrum, 38 (6):39–40, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2002:BSC

[Jon02a] W. D. Jones. Building safer cars. IEEE Spectrum, 39(1): 82–85, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2002:GH

[Jon02b] W. D. Jones. Gifts for the holidays. IEEE Spectrum, 39(11): 44–49, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2003:BTL

[Jon03a] W. D. Jones. Blackout a turn-on for long island cable. IEEE Spectrum, 40(10):11, October 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2003:CCP

[Jon03b] W. D. Jones. Cable controversy pits power haves against have-nots. IEEE Spectrum, 40 (5):15–16, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2003:FEL

[Jon03c] W. D. Jones. Fcc embraces localism. IEEE Spectrum, 40 (10):16, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2003:GH

[Jon03d] W. D. Jones. Gifts for the holidays. IEEE Spectrum, 40(11): 42–47, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2003:HRH

[Jon03e]  W. D. Jones.  Hybrids to

the rescue [hybrid electric vehicles]. IEEE Spectrum, 40 (1):70–71, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2003:RPG

[Jon03f] W. D. Jones. Radio pirates grapple with U.S. regulators. IEEE Spectrum, 40(7):19, July 2003.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2003:SS

[Jon03g]  W. D. Jones.  Share and

share not. IEEE Spectrum, 40 (4):19–21, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2004:ARG

[Jon04a] W. D. Jones. Amped up and ready to go [broadband over power lines]. IEEE Spectrum, 41 (9):14–16, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2004:BBG

[Jon04b] W. D. Jones. Black boxes get green light [car crash data recorders]. IEEE Spectrum, 41 (12):14–16, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2004:BRT

[Jon04c] W. D. Jones. Boeing runs into turbulence. IEEE Spectrum, 41(7):10–13, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2004:ECT

[Jon04d] W. D. Jones. Eyeglasses on the cheap [tabletop lens molding]. IEEE Spectrum, 41(11): 22, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2004:SSE

[Jon04e] W. D. Jones. Sight for sore ears [visual aids for the blind]. IEEE Spectrum, 41(2):13–14, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2004:TTG

[Jon04f] W. D. Jones. Tools toys: Gifts for the holidays. IEEE Spectrum, 41(11):86–94, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2004:UGU

[Jon04g] W. D. Jones. Ultrawide gap on ultrawideband. IEEE Spectrum, 41(1):30, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:AVB

[Jon05a] W. D. Jones. Airbus vs. Boeing: some battle stats. IEEE Spectrum, 42(8):13, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:BTB

[Jon05b] W. D. Jones. A Bronx tale:

broadband TV concern hooks up the ’hood. IEEE Spectrum, 42 (9):10–12, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:FBB

[Jon05c] W. D. Jones. Fiber to the brain [brain surgery]. IEEE Spectrum, 42(10):20, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:HDR

[Jon05d] W. D. Jones. Hard drive — robotic cars impress in rough road race. IEEE Spectrum, 42 (12):13, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:NPH

[Jon05e] W. D. Jones. No place to hide [radar devices]. IEEE Spectrum, 42(11):20–21, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:TCP

[Jon05f] W. D. Jones. Take this car and plug it [plug-in hybrid vehicles]. IEEE Spectrum, 42(7):10– 13, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:TTS

[Jon05g]  W. D. Jones.  Tools toys:

Stargazing made easy. IEEE Spectrum, 42(1):68, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2005:WWW

[Jon05h] W. D. Jones. Wringing watts from waves: a high schooler’s invention turns water motion into “juice”. IEEE Spectrum, 42(4):14, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:BT

[Jon06a] W. D. Jones. Blood test. IEEE Spectrum, 43(11):16–18, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:HT

[Jon06b] W. D. Jones. Hydrogen on track. IEEE Spectrum, 43 (8):10–13, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:IGP

[Jon06c] W. D. Jones. I’ve got the power.

IEEE Spectrum, 43(10):18, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:MGV

[Jon06d] W. D. Jones. Microsoft and Google vie for virtual world domination. IEEE Spectrum, 43(7):16–18, July 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:MHL

[Jon06e]  W. D. Jones.  More heat,

less sag [power cable upgrades]. IEEE Spectrum, 43(6):16–18, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:NSA

[Jon06f] W. D. Jones. No strings attached [wireless usb]. IEEE Spectrum, 43(4):16–18, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:P

[Jon06g] W. D. Jones. Parking 2.0. IEEE Spectrum, 43(12):20, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:PTA

[Jon06h] W. D. Jones. Plane talk about cellphones [spectral lines]. IEEE Spectrum, 43(3):13, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2006:TBT

[Jon06i] W. D. Jones. Taking body temperature, inside out [body temperature monitoring]. IEEE Spectrum, 43(1):13–15, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:HSH

[Jon07a] W. D. Jones. Helmets sense the hard knocks [news]. IEEE Spectrum, 44(10):10–12, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:HBM

[Jon07b] W. D. Jones. How to build a mile-high skyscraper. IEEE Spectrum, 44(6):52–53, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:LPP

[Jon07c] W. D. Jones. Low power to the people. IEEE Spectrum, 44 (8):16–18, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:MMM

[Jon07d] W. D. Jones. Made-to-measure mass transit. IEEE Spectrum, 44(6):14–16, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:OQC

[Jon07e] W. D. Jones. Ontario quits coal.

IEEE Spectrum, 44(3):12–13, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:OMT

[Jon07f] W. D. Jones. Oslo Metro takes greener track [news]. IEEE Spectrum, 44(11):16, November 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:PEW

[Jon07g] W. D. Jones. Putting electricity where the rubber meets the road [news]. IEEE Spectrum, 44(7):18–20, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:TYT

[Jon07h]  W. D. Jones.  Termites in

your tank. IEEE Spectrum, 44(4):16–17, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2007:TEL

[Jon07i] W. D. Jones. Tulane engineering is latest katrina victim. IEEE Spectrum, 44(5):11, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2008:RNR

[Jon08a] R. A. L. Jones. Rupturing the nanotech rapture. IEEE Spectrum, 45(6):64–67, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2008:CCS

[Jon08b] W. D. Jones. Colon cancer screening, the easy way — [update]. IEEE Spectrum, 45(10): 14–15, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2008:EGN

[Jon08c] W. D. Jones. Engineers graduate from New Afghan Military Academy. IEEE Spectrum, 45(1):20, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2008:SSF

[Jon08d] W. D. Jones. Silicon siivers for flexible circuits. IEEE Spectrum, 45(3):15, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2008:UOP

[Jon08e] W. D. Jones. Update — ocean power catches a wave. IEEE Spectrum, 45(7):14, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2008:VRT

[Jon08f] W. D. Jones. Virtual-reality test reveals hidden concussion damage [update]. IEEE Spectrum, 45(9):14, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2008:WMP

[Jon08g] W. D. Jones. World’s most powerful magnet under construction [update]. IEEE Spectrum, 45 (9):13, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:AAC

[Jon09a] W. D. Jones. Automated auditors to chase down cheats. IEEE Spectrum, 46(4):13, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:FCC

[Jon09b] W. D. Jones. Fuel cells could power a streetcar revival. IEEE Spectrum, 46(9):15–16, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:FFS

[Jon09c] W. D. Jones. Fusion factory starts up. IEEE Spectrum, 46 (3):11–12, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:MM

[Jon09d] W. D. Jones. Musical microfluidics. IEEE Spectrum, 46 (9):17, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:PSF

[Jon09e] W. D. Jones. Plastic surgery 1, face recognition 0. IEEE Spectrum, 46(9):17, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:SWB

[Jon09f] W. D. Jones. Sound waves for brain waves — [update]. IEEE Spectrum, 46(1):16–17, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:TLD

[Jon09g] W. D. Jones. ‘Til lawsuits do us part’. IEEE Spectrum, 46(4):23, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:2009:TSF

[Jon09h] W. D. Jones. Touch screens with feeling. IEEE Spectrum, 46(5):15, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jain:2006:TMB

[JP06] A. K. Jain and S. Pankanti. A touch of money [biometric authentication systems]. IEEE Spectrum, 43(7):22–27, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Javor:2008:F

[JS08] K. Javor and J. Spencer. Forum. IEEE Spectrum, 45(3):10, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

K:2008:MSN

[K.08]  J. K.  Mark Schubin —

nights at the opera [dream jobs 2008]. IEEE Spectrum, 45(2): 43–44, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2000:WW

[KA00] J. Kumagai and J. Adam. Washington watch. IEEE Spectrum, 37 (8):20I2–20I4, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:WWB

[KA01] J. Kumagai and A. Applewhite.

Who wants to be a billionaire? [careers]. IEEE Spectrum, 38(7):74–77, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaner:2002:UDP

[Kan02] C. Kaner. UCITA: a disaster in progress. IEEE Spectrum, 39 (8):13–15, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:2000:IET

[Kap00] Gadi Kaplan. Industrial electronics [technology 2000 analysis and forecast]. IEEE Spectrum, 37 (1):104–109, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:2001:EWW

[Kap01a] Gadi Kaplan. Ethernet’s winning ways. IEEE Spectrum, 38 (1):113–115, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:2001:SN

[Kap01b] Gadi Kaplan. Simulating networks. IEEE Spectrum, 38 (1):74–76, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kapustka:2002:WUE

[Kap02] Paul Kapustka. Will the U.S. economy get a second wind any time soon? IEEE Spectrum, 39(7):17–19, July 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2001:ERU

[Kar01a] S. Kariya. Effective resumes for uncertain times [careers]. IEEE Spectrum, 38(10):62–64, October 2001. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2001:HTH

[Kar01b] S. Kariya. The how-tos of hiring. IEEE Spectrum, 38(12): 55–56, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2001:TPD

[Kar01c] S. Karlin. Take off, plug in, dial up [aircraft personal communication]. IEEE Spectrum, 38 (8):52–59, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:DH

[Kar02a] S. Kariya. Dealing with headhunters. IEEE Spectrum, 39 (6):69–71, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:FBC

[Kar02b] S. Kariya. Faster, better — costlier? [us patent trademark office’s strategic plan]. IEEE Spectrum, 39(12):52–54, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:GM

[Kar02c] S. Kariya. Getting the message. IEEE Spectrum, 39(8): 56–57, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:MDI

[Kar02d] S. Kariya. A million dollar ideaand your next job [intellectual property engineers]. IEEE Spectrum, 39(4):67–69, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:WSW

[Kar02e] S. Kariya. “So why should we hire you?” [interview techniques]. IEEE Spectrum, 39 (2):68–71, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:WDL

[Kar02f] S. Kariya. “What I did last summer” [internship]. IEEE Spectrum, 39(7):48–49, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:WJ

[Kar02g] S. Kariya. Where the jobs are. IEEE Spectrum, 39(1): 89–91, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2002:EE

[Kar02h] S. Karlin. Engineering entertainment. IEEE Spectrum, 39 (3):69–70, March 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2002:SHL

[Kar02i] S. Karlin. From squirts to Hertz [Lonnie Johnson, inventor]. IEEE Spectrum, 39(7):46– 48, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2002:BSB

[Kar02j] S. Karlin. Is business school best for you? [electrical engineering careers]. IEEE Spectrum, 39 (2):64–67, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2003:MDA

[Kar03a] S. Kariya. Missile defense for airliners finds growing support in United States. IEEE Spectrum, 40(8):14–15, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2003:OEE

[Kar03b] S. Kariya. Online education expands and evolves. IEEE Spectrum, 40(5):49–51, May 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2003:PSC

[Kar03c] S. Kariya. Patriot’s second chance at glory. IEEE Spectrum, 40(5):17–18, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2003:WJ

[Kar03d] S. Kariya. Where the jobs are. IEEE Spectrum, 40(1): 91–94, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2003:ALS

[Kar03e] S. Karlin. Aural laureate [Scott Jones biography]. IEEE Spectrum, 40(3):68–69, March 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2003:CBM

[Kar03f] S. Karlin. Careers: Boosting morale in an iffy economy. IEEE Spectrum, 40(11):52, November 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2003:ID

[Kar03g] S. Karlin. Invention on demand. IEEE Spectrum, 40(9): 46–48, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2004:CDV

[Kar04a] S. Karlin. Careers — from Dilbert to Da Vinci. IEEE Spectrum, 41(11):96–98, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2004:YFM

[Kar04b] S. Karlin. Not your father’s MBA [technology management degree programs]. IEEE Spectrum, 41 (9):55–60, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2004:YIW

[Kar04c] S. Karlin. Young inventors of the world unite [innovators international network]. IEEE Spectrum, 41(3):53–55, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2005:ECS

[Kar05a] S. Karlin. Engineering a career switch. IEEE Spectrum, 42(5):58–61, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2005:HCG

[Kar05b] S. Karlin. How to choose a grad school [graduate studies]. IEEE Spectrum, 42(9):59–60, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2005:NKH

[Kar05c] S. Karlin. Networking knowhow [online networking]. IEEE Spectrum, 42(11):68–70, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2006:CCU

[Kar06a] S. Karlin. Careers: Certification uncertainty. IEEE Spectrum, 43 (11):58–61, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2006:CSBa

[Kar06b] S. Karlin. Careers: Shaman, bless this lab. IEEE Spectrum, 43(10):49–54, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2006:CSBb

[Kar06c] S. Karlin. Careers: Shaman, bless this lab. IEEE Spectrum, 43(10):49–51, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2007:CRD

[Kar07a] S. Karlin. Careers: A recipe for development. IEEE Spectrum, 44(5):49–54, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2007:CMV

[Kar07b] S. Karlin. Careers: Mom vs. bomb. IEEE Spectrum, 44 (3):45–47, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2007:EW

[Kar07c] S. Karlin. Engineers at war. IEEE Spectrum, 44(9):59–63, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:BSM

[Kar08a] S. Karlin. Bob Saget: Mac Nut [mini-profile]. IEEE Spectrum, 45(2):24, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:CDB

[Kar08b] S. Karlin. Careers: Denee busby: technical actress. IEEE Spectrum, 45(6):25, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:COW

[Kar08c] S. Karlin. Careers [oscar-winning software]. IEEE Spectrum, 45 (3):21–22, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:CQU

[Kar08d] S. Karlin. Careers: Qatar University opens EE doors to women. IEEE Spectrum, 45 (8):22, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:CYW

[Kar08e] S. Karlin. Careers: So you want to be an expert witness. IEEE Spectrum, 45(4):22–23, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:ETJ

[Kar08f] S. Karlin. Engineer turned jet jockey [careers]. IEEE Spectrum, 45(1):26–27, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:EEC

[Kar08g] S. Karlin. Extremist engineers [careers]. IEEE Spectrum, 45 (9):22–23, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:GGC

[Kar08h] S. Karlin. Generation GPS — [careers]. IEEE Spectrum, 45 (10):24, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:HTH

[Kar08i] S. Karlin. High-tech haunting — [careers]. IEEE Spectrum, 45(10):24, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:JBA

[Kar08j] S. Karlin. James Brown — above beyond [dream jobs 2008]. IEEE Spectrum, 45(2):35–36, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:MKR

[Kar08k] S. Karlin. Math kisser (review of Kiss My Math: Showing Pre-Algebra Who’s Boss by D. McKellar; 2008) [books]. IEEE Spectrum, 45(9):24, September 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:MPG

[Kar08l] S. Karlin. Mini-profile — Gordon Clapp: the voice in your computer after starring turns. IEEE Spectrum, 45(6):25, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:MPJb

[Kar08m] S. Karlin. Mini-profile — jackie martling: Joke man. IEEE Spectrum, 45(7):21, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:MPJa

[Kar08n] S. Karlin. Mini-profile: Jeannine mosely: Paper sculptor. IEEE Spectrum, 45(5):21, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:MPJc

[Kar08o] S. Karlin. Mini-profile: Joseph gatt: motion-capture actor. IEEE Spectrum, 45(8):22, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:MPM

[Kar08p] S. Karlin. Mini-profile: Mike daisey: One-man show. IEEE Spectrum, 45(4):25, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2008:RNJ

[Kar08q] S. Karlin. A rock named judy.

IEEE Spectrum, 45(12):24, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:AFF

[Kar09a] S. Karlin. Art fraud forensics. IEEE Spectrum, 46(7):23, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:D

[Kar09b] S. Karlin. The doodlemeister.

IEEE Spectrum, 46(2):17, February 2009. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:DDR

[Kar09c] S. Karlin. Dumb and dumber [review of “The Dumbest Generation: How the Digital Age Stupifies Young Americans and Jeopardizes Our Future, Or Don’t Trust Anyone Under 30” (Bauerlein, M.; 2008)]. IEEE Spectrum, 46(2):19, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:FA

[Kar09d] S. Karlin. Foreign affair. IEEE Spectrum, 46(3):22, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:GGI

[Kar09e] S. Karlin. Getting getting-it-done done. IEEE Spectrum, 46(6):24, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:MN

[Kar09f] S. Karlin. Money for nothing. IEEE Spectrum, 46(4):22, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:PB

[Kar09g] S. Karlin. Phone-y brew. IEEE Spectrum, 46(7):22, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:QR

[Kar09h] S. Karlin. Quantum rock. IEEE Spectrum, 46(8):22, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:R

[Kar09i] S. Karlin. Robogolf. IEEE Spectrum, 46(11):19, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:SS

[Kar09j]  S. Karlin.  Scientific sages.

IEEE Spectrum, 46(2):21, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:YBG

[Kar09k] S. Karlin. This is your brain on Google [review of iBrain: Surviving the Technological Alteration of the Modern Mind (Small, G. and Vorgan, G.; 2008)]. IEEE Spectrum, 46(3):23, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:TPC

[Kar09l] S. Karlin. Throwing physics a curve. IEEE Spectrum, 46(3):23, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karlin:2009:WRP

[Kar09m] S. Karlin. Where rap and physics collide [YouTube rap video review]. IEEE Spectrum, 46 (2):19, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kasevich:2002:CRH

[Kas02] R. S. Kasevich. Cellphones, radars, and health. IEEE Spectrum, 39(8):15–16, August 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Katz:2009:TTB

[Kat09] R. H. Katz. Tech titans building boom. IEEE Spectrum, 46 (2):40–54, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kinard:2001:VCV

[KB01] B. Kinard and J. Balbach. A venture capitalist’s view of engineers [speakout]. IEEE Spectrum, 38 (12):14–15, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:SS

[KC04] J. Kumagai and S. Cherry. Sensors and sensibility. IEEE Spectrum, 41(7):22–26, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2008:G

[KDG^{+}08] J. Kumagai, D. P. Dern, Erico Guizzo, S. Sontag, S. Cherry, and Tekla S. Perry. Gifts. IEEE Spectrum, 45(11):25–30, November 2008. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Kelly:2001:TCS

[Kel01] C. Kelly. Teaching from a clean slate. IEEE Spectrum, 38(9): 58–61, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kent:2000:TII

[Ken00] S. Kent. On the trail of intrusions into information systems. IEEE Spectrum, 37(12):52–56, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koshima:2000:PLS

[KH00] H. Koshima and J. Hoshen. Personal locator services emerge. IEEE Spectrum, 37(2):41–48, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kimura:2005:PEH

[Kim05] J. Kimura. Photo essay: Hold me closer, tinny dancer. IEEE Spectrum, 42(9):22–23, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kitamura:2005:PER

[Kit05] T. Kitamura. Photo essay:

Red LEDs for green groceries. IEEE Spectrum, 42(4):22–23, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kariya:2002:NTT

[KK02] S. Kariya and P. Kaufman. New technology transforms tactics in Afghanistan [warfare technology]. IEEE Spectrum, 39 (4):30–33, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kroll:2008:IPD

[KKG08] M. W. Kroll, K. Kroll, and B. Gilman. Idiot-proofing the defibrillator. IEEE Spectrum, 45 (11):40–45, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Klemens:2005:NLC

[Kle05a] B. Klemens. New legal code [software patents]. IEEE Spectrum, 42(8):60–62, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Klemens:2005:SPD

[Kle05b] B. Klemens. Software patents don’t compute: no clear boundary between math and software exists. IEEE Spectrum, 42(7):56– 58, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:NA

[KLS^{+}01] J. Kumagai, P. S. Leven, M. Schrope, J. Mullins, E. A. Bretz, R. M. Sherwin, S. Barlas, G. Weiss, Samuel K. Moore, S. R. Das, and A. Hellemans. News analysis. IEEE Spectrum, 38(5):18–40, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karn:2007:NPM

[KM07] B. Karn and H. S. Matthews.

Nano particles without macroproblems. IEEE Spectrum, 44 (9):55–58, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kassakian:2000:AEP

[KMT00] J. G. Kassakian, J. M. Miller, and N. Traub. Automotive electronics power up. IEEE Spectrum, 37(5):34–39, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kolesidis:2004:PES

[Kol04] J. Kolesidis. Photo essay: Smart bridge. IEEE Spectrum, 41 (7):20–21, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koomey:2003:SWN

[Koo03] J. G. Koomey. Sorry, wrong number — how to separate fact from fiction in the information age. IEEE Spectrum, 40(6):11–12, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kornbluh:2000:TST

[Kor00] Ken Kornbluh. Technical software [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1):58–62, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Korte:2006:BPF

[Kor06] F. Korte. The big picture: Four (million) eyes. IEEE Spectrum, 43(6):22–23, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kruft:2000:CEA

[KP00] E. Kruft and P. Packebush. Calibration ensures accuracy. IEEE Spectrum, 37(11):66–69, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koch:2009:UH

[KP09] P. Koch and R. Prasad. The universal handset. IEEE Spectrum, 46(4):36–41, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kraemer:2001:TSB

[Kra01a] A. Kraemer. Two speakers are better than 5.1 [surround sound]. IEEE Spectrum, 38 (5):70–74, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kraft:2001:FFB

[Kra01b] C. Kraft. Fights and flights [book review]. IEEE Spectrum, 38(6):80–82, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kraft:2002:NCN

[Kra02] C. C. Kraft. A new course for nasa. IEEE Spectrum, 39 (1):14–15, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Krein:2005:F

[Kre05] P. Krein. Forum. IEEE Spectrum, 42(1):6, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Krieger:2006:HNM

[Kri06] K. Krieger. Heart of a new machine [robot that can express different moods]. IEEE Spectrum, 43(7):48–51, July 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kroll:2007:CPS

[Kro07] M. W. Kroll. Crafting the perfect shock. IEEE Spectrum, 44 (12):27–31, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:EAR

[KS04] J. Kumagai and W. Sweet. East Asia rising. IEEE Spectrum, 41 (10):42–47, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2005:SRC

[KSC^{+}05] J. Kumagai, W. Sweet, S. Cherry, P. Fairley, L. Geppert, M. Hood, Jen Lin-Liu, Tekla S. Perry, and P. P. Predd. Special report: China’s tech revolution. IEEE Spectrum, 42(6):19, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koch:2008:CMC

[KT08] C. Koch and G. Tononi. Can machines be conscious? IEEE Spectrum, 45(6):55–59, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2000:CDR

[Kum00a] J. Kumagai. Chip detectives [reverse engineering]. IEEE Spectrum, 37(11):43–48, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2000:TPW

[Kum00b] J. Kumagai. Technology priorities for the white house. IEEE Spectrum, 37(9):52–56, September 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:DMC

[Kum01a] J. Kumagai. Disciplined minds: a critical look at salaried professionals and the soul-battering system that shapes their lives, Jeff Schmidt, Rowman & Littlefield, Lanhan, MD, $26.95 [book review]. IEEE Spectrum, 38 (11):66, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:CMT

[Kum01b] J. Kumagai. Choosing management tools that work [careers]. IEEE Spectrum, 38 (8):60–62, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:FSH

[Kum01c] J. Kumagai. Fighting in the streets [high-tech military equipment]. IEEE Spectrum, 38(2): 68–71, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:LYO

[Kum01d] J. Kumagai. Leaving on your own terms. IEEE Spectrum, 38 (3):75–76, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:OHC

[Kum01e] J. Kumagai. Occupational hazard [careers]. IEEE Spectrum, 38 (2):73–75, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:RFF

[Kum01f] J. Kumagai. Rambus: friend or foe? [high-speed memory interface technology manufacture]. IEEE Spectrum, 38 (5):42–47, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:RDW

[Kum01g] J. Kumagai. Redefining diversity [workplace minorities representation]. IEEE Spectrum, 38 (12):56–57, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:TWE

[Kum01h] J. Kumagai. Tapping into a world of expertise [careers].

IEEE Spectrum, 38(1):123– 126, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:WWC

[Kum01i] J. Kumagai. The web as weapon [cyber warfare]. IEEE Spectrum, 38(1):118–121, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2001:WDS

[Kum01j] J. Kumagai. When disaster strikes. IEEE Spectrum, 38(11): 63–64, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2002:PMI

[Kum02a] J. Kumagai. On point for the military [integrated fighting ensemble combat outfit]. IEEE Spectrum, 39(4):65–67, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2002:TM

[Kum02b] J. Kumagai. Talk to the machine. IEEE Spectrum, 39(9): 60–64, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2002:TCP

[Kum02c] J. Kumagai. Techno cops [police robotic and electronic technology]. IEEE Spectrum, 39(12): 34–39, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2002:DWA

[Kum02d] J. Kumagai. To disarm [weapons abolition]. IEEE Spectrum, 39 (1):30–37, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2003:WSY

[Kum03a] J. Kumagai. Are we safe yet?

[nuclear and biological terrorism threats]. IEEE Spectrum, 40 (1):63–64, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2003:CT

[Kum03b] J. Kumagai. It’s cold out there. IEEE Spectrum, 40(7):13– 15, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2003:MIF

[Kum03c] J. Kumagai. Mission impossible? [fbi computer network]. IEEE Spectrum, 40(4):26–31, April 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2003:PPU

[Kum03d] J. Kumagai. Problems persist with U.S. visas. IEEE Spectrum, 40(12):12–14, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2003:RFE

[Kum03e] J. Kumagai. Reversal of fortune [engineer career prospects].

IEEE Spectrum, 40(1):74, January 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2003:WUS

[Kum03f] J. Kumagai. Will U.S. sanctions have chilling effect on scholarly publishing? IEEE Spectrum, 40 (11):12–15, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:CBA

[Kum04a] J. Kumagai. Careers: A balancing act. IEEE Spectrum, 41 (3):60–61, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:JNC

[Kum04b] J. Kumagai. Jim Newman: call of the sea. IEEE Spectrum, 41 (2):35–36, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:LBW

[Kum04c] J. Kumagai. Life of birds [wireless sensor network for bird study]. IEEE Spectrum, 41 (4):42–49, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:RBI

[Kum04d] J. Kumagai. Recycling a bad idea. IEEE Spectrum, 41(1): 48–49, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:ST

[Kum04e] J. Kumagai. Sand trap. IEEE Spectrum, 41(6):44–50, June 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2004:WBD

[Kum04f] J. Kumagai. The whistleblower’s dilemma. IEEE Spectrum, 41(4):53–55, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2005:CNE

[Kum05a] J. Kumagai. Craig Nance: Engineer to the stars. IEEE Spectrum, 42(2):31–32, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2005:KDT

[Kum05b] J. Kumagai. Keith Donovan:

The tintinnabulator [clock and bell tower automation]. IEEE Spectrum, 42(2):32, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2005:PCN

[Kum05c] J. Kumagai. The panda connection: a new wireless network helps panda researchers and school kids, too. IEEE Spectrum, 42(6):61, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagal:2005:SMS

[Kum05d] J. Kumagal. Space mountain [submillimeter radio telescopes]. IEEE Spectrum, 42(12): 12–14, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2006:DS

[Kum06a] J. Kumagai. Drowning in sound. IEEE Spectrum, 43 (4):54–60, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2006:HMN

[Kum06b] J. Kumagai. Halfway to Mars [next generation planetary rover]. IEEE Spectrum, 43(3):38–43, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2006:MGB

[Kum06c] J. Kumagai. Mars gets broadband connection. IEEE Spectrum, 43(2):15, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2006:UDD

[Kum06d] J. Kumagai. U.S. defense dollars for computer science plunge [computer science research]. IEEE Spectrum, 43 (2):12–15, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2007:AA

[Kum07a] J. Kumagai. AT&T’s attic.

IEEE Spectrum, 44(3):30–37, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2007:HMS

[Kum07b] J. Kumagai. How to master a seismic disaster. IEEE Spectrum, 44(6):48–51, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2007:RR

[Kum07c] J. Kumagai. Radio revolutionaries. IEEE Spectrum, 44 (1):28–32, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2007:RSK

[Kum07d] J. Kumagai. A robotic sentry for Korea’s Demilitarized Zone. IEEE Spectrum, 44 (3):16–17, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2007:SE

[Kum07e] J. Kumagai. Swimming to europa. IEEE Spectrum, 44(9): 33–40, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2008:ANE

[Kum08a] J. Kumagai. Ash nehro — everything is illuminated [dream jobs 2008]. IEEE Spectrum, 45 (2):40–42, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2008:UNT

[Kum08b] J. Kumagai. Update — NASA touches down in Moses Lake. IEEE Spectrum, 45(7):16, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2009:KKR

[Kum09a] J. Kumagai. Kunio koike real time. IEEE Spectrum, 46(2): 35–36, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumagai:2009:QL

[Kum09b] J. Kumagai. Quantum leap. IEEE Spectrum, 46(1):51–54, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kunz:2001:WI

[Kun01] T. Kunz. Is the WAP a inner? IEEE Spectrum, 38(2): 78–79, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kuniholm:2009:OA

[Kun09] J. Kuniholm. Open arms. IEEE Spectrum, 46(3):36–41, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kurzweil:2008:SLW

[Kur08] R. Kurzweil. The singularity: The last word — [forum].

IEEE Spectrum, 45(10):10, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2002:DYD

[Kus02a] D. Kushner. 21 dog years — doing time @amazon.com [book review]. IEEE Spectrum, 39(7):43– 44, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2002:LL

[Kus02b] D. Kushner. Living in the limelight. IEEE Spectrum, 39(9): 76–79, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2002:WWW

[Kus02c] D. Kushner. Watching those who watch us. IEEE Spectrum, 39(6):67–68, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2002:WIV

[Kus02d] D. Kushner. The wizardry of id [video games]. IEEE Spectrum, 39(8):42–47, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2003:AFC

[Kus03a] D. Kushner. The audio frontier [computer audio engineers’ retreat]. IEEE Spectrum, 40 (2):42–47, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2003:CWG

[Kus03b] D. Kushner. Careers: The wandering gadgeteer. IEEE Spectrum, 40(12):50–52, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2003:DDF

[Kus03c] D. Kushner. Digital decoys [fake MP3 song files to deter music pirating]. IEEE Spectrum, 40(5):27, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2003:MMW

[Kus03d] D. Kushner. It’s a mod, mod world [computer games, copyrighted material modification]. IEEE Spectrum, 40(2): 56–57, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2003:SLE

[Kus03e] D. Kushner. The social life of engineers. IEEE Spectrum, 40(10):52, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2003:WYG

[Kus03f] D. Kushner. Who ya gonna call? [technical support]. IEEE Spectrum, 40(1):94–96, January 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2003:WMP

[Kus03g] D. Kushner. Why are so many patents pending? IEEE Spectrum, 40(4):52–54, April 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2004:BBW

[Kus04a] D. Kushner. Big brother at work [workplace surveillance]. IEEE Spectrum, 41(12):57–58, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2004:LUE

[Kus04b] D. Kushner. Let us entertain you [entertainment engineering and technology]. IEEE Spectrum, 41 (2):58–60, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2005:EEO

[Kus05a] D. Kushner. Engineering EverQuest: online gaming demands heavyweight data centers. IEEE Spectrum, 42(7):34– 39, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2005:MIH

[Kus05b] D. Kushner. Media: Inside hollywood. IEEE Spectrum, 42(12): 53–56, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2005:OSM

[Kus05c] D. Kushner. Open-source media [mobile multimedia software]. IEEE Spectrum, 42(2): 50–51, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2006:BSO

[Kus06a] D. Kushner. Beating Sony at its own game [copyright protection]. IEEE Spectrum, 43 (2):57–59, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2006:FK

[Kus06b] D. Kushner. The Firefox Kid. IEEE Spectrum, 43(11): 26–31, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2006:I

[Kus06c] D. Kushner. The insomniacs. IEEE Spectrum, 43(12): 24–29, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2006:LLL

[Kus06d] D. Kushner. Location, location, location [location-based games]. IEEE Spectrum, 43 (1):62–67, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2006:ULC

[Kus06e] D. Kushner. Uncommon law [copyright laws]. IEEE Spectrum, 43(3):57–58, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2006:V

[Kus06f] D. Kushner. Vegas 911. IEEE Spectrum, 43(4):44–49, April 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2007:PD

[Kus07a] D. Kushner. Playing dirty. IEEE Spectrum, 44(12):32–37, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2007:RP

[Kus07b] D. Kushner. Resources — the player. IEEE Spectrum, 44(12): 57–60, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2007:SS

[Kus07c] D. Kushner. The slashdot supremacy. IEEE Spectrum, 44 (11):34–38, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2008:ES

[Kus08a] D. Kushner. Engineering spore.

IEEE Spectrum, 45(9):36–40, September 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2008:MMM

[Kus08b] D. Kushner. Machinima’s movie moguls. IEEE Spectrum, 45 (7):36–41, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2008:WSM

[Kus08c] D. Kushner. Winner: Software — make your very own virtual world with olive. IEEE Spectrum, 45(1):45–47, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2009:MBR

[Kus09a] D. Kushner. The making of the beatles: Rock band. IEEE Spectrum, 46(9):30–35, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kushner:2009:TIV

[Kus09b] D. Kushner. Texas Instruments vs. the calculator hackers. IEEE Spectrum, 46(12):14, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Khuri-yakub:2009:NGU

[KyOK09] B. T. Khuri-yakub, O. Oralkan, and M. Kupnik. Next-gen ultrasound. IEEE Spectrum, 46 (5):44–54, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laird:2003:SRM

[Lai03] C. Laird. Software reviews: Microsoft hits a new programming note. IEEE Spectrum, 40(11): 50–51, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Langley:2002:HTG

[Lan02] M. Langley. Hands off telecom: give deregulation a chance. IEEE Spectrum, 39(6):10–12, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lang:2003:HTE

[Lan03a] C. Lang. Harnessing tidal energy takes new turn. IEEE Spectrum, 40(9):13, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lang:2003:PBO

[Lan03b] C. Lang. President Bush orders government spectrum review. IEEE Spectrum, 40(8):15, August 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Larson:2007:TT

[Lar07] K. Larson. The technology of text. IEEE Spectrum, 44 (5):26–31, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Layer:2001:DRT

[Lay01] D. H. Layer. Digital radio takes to the road. IEEE Spectrum, 38(7):40–46, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lazarus:2009:RRR

[Laz09] M. Lazarus. Radio’s regulatory roadblocks. IEEE Spectrum, 46 (9):42–56, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lal:2004:DDN

[LB04] A. Lal and J. Blanchard. Daintiest dynamos [nuclear microbatteries]. IEEE Spectrum, 41(9): 36–41, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

LaMarca:2004:TTF

[LCS04] A. LaMarca, Y. Chawathe, and I. Smith. Tools toys: Finding yourself. IEEE Spectrum, 41 (12):51–54, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lee:2005:F

[Lee05] J. C. M. Lee. Forum. IEEE Spectrum, 42(8):8, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Leventon:2002:SS

[Lev02] W. Leventon. Synthetic skin. IEEE Spectrum, 39(12):28–33, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levy:2008:GTB

[Lev08] S. Levy. 10 great tech books. IEEE Spectrum, 45(7):42–49, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lewis:2001:DSB

[Lew01] J. S. Lewis. Designs on space [book review]. IEEE Spectrum, 38(4):84, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lloyd-Hart:2003:TTS

[LH03] M. Lloyd-Hart. Taking the twinkle out of starlight. IEEE Spectrum, 40(12):22–29, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Li:2001:WYM

[Li01] Ambrose Li. Why your e-mail may never arrive after all. IEEE Spectrum, 38(4):15, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Linden:2008:PWR

[Lin08] G. Linden. People who read this article also read . . .. IEEE Spectrum, 45(3):46–60, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lipson:2005:HFT

[Lip05] H. Lipson. Homemade [fabrication technology]. IEEE Spectrum, 42(5):24–31, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lang:2003:VMS

[LJ03] C. Lang and W. D. Jones. Virginia Maglev system off to shaky start. IEEE Spectrum, 40(11): 14–15, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levin:2002:CMS

[LL02a] P. L. Levin and R. Ludwig. Crossroads for mixedsignal chips. IEEE Spectrum, 39(3):38–43, March 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2002:NKN

[LL02b] Jen Lin-Liu. North Korea’s nuclear revelation puts spotlight on China. IEEE Spectrum, 39(12): 22–23, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2002:OIP

[LL02c] Jen Lin-Liu. Offshore investors place their chips on Chinese foundries. IEEE Spectrum, 39 (11):15–16, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2003:HCT

[LL03a]  Jen Lin-Liu.  Huawei–Cisco

tests China’s respect for property rights. IEEE Spectrum, 40 (8):10–12, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2003:MSC

[LL03b] Jen Lin-Liu. Microsoft suffers Chinese triple whammy. IEEE Spectrum, 40(10):18, October 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2003:WWG

[LL03c] Jen Lin-Liu. Who will get China’s 3G dowry? IEEE Spectrum, 40(4):16–18, April 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2004:CTG

[LL04a] J. Lin-Liu. A Chinese town’s get-rich scheme: e-trash to cash. IEEE Spectrum, 41(8):13–16, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2004:TUO

[LL04b] Jen Lin-Liu. Top U.S. officials successfully defend IEEE’s Wi-Fi standard. IEEE Spectrum, 41(6):16, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2005:MAS

[LL05a] J. Lin-Liu. Management American style: a US university teaches Western business ways to Chinese students. IEEE Spectrum, 42(6):36, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2005:DTT

[LL05b] Jen Lin-Liu. Digital TV trial brews in China. IEEE Spectrum, 42(12):20–21, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin-Liu:2006:CJH

[LL06] J. Lin-Liu. China’s e-junkies head for rehab [Internet addicts treatment]. IEEE Spectrum, 43 (2):19, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lave:2001:HVW

[LM01] L. B. Lave and H. L. Maclean. Are hybrid vehicles worth it? IEEE Spectrum, 38(3):47–50, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Loebner:2006:DIY

[Loe06]  H. Loebner.  Do-it-yourself

patents. IEEE Spectrum, 43 (4):68–70, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Loui:2000:IEB

[Lou00] M. C. Loui. Internet ethics [books]. IEEE Spectrum, 37(11): 15–16, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lahey:2005:BPO

[LTN05] R. T. Lahey, R. P. Taleyarkhan, and R. I. Nigmatulin. Bubble power [other sources of nuclear energy]. IEEE Spectrum, 42(5):38–43, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lubell:2000:CAD

[Lub00] P. D. Lubell. A coming attraction: digital cinema. IEEE Spectrum, 37(3):72–78, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Luck:2000:R

[Luc00a] R. W. Luck. Reflections. IEEE Spectrum, 37(7):25, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2000:MBR

[Luc00b] R. W. Lucky. Monkey business [reflections]. IEEE Spectrum, 37 (9):37, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2000:Ra

[Luc00c]  R. W. Lucky.  Reflections.

IEEE Spectrum, 37(1):25, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2000:Rb

[Luc00d] R. W. Lucky. Reflections. IEEE Spectrum, 37(3):27, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2000:Rc

[Luc00e] R. W. Lucky. Reflections. IEEE Spectrum, 37(5):25, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2000:Rd

[Luc00f] R. W. Lucky. Reflections. IEEE Spectrum, 37(11):33, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2001:TMR

[Luc01a] R. W. Lucky. 2001: Technology and the movie [reflections]. IEEE Spectrum, 38 (3):102, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2001:CTR

[Luc01b] R. W. Lucky. Communications triage [reflections]. IEEE Spectrum, 38(1):162, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2001:FFR

[Luc01c] R. W. Lucky. Fact or fiction? [reflections]. IEEE Spectrum, 38(7): 96, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2001:PRS

[Luc01d] R. W. Lucky. The precious radio spectrum [reflections]. IEEE Spectrum, 38(9):90, September 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2001:TDA

[Luc01e] R. W. Lucky. The telecom dark ages [reflections]. IEEE Spectrum, 38(11):96, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2001:WGW

[Luc01f] R. W. Lucky. When giants walked the earth [reflections]. IEEE Spectrum, 38(4):104, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2002:FER

[Luc02a] R. W. Lucky. The future of engineering [reflections]. IEEE Spectrum, 39(9):86, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2002:PFR

[Luc02b] R. W. Lucky. A paramecium, a fish, and a rat [reflections]. IEEE Spectrum, 39 (11):64, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2002:WDY

[Luc02c] R. W. Lucky. What do you do at work? [reflections]. IEEE Spectrum, 39(7):56, July 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2002:WRI

[Luc02d] R. W. Lucky. Who runs the Internet? [reflections]. IEEE Spectrum, 39(3):80, March 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2003:DFR

[Luc03a] R. W. Lucky. Down into darkness or up into fog [reflections]. IEEE Spectrum, 40(3):76, March 2003.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2003:EAR

[Luc03b] R. W. Lucky. Engineering appearance [reflections]. IEEE Spectrum, 40(11):60, November 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2003:FNR

[Luc03c] R. W. Lucky. Flames from the net [reflections]. IEEE Spectrum, 40(1):104, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2003:IRO

[Luc03d] R. W. Lucky. Is “industrial research” an oxymoron? [reflections]. IEEE Spectrum, 40 (9):56, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2003:ODR

[Luc03e] R. W. Lucky. Order and disorder [reflections]. IEEE Spectrum, 40(7):52, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2003:WWW

[Luc03f] R. W. Lucky. What were we thinking? [reflections]. IEEE Spectrum, 40(5):56, May 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2004:DGL

[Luc04a] R. W. Lucky. Does Google like me? [reflections]. IEEE Spectrum, 41(11):136, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2004:EER

[Luc04b] R. W. Lucky. Engineering expressiveness [reflections]. IEEE Spectrum, 41(9):80, September 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2004:GMR

[Luc04c] R. W. Lucky. Goodbye to e-mail [reflections]. IEEE Spectrum, 41(1):120, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2004:IKR

[Luc04d] R. W. Lucky. The impermanence of knowledge [reflections]. IEEE Spectrum, 41(3):80, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2004:ROR

[Luc04e] R. W. Lucky. Retrospective organization [reflections]. IEEE Spectrum, 41(7):72, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2004:WVT

[Luc04f] R. W. Lucky. Where is the vision for telecom? [reflections]. IEEE Spectrum, 41(5):72, May 2004.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2005:DR

[Luc05a] R. W. Lucky. The dream [reflections]. IEEE Spectrum, 42(7):76, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2005:LSR

[Luc05b] R. W. Lucky. Lab for sale [reflections]. IEEE Spectrum, 42 (9):92, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2005:LBR

[Luc05c] R. W. Lucky. Life bits [reflections]. IEEE Spectrum, 42(5):76, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2005:MNL

[Luc05d] R. W. Lucky. My new laptop [reflections]. IEEE Spectrum, 42 (11):104, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2005:RRR

[Luc05e] R. W. Lucky. Robert’s Rules [reflections]. IEEE Spectrum, 42(3):80, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2005:WR

[Luc05f] R. W. Lucky. Wordsmithing [reflections]. IEEE Spectrum, 42(1):96, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2006:DPR

[Luc06a] R. W. Lucky. Digital photographs [reflections]. IEEE Spectrum, 43(5):84, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2006:FPR

[Luc06b] R. W. Lucky. Famous people [reflections]. IEEE Spectrum, 43(7): 68, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2006:ITR

[Luc06c] R. W. Lucky. Impossible tradeoffs [reflections]. IEEE Spectrum, 43(11):84, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2006:RCR

[Luc06d] R. W. Lucky. Remote control [reflections]. IEEE Spectrum, 43(3):84, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2006:TCR

[Luc06e] R. W. Lucky. Tragedy of the commons [reflections]. IEEE Spectrum, 43(1):88, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2006:UER

[Luc06f] R. W. Lucky. Unsystematic engineering [reflections]. IEEE Spectrum, 43(9):84, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2007:YAR

[Luc07a] R. W. Lucky. Are you aware? [reflections]. IEEE Spectrum, 44(3):72, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2007:BAR

[Luc07b] R. W. Lucky. A billion amateurs [reflections]. IEEE Spectrum, 44 (11):96, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2007:CRR

[Luc07c] R. W. Lucky. Computer rot [reflections]. IEEE Spectrum, 44(7): 68, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2007:GTR

[Luc07d] R. W. Lucky. Great thoughts [reflections]. IEEE Spectrum, 44(1):84, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2007:MBR

[Luc07e] R. W. Lucky. Math blues [reflections]. IEEE Spectrum, 44 (9):80, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2007:WYR

[Luc07f] R. W. Lucky. Who are you?

[reflections]. IEEE Spectrum, 44(5):68, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2008:BLR

[Luc08a] R. W. Lucky. The blinking light [reflections]. IEEE Spectrum, 45 (11):24, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2008:EFR

[Luc08b] R. W. Lucky. The elusive future [reflections]. IEEE Spectrum, 45 (9):20, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2008:TPI

[Luc08c] R. W. Lucky. Technical publications and the Internet [reflections]. IEEE Spectrum, 45 (1):25, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2008:UEF

[Luc08d] R. W. Lucky. US engineers and the flat Earth [reflections]. IEEE Spectrum, 45(3):19, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2008:ZPR

[Luc08e] R. W. Lucky. Zero privacy [reflections]. IEEE Spectrum, 45(7): 20, July 2008. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2008:ZDR

[Luc08f] R. W. Lucky. Zipf drive [reflections]. IEEE Spectrum, 45(5):18, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2009:CCR

[Luc09a] R. W. Lucky. Cloud computing [reflections]. IEEE Spectrum, 46(5):27, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2009:DCU

[Luc09b] R. W. Lucky. The dreaded computer upgrade [reflections]. IEEE Spectrum, 46(3):25, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2009:EAT

[Luc09c]  R. W. Lucky.  Engineering

achievements: the two lists [reflections]. IEEE Spectrum, 46 (11):23, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2009:FBE

[Luc09d] R. W. Lucky. The first book of electronics [reflections]. IEEE Spectrum, 46(9):29, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2009:TTR

[Luc09e] R. W. Lucky. To twitter or not to twitter? [reflections]. IEEE Spectrum, 46(1):22, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lucky:2009:WPR

[Luc09f] R. W. Lucky. Wicked problems [reflections]. IEEE Spectrum, 46(7):27, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lukaszewski:2009:MYB

[Luk09] J. E. Lukaszewski. Managing your boss. IEEE Spectrum, 46 (2):17, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lauterbach:2000:AJI

[LW00] M. Lauterbach and T. Wey. Analyze jitter to improve highspeed design. IEEE Spectrum, 37(7):62–67, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MacGregor:2001:ERB

[Mac01] T. MacGregor. Electricity restructuring in britain: not a model to follow. IEEE Spectrum, 38(6):15–19, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Macedonia:2002:GSP

[Mac02] M. Macedonia. Games soldiers play. IEEE Spectrum, 39 (3):32–37, March 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mack:2008:SD

[Mac08] C. A. Mack. Seeing double. IEEE Spectrum, 45(11):46–51, November 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maize:2003:OET

[Mai03] K. Maize. Opening up energy trading. IEEE Spectrum, 40 (1):54–58, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Makansi:2001:RBG

[Mak01] J. Makansi. Reactorland: A board game. IEEE Spectrum, 38 (11):42–43, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Makansi:2002:SSE

[Mak02] J. Makansi. Sham? Shame! The electric power industry [book review]. IEEE Spectrum, 39(4):59, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mandell:2000:LUS

[Man00] M. Mandell. 120000 leagues under the sea [undersea optical cables]. IEEE Spectrum, 37(4):50–54, April 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mandelbaum:2002:RWW

[Man02] R. Mandelbaum. Reap the wild wind [offshore wind farm]. IEEE Spectrum, 39(10):34–39, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mandelbaum:2008:BWC

[Man08]  R. Mandelbaum.  Brake-by-

wire comes to freight trains. IEEE Spectrum, 45(2):18, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mandelbaum:2009:ST

[Man09] R. Mandelbaum. Stop that train. IEEE Spectrum, 46 (3):46–49, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:2000:MDD

[Mar00] C. D. Martin. Mad for digital or digital madness? [book review]. IEEE Spectrum, 37(9): 14–16, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:2001:ACA

[Mar01] B. Martin. Automation comes to analog. IEEE Spectrum, 38(6):70–75, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:2002:ESA

[Mas02] J. F. Mason. Euro space agency opens ’02 with a bang. IEEE Spectrum, 39(4):18–20, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mosher:2000:MRN

[MBP00] D. E. Mosher, O. Bukharin, and T. E. Perry. Minding Russia’s nuclear store. IEEE Spectrum, 37(3):44–50, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mouli:2007:FF

[MC07] C. Mouli and W. Carriker. Future fab. IEEE Spectrum, 44 (3):38–43, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McCarthy:2005:LLJ

[McC05] W. McCarthy. This looks like a job for superatoms [atom lasers]. IEEE Spectrum, 42 (8):42–47, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McDowell:2004:DSV

[McD04] J. C. McDowell. Downloading the sky [virtual observatory]. IEEE Spectrum, 41 (8):35–39, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2002:FG

[McF02a] Paul McFedries. Filling the gaps. IEEE Spectrum, 39(6):80, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2002:TPS

[McF02b] Paul McFedries. Tall poppy syndrome dot-com. IEEE Spectrum, 39(12):68, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2002:TMC

[McF02c] Paul McFedries. Too many chiefs. IEEE Spectrum, 39(8):64, August 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2003:MRU

[Mcf03a] P. Mcfedries. Mobs R us.

IEEE Spectrum, 40(10):56, October 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2003:AHT

[McF03b] Paul McFedries. The age of high (tech) anxiety. IEEE Spectrum, 40(6):56, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2003:G

[McF03c] Paul McFedries. Google this.

IEEE Spectrum, 40(2):68, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2003:IE

[McF03d] Paul McFedries. The Internet ecology. IEEE Spectrum, 40(4):68, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2003:OPL

[McF03e] Paul McFedries. Our posthuman lingo. IEEE Spectrum, 40(8):52, August 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2003:TSB

[McF03f] Paul McFedries. Technically speaking: Blah, blah, blog. IEEE Spectrum, 40(12):60, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2004:HUR

[McF04a] Paul McFedries. Hacking unplugged [review of hacking activities and terminology]. IEEE Spectrum, 41(2):80, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2004:TSN

[McF04b] Paul McFedries. Technically speaking: No “I” in “internet”? IEEE Spectrum, 41(12):92, December 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2004:TSS

[McF04c] Paul McFedries. Technically speaking: Slicing the ham from the spam. IEEE Spectrum, 41(4):72, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2004:TSH

[McF04d] Paul McFedries. Technically speaking: The hidden plaque. IEEE Spectrum, 41(10):88, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2004:TSP

[McF04e] Paul McFedries. Technically speaking: The (pre) fix is in. IEEE Spectrum, 41(8):64, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2005:BC

[McF05a] Paul McFedries. Bluetooth cavities. IEEE Spectrum, 42(6):88, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2005:TSC

[McF05b] Paul McFedries. Technically speaking: Call me, ishmael. IEEE Spectrum, 42(10):92, October 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2005:TSI

[McF05c] Paul McFedries. Technically speaking: The iPod people. IEEE Spectrum, 42(2):76, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2005:TSS

[McF05d] Paul McFedries. Technically speaking: The spyware nightmare. IEEE Spectrum, 42(8):72, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2005:TSW

[McF05e] Paul McFedries. Technically speaking: Watchwords. IEEE Spectrum, 42(12):88, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2005:WWN

[McF05f] Paul McFedries. Words in the wind [new tech terms]. IEEE Spectrum, 42(4):80, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2006:TSC

[McF06a] Paul McFedries. Technically speaking: Changing climate, changing language. IEEE Spectrum, 43(8):60, August 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2006:TSF

[McF06b] Paul McFedries. Technically speaking: Folk wisdom. IEEE Spectrum, 43(2):80, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2006:TSG

[McF06c] Paul McFedries. Technically speaking: Gone phishin’. IEEE Spectrum, 43(4):80, April 2006.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2006:TSWc

[McF06d] Paul McFedries. Technically speaking: It’s a wiki, wiki world. IEEE Spectrum, 43(12):88, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2006:TSWa

[McF06e] Paul McFedries. Technically speaking the web, take two. IEEE Spectrum, 43(6):68, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2006:TSWb

[McF06f] Paul McFedries. Technically speaking: When good clicks go bad. IEEE Spectrum, 43 (10):88, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2007:TS

[McF07a] Paul McFedries. Technically speaking. IEEE Spectrum, 44 (12):96, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2007:TSA

[McF07b] Paul McFedries. Technically speaking: All a-twitter. IEEE Spectrum, 44(10):84, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2007:TSG

[McF07c] Paul McFedries. Technically speaking: Game on. IEEE Spectrum, 44(2):72, February 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2007:TSS

[McF07d] Paul McFedries. Technically speaking: Social-isms. IEEE Spectrum, 44(4):68, April 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2007:TSH

[McF07e] Paul McFedries. Technically speaking: The hobbyist renaissance. IEEE Spectrum, 44(6):88, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McFedries:2007:TST

[McF07f] Paul McFedries. Technically speaking: Tired vs. wired. IEEE Spectrum, 44(8):68, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2008:CWT

[Mcf08a] P. Mcfedries. E-cycling ewaste — [technically speaking]. IEEE Spectrum, 45(10):21, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2008:SNC

[Mcf08b] P. Mcfedries. Snowclone is the new cliche [technically speaking]. IEEE Spectrum, 45(2):27, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2008:TSH

[Mcf08c] P. Mcfedries. Technically speaking: Homo nerdus. IEEE Spectrum, 45(6):24, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2008:TSG

[Mcf08d] P. Mcfedries. Technically speaking: Of geeks, modders, and overclockers. IEEE Spectrum, 45(4):20, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2008:TW

[Mcf08e] P. Mcfedries. Trafficking in words. IEEE Spectrum, 45 (12):22, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfredries:2008:TSC

[Mcf08f] P. Mcfredries. Technically speaking: The cloud is the computer. IEEE Spectrum, 45(8):20, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2009:BNW

[Mcf09a] P. Mcfedries. Brave neuro world.

IEEE Spectrum, 46(8):25, August 2009. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2009:HD

[Mcf09b] P. Mcfedries. Hacking DNA. IEEE Spectrum, 46(10):23, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2009:IYF

[Mcf09c] P. Mcfedries. Information at your fingertips. IEEE Spectrum, 46(4):24, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfedries:2009:SIS

[Mcf09d] P. Mcfedries. Suffix it to say. IEEE Spectrum, 46(6):25, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mcfredries:2009:OFL

[Mcf09e] P. Mcfredries. The office-free lifestlye. IEEE Spectrum, 46 (12):23, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McNew:2005:PEL

[McN05] D. McNew. Photo essay: Lookin’ for love photograph. IEEE Spectrum, 42(8):22–23, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McNew:2007:BPB

[McN07] D. McNew. The big picture: Breakthrough. IEEE Spectrum, 44(2):20–21, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meinel:2007:LG

[Mei07]  C. Meinel.  For love of a

gun. IEEE Spectrum, 44(7):40– 46, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Merkle:2001:NWW

[Mer01] R. C. Merkle. Nanotechnology:

what will it mean? IEEE Spectrum, 38(1):19–21, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mercuri:2002:BBB

[Mer02] R. Mercuri. A better ballot box? IEEE Spectrum, 39 (10):46–50, October 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meres:2004:PER

[Mer04] A. Meres. Photo essay: Rock and rile. IEEE Spectrum, 41 (9):28–29, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Molinski:2000:SGS

[MFD00] T. S. Molinski, W. E. Feero, and B. L. Damsky. Shielding grids from solar storms [power system protection]. IEEE Spectrum, 37 (11):55–60, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2000:Ia

[MG00a] Samuel K. Moore and L. Geppert. Innovations. IEEE Spectrum, 37(6):28–29, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2000:SIM

[MG00b] Samuel K. Moore and L. Geppert. Silicon IC models brain activity [innovations]. IEEE Spectrum, 37(9):18–20, September 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2007:MN

[MG07] Samuel K. Moore and A. Gardner. Megacities by the numbers. IEEE Spectrum, 44(6):24–25, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Matthews:2000:S

[MHLS00] H. S. Matthews, C. Hendrickson, L. Lave, and A. J. Schwab. Speakout. IEEE Spectrum, 37 (11):20–26, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:2001:PCF

[Mil01] S. Miller. Pragmatic concerns fuel nuclear support. IEEE Spectrum, 38(11):34–36, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Minkel:2003:CTT

[Min03] J. R. Minkel. Could trapping tiny ions crack the toughest codes? IEEE Spectrum, 40(3):31–32, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Minkel:2006:CIF

[Min06a] J. R. Minkel. Could the Internet fragment? [domain name system]. IEEE Spectrum, 43 (6):20–21, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Minkel:2006:QLQ

[Min06b] J. R. Minkel. Quantum leap for quantum computing. IEEE Spectrum, 43(3):17–18, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Minkel:2006:SBT

[Min06c] J. R. Minkel. A smashing bad time for the United States. IEEE Spectrum, 43(8):12–13, August 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Minkel:2006:TUM

[Min06d] J. R. Minkel. A tabletop UV microscope. IEEE Spectrum, 43 (9):16, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mindell:2009:ECA

[Min09] D. A. Mindell. The end of the cult of the astronaut. IEEE Spectrum, 46(6):68–69, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:2000:LNH

[Mir00a] Mark S. Mirotznik. ‘Look, no hands’ signal analysis [software reviews]. IEEE Spectrum, 37 (2):24, February 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mirotznik:2000:SR

[Mir00b] Mark S. Mirotznik. Software reviews. IEEE Spectrum, 37 (8):19–20, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mbarika:2006:ACA

[MM06] V. W. A. Mbarika and I. Mbarika. Africa calling [African wireless connection]. IEEE Spectrum, 43(5):56–60, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2001:NA

[MMG^{+}01] Samuel K. Moore, J. Mullins, P. Gwynne, R. Teitelman, Y. Chiu, Stephen Cass, and V. Singh. News analysis. IEEE Spectrum, 38(1):24–40, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:2001:NA

[MMH^{+}01] E. Murphy, S. Miller, A. Hellemans, P. Gwynne, P. R. Kaufman, and P. Harclerode. News analysis. IEEE Spectrum, 38 (12):16–27, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marculescu:2003:RW

[MMPJ03] D. Marculescu, R. Marculescu, S. Park, and S. Jayaraman. Ready to ware. IEEE Spectrum, 40(10):28–32, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Madani:2005:GGG

[MN05] V. Madani and D. Novosel. Getting a grip on the grid. IEEE Spectrum, 42(12):42–47, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mueller:2000:ADT

[MO00] J. Mueller and R. Oblad. Architecture drives test system standards. IEEE Spectrum, 37(9): 68–73, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mock:2004:ME

[Moc04] T. Mock. Music everywhere.

IEEE Spectrum, 41(9):42–47, September 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mockapetris:2006:TNA

[Moc06] P. V. Mockapetris. Telephony’s next act. IEEE Spectrum, 43 (4):28–32, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Montgomery:2009:HGG

[Mon09] M. Montgomery. The hundredgrand grand. IEEE Spectrum, 46(11):18, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2000:Ib

[Moo00a] Samuel K. Moore. Innovations. IEEE Spectrum, 37(7):75, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2000:Ic

[Moo00b] Samuel K. Moore. Innovations. IEEE Spectrum, 37(10): 18–19, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2000:UHG

[Moo00c] Samuel K. Moore. Understanding the human genome. IEEE Spectrum, 37(11):33–35, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moody:2001:HRB

[Moo01a] G. Moody. The hidden revolution [book review]. IEEE Spectrum, 38(3):78–80, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moor:2001:MCC

[Moo01b] S. K. Moor. Microfluidics for complex computation? IEEE Spectrum, 38(6):28–29, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2001:BBC

[Moo01c] Samuel K. Moore. Better breast cancer detection. IEEE Spectrum, 38(5):50–54, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2001:CMF

[Moo01d] Samuel K. Moore. CO_{2} makes fillings for chip cavities. IEEE Spectrum, 38(11):28–29, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2001:HDS

[Moo01e] Samuel K. Moore. Harmonizing data, setting standards [genomics, information sets]. IEEE Spectrum, 38(1):111–112, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2001:MCP

[Moo01f] Samuel K. Moore. Making chips to probe genes. IEEE Spectrum, 38(3):54–60, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2001:UMW

[Moo01g] Samuel K. Moore. Unhooking medicine [wireless networking]. IEEE Spectrum, 38(1): 107–108, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2002:BUI

[Moo02a] Samuel K. Moore. The brain as user interface [remotely guided rats]. IEEE Spectrum, 39(8):29, August 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2002:EHR

[Moo02b] Samuel K. Moore. Extending healthcare’s reach. IEEE Spectrum, 39(1):66–71, January 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2002:JOW

[Moo02c] Samuel K. Moore. Just one word — plastics [organic semiconductors]. IEEE Spectrum, 39(9): 55–59, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2003:DFN

[Moo03a] Samuel K. Moore. Danes forge nanotools. IEEE Spectrum, 40(7):15, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2003:ESS

[Moo03b] Samuel K. Moore. Extreme solar storm strikes earth. IEEE Spectrum, 40(12):15–16, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2004:ERT

[Moo04a] Samuel K. Moore. Electric rainmaking technology gets Mexico’s blessing. IEEE Spectrum, 41 (4):26–27, April 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2004:PFM

[Moo04b] Samuel K. Moore. Price fixing in the memory market [dram prices]. IEEE Spectrum, 41(12): 18–20, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2005:BBM

[Moo05a] Samuel K. Moore. Back from the brink: a made-for-TV technology is finally ready for prime time. IEEE Spectrum, 42(5):18, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2005:NPT

[Moo05b] Samuel K. Moore. The nanotech patent trap. IEEE Spectrum, 42(7):18, July 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2005:ZBP

[Moo05c] Samuel K. Moore. Zapping away the blues: a pacemaker-like device to treat depression takes a giant step forward. IEEE Spectrum, 42(5):16–17, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:BSI

[Moo06a] Samuel K. Moore. Barbados has a sweet idea [power generation from sugarcane breeding]. IEEE Spectrum, 43(1): 16–17, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:CYP

[Moo06b] Samuel K. Moore. Calm in your palm: biofeedback device promises to reduce stress. IEEE Spectrum, 43(3):60, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:CCN

[Moo06c] Samuel K. Moore. Cheap chips for next wireless frontier. IEEE Spectrum, 43(6):12–13, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:FIW

[Moo06d] Samuel K. Moore. Fingerprint ID for wireless keys. IEEE Spectrum, 43(1):53, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:LS

[Moo06e] Samuel K. Moore. Laser on silicon. IEEE Spectrum, 43 (11):18, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:MMS

[Moo06f] Samuel K. Moore. Multimedia monster [supercomputer on a single chip]. IEEE Spectrum, 43 (1):20–23, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:PPP

[Moo06g] Samuel K. Moore. Poky plastic perks up [polymer semiconductor]. IEEE Spectrum, 43 (5):20–21, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2006:PSN

[Moo06h] Samuel K. Moore. Psychiatry’s shocking new tools [brain stimulation techniques]. IEEE Spectrum, 43(3):24–31, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2007:CF

[Moo07a] Samuel K. Moore. Cold fission.

IEEE Spectrum, 44(2):16, February 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2007:CQK

[Moo07b] Samuel K. Moore. Commercializing quantum keys. IEEE Spectrum, 44(3):15–17, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2007:FTN

[Moo07c] Samuel K. Moore. Fairchild turns 50 [news]. IEEE Spectrum, 44 (10):12, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2007:HMC

[Moo07d] Samuel K. Moore. How to measure a city’s metabolism. IEEE Spectrum, 44(6):30–31, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2007:MM

[Moo07e] Samuel K. Moore. Masters of memory. IEEE Spectrum, 44 (1):45–49, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2008:IMB

[Moo08a] Samuel K. Moore. Intel makes a big jump in computer math. IEEE Spectrum, 45(2): 14–15, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2008:MBN

[Moo08b] Samuel K. Moore. Multicore is bad news for supercomputers. IEEE Spectrum, 45 (11):15, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2008:SRS

[Moo08c] Samuel K. Moore. Supply risk, scarcity, and cellphones [the data]. IEEE Spectrum, 45(3):76, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2008:WGG

[Moo08d] Samuel K. Moore. Water goes off the grid [update]. IEEE Spectrum, 45(11):18, November 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2009:BMM

[Moo09a] Samuel K. Moore. Bailouts for memory makers. IEEE Spectrum, 46(2):14, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2009:FLP

[Moo09b] Samuel K. Moore. Future of lowpower chips. IEEE Spectrum, 46(8):16, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2009:IVC

[Moo09c] Samuel K. Moore. Ingres and VectorWise claim database speedup. IEEE Spectrum, 46 (10):18, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2009:SBS

[Moo09d] Samuel K. Moore. Some bright spots in the gloom. IEEE Spectrum, 46(5):14, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:2000:JRC

[Mor00] S. Morgan. Jini to the rescue [computer network interconnection technology]. IEEE Spectrum, 37(4):44–49, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:2001:CKP

[Mor01a] G. Morgan. The collapse of the Kyoto Protocol and the struggle to slow global warming [books]. IEEE Spectrum, 38(7): 78, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morris:2001:STN

[Mor01b] A. S. Morris. In search of transparent networks. IEEE Spectrum, 38(10):47–51, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morrison:2001:HOP

[Mor01c] C. Morrison. Hitachi, Oracle are partners in protein-mapping megaproject. IEEE Spectrum, 38(6):22–24, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moraes:2006:BPD

[Mor06] S. Moraes. The big picture: Dude, you’ve got mail. Photograph by Sergio Moraesg/ Reuters/Landov. IEEE Spectrum, 43(2):20–21, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morgan:2007:FIR

[Mor07] M. G. Morgan. Fire and ice [review of “What We Know About Climate Change” (Emanuel, K.; 2007)]. IEEE Spectrum, 44(12): 64–66, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2009:NPD

[MS09] Samuel K. Moore and Neil Savage. The Nobel Prize and its discontents. IEEE Spectrum, 46 (12):11–12, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mueller:2002:WIC

[Mue02] M. Mueller. Why ICANN can’t [failings of Internet Corporation for Assigned Names and Numbers]. IEEE Spectrum, 39(7):15– 16, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2001:BPM

[Mul01a] J. Mullins. Britain’s Payne may have key to all-optical routers. IEEE Spectrum, 38(11): 19–21, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2001:TTW

[Mul01b] J. Mullins. The topsy-turvy world of quantum computing. IEEE Spectrum, 38(2): 42–49, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2002:MUC

[Mul02a] J. Mullins. Making unbreakable code. IEEE Spectrum, 39(5):40–45, May 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2002:UUF

[Mul02b] J. Mullins. Using unusable frequencies [solid-state terahertz laser]. IEEE Spectrum, 39(7):22– 23, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2004:BWM

[Mul04a] J. Mullins. A better way of making blue laser diodes? IEEE Spectrum, 41(4):13–14, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2004:CFB

[Mul04b] J. Mullins. Cold fusion back from the dead. IEEE Spectrum, 41(9):22, September 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2005:CSM

[Mul05a] J. Mullins. Could simple mixed materials make flat lenses cheap? IEEE Spectrum, 42(12):19, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2005:LBU

[Mul05b] J. Mullins. London broil? The UK’s capital prepares for rising tides and temperatures. IEEE Spectrum, 42(3):10–12, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2005:SLO

[Mul05c] J. Mullins. Shedding light on organic transistors: insights could lead to brighter and cheaper displays. IEEE Spectrum, 42 (8):15–16, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2006:BFL

[Mul06a] J. Mullins. Ben Franklin’s London house opens as a landmark. IEEE Spectrum, 43(3):16, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2006:BEL

[Mul06b] J. Mullins. Butterfly effect [light emitting diodes]. IEEE Spectrum, 43(2):14, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2006:CEI

[Mul06c] J. Mullins. The case of Earth’s incredible shrinking field. IEEE Spectrum, 43(11):22, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2006:CC

[Mul06d] J. Mullins. Chaotic communication. IEEE Spectrum, 43 (1):11–12, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2006:RSI

[Mul06e] J. Mullins. Ring of steel II — New York City gets set to replicate London’s high-security zone. IEEE Spectrum, 43(7):12– 13, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2007:LCB

[Mul07] J. Mullins. LEDs could borrow from beetles. IEEE Spectrum, 44(6):18–20, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:2008:CM

[Mul08] J. Mullins. The church of microsoft. IEEE Spectrum, 45 (3):11–12, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:2007:SLH

[Mur07]  K. Murphy.  Spectral lines

— happy birthday, fairchild. IEEE Spectrum, 44(12):7, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murphy:2009:WPR

[Mur09] K. Murphy. World power (review of The Power Makers: Steam, Electricity, and the Men Who Invented Modern America by M. Klein; 2008) [books]. IEEE Spectrum, 46(8):23, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Musk:2009:RB

[Mus09] E. Musk. Risky business. IEEE Spectrum, 46(6):40–41, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nahin:2001:TTE

[Nah01] Paul J. Nahin. Time travel in Einstein’s universe, J. Richart Gott, Houghton Mifflin, New York, 2001. ISBN: 0-395-95563-7, 291 pp., $25 [books]. IEEE Spectrum, 38(7):79–80, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nakao:2004:PET

[Nak04] Y. Nakao. Photo essay: Taking Tokyo by swarm. IEEE Spectrum, 41(6):26–27, June 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nucci:2007:CCI

[NB07] A. Nucci and S. Bannerman.

Controlled chaos [Internet security]. IEEE Spectrum, 44(12): 42–48, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nebeker:2000:EC

[Neb00] F. Nebeker. The electric century.

IEEE Spectrum, 37(6):68–74, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nebeker:2003:MTT

[Neb03a] F. Nebeker. More treasured texts [educational textbooks]. IEEE Spectrum, 40(7):31–36, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nebeker:2003:TT

[Neb03b] F. Nebeker. Treasured texts. IEEE Spectrum, 40(4):44–49, April 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nevens:2001:FLD

[Nev01] T. M. Nevens. Fast lines at digital high. IEEE Spectrum, 38(4):16–17, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nguyen:2009:MR

[Ngu09] Clark T.-C. Nguyen. Mechanical radio. IEEE Spectrum, 46 (12):30–35, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Numagai:2005:CTR

[NH05] J. Numagai and M. Hood. China’s tech revolution: how technology is driving the country’s economic boom and what that means for the World. IEEE Spectrum, 42(6):28–35, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nordmann:2008:SS

[Nor08] A. Nordmann. Singular simplicity. IEEE Spectrum, 45(6):60–63, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2000:NBP

[Obe00a] J. Oberg. NASA’s big push for the Space Station. IEEE Spectrum, 37(11):49–54, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2000:RSI

[Obe00b] J. Oberg. Russia’s sorry infrastructure. IEEE Spectrum, 37 (12):47–51, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2000:SMR

[Obe00c] J. Oberg. Saving Mir with a rope trick. IEEE Spectrum, 37(7):32– 37, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2001:TPL

[Obe01a] J. Oberg. Taikonauts prepare for liftoff [chinese manned space programme]. IEEE Spectrum, 38 (12):42–47, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2001:YRC

[Obe01b] J. Oberg. Year of the rocket [chinese space research]. IEEE Spectrum, 38(5):62–68, May 2001.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2002:PCF

[Obe02a] J. Oberg. Paperboy: confessions of a future engineer [book review]. IEEE Spectrum, 39(8):51, August 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2002:RFA

[Obe02b] J. Oberg. A rocket family album [book review]. IEEE Spectrum, 39(2):59, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2003:SSB

[Obe03a] J. Oberg. For spacious skies: [book review]. IEEE Spectrum, 40(2):49–50, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2003:SPS

[Obe03b] J. Oberg. The Shuttle puzzle [Space Suttle Columbia disaster]. IEEE Spectrum, 40 (3):22–24, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2003:SMS

[Obe03c] J. Oberg. Soyuz mishap sounds alarms. IEEE Spectrum, 40 (6):16, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2004:GLC

[Obe04] J. Oberg. A giant leap for commercial space travel. IEEE Spectrum, 41(8):10–12, August 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2005:WCC

[Obe05] J. Oberg. The watchman: as commander of Cheyenne Mountain Operations Center, Duane Deal guards the skies. IEEE Spectrum, 42(6):69–70, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2006:CSE

[Obe06] J. Oberg. Careers: Stellar engineer. IEEE Spectrum, 43 (4):61–64, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2007:LSR

[Obe07a] J. Oberg. Lost in space [a review of ISScapades: The Crippling of America’s Space Program (Beattie, D. A.)]. IEEE Spectrum, 44(9):64, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2007:NE

[Obe07b] J. Oberg. No exit. IEEE Spectrum, 44(9):22, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2008:CNM

[Obe08] J. Oberg. Copying NASA’s mistakes (a review of Energiya-Buran: The Soviet Space Shuttle by B. Hendricks and B. Vis; 2007) [books]. IEEE Spectrum, 45(3):22, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2009:CCS

[Obe09a] J. Oberg. Commercial communications satellites for the moon. IEEE Spectrum, 46(2):11, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2009:CCG

[Obe09b] J. Oberg. Could China get to Mars first? IEEE Spectrum, 46(6):50–51, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

OConnor:2001:NTY

[O’C01] P. D. T. O’Connor. Neglect testing at your peril. IEEE Spectrum, 38(7):18, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

ODonovan:2006:GC

[O’D06]  P. O’Donovan.  Goodbye,

crt. IEEE Spectrum, 43(11): 38–43, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

OKeefe:2002:NMP

[O’K02] S. O’Keefe. NASA’s master plan. IEEE Spectrum, 39(11):

11–12, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliphant:2000:RIM

[Oli00] M. W. Oliphant. Radio interfaces make the difference in 3G cellular systems. IEEE Spectrum, 37 (10):53–58, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliphant:2002:TGW

[Oli02]  M. W. Oliphant.  Third-

generation wireless must wait for services to catch up. IEEE Spectrum, 39(6):14, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

ONeill:2003:MBR

[O’N03] A. O’Neill. Microchip [book review]. IEEE Spectrum, 40 (6):40–41, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ortiz:2009:HSF

[Ort09] R. Ortiz. How software found the Air France wreckage. IEEE Spectrum, 46(8):18, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Overbye:2001:VEG

[OW01] T. J. Overbye and J. D. Weber. Visualizing the electric grid. IEEE Spectrum, 38(2): 52–58, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oberg:2003:SS

[OZC03] J. Oberg, A. Zak, and Stephen Cass. Saving the station. IEEE Spectrum, 40(10):20–27, October 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

P:2008:RHM

[P.08]  T. S. P.  Roger Hill —

marches with penguins [dream jobs 2008]. IEEE Spectrum, 45 (2):37–39, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:MNH

[PA02] T. C. Perry and C. R. Abraham. Money for nothing [hanford nuclear waste reservation]. IEEE Spectrum, 39(7):12–14, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parloff:2002:MAV

[Par02a] R. Parloff. Microsoft’s anticlimactic victory. IEEE Spectrum, 39(12):17–18, December 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parloff:2002:WDM

[Par02b] R. Parloff. What a difference 18 months make. IEEE Spectrum, 39(9):24–26, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parloff:2002:WMV

[Par02c] R. Parloff. Will Microsoft verdict miss mark? IEEE Spectrum, 39(6):26–28, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parloff:2003:ECT

[Par03a] R. Parloff. Executioner’s current: Thomas Edison, George Westinghouse, and the invention of the electric chair [book review]. IEEE Spectrum, 40(3):56, March 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parloff:2003:MF

[Par03b] R. Parloff. Morpheus falling?

IEEE Spectrum, 40(12):18–19, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parloff:2003:TWR

[Par03c] R. Parloff. Technical win for Rambus in patents case. IEEE Spectrum, 40(4):22–24, April 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parviz:2009:YEO

[Par09] B. A. Parviz. For your eye only. IEEE Spectrum, 46(9): 36–41, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pascoe:2001:BNF

[Pas01] R. Pascoe. Building networks on the fly [cellular radio]. IEEE Spectrum, 38(3):61–65, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:ANG

[Pat09a] P. Patel. Advance in nanopore gene sequencing. IEEE Spectrum, 46(6):14, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:AJF

[Pat09b] P. Patel. Aerospace job forecast: skies are clearing. IEEE Spectrum, 46(11):22, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:BMI

[Pat09c] P. Patel. The brain-machine interface, unplugged. IEEE Spectrum, 46(10):13–14, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:D

[Pat09d] P. Patel. The data. IEEE Spectrum, 46(12):68, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:DSM

[Pat09e]  P. Patel.  Despite stimulus

money, most U.S. bridges might stay dumb. IEEE Spectrum, 46(4):11–12, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:EMV

[Pat09f] P. Patel. Engineers map volcanic lightning. IEEE Spectrum, 46(4):15, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:EER

[Pat09g] P. Patel. Experts expect résumé fraud to rise. IEEE Spectrum, 46(6):24, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:ORS

[Pat09h]  P. Patel.  It’s only rocket

science. IEEE Spectrum, 46 (6):58–61, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:MQW

[Pat09i] P. Patel. Math quiz: why do men predominate. IEEE Spectrum, 46 (12):22, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:RFQ

[Pat09j] P. Patel. The rise and fall of the quants. IEEE Spectrum, 46 (8):20–21, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel:2009:WCT

[Pat09k] P. Patel. The well-connected traveler. IEEE Spectrum, 46 (10):22, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pavey:2007:RTC

[Pav07] E. Pavey. Ready for their close-ups. IEEE Spectrum, 44 (4):42–47, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:SD

[PC05] P. P. Predd and Stephen Cass. Showdown on the desktop. IEEE Spectrum, 42(4):19–20, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peck:2008:USB

[Pec08] Morgan E. Peck. Update: Standardizing the brain-machine interface. IEEE Spectrum, 45 (4):16, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peha:2001:BDB

[Peh01a] J. M. Peha. Bridging the divide between technologists and policy-makers. IEEE Spectrum, 38(3):15–19, March 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peha:2001:CNN

[Peh01b] J. M. Peha. Congress needs nonpartisan advice on science, technology [speakout]. IEEE Spectrum, 38(9):19–20, September 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2000:CET

[Per00a] Tekla S. Perry. Consumer electronics [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1):51–56, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2000:ETA

[Per00b] Tekla S. Perry. The environment [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1):81–85, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2000:JRB

[Per00c] Tekla S. Perry. Jon rubinstein [biography]. IEEE Spectrum, 37 (2):52–57, February 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2000:PSK

[Per00d] Tekla S. Perry. Profile: Steve Kirch. IEEE Spectrum, 37 (8):53–57, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2000:TWF

[Per00e]  Tekla S. Perry.  Tracking

weather’s flight path. IEEE Spectrum, 37(9):38–45, September 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2000:WC

[Per00f] Tekla S. Perry. Watching the clocks. IEEE Spectrum, 37(12):

56–60, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2001:OGL

[Per01a]  Tekla S. Perry.  And the

Oscar goes to … [computer graphics]. IEEE Spectrum, 38 (4):42–49, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2001:FIC

[Per01b] Tekla S. Perry. Fueling the Internet [computer power consumption]. IEEE Spectrum, 38 (1):80–82, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2001:NPG

[Per01c] Tekla S. Perry. Nuclear power gets a second look. IEEE Spectrum, 38(11):32–33, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2001:OND

[Per01d] Tekla S. Perry. Operating in neutral [digital music legislation]. IEEE Spectrum, 38(10): 52–57, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2001:PST

[Per01e] Tekla S. Perry. Planting seeds of technology [business start-up management]. IEEE Spectrum, 38(7):68–73, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2001:STN

[Per01f] Tekla S. Perry. Service takes over in the networked world. IEEE Spectrum, 38(1):102–106, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:CCC

[Per02a] Tekla S. Perry. Capturing climate change. IEEE Spectrum, 39(1):58–65, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:JBS

[Per02b] Tekla S. Perry. Not just blue sky [herbert kroemer]. IEEE Spectrum, 39(6):32–37, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:RM

[Per02c] Tekla S. Perry. RISC maker.

IEEE Spectrum, 39(11):33–37, November 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:SAR

[Per02d] Tekla S. Perry. State-of-the-art recycling plant. IEEE Spectrum, 39(12):19, December 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:VNV

[Per02e] Tekla S. Perry. Vying for next video player crown. IEEE Spectrum, 39(6):19–20, June 2002.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2003:AG

[Per03a] Tekla S. Perry. All in the game. IEEE Spectrum, 40(11): 31–35, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2003:CW

[Per03b] Tekla S. Perry. The copyright wars. IEEE Spectrum, 40(5):20–23, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2003:RHL

[Per03c] Tekla S. Perry. Red hot [light emitting diodes]. IEEE Spectrum, 40(6):26–29, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:BFS

[Per04a] Tekla S. Perry. Building a in a flash supercomputer. IEEE Spectrum, 41(6):24–25, June 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:DFH

[Per04b] Tekla S. Perry. Dave Fruehling: His job rocks! IEEE Spectrum, 41(2):40–41, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:DST

[Per04c]  Tekla S. Perry.  Dennis

Shelden: turning dreams into reality. IEEE Spectrum, 41(2): 39–40, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:DJ

[Per04d] Tekla S. Perry. Dream jobs 2004.

IEEE Spectrum, 41(2):28, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:GMH

[Per04e] Tekla S. Perry. General motors on the hy-wire. IEEE Spectrum, 41(1):64–65, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:JGN

[Per04f] Tekla S. Perry. John Gage: he is the network. IEEE Spectrum, 41(2):32–33, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:MC

[Per04g] Tekla S. Perry. Master of convergence. IEEE Spectrum, 41 (6):34–37, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:SSS

[Per04h] Tekla S. Perry. Steve sullivan:

shaping the future of the movies [special effects]. IEEE Spectrum, 41(2):29–30, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:TTG

[Per04i] Tekla S. Perry. Tomorrow’s TV — the grating light valve. IEEE Spectrum, 41(4):38–41, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:TTS

[Per04j] Tekla S. Perry. Tools toys: The solution was sake. IEEE Spectrum, 41(4):51–52, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:BWI

[Per05a] Tekla S. Perry. Bill Woodcock: on an Internet odyssey. IEEE Spectrum, 42(2):26–28, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:DGM

[Per05b] Tekla S. Perry. Dale Gardener: The modern mariner. IEEE Spectrum, 42(2):30–31, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:DTM

[Per05c] Tekla S. Perry. Digital TV’s 100meter dash: China’s huge TV industry faces a 2008 deadline. IEEE Spectrum, 42(6):46–51, June 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:DJ

[Per05d] Tekla S. Perry. Dream jobs 2005. IEEE Spectrum, 42(2): 20–21, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:DCP

[Per05e] Tekla S. Perry. DVD copy protection: take 2. IEEE Spectrum, 42(1):38–39, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:HTF

[Per05f]  Tekla S. Perry.  High-tech

faberge. IEEE Spectrum, 42(7): 11, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:HWH

[Per05g] Tekla S. Perry. Hurricane watchers hit their mark [hurricane forecasting]. IEEE Spectrum, 42 (12):14–15, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:ICP

[Per05h] Tekla S. Perry. Ian caven: he had to be in pictures [image processing techniques]. IEEE Spectrum, 42(2):24–25, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:MMU

[Per05i] Tekla S. Perry. Of modes men [user interfaces]. IEEE Spectrum, 42(8):48–53, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:RBB

[Per05j] Tekla S. Perry. Recycling behind bars: prison practices blacken a green endeavor. IEEE Spectrum, 42(6):10–13, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:SFD

[Per05k] Tekla S. Perry. The sky is falling [digital theaters]. IEEE Spectrum, 42(11):14, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:SSS

[Per05l] Tekla S. Perry. Sultan of sound [speech technology]. IEEE Spectrum, 42(5):44–48, May 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:TTB

[Per05m] Tekla S. Perry. Tools toys: A better light bulb? IEEE Spectrum, 42(2):49, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:TTA

[Per05n] Tekla S. Perry. Tools toys: Akimbo in limbo. IEEE Spectrum, 42(3):65–66, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:TTH

[Per05o] Tekla S. Perry. Tools toys:

Home automation for the rest of us. IEEE Spectrum, 42(9): 68–72, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2005:TGW

[Per05p] Tekla S. Perry. Tracking a great white [electronic tracking tag]. IEEE Spectrum, 42 (8):20–21, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perrow:2006:ST

[Per06a] C. Perrow. Shrink the targets. IEEE Spectrum, 43(9): 46–49, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2006:BWR

[Per06b] Tekla S. Perry. Black, white and readable [nanochromics display]. IEEE Spectrum, 43(1): 38–41, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2006:DWD

[Per06c] Tekla S. Perry. Doing well by doing good. IEEE Spectrum, 43 (12):30–35, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2006:MGR

[Per06d] Tekla S. Perry. Media: Geek rhythms. IEEE Spectrum, 43 (9):57–59, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2006:WPW

[Per06e] Tekla S. Perry. Who pays for e-waste? [electronic waste recycling]. IEEE Spectrum, 43(7):14– 15, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2006:WWE

[Per06f] Tekla S. Perry. Wizard of watts [electronics engineering education]. IEEE Spectrum, 43 (6):32–37, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2007:C

[Per07a] Tekla S. Perry. The codemaker. IEEE Spectrum, 44 (10):36–41, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2007:CIT

[Per07b] Tekla S. Perry. Commentary: Imagine there’s no DRM … I wonder if you can [news]. IEEE Spectrum, 44(7):14–18, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2007:LC

[Per07c] Tekla S. Perry. The laptop crusade. IEEE Spectrum, 44 (4):28–33, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2007:NZZ

[Per07d] Tekla S. Perry. News: The zero– zero hero. IEEE Spectrum, 44 (9):12–14, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2007:SYL

[Per07e] Tekla S. Perry. Sizing up your LCD TV. IEEE Spectrum, 44 (6):75–77, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2007:TK

[Per07f] Tekla S. Perry. Thomas Kailath.

IEEE Spectrum, 44(5):44–47, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:DDM

[Per08a] Tekla S. Perry. David Downey — Marathon Man [dream jobs 2008]. IEEE Spectrum, 45(2): 33–35, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:DDS

[Per08b] Tekla S. Perry. Digital dilemma — [spectral lines]. IEEE Spectrum, 45(10):9, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:PPP

[Per08c] Tekla S. Perry. From podfather to Palm’s Pilot. IEEE Spectrum, 45(9):48–58, September 2008. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:GMN

[Per08d] Tekla S. Perry. Gordon Moore’s next act. IEEE Spectrum, 45 (5):40–43, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:LLI

[Per08e] Tekla S. Perry. The lady and the li-ion. IEEE Spectrum, 45 (3):30–36, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:LCS

[Per08f] Tekla S. Perry. Loser: Consumer — satellite TV soccer moms will hate. IEEE Spectrum, 45(1):43–44, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:SLA

[Per08g] Tekla S. Perry. Spectral lines: Airlines: got fuel? IEEE Spectrum, 45(8):9, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:TTL

[Per08h] Tekla S. Perry. Tools toys — let ’er rip. IEEE Spectrum, 45(7): 22, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:TTP

[Per08i] Tekla S. Perry. Tools toys: Pleo, the poop-free pet. IEEE Spectrum, 45(4):24, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2008:UAM

[Per08j] Tekla S. Perry. Update: Ausra makes solar thermal simple and cheap. IEEE Spectrum, 45(5):15, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:BHC

[Per09a] Tekla S. Perry. Big house, cheap audio. IEEE Spectrum, 46(7):26, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:F

[Per09b] Tekla S. Perry. Fruitless. IEEE Spectrum, 46(1):55–58, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:GBG

[Per09c] Tekla S. Perry. Go big or go home. IEEE Spectrum, 46(12): 42–46, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:IMM

[Per09d] Tekla S. Perry. The iphone’s music man. IEEE Spectrum, 46 (9):24, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:KKZ

[Per09e] Tekla S. Perry. Kenyon kluge zero to 60. IEEE Spectrum, 46 (2):28–29, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:KCG

[Per09f] Tekla S. Perry. The kindlers are coming to get me. IEEE Spectrum, 46(8):8, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:LMA

[Per09g] Tekla S. Perry. Little mass appeal for Intel’s mobile Internet device. IEEE Spectrum, 46(7): 15, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:SBT

[Per09h]  Tekla S. Perry.  Shannon

Bruzelius toy story. IEEE Spectrum, 46(2):26–27, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2009:ZMF

[Per09i] Tekla S. Perry. Zink: A modern fairy tale. IEEE Spectrum, 46 (11):24–56, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pescovitz:2002:HCH

[Pes02] D. Pescovitz. Helping computers help themselves. IEEE Spectrum, 39(9):49–53, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pettis:2009:BAB

[Pet09] B. Pettis. Barbot, the automated bartender. IEEE Spectrum, 46 (12):20–21, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pfleeger:2000:ASD

[Pfl00] S. L. Pfleeger. Adaptive software development [books]. IEEE Spectrum, 37(5):10–14, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pierse:2005:PER

[Pie05] R. Pierse. Photo essay: Rocket man. IEEE Spectrum, 42(11): 22–23, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Panicia:2005:SSS

[PK05] M. Panicia and S. Koehl. The silicon solution [silicon photonics]. IEEE Spectrum, 42(10): 38–43, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pollack:2006:DDS

[Pol06] J. Pollack. Displays of a different stripe. IEEE Spectrum, 43(8):40–44, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poston:2000:AGO

[Pos00] J. W. Poston. The angry genie: one man’s walk through the nuclear age [books]. IEEE Spectrum, 37(2):12–14, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2005:RLE

[PP05a] P. Patel-Predd. Resources — a league of extraordinary women [engineering profession]. IEEE Spectrum, 42(10):59–61, October 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predo:2005:JSC

[PP05b] P. Patel-Predo. Juice from solar concentrate [photovoltaic collector]. IEEE Spectrum, 42 (10):15–16, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2006:CWP

[PP06a] P. Patel-Predd. Careers: What’s up, PostDoc? IEEE Spectrum, 43(9):60–65, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predo:2006:TLM

[PP06b] P. Patel-Predo. Traveling light — micro fuel cells could give soldiers less weight to carry. IEEE Spectrum, 43(7):20–21, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2007:CE

[PP07a] P. Patel-Predd. Careers: The enabler. IEEE Spectrum, 44 (1):51–54, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2007:RNW

[PP07b] P. Patel-Predd. Resources: From nerd to wonk. IEEE Spectrum, 44(10):57–60, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:BCH

[PP08a] P. Patel-Predd. A batterycapacitor hybrid for hybrids. IEEE Spectrum, 45(12):15, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:BCC

[PP08b] P. Patel-Predd. The biggest career choice of all is when to start [careers]. IEEE Spectrum, 45 (9):21, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:CUE

[PP08c] P. Patel-Predd. Careers: U.S.

engineering salaries rise sharply. IEEE Spectrum, 45(8):21, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:DBD

[PP08d] P. Patel-Predd. Data breaches [the data]. IEEE Spectrum, 45 (9):72, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:EDW

[PP08e] P. Patel-Predd. Engineers are doing well by doing good. IEEE Spectrum, 45(8):64, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:UCN

[PP08f]  P. Patel-Predd.  Update:

Carbon-nanotube wiring gets real. IEEE Spectrum, 45(4):14, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:UEE

[PP08g] P. Patel-Predd. Update: Energyefficient ethernet. IEEE Spectrum, 45(5):13, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2008:UTS

[PP08h] P. Patel-Predd. Update: Transistors in space. IEEE Spectrum, 45(8):17, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2009:ASE

[PP09a] P. Patel-Predd. The all-seeing employer — [careers]. IEEE Spectrum, 46(1):23, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2009:CS

[PP09b] P. Patel-Predd. Cellphones for science. IEEE Spectrum, 46 (2):16, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2009:GCD

[PP09c] P. Patel-Predd. Generation cogeneration [the data]. IEEE Spectrum, 46(3):68, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patel-Predd:2009:TTS

[PP09d]  P. Patel-Predd.  The trou-

ble with touch screens — [update]. IEEE Spectrum, 46(1): 11–12, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Prather:2001:MGF

[Pra01]  D. W. Prather.  Mathcad

gets its feet webbed [software]. IEEE Spectrum, 38(7):81–82, July 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2004:CLB

[Pre04] P. P. Predd. Cellphones linked to brain tumors. IEEE Spectrum, 41 (12):20–22, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:CHB

[Pre05a] P. P. Predd. Cheaper hydrogen beckons [hydrogen energy]. IEEE Spectrum, 42(3):16, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:GEE

[Pre05b] P. P. Predd. The gay engineer [equal opportunities]. IEEE Spectrum, 42(2):47–48, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:HSD

[Pre05c] P. P. Predd. How a strand of DNA launched a career. IEEE Spectrum, 42(4):57–58, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:HPY

[Pre05d]  P. P. Predd.  How PC is

your PC? [personal computer lifespan]. IEEE Spectrum, 42 (1):11, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:PPS

[Pre05e] P. P. Predd. Perils of plutonium [spent nuclear fuel storage]. IEEE Spectrum, 42(7):16– 17, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:PWC

[Pre05f] P. P. Predd. The price is wrong [consumer electronics]. IEEE Spectrum, 42(3):50–51, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2005:SCF

[Pre05g] P. P. Predd. Stealing celestial fire [artificial lightning]. IEEE Spectrum, 42(4):21, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2007:BB

[Pre07a] P. P. Predd. Beyond blue. IEEE Spectrum, 44(3):14, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Predd:2007:PPH

[Pre07b] P. P. Predd. A power plant for the home. IEEE Spectrum, 44(4):14–15, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Prevelakis:2007:AA

[PS07] V. Prevelakis and D. Spinellis.

The Athens Affair. IEEE Spectrum, 44(7):26–33, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pecht:2006:BEM

[PT06] M. Pecht and S. Tiku. Bogus: electronic manufacturing and consumers confront a rising tide of counterfeit electronics. IEEE Spectrum, 43(5):37–46, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pustai:2000:SR

[Pus00] D. Pustai. Software reviews.

IEEE Spectrum, 37(4):21–22, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Putzeys:2003:DAF

[Put03] B. Putzeys. Digital audio’s final frontier. IEEE Spectrum, 40(3):34–41, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Qu:2005:NSB

[QP05] P. Qu and C. Polley. The new standard-bearer. IEEE Spectrum, 42(12):48–52, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ragsdale:2009:DSV

[Rag09] R. D. Ragsdale. DIY streetview camera. IEEE Spectrum, 46 (10):20–21, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rajlich:2008:GKR

[Raj08] V. Rajlich. Game of kings (review of King’s Gambit: A Son, a Father, and the World’s Most Dangerous Game by P. Hoffman; 2007) [books]. IEEE Spectrum, 45(2):26, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ramana:2000:INB

[Ram00] M. V. Ramana. India’s Nuclear Bomb: The Impact on Global Proliferation [books]. IEEE Spectrum, 37(3):14–17, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ramana:2007:MMT

[Ram07] M. V. Ramana. More missiles than megawatts. IEEE Spectrum, 44(7):34–39, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rast:2005:DDT

[Ras05] R. M. Rast. The dawn of digital TV. IEEE Spectrum, 42 (10):26–31, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Raymond:2001:WMS

[Ray01] E. S. Raymond. Why Microsoft smears — and fears — open source. IEEE Spectrum, 38 (8):14–15, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Razavi:2008:GGG

[Raz08] D. Razavi. Gadgets gab at 60 ghz. IEEE Spectrum, 45 (2):46–58, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTh

[RB00a] M. J. Riezenman and E. A.

Bretz. EEs’ tools toys. IEEE Spectrum, 37(11):96–97, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTi

[RB00b] M. J. Riezenman and E. A. Bretz. EEs’ tools toys. IEEE Spectrum, 37(12):89–93, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTc

[RBJ00a] M. J. Riezenman, E. A. Bretz, and W. D. Jones. EEs’ tools toys. IEEE Spectrum, 37(6):88–91, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTd

[RBJ00b] M. J. Riezenman, E. A. Bretz, and W. D. Jones. EEs’ tools toys. IEEE Spectrum, 37(7):76– 81, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reese:2003:ELD

[Ree03] R. A. Reese. Extreme lawsuits [digital copyright]. IEEE Spectrum, 40(5):23–25, May 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reed:2009:FOO

[Ree09] D. P. Reed. Fouling our own net [review of “Future of the Internet and How to Stop It” (Zittrain, J.; 2008)]. IEEE Spectrum, 46 (2):18–19, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reis:2002:CC

[Rei02] R. Reis. Critical challenges 2002.

IEEE Spectrum, 39(1):9, January 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosen:2006:DD

[RH06] J. Rosen and B. Hannaford. Doc at a distance. IEEE Spectrum, 43 (10):34–39, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Richards:2007:BPC

[Ric07] M. Richards. The big picture:

Core memories. IEEE Spectrum, 44(7):24–25, July 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:CSB

[Rie00a] M. J. Riezenman. Cellular security: better, but foes still lurk. IEEE Spectrum, 37(6):39–42, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:CTA

[Rie00b] M. J. Riezenman. Communications [technology 2000 analysis and forecast]. IEEE Spectrum, 37(1):33–39, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTa

[Rie00c] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 37(4): 61–62, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTb

[Rie00d] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 37(5): 73–75, May 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTe

[Rie00e] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 37(8): 74–75, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTf

[Rie00f] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 37(9): 90–92, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:ETTg

[Rie00g] M. J. Riezenman. EEs’ tools toys. IEEE Spectrum, 37(10): 75–78, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:EW

[Rie00h] M. J. Riezenman. EV watch. IEEE Spectrum, 37(3):25–26, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:FF

[Rie00i] M. J. Riezenman. Faults failures. IEEE Spectrum, 37(5):23–24, May 2000. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:IQR

[Rie00j] M. J. Riezenman. Instrumentation: the quiet revolution. IEEE Spectrum, 37(2):49–51, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:TWa

[Rie00k] M. J. Riezenman. Technology watch. IEEE Spectrum, 37(6):24–25, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2000:TWb

[Rie00l] M. J. Riezenman. Technology watch. IEEE Spectrum, 37 (11):32, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:BIE

[Rie01a] M. J. Riezenman. Beneath the Internet: explosive growth drives improvements to the infrastructure. IEEE Spectrum, 38(1):54, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:DBR

[Rie01b] M. J. Riezenman. The disposable battery [review — power to the phones]. IEEE Spectrum, 38 (8):68–69, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:FCL

[Rie01c] M. J. Riezenman. Fuel cells for the long haul, batteries for the spurts [electric vehicles]. IEEE Spectrum, 38(1): 95–97, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:HUT

[Rie01d] M. J. Riezenman. Handheld unit tests network wiring and performance. IEEE Spectrum, 38 (12):70–71, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:MFC

[Rie01e] M. J. Riezenman. Metal fuel cells [zn-air fuel cells]. IEEE Spectrum, 38(6):55–59, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:ONB

[Rie01f]  M. J. Riezenman.  Optical

nets brace for even heavier traffic. IEEE Spectrum, 38(1): 44–46, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:RR

[Rie01g] M. J. Riezenman. The rebirth of radio. IEEE Spectrum, 38 (1):62–64, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:TTR

[Rie01h] M. J. Riezenman. Tools toys [resources]. IEEE Spectrum, 38 (4):88–91, April 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2001:UUB

[Rie01i] M. J. Riezenman. The universe unveiled [book review]. IEEE Spectrum, 38(5):92–93, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riebeek:2002:BHA

[Rie02a]  H. Riebeek.  Brazil holds

all-electronic national election. IEEE Spectrum, 39(11):25–26, November 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2002:ERS

[Rie02b] M. J. Riezenman. Over and over and out [emergency radio systems]. IEEE Spectrum, 39(9): 41–42, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2002:RSO

[Rie02c] M. J. Riezenman. Robots stand on own two feet. IEEE Spectrum, 39(8):24–25, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riebeck:2003:SCC

[Rie03a] H. Riebeck. A second coffin for chernobyl. IEEE Spectrum, 40(3):30–31, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riebeek:2003:BTW

[Rie03b] H. Riebeek. Brazil tests world’s largest environmental monitoring system. IEEE Spectrum, 40 (9):10–12, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riebeek:2003:KRJ

[Rie03c]  H. Riebeek.  Key research

journals embrace self-censorship. IEEE Spectrum, 40(4):17, April 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riebeek:2003:TC

[Rie03d]  H. Riebeek.  The ticket

chase. IEEE Spectrum, 40(1): 72–73, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2003:EBR

[Rie03e] M. J. Riezenman. Extending broadband’s reach. IEEE Spectrum, 40(3):20–21, March 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riezenman:2003:MMF

[Rie03f]  M. J. Riezenman.  Mighty

mites [fuel cell power supplies]. IEEE Spectrum, 40(6):30–33, June 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2002:DAR

[Rio02] Michael Riordan. To the digital age: research labs, start-up companies, and the rise of MOS technology [book review]. IEEE Spectrum, 39(11):53–54, November 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2004:LHT

[Rio04] Michael Riordan. The lost history of the transistor. IEEE Spectrum, 41(5):44–49, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2005:EAM

[Rio05a] Michael Riordan. The end of AT&T: Ma Bell may be gone, but its innovations are everywhere. IEEE Spectrum, 42 (7):46–51, July 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2005:HEM

[Rio05b] Michael Riordan. How Europe missed the transistor. IEEE Spectrum, 42(11):52–57, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2006:BMW

[Rio06a] Michael Riordan. Books: The men who made the microchip. IEEE Spectrum, 43(4):64–68, April 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2006:HBL

[Rio06b] Michael Riordan. How Bell Labs missed the microchip. IEEE Spectrum, 43(12):36–41, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2007:BTN

[Rio07a] Michael Riordan. Books: Tales of Nakamura [review of Brilliant! Shuji Nakamura and the Revolution in Lighting Technology (Johnstone, B.; 2007)]. IEEE Spectrum, 44(5):56–58, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2007:SDS

[Rio07b] Michael Riordan. The silicon dioxide solution. IEEE Spectrum, 44(12):51–56, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riebeek:2003:BUM

[RM03] H. Riebeek and Samuel K. Moore. Building unable matter. IEEE Spectrum, 40(8):16–17, August 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robb:2000:DRS

[Rob00] D. Robb. Defragmenting really speeds up Windows NT machines. IEEE Spectrum, 37(9): 74–77, September 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robinson:2003:WGT

[Rob03a] B. T. Robinson. Who goes there? IEEE Spectrum, 40 (10):42–47, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robinson:2003:WWE

[Rob03b] M. Robinson. Watts on way for Ethernet. IEEE Spectrum, 40(3):21, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roberts:2009:RNR

[Rob09a] L. G. Roberts. A radical new router. IEEE Spectrum, 46 (7):34–39, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robinson:2009:MFM

[Rob09b] K. S. Robinson. My 10 favorite Mars novels. IEEE Spectrum, 46(6):20–21, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rodriguez:2003:UFI

[Rod03] R. Rodriguez. Unite or face irrelevance. IEEE Spectrum, 40(7):9– 10, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Romero:2007:CNT

[Rom07] J. J. Romero. Carbon nanotubes take the heat off chips [news]. IEEE Spectrum, 44(12): 12–14, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Romero:2008:SS

[Rom08] J. J. Romero. Searching the sky. IEEE Spectrum, 45(1): 13–14, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Romero:2009:MRU

[Rom09a] J. J. Romero. Mars for the rest of us. IEEE Spectrum, 46(6):64–78, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Romero:2009:SMM

[Rom09b] J. J. Romero. A stowaway mission to the moon. IEEE Spectrum, 46(3):15, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roos:2000:MET

[Roo00] M. Roos. Medical electronics [technology 2000 analysis and forecast]. IEEE Spectrum, 37 (1):110–115, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:2000:CBE

[Ros00] A. Rosenblatt. The code book: the evolution of secrecy from Mary Queen of Scots to quantum cryptography [books]. IEEE Spectrum, 37(10):10–14, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:2001:BSB

[Ros01a] A. Rosenblatt. Body of secrets [book review]. IEEE Spectrum, 38(11):65, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:2001:IDL

[Ros01b] A. Rosenblatt. Inviting disaster lessons from the edge of technology [book review]. IEEE Spectrum, 38(11):65, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:2002:RAV

[Ros02] A. Rosenblatt. Riding on air in Virginia [Maglev train]. IEEE Spectrum, 39(10):20–21, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2003:TTC

[Ros03a] P. Ross. Top 10 techno-cool cars. IEEE Spectrum, 40(2): 30–35, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2003:CTL

[Ros03b] P. E. Ross. 5 commandments [technology laws and rules of thumb]. IEEE Spectrum, 40(12): 30–35, December 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2003:SSM

[Ros03c] P. E. Ross. Silicon shows its mettle [computer chess]. IEEE Spectrum, 40(3):24–26, March 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2003:SYM

[Ros03d] P. E. Ross. Start your motors! IEEE Spectrum, 40 (10):37–41, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenblatt:2004:ATE

[Ros04a] A. Rosenblatt. The attractions of technology [engineering/technology survey — professionals and students]. IEEE Spectrum, 41(2):22–25, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2004:TCN

[Ros04b] P. E. Ross. 10 tech companies for the next 10 years. IEEE Spectrum, 41(11):52, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2004:CCC

[Ros04c] P. E. Ross. Carbon-crunching critters. IEEE Spectrum, 41 (1):79, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2004:MCT

[Ros04d] P. E. Ross. Managing care through the air [remote health monitoring]. IEEE Spectrum, 41 (12):26–31, December 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2004:PCC

[Ros04e]  P. E. Ross.  Psyching out

computer chess players. IEEE Spectrum, 41(2):14–15, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2004:YBO

[Ros04f] P. E. Ross. This is your brain online. IEEE Spectrum, 41 (1):76–78, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2005:ESB

[Ros05a] P. E. Ross. The exterminators [software bugs]. IEEE Spectrum, 42(9):36–41, September 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2005:PNE

[Ros05b] P. E. Ross. Passport to nowhere [electronic passport with biometric information]. IEEE Spectrum, 42(1):54–55, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2005:WWN

[Ros05c] P. E. Ross. Waiting and waiting for the next killer wave [tsunami early warning system]. IEEE Spectrum, 42(3):17, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2006:NS

[Ros06a] M. Ross. The new search for E.T. IEEE Spectrum, 43(11): 32–37, November 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2006:HL

[Ros06b] P. E. Ross. The 8-hour laptop. IEEE Spectrum, 43(1):52, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2006:BPL

[Ros06c] P. E. Ross. Broadband over power liens. IEEE Spectrum, 43(1):54, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2006:MSG

[Ros06d] P. E. Ross. Microsoft to spammers: go phish [e-mail security]. IEEE Spectrum, 43(1): 48–49, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2006:MBD

[Ros06e] P. E. Ross. Motorola becomes a disk jockey [cellphone music]. IEEE Spectrum, 43(1): 36–37, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2006:SP

[Ros06f] P. E. Ross. Silicon photonics. IEEE Spectrum, 43(1):54, January 2006. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2007:WLD

[Ros07] P. E. Ross. What’s the Latin for “Delusional”? IEEE Spectrum, 44(1):49–50, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2008:LNS

[Ros08a] P. E. Ross. Loser: Nanotechnology — still waiting for nanotube memory chip. IEEE Spectrum, 45(1):61–62, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2008:TTT

[Ros08b] P. E. Ross. Tools toys — on time and under budget. IEEE Spectrum, 45(7):23, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2008:UCC

[Ros08c] P. E. Ross. Update: A computer for the clouds. IEEE Spectrum, 45(8):11–12, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2008:WCF

[Ros08d] P. E. Ross. Why CPU frequency stalled. IEEE Spectrum, 45(4):72, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2009:CCK

[Ros09a] P. E. Ross. Cloud computing’s killer app: Gaming. IEEE Spectrum, 46(3):14, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2009:CR

[Ros09b] P. E. Ross. Computer reborn. IEEE Spectrum, 46(11): 42–47, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2009:TAS

[Ros09c] P. E. Ross. Tsunami alert system starts up in Indonesia — [update]. IEEE Spectrum, 46 (1):18–19, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2009:WLV

[Ros09d] P. E. Ross. Winners losers VI. IEEE Spectrum, 46(1): 31–32, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rotberg:2001:VVG

[Rot01a] E. Rotberg. Vintage video games [book review]. IEEE Spectrum, 38(12):60–61, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rothke:2001:PDI

[Rot01b] B. Rothke. The pipe dream of Internet voting. IEEE Spectrum, 38(2):14, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roush:2001:EER

[Rou01] W. Roush. EUV edging out rivals as next-generation IC fab tool. IEEE Spectrum, 38(6):25–26, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roy:2000:FFO

[Roy00] A. D. Roy. The first 100 feet: options for Internet and broadband access [books]. IEEE Spectrum, 37(4):9–12, April 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Raghavan:2000:IPI

[RSS00] G. Raghavan, M. Sokolich, and W. E. Stanchina. Indium phosphide ICs unleash the highfrequency spectrum. IEEE Spectrum, 37(10):47–52, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubenstein:2007:RGS

[Rub07]  R. Rubenstein.  Radios get

smart. IEEE Spectrum, 44(2): 46–50, February 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Riordan:2005:HHW

[RWZ05] J. Riordan, A. Wespi, and D. Zamboni. How to hook worms [computer network security]. IEEE Spectrum, 42 (5):32–36, May 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siang:2000:WWa

[SA00a] S. Siang and J. Adam. Washington watch. IEEE Spectrum, 37(10):32I2–32I4, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siang:2000:WWb

[SA00b] S. Siang and J. Adam. Washington watch. IEEE Spectrum, 37 (11):32I2–32I8, November 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:WW

[SA00c] W. Sweet and J. Adam. Washington watch. IEEE Spectrum, 37(7):20I2–20I4, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:2009:MMS

[SA09] B. Santo and S. Adee. Multicore made simple. IEEE Spectrum, 46(1):32–36, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Samuelson:2003:CCD

[Sam03] P. Samuelson. The Congressional corral [digital media copying prevention]. IEEE Spectrum, 40(5):33–35, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:2001:EBR

[San01] S. Santo. Embedded battle royal [embedded operating systems]. IEEE Spectrum, 38(12):

36–41, December 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sandell:2003:FBJ

[San03] P. Sandell. First a broken jaw, now a beautiful partnership. IEEE Spectrum, 40(2):11, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:2006:AA

[San06] B. R. Santo. Acronym addiction. IEEE Spectrum, 43 (10):28–33, October 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:2007:PNG

[San07] B. Santo. Plans for next-gen chips imperiled. IEEE Spectrum, 44(8):12–14, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sandborn:2008:TTE

[San08] P. Sandborn. Trapped on technology’s edge. IEEE Spectrum, 45(4):42–58, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:2009:MSW

[San09] B. R. Santo. 25 microchips that shook the world. IEEE Spectrum, 46(5):34–43, May 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sarpeshkar:2006:BPB

[Sar06] R. Sarpeshkar. Brain power — borrowing from biology makes for low power computing [bionic ear]. IEEE Spectrum, 43(5): 24–29, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2001:HKB

[Sav01] Neil Savage. Herwig Kogelnik [biography]. IEEE Spectrum, 38(6):42–47, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2002:LLH

[Sav02a]  Neil Savage.  Linking with

light [high-speed optical interconnects]. IEEE Spectrum, 39 (8):32–36, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2002:NYS

[Sav02b] Neil Savage. New York State wins top semiconductor R&D lab. IEEE Spectrum, 39(9): 26–30, September 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2002:BNI

[Sav02c] P. R. Savage. Betting on a new idea [organic leds]. IEEE Spectrum, 39(6):63–64, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2003:RCT

[Sav03a] Neil Savage. A revolutionary chipmaking technique? IEEE Spectrum, 40(11):18, November 2003. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2003:PHD

[Sav03b] P. Savage. The perfect handheld:

dream on. IEEE Spectrum, 40 (1):44–46, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2003:DA

[Sav03c] P. R. Savage. The data alchemists. IEEE Spectrum, 40 (7):38–39, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2004:LSS

[Sav04] Neil Savage. Light from silicon [silicon-based lasers]. IEEE Spectrum, 41(1):59, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2005:IPWa

[Sav05a] P. R. Savage. Invention: Patent watch. IEEE Spectrum, 42 (1):70–71, January 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2005:IPWb

[Sav05b] P. R. Savage. Inventions: Patent watch. IEEE Spectrum, 42(3):64, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2005:PWE

[Sav05c] P. R. Savage. Patent watch [electronics and communications devices]. IEEE Spectrum, 42 (9):62, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2008:URP

[Sav08a] Neil Savage. Update: Researchers pencil in graphene transistors. IEEE Spectrum, 45(6):13–14, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2008:WCM

[Sav08b] Neil Savage. A weaker, cheaper MRI. IEEE Spectrum, 45 (1):21, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2009:GMT

[Sav09a] Neil Savage. Graphene makes transistors tunable. IEEE Spectrum, 46(9):20, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savage:2009:GC

[Sav09b] Neil Savage. Grid on a chip. IEEE Spectrum, 46(12):15, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2001:EWU

[SB01] W. Sweet and E. A. Bretz. Energy woes [us national energy policy]. IEEE Spectrum, 38(7):48–53, July 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2002:HMD

[SB02] W. Sweet and E. A. Bretz. How to make deregulation work. IEEE Spectrum, 39(1):50–56, January 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schaeffer:2008:BGS

[SBB08] J. Schaeffer, V. Bulitko, and M. Buro. Bots get smart. IEEE Spectrum, 45(12):48–56, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Snoble:2000:S

[SC00] R. Snoble and W. Cox. Speakout. IEEE Spectrum, 37(10): 20–26, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stauffer:2006:NCT

[SC06a] B. Stauffer and B. Christie.

Nine cautionary tales. IEEE Spectrum, 43(9):36–45, September 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2006:FRE

[SC06b] W. Sweet and S. Cherry. A French regulatory edge [Internet regulations]. IEEE Spectrum, 43(5):14, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2007:HFC

[SC07] W. Sweet and Stephen Cass.

How to fight crime in real time. IEEE Spectrum, 44 (6):54–57, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneiderman:2000:BSD

[Sch00] R. Schneiderman. Bluetooth’s slow dawn. IEEE Spectrum, 37 (11):61–65, November 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scheibelhofer:2001:MPB

[Sch01] W. K. Scheibelhofer. The making of the pacemaker [book review]. IEEE Spectrum, 38(4):83, April 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schellenberg:2003:LLM

[Sch03a]  F. Schellenberg.  A little

light magic [optical lithography]. IEEE Spectrum, 40(9): 34–39, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schiff:2003:MGA

[Sch03b] D. Schiff. Mind the gap [age gap problems between employees]. IEEE Spectrum, 40(2): 60–61, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schewe:2006:EI

[Sch06] P. F. Schewe. Electric idyll. IEEE Spectrum, 43(10):40– 47, October 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schindall:2007:CU

[Sch07] J. Schindall. The charge of the ultracapacitors. IEEE Spectrum, 44(11):42–46, November 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2008:FP

[Sch08a] D. Schneider. Fresh phish. IEEE Spectrum, 45(10):34–38, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2008:ICH

[Sch08b] D. Schneider. The innovators club — [hands on]. IEEE Spectrum, 45(10):22–23, October 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneier:2008:HCH

[Sch08c] B. Schneier. Here comes here comes everybody (review of Here Comes Everybody: The Power of Organizing Without Organizations by C. Shirky) [books]. IEEE Spectrum, 45(9): 23–24, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:AYH

[Sch09a] D. Schneider. Automate your home shop. IEEE Spectrum, 46(6):22–23, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:ER

[Sch09b] D. Schneider. Easy rider. IEEE Spectrum, 46(9):26–27, September 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:FEG

[Sch09c] D. Schneider. A fuel-economy gauge for the rest of us. IEEE Spectrum, 46(4):20–21, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:IYB

[Sch09d] D. Schneider. If you build it, Santa will come. IEEE Spectrum, 46(11):20–21, November 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:MBP

[Sch09e] D. Schneider. Is this the moment for broadband over power lines? IEEE Spectrum, 46(7): 17, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:RES

[Sch09f] D. Schneider. Radio eye in the sky. IEEE Spectrum, 46 (1):58–62, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:RMR

[Sch09g] D. Schneider. Robin Murphy: Roboticist to the rescue. IEEE Spectrum, 46(2):36–37, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:RRS

[Sch09h] D. Schneider. Room for rent, stellar view. IEEE Spectrum, 46 (12):13, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:SC

[Sch09i] D. Schneider. The silence of the cellphones. IEEE Spectrum, 46(4):14, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scott:2000:AMT

[Sco00] P. Scott. Aerospace and military [technology 2000 analysis and forecast]. IEEE Spectrum, 37 (1):97–102, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2000:TSa

[Sel00a] K. Self. Technically speaking.

IEEE Spectrum, 37(2):25, February 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2000:TSb

[Sel00b] K. Self. Technically speaking. IEEE Spectrum, 37(4):25, April 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2000:TSc

[Sel00c] K. Self. Technically speaking. IEEE Spectrum, 37(6):33, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2000:TSd

[Sel00d] K. Self. Technically speaking. IEEE Spectrum, 37(8):22, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2000:TSe

[Sel00e] K. Self. Technically speaking. IEEE Spectrum, 37(10):33, October 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2000:TSf

[Sel00f] K. Self. Technically speaking. IEEE Spectrum, 37(12):30, December 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2001:AHS

[Sel01a] K. Self. About hobgoblins and @ signs. IEEE Spectrum, 38(6):98, June 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2001:CS

[Sel01b] K. Self. Crewless in space. IEEE Spectrum, 38(12):96, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2001:LSU

[Sel01c] K. Self. A lighter side to unit conversions [technically speaking]. IEEE Spectrum, 38(2): 104, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2001:MID

[Sel01d] K. Self. Measurement by imperial decree. IEEE Spectrum, 38(8):80, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2001:TLD

[Sel01e] K. Self. Technical language for deaf ears [technically speaking]. IEEE Spectrum, 38(10):88, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2002:AWE

[Sel02a] K. Self. All’s well that ends well. IEEE Spectrum, 39(4):82, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Self:2002:TON

[Sel02b] K. Self. A transmitter by any other name. IEEE Spectrum, 39(2):80, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2003:SYD

[Sel03] C. Selinger. Stuff you don’t learn in engineering school. IEEE Spectrum, 40(9):49–52, September 2003. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2004:AO

[Sel04a] C. Selinger. Anatomy of the office. IEEE Spectrum, 41(1): 90–93, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2004:CEW

[Sel04b] C. Selinger. The creative engineer: what can you do to spark new ideas? IEEE Spectrum, 41(8):47–49, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2005:AD

[Sel05a] C. Selinger. The art of delegating. IEEE Spectrum, 42 (8):62–63, August 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2005:PWO

[Sel05b] C. Selinger. From peace to war [office conflict]. IEEE Spectrum, 42(3):60–62, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2005:RM

[Sel05c] C. Selinger. Recovering from mistakes. IEEE Spectrum, 42 (12):56–57, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2006:CGY

[Sel06a] C. Selinger. Careers: Gilding your golden years. IEEE Spectrum, 43(8):51–53, August 2006.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2006:CMC

[Sel06b] C. Selinger. Careers: Making committees work. IEEE Spectrum, 43(6):53–55, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2006:CWD

[Sel06c] C. Selinger. Careers: Who dares, wins. IEEE Spectrum, 43(12): 49–52, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2006:LWJ

[Sel06d] C. Selinger. Learning where the jobs are [career growth]. IEEE Spectrum, 43(2):59–61, February 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2007:CBY

[Sel07a] C. Selinger. Careers: Becoming your own boss. IEEE Spectrum, 44(4):48–52, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2007:DM

[Sel07b] C. Selinger. Dealing with the media. IEEE Spectrum, 44 (8):58–59, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2007:W

[Sel07c] C. Selinger. Workaholism. IEEE Spectrum, 44(6):71–72, June 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2008:CYM

[Sel08a] C. Selinger. Careers — your mentor and you. IEEE Spectrum, 45(7):21, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2008:CSO

[Sel08b] C. Selinger. Careers: Survivor: The office. IEEE Spectrum, 45(5):20, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2008:DDP

[Sel08c] C. Selinger. Dealing with difficult people. IEEE Spectrum, 45(12):24, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2008:HTL

[Sel08d]  C. Selinger.  How to talk

like a salesman. IEEE Spectrum, 45(6):26, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2008:HD

[Sel08e] C. Selinger. Humor dos and don’ts. IEEE Spectrum, 45 (10):25, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selinger:2008:TTC

[Sel08f] C. Selinger. Travel tips [careers].

IEEE Spectrum, 45(2):23, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shenai:2000:MOP

[She00] K. Shenai. Made-to-order power.

IEEE Spectrum, 37(7):50–55, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sherwin:2001:TTR

[She01a] R. Sherwin. Tools toys [resources]. IEEE Spectrum, 38 (5):96–98, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sherwin:2001:MM

[She01b] R. M. Sherwin. Memory on the move. IEEE Spectrum, 38(5):55–59, May 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sheppard:2002:YMM

[She02] L. M. Sheppard. Not your mother’s mammography [breast cancer detection]. IEEE Spectrum, 39(10):56–57, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sheehan:2004:PEG

[She04] D. Sheehan. Photo essay: Greetings Earthlings! IEEE Spectrum, 41(8):20–21, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shorrock:2005:CAD

[Sho05a] T. Shorrock. Crash ABM deployment. IEEE Spectrum, 42 (2):11, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shorrock:2005:CI

[Sho05b] T. Shorrock. Cyber insecurity? IEEE Spectrum, 42 (3):15–16, March 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shorrock:2005:UDM

[Sho05c] T. Shorrock. U.S. deploys missile defense system. IEEE Spectrum, 42(2):10–13, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shrope:2001:FHT

[Shr01] M. Shrope. Florida’s high-tech haven. IEEE Spectrum, 38(11): 23–24, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shreeves:2009:WPD

[Shr09] K. Shreeves. Winner: Poseidon discovery — [tools toys]. IEEE Spectrum, 46(1):28–29, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siang:2001:BCC

[Sia01] Sanyin Siang. Bush and Congress converge on anti-terrorist legislation. IEEE Spectrum, 38(11):

27–28, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simpson:2001:CSE

[Sim01]  T. R. Simpson.  Chang-

ing the strategy for exploring space. IEEE Spectrum, 38(2): 16–17, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simonson:2007:BPA

[Sim07] E. Simonson. The big picture: Around the world in 65 days. IEEE Spectrum, 44(1): 18–19, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2002:IUG

[Sin02] S. Singh. India to use global positioning in civil air navigation system. IEEE Spectrum, 39(4):27–28, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2003:ICS

[Sin03a]  S. Singh.  Indian company

stretches fixed-wireless technology to create cellular network. IEEE Spectrum, 40(7):16–17, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2003:PET

[Sin03b] S. Singh. A passage of engineering talent back to india. IEEE Spectrum, 40(1): 22–25, January 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2004:VBF

[Sin04a] S. Singh. Venkataraman balaji: fighting drought with technology. IEEE Spectrum, 41(2): 36–37, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2004:ING

[Sin04b] V. Singh. India’s new government pledges to continue economic reforms — but investors worry a fractious coalition will be bad for business. IEEE Spectrum, 41(7):14, July 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2005:IMR

[Sin05] V. Singh. India may resuscitate disputed Enron plant: summer shortages give new life to India’s biggest power project. IEEE Spectrum, 42(8):18–19, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2006:BPC

[Sin06a] S. Singh. Big players in chip design buy into India. IEEE Spectrum, 43(3):14–17, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2006:ICW

[Sin06b] S. Singh. India connects to wireless Internet. IEEE Spectrum, 43(9):20, September 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2006:ISM

[Sin06c] S. Singh. India shoots for the Moon. IEEE Spectrum, 43 (12):16, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2006:IYC

[Sin06d] S. Singh. India’s year of computing inexpensively. IEEE Spectrum, 43(1):14, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2007:DDS

[Sin07a] S. Singh. Delhi’s defense spending spree. IEEE Spectrum, 44 (6):10–12, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2007:ISU

[Sin07b]  S. Singh.  Indian start-ups

lure Silicon Valley cash. IEEE Spectrum, 44(9):14–15, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2008:UTH

[Sin08] S. Singh. Update: Tata hopes its supercomputer is a money machine. IEEE Spectrum, 45(6):18, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2005:HGS

[Six05a] J. M. Six. Hidden gems [science and technology museums]. IEEE Spectrum, 42(1):70–71, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2005:STMa

[Six05b] J. M. Six. Sci… tech museums. IEEE Spectrum, 42(2):55, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2005:STMd

[Six05c] J. M. Six. Sci/tech museums. IEEE Spectrum, 42(7):62, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2005:STMe

[Six05d] J. M. Six. Sci/tech museums. IEEE Spectrum, 42(12): 60–61, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2005:STMb

[Six05e] J. M. Six. Sci/tech museums: Ann Arbor Hands-On Museum. IEEE Spectrum, 42(4):65, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2005:STMc

[Six05f] J. M. Six. Sci/tech museums: Museu de la Ciencia I de la Technica de Catalunya. IEEE Spectrum, 42(4):66, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2007:STMa

[Six07a] J. M. Six. Sci/tech museums.

IEEE Spectrum, 44(1):60, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Six:2007:STMb

[Six07b] J. M. Six. Sci/tech museums.

IEEE Spectrum, 44(3):54–56, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:TSN

[SK00] W. Sweet and J. Kumagai. The troubling state of nuclear controls. IEEE Spectrum, 37 (3):28–30, March 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Skolnikoff:2003:SS

[Sko03] Eugene B. Skolnikoff. Security and sanity. IEEE Spectrum, 40(4):13–14, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sorid:2000:VSV

[SM00] D. Sorid and Samuel K. Moore.

The virtual surgeon [virtual reality trainer]. IEEE Spectrum, 37(7):26–31, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strauss:2006:UA

[SMAS06] B. Strauss, M. G. Morgan, J. Apt, and D. D. Stancil. Unsafe at any airspeed? IEEE Spectrum, 43(3):44–49, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2003:FTT

[Smi03] B. Smith. Fort TV [TV show e-mail transmission prevention]. IEEE Spectrum, 40(5):30–32, May 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:2005:EBD

[Smi05] M. Smith. Eternal bits [digital files preservation]. IEEE Spectrum, 42(7):22–27, July 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schoenbach:2006:ZEV

[SNB06] K. H. Schoenbach, R. Nuccitelli, and S. J. Beebe. Zap [extreme voltage for fighting diseases]. IEEE Spectrum, 43 (8):20–26, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Snider:2000:FAM

[Sni00] A. D. Snider. Fourier analysis for the masses: [A First Course in Fourier Analysis, David W. Kammler, Prentice-Hall, Upper Saddle River, NJ, 2000, ISBN 0-13-578782-3, 830 pp., $84] [books]. IEEE Spectrum, 37(12): 15–16, December 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Snoonian:2003:SB

[Sno03]  D. Snoonian.  Smart build-

ings. IEEE Spectrum, 40(8): 18–23, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sontag:2008:JLP

[Son08a] S. Sontag. The jurassic, live and in person. IEEE Spectrum, 45 (12):25, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sontag:2008:TTR

[Son08b] S. Sontag. Tools toys — rekindling a love of books. IEEE Spectrum, 45(7):25, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sontag:2008:VDT

[Son08c] S. Sontag. The voltage the Dell [tools toys]. IEEE Spectrum, 45 (9):25, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sontag:2009:TWN

[Son09a] S. Sontag. Time waits for no engineer. IEEE Spectrum, 46(5):25, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sontag:2009:U

[Son09b] S. Sontag. Unsqueezed. IEEE Spectrum, 46(5):26, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sopensky:2003:BCC

[Sop03] E. Sopensky. Bush country cleans up its energy act. IEEE Spectrum, 40(2):26–29, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Skala:2008:F

[SOR08] M. Skala, P. O’Connor, and L. Rolland. Forum. IEEE Spectrum, 45(7):10, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spruth:2002:FDD

[Spr02] W. G. Spruth. Forum — digital divisiveness. IEEE Spectrum, 39(4):10–12, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sipper:2000:NSH

[SR00] M. Sipper and E. M. A. Ronald. A new species of hardware. IEEE Spectrum, 37(3):59–64, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schrick:2002:WBB

[SR02] B. Schrick and M. J. Riezenman. Wireless broadband in a box. IEEE Spectrum, 39 (6):38–43, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2009:ANA

[SR09] R. Schneider and J. Ross. Antennas for the new airwaves. IEEE Spectrum, 46(2):44–49, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2007:RHG

[SRJ^{+}07] W. Sweet, P. E. Ross, W. D.

Jones, Erico Guizzo, P. Bhattacharyya, Glenn Zorpette, and C. J. James. Resources: Holiday gifts. IEEE Spectrum, 44 (11):49–60, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Salvador:2005:TIP

[SS05] T. Salvador and J. Sherry. Taking the Internet to the people. IEEE Spectrum, 42(10): 32–35, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Santo:2008:PMT

[SS08] B. Santo and M. Santo. A plugin motorcycle [tools toys]. IEEE Spectrum, 45(2):24–25, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siang:2001:NA

[SSB^{+}01] S. Siang, W. Sweet, E. A.

Bretz, J. Makansi, B. Klein, and Samuel K. Moore. News analysis. IEEE Spectrum, 38 (2):18–32, February 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stick:2007:TT

[SSM07] D. Stick, J. D. Stern, and C. Monroe. The trap technique. IEEE Spectrum, 44 (8):36–43, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2005:CMS

[ST05] R. Singh and R. Thakur. Chip making’s singular future. IEEE Spectrum, 42(2):40–45, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Slakey:2009:WAN

[ST09] F. Slakey and B. Tannenbaum. What about the nukes? IEEE Spectrum, 46(3):26–52, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stahlkopf:2001:TOS

[Sta01] K. Stahlkopf. Technology offers solutions to the current power crisis. IEEE Spectrum, 38(6):14–15, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stevenson:2008:FSQ

[Ste08] R. Stevenson. First solar: quest for the $1 watt. IEEE Spectrum, 45(8):26–31, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stevenson:2009:LDS

[Ste09] R. Stevenson. The LED’s dark secret. IEEE Spectrum, 46 (8):26–31, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stirr:2003:ICI

[Sti03] B. Stirr. Invention: Caveat inventor. IEEE Spectrum, 40 (11):49, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stroh:2003:WMU

[Str03] S. Stroh. Wideband: multimedia unplugged. IEEE Spectrum, 40 (9):23–27, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sullivan:2005:DND

[Sul05] A. Sullivan. 3-deep: new displays render images you can almost reach out and touch. IEEE Spectrum, 42(4):30–35, April 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Svoboda:2004:FWE

[Svo04] E. Svoboda. Finding work the eBay way. IEEE Spectrum, 41 (8):16–17, August 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Svoboda:2006:OCC

[Svo06a] E. Svoboda. One-click content, no guarantees [online encyclopedia reliability]. IEEE Spectrum, 43(5):64–65, May 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Svoboda:2006:SDN

[Svo06b]  E. Svoboda.  Shot in the

dark [nuclear weapon testing]. IEEE Spectrum, 43(5):12–14, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Svoboda:2007:FSL

[Svo07] E. Svoboda. Fresh spin on logic. IEEE Spectrum, 44(1):15, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Staple:2004:ESS

[SW04] G. Staple and K. Werbach. The end of spectrum scarcity [spectrum allocation and utilization]. IEEE Spectrum, 41 (3):48–52, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stuck:2005:HVC

[SW05] B. Stuck and M. Weingarten. How venture capital thwarts innovation. IEEE Spectrum, 42 (4):50–55, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:TDT

[Swe00a] W. Sweet. The Telcordia Directory of Telecommunications Forums, Standards Bodies and Associations, Telcordia Technologies/SAIC, Morristown, NJ (Special Report SR-3776), September 1999, 332 pp., $99.95 (or $150 for book plus CD for one user) [books]. IEEE Spectrum, 37(8):14, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:CPA

[Swe00b] W. Sweet. Cell phones answer Internet’s call. IEEE Spectrum, 37(8):42–46, August 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:EIE

[Swe00c] W. Sweet. Electrical injuries.

Engineering, medical, and legal aspects [books]. IEEE Spectrum, 37(7):12, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:GNS

[Swe00d] W. Sweet. Going nuclear, Soviet style [books]. IEEE Spectrum, 37(5):14–15, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:MRW

[Swe00e] W. Sweet. The multifaceted Richard W. Sonnenfeldt. IEEE Spectrum, 37(7):56–61, July 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:PET

[Swe00f] W. Sweet. Power and energy [technology 2000 analysis and forecast]. IEEE Spectrum, 37 (1):86–90, January 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:RTS

[Swe00g] W. Sweet. Restructuring the thin-stretched [us electricity supply industry]. IEEE Spectrum, 37(6):43–49, June 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2000:TW

[Swe00h] W. Sweet. Technology watch. IEEE Spectrum, 37(8):21, August 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2001:CBD

[Swe01a] W. Sweet. China’s big dams. IEEE Spectrum, 38(9):46–51, September 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2001:GTA

[Swe01b] W. Sweet. Getting tough on antitrust. IEEE Spectrum, 38 (1):51–53, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2001:NAD

[Swe01c] W. Sweet. Networking assets [distributed generation]. IEEE Spectrum, 38(1):84–86, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2002:NIC

[Swe02] W. Sweet. Nuclear incident could prompt more costly inspections [power plants]. IEEE Spectrum, 39(6):21–22, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2003:NKN

[Swe03a] W. Sweet. Next up: Korean nuclear crisis. IEEE Spectrum, 40(5):18–19, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2003:TRR

[Swe03b] W. Sweet. Tighter regional regulation needed of power grids. IEEE Spectrum, 40(10): 12–13, October 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2003:TID

[Swe03c] W. Sweet. Tritium on ice: The dangerous new alliance of nuclear weapons and nuclear power [book review]. IEEE Spectrum, 40(5):42–43, May 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2003:UEL

[Swe03d] W. Sweet. U.S. energy legislation, much maligned, hangs fire. IEEE Spectrum, 40(11): 20–22, November 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2004:BRE

[Swe04a] W. Sweet. Book review: Energy at the crossroads. IEEE Spectrum, 41(5):58, May 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2004:BAS

[Swe04b] W. Sweet. Bush Administration’s “science” is under fire.

IEEE Spectrum, 41(4):12–13, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2004:DWT

[Swe04c] W. Sweet. Danish wind turbines take unfortunate turn. IEEE Spectrum, 41(11):30, November 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2004:DTW

[Swe04d] W. Sweet. Digital television wins holiday sweepstakes. IEEE Spectrum, 41(2):11, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2004:EAB

[Swe04e] W. Sweet. Energy answer — blowing in the wind. IEEE Spectrum, 41(2):10–12, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2004:INP

[Swe04f] W. Sweet. Iran’s nuclear program reaches critical juncture. IEEE Spectrum, 41(6):12–15, June 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2004:SSF

[Swe04g] W. Sweet. Self-sustaining fusion. IEEE Spectrum, 41(1): 50–51, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweeney:2005:TTM

[Swe05a] D. Sweeney. Tools toys: Mixed marriage. IEEE Spectrum, 42 (6):72–73, June 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2005:AEA

[Swe05b] W. Sweet. The atomic energy agency’s peace prize. IEEE Spectrum, 42(12):16–19, December 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2005:UEB

[Swe05c] W. Sweet. A U.S. energy bill at last. IEEE Spectrum, 42 (9):11, September 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2005:VZG

[Swe05d] W. Sweet. Victor Zagorodnov:

Getting high on glaciers [drilling equipment]. IEEE Spectrum, 42 (2):23–24, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2006:AGS

[Swe06a] W. Sweet. Adrenaline for the grid [superconducting synchronous condenser]. IEEE Spectrum, 43(1):44–47, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2006:GRQ

[Swe06b] W. Sweet. Get-rich-quick scheme [innovative electricity conservation]. IEEE Spectrum, 43(1): 10–12, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2006:MCE

[Swe06c] W. Sweet. Media: Climate emergency. IEEE Spectrum, 43 (8):49–50, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2006:SES

[Swe06d] W. Sweet. Security expert slams Bush’s surveillance program. IEEE Spectrum, 43 (4):23–25, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2006:YPU

[Swe06e] W. Sweet. Yangtze’s power is unleashed. IEEE Spectrum, 43 (8):14–15, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2007:CCC

[Swe07a] W. Sweet. Catastropic climate change from outer space [news]. IEEE Spectrum, 44(12):18, December 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2007:CC

[Swe07b] W. Sweet. Corn-o-copia. IEEE Spectrum, 44(1):26–28, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2007:KSI

[Swe07c] W. Sweet. Keeping score in the IP game. IEEE Spectrum, 44 (11):39–41, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2007:PBE

[Swe07d] W. Sweet. Protecting the big easy from the next big one. IEEE Spectrum, 44(3):10–12, March 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2007:SMP

[Swe07e]  W. Sweet.  Superconductor

maker in political crosshairs [news]. IEEE Spectrum, 44(11): 18–20, November 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2008:DAR

[Swe08a] W. Sweet. A disarming argument (review of Why Nuclear Disarmament Matters by H. Blix; 2008) [books]. IEEE Spectrum, 45(10):26, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2008:GGT

[Swe08b] W. Sweet. Greenhouse gas trends. IEEE Spectrum, 45 (1):88, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2008:LGN

[Swe08c] W. Sweet. Loser: Geopositioning — no payoff for Galileo Navigation System. IEEE Spectrum, 45(1):48–49, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2008:RRE

[Swe08d] W. Sweet. Russia rules Eurasian gas roost [the data]. IEEE Spectrum, 45(11):100, November 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2008:UCF

[Swe08e] W. Sweet. Update: A Connecticut fuel cell in South Korea’s grid. IEEE Spectrum, 45(8):16, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2008:WCC

[Swe08f] W. Sweet. Winner: Clean coal — restoring coal’s sheen. IEEE Spectrum, 45(1):57–60, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2009:K

[Swe09a] W. Sweet. Kyoto 2.0. IEEE Spectrum, 46(12):16, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2009:M

[Swe09b] W. Sweet. Moonstruck. IEEE Spectrum, 46(6):52–57, June 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sweet:2009:PMS

[Swe09c] W. Sweet. £10 million sea power challenge. IEEE Spectrum, 46(8):17, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tang:2009:MFL

[Tan09] H. X. Tang. May the force of light be with you. IEEE Spectrum, 46 (10):46–51, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Theberge:2006:GSS

[TD06] M. Theberge and G. Dudek.

Gone swimmin’ [seagoing robots]. IEEE Spectrum, 43(6):38–43, June 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teasley:2008:TTS

[Tea08] H. Teasley. Tools toys: Securing your laptop. IEEE Spectrum, 45(6):30, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teasley:2009:LEW

[Tea09] H. Teasley. Loser: eyes wide shut. IEEE Spectrum, 46 (1):29, January 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teitelman:2000:AG

[Tei00]  R. Teitelman.  Up against

gates. IEEE Spectrum, 37 (5):59–63, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teitelman:2001:CAP

[Tei01] R. Teitelman. Cisco’s accounting proved vulnerable to human error. IEEE Spectrum, 38(6):32–34, June 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teitelman:2002:ETL

[Tei02] R. Teitelman. Enron: The tale less trumpeted [management]. IEEE Spectrum, 39 (8):25–28, August 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tenner:2009:EE

[Ten09] E. Tenner. The effect effect. IEEE Spectrum, 46(2):22, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Terren:2007:BPS

[Ter07] P. Terren. The big picture: Step away from the vehicle! IEEE Spectrum, 44(5):18–19, May 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2003:PPP

[Tes03a] K. Teska. The problem with provisionals [patents]. IEEE Spectrum, 40(6):44, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Testa:2003:LPN

[Tes03b] B. M. Testa. Legacy phone networks get new lease on life: a telecom evolution caters to data demands. IEEE Spectrum, 40 (3):18–20, March 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Testa:2003:UPC

[Tes03c] B. M. Testa. US phone companies set stage for fiber to the curb. IEEE Spectrum, 40(9): 14–15, September 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Testa:2004:MPS

[Tes04] B. M. Testa. Mobile phone system passes Texas test. IEEE Spectrum, 41(2):18–21, February 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2006:IPO

[Tes06] K. Teska. Invention: Patently obvious. IEEE Spectrum, 43 (12):56–58, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2007:FPM

[Tes07a] K. Teska. The foreign patent money trap [resources: Invention]. IEEE Spectrum, 44(12): 62–63, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2007:RHH

[Tes07b] K. Teska. Resources: How high is the patent bar now? IEEE Spectrum, 44(10):60–63, October 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2008:HII

[Tes08a] K. Teska. Hired to invent [invention]. IEEE Spectrum, 45 (2):25–26, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2008:ISP

[Tes08b] K. Teska. Invention [software patents 101]. IEEE Spectrum, 45(3):24, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2008:IPP

[Tes08c] K. Teska. Invention: the pitfalls of patent searches knowing about. IEEE Spectrum, 45(6):29, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2008:IPM

[Tes08d] K. Teska. Invention: the poor man’s patent. IEEE Spectrum, 45(8):23, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2009:NPP

[Tes09a] K. Teska. A new patent perspective. IEEE Spectrum, 46 (10):22, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2009:SP

[Tes09b] K. Teska. Strategic patenting. IEEE Spectrum, 46(2):20, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teska:2009:WMY

[Tes09c] K. Teska. What makes you an inventor? IEEE Spectrum, 46(8):24, August 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tietjen:2004:WFW

[Tie04] J. S. Tietjen. Why so few women, still? IEEE Spectrum, 41(10):57–58, October 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thomsen:2000:ITG

[TJ00] G. Thomsen and Y. Jani. Internet telephony: going like crazy. IEEE Spectrum, 37 (5):52–58, May 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tamada:2002:KWG

[TLT02] J. A. Tamada, M. Lesho, and M. J. Tierney. Keeping watch on glucose. IEEE Spectrum, 39(4):52–57, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trope:2006:LMV

[TP06] R. L. Trope and M. Power.

The lessons of MGM v Grokster [copyright-infringement]. IEEE Spectrum, 43(1):57–60, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Treleaven:2004:SU

[Tre04] P. Treleaven. Sizing us up.

IEEE Spectrum, 41(4):28–31, April 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Trope:2004:GAT

[Tro04] R. L. Trope. Guarding against terrorism and liability. IEEE Spectrum, 41(1):85–86, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tredennick:2001:EVV

[TS01] N. Tredennick and B. Shimamoto. An engineer’s view of venture capitalists [speakout]. IEEE Spectrum, 38(9): 13–19, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tredennick:2003:GRP

[TS03] N. Tredennick and B. Shimamoto. Go reconfigure [programmable logic in handheld devices]. IEEE Spectrum, 40(12): 36–40, December 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tummala:2006:MLM

[Tum06] R. R. Tummala. Moore’s Law meets its match (system-on-package). IEEE Spectrum, 43 (6):44–49, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turner:2008:HHO

[Tur08] J. Turner. Hands-on: hacking the olpc. IEEE Spectrum, 45 (8):24–25, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turner:2009:PP

[Tur09a] J. Turner. The projector project.

IEEE Spectrum, 46(5):22–23, May 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turner:2009:TYS

[Tur09b] J. Turner. Time is on your side. IEEE Spectrum, 46(7):24– 25, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ubell:2000:ETL

[Ube00] R. Ubell. Engineers turn to elearning. IEEE Spectrum, 37 (10):59–63, October 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ullman:2003:BRB

[Ull03] E. Ullman. Book review: The bug. IEEE Spectrum, 40(11): 48, November 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2006:SEF

[Ups06] Sandra Upson. Showdown on the energy frontier. IEEE Spectrum, 43(12):12–15, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:GSE

[Ups07a] Sandra Upson. A global search engine for geospatial data. IEEE Spectrum, 44(8):20, August 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:GG

[Ups07b] Sandra Upson. The greening of google. IEEE Spectrum, 44 (10):24–28, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:G

[Ups07c] Sandra Upson. Grounded. IEEE Spectrum, 44(1):32–34, January 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:HTW

[Ups07d] Sandra Upson. How a taser works. IEEE Spectrum, 44(12): 24–31, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:HSU

[Ups07e] Sandra Upson. How to see the unseen city. IEEE Spectrum, 44(6):36–37, June 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:SLR

[Ups07f] Sandra Upson. Science for the layman [review of The Canon: A Whirligig Tour of the Beautiful Basics of Science (Angier, N.; 2007)]. IEEE Spectrum, 44 (8):59–60, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:TV

[Ups07g] Sandra Upson. Tongue vision. IEEE Spectrum, 44(1): 44–45, January 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:UES

[Ups07h] Sandra Upson. U.S. Earthsensing satellites left out in the cold. IEEE Spectrum, 44 (4):10–12, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2007:WW

[Ups07i] Sandra Upson. Wiretapping woes. IEEE Spectrum, 44 (5):10–12, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:AOP

[Ups08a] Sandra Upson. Across the outback on photon alone. IEEE Spectrum, 45(2):50–60, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:DHF

[Ups08b] Sandra Upson. The data: How free is solar energy? When solar cells break even; winning back the watts; global warming potential of energy options. IEEE Spectrum, 45(2):72, February 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:LWO

[Ups08c] Sandra Upson. A less well-oiled war machine. IEEE Spectrum, 45 (10):28–60, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:LCE

[Ups08d] Sandra Upson. Loser: Climate engineering algae bloom climatechange scheme doomed. IEEE Spectrum, 45(1):37–38, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:LSN

[Ups08e] Sandra Upson. Loser: Social networking — social networking goes to the dogs. IEEE Spectrum, 45(1):55–56, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:SCM

[Ups08f] Sandra Upson. Steven camilleri — motor maniac [dream jobs 2008]. IEEE Spectrum, 45(2): 36–37, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:UDT

[Ups08g] Sandra Upson. Update: A difficult time for depression devices. IEEE Spectrum, 45(5):14, May 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:UPW

[Ups08h] Sandra Upson. Update: Putting wireless power to work smart sensors harvest. IEEE Spectrum, 45(6):16, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2008:URS

[Ups08i] Sandra Upson. Update: Radiation sensor fine-tunes cancer treatments. IEEE Spectrum, 45(6):22, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2009:AGB

[Ups09a] Sandra Upson. Arieta gonelevu bright lights, little islands. IEEE Spectrum, 46(2):29–30, February 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2009:FNW

[Ups09b] Sandra Upson. Finland’s nuclear waste solution. IEEE Spectrum, 46(12):25–49, December 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2009:HTO

[Ups09c] Sandra Upson. A high-tech oasis in Saudi Arabia. IEEE Spectrum, 46(9):18, September 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2009:HR

[Ups09d] Sandra Upson. Hot rocks. IEEE Spectrum, 46(1):39–41, January 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2009:JRW

[Ups09e] Sandra Upson. Jordan’s radioactive water problem. IEEE Spectrum, 46(8):13–15, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2009:JRS

[Ups09f] Sandra Upson. Jordan’s Red Sea desalination plan. IEEE Spectrum, 46(7):18, July 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Upson:2009:RRP

[Ups09g] Sandra Upson. Rockets for the red planet. IEEE Spectrum, 46(6):42–47, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ulrich:2003:PPT

[US03] R. Ulrich and L. Schaper. Putting passives in their place [pcb integration]. IEEE Spectrum, 40(7):26–30, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vandersypen:2007:DDD

[Van07] L. Vandersypen. Dot-to-dot design. IEEE Spectrum, 44(9): 42–47, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vatis:2001:CMP

[Vat01]  M. Vatis.  A counterterror-

ist Manhattan Project. IEEE Spectrum, 38(11):14–15, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Venkat:2002:DDD

[Ven02] K. Venkat. Delving into the digital divide. IEEE Spectrum, 39(2):14, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Verkest:2003:MCH

[Ver03] D. Verkest. Machine chameleon [handheld devices]. IEEE Spectrum, 40(12):41–46, December 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vinge:2004:SS

[Vin04] V. Vinge. Synthetic serendipity. IEEE Spectrum, 41(7):35– 44, July 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vinge:2008:SS

[Vin08] V. Vinge. Signs of the singularity. IEEE Spectrum, 45(6):76–82, June 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vaughan-Nichols:2003:SPM

[VN03] S. J. Vaughan-Nichols. Saving private e-mail. IEEE Spectrum, 40(8):40–44, August 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2004:TTC

[Voe04] J. Voelcker. Top 10 tech cars. IEEE Spectrum, 41(3):28–35, March 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2005:TTC

[Voe05] J. Voelcker. Top 10 tech cars [hybrid electric vehicles]. IEEE Spectrum, 42(3):22–30, March 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2006:ECE

[Voe06a] J. Voelcker. Electric cars for enlightened stars. IEEE Spectrum, 43(11):14–16, November 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2006:SSA

[Voe06b] J. Voelcker. Stalked by satellite — an alarming rise in GPSenabled harassment. IEEE Spectrum, 43(7):15–16, July 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2006:TTC

[Voe06c] J. Voelcker. Top 10 tech cars [fuel efficient cars]. IEEE Spectrum, 43(4):34–35, April 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2007:CGS

[Voe07a] J. Voelcker. Cars get street smart [news]. IEEE Spectrum, 44 (10):16–18, October 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2007:LBT

[Voe07b] J. Voelcker. Lithium batteries take to the road. IEEE Spectrum, 44(9):26–31, September 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2007:TTC

[Voe07c]  J. Voelcker.  Top 10 tech

cars 2007. IEEE Spectrum, 44 (4):34–41, April 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2008:BC

[Voe08a] J. Voelcker. Battery czar. IEEE Spectrum, 45(8):32–37, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2008:PP

[Voe08b] J. Voelcker. Plugging away in a prius. IEEE Spectrum, 45(5):30–48, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2008:SNM

[Voe08c] J. Voelcker. The soul of a new mercedes. IEEE Spectrum, 45 (12):36–41, December 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2008:SLA

[Voe08d] J. Voelcker. Spectral lines: An almost all-electric car? IEEE Spectrum, 45(4):9, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2008:TTC

[Voe08e] J. Voelcker. Top 10 tech cars it’s the environment, stupid. IEEE Spectrum, 45(4):26–35, April 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2009:HGM

[Voe09a] J. Voelcker. How green is my plug-in? IEEE Spectrum, 46 (3):42–58, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2009:TTC

[Voe09b] J. Voelcker. Top 10 tech cars. IEEE Spectrum, 46(4):42–49, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:2009:WWK

[Voe09c] J. Voelcker. Will Washington kick-start the U.S. battery biz? IEEE Spectrum, 46 (3):11–12, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

vonLohmann:2006:DD

[vS06] F. von Lohmann and W. Seltzer. Death by DMCA. IEEE Spectrum, 43(6):24–30, June 2006.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Walker:2000:MWD

[Wal00] M. G. Walker. Modeling the wiring of deep submicron ics. IEEE Spectrum, 37(3):65–71, March 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2001:ACW

[Wal01a] P. Wallich. Amiga: the computer that wouldn’t die. IEEE Spectrum, 38(3):40–46, March 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2001:MJK

[Wal01b] P. Wallich. Mindstorms: not just a kid’s toy. IEEE Spectrum, 38 (9):52–57, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2002:DH

[Wal02a] P. Wallich. Digital hubbub.

IEEE Spectrum, 39(7):26–31, July 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2002:GCP

[Wal02b] P. Wallich. The ghosts of computers past. IEEE Spectrum, 39 (11):38–43, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2003:GM

[Wal03] P. Wallich. Getting the message. IEEE Spectrum, 40 (4):38–42, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2004:EEI

[Wal04] P. Wallich. Electrical engineering’s identity crisis — when does a vast and vital profession become unrecognizably diffuse? IEEE Spectrum, 41(11): 66–73, November 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2005:TTR

[Wal05] P. Wallich. Tools toys: I, roboticist. IEEE Spectrum, 42 (10):63–65, October 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2007:BYE

[Wal07a] P. Wallich. Build yourself an electric gun [resources: Tools toys]. IEEE Spectrum, 44(12): 60–61, December 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2007:DS

[Wal07b] P. Wallich. Deeply superficial. IEEE Spectrum, 44(8): 56–57, August 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2007:KDT

[Wal07c] P. Wallich. Kids, do try this at home. IEEE Spectrum, 44 (9):62–64, September 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2008:TTP

[Wal08a] P. Wallich. Tools toys — playing with cyc. IEEE Spectrum, 45(7):24–25, July 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2008:TTH

[Wal08b] P. Wallich. Tools toys: Hacking the Nokia N800. IEEE Spectrum, 45(4):25, April 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2008:TTN

[Wal08c] P. Wallich. Tools toys: N800 fights the bad guys. IEEE Spectrum, 45(5):23, May 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2009:FBa

[Wal09a] P. Wallich. Fusion on a budget.

IEEE Spectrum, 46(3):20–21, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2009:FBb

[Wal09b] P. Wallich. The future’s so bright. IEEE Spectrum, 46(8):22, August 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallich:2009:SGP

[Wal09c] P. Wallich. The supercomputer goes personal. IEEE Spectrum, 46(4):64, April 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wang:2002:CSP

[Wan02] G. Wang. Ch solves portability headaches. IEEE Spectrum, 39(6):59, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Webb:2004:CS

[Web04] C. E. Webb. Chip shots. IEEE Spectrum, 41(10):48–53, October 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Webb:2005:BS

[Web05]  C. E. Webb.  The body

shops. IEEE Spectrum, 42(2): 34–39, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weeks:2005:PEL

[Wee05] H. Weeks. Photo essay: The long and winding road. IEEE Spectrum, 42(2):18–19, February 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weil:2001:PBD

[Wei01] J. Weil. Pebble-bed design returns [nuclear reactor]. IEEE Spectrum, 38(11):37–40, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:2002:EFU

[Wei02a] G. Weiss. Europe favors United Air Traffic Control System. IEEE Spectrum, 39(11): 22–23, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:2002:WAD

[Wei02b] G. Weiss. Welcome to the (almost) digital hospital. IEEE Spectrum, 39(3):44–49, March 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:2003:WPB

[Wei03] G. Weiss. Who’s politically biased? IEEE Spectrum, 40 (4):20–21, April 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:2004:AWS

[Wei04] G. Weiss. Around the world in a solar plane. IEEE Spectrum, 41(3):12–14, March 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:2007:CDD

[Wei07] G. Weiss. CERN’s discerning detectors. IEEE Spectrum, 44(5):12–14, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weinberger:2009:PG

[Wei09a] S. Weinberger. Powerless in Gaza. IEEE Spectrum, 46(12): 36–41, December 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:2009:PLC

[Wei09b] G. Weiss. Philippe Lauper chase the Sun. IEEE Spectrum, 46 (2):38, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Westly:2009:SLC

[Wes09] E. Westly. Sign language by cellphone. IEEE Spectrum, 46(3):16, March 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Willebrand:2001:FOF

[WG01] H. A. Willebrand and B. S. Ghuman. Fiber optics without fiber. IEEE Spectrum, 38 (8):40–45, August 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Williams:2008:HWF

[Wil08] R. S. Williams. How we found the missing memristor. IEEE Spectrum, 45(12):28–35, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wintner:2006:BBS

[Win06] R. Wintner. Bits on the big screen. IEEE Spectrum, 43(12): 42–48, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Warrior:2003:TKW

[WMM03] J. Warrior, E. McHenry, and K. McGee. They know where you are [location detection]. IEEE Spectrum, 40(7):20–25, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wohl:2005:ILU

[Woh05] A. D. Wohl. Interface lift [user interface]. IEEE Spectrum, 42 (11):32–37, November 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolfe:2002:OSP

[Wol02a] A. Wolfe. An open-source platform to tackle scientific computing. IEEE Spectrum, 39 (6):22–24, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolfe:2002:PPS

[Wol02b] A. Wolfe. Putting pen to screen on tablet pcs. IEEE Spectrum, 39(10):16–18, October 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolfe:2003:HRL

[Wol03] A. Wolfe. HDTV — ready for the long drive? IEEE Spectrum, 40(6):13–15, June 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolf:2009:KPW

[Wol09] D. A. Wolf. The kind of people who will go to Mars. IEEE Spectrum, 46(6):38, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wood:2008:FRF

[Woo08] R. Wood. Fly, robot, fly. IEEE Spectrum, 45(3):25–29, March 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wehner:2009:RCC

[WOS09] M. Wehner, L. Oliker, and J. Shalf. A real cloud computer. IEEE Spectrum, 46(10): 24–29, October 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wouk:2001:LP

[Wou01] V. Wouk. Lessons of the past. IEEE Spectrum, 38(2):77, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wright:2005:BEB

[Wri05] J. Wright. Blades have the edge [blade servers]. IEEE Spectrum, 42(4):24–29, April 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Whitehorn-Umphres:2001:HHR

[WU01] D. Whitehorn-Umphres. Hackers, hot rods, and the information drag strip. IEEE Spectrum, 38(10):14–17, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wu:2008:CEW

[Wu08a] C. Wu. Careers: An engineer walks into a comedy club . . .. IEEE Spectrum, 45(4):23, April 2008. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Wu:2008:TTR

[Wu08b] C. Wu. Tools toys [robotic photographers]. IEEE Spectrum, 45(3):23, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yang:2008:RF

[YDLC08] L. Yang, R. P. Dick, H. Lekatsas, and S. Chakradhar. RAM for free. IEEE Spectrum, 45 (8):38–43, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ye:2008:BSE

[Ye08] P. D. Ye. Beyond silicon’s elemental logic. IEEE Spectrum, 45 (9):42–47, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yeung:2006:BPJ

[Yeu06] P. Yeung. The big picture: Just charge it. IEEE Spectrum, 43 (12):22–23, December 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yim:2002:MR

[YZD02] M. Yim, Ying Zhang, and D. Duff. Modular robots. IEEE Spectrum, 39(2):30–34, February 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zachary:2005:AHS

[Zac05] G. P. Zachary. The African hacker [software development].

IEEE Spectrum, 42(8):24–29, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zachary:2007:TP

[Zac07]  G. P. Zachary.  Thirst for

power. IEEE Spectrum, 44 (5):32–37, May 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zak:2002:TTR

[Zak02a] A. Zak. Taking tourists for a ride (in space). IEEE Spectrum, 39(6):30–31, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zak:2002:URR

[Zak02b] A. Zak. US rocket with Russian engine gets to work. IEEE Spectrum, 39(10):23, October 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zak:2003:RGI

[Zak03] A. Zak. Russia gives International Space Station a lift. IEEE Spectrum, 40(2):28–29, February 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zak:2005:EJR

[Zak05a] A. Zak. Europe to join Russia in building next space shuttle. IEEE Spectrum, 42(8):12–14, August 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zak:2005:RPN

[Zak05b] A. Zak. Russians propose a new space shuttle. IEEE Spectrum, 42 (2):13–14, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zak:2009:RRM

[Zak09] A. Zak. A Russian return to a Martian moon. IEEE Spectrum, 46(6):36–37, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zbikowski:2005:FLF

[Zbi05]  R. Zbikowski.  Fly like a

fly [micro-air vehicle]. IEEE Spectrum, 42(11):46–51, November 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zeldes:2009:I

[Zel09] N. Zeldes. Infoglut. IEEE Spectrum, 46(10):30–55, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zimmerman:2000:RFC

[Zim00] D. Zimmerman. Radar’s first comprehensive history [books]. IEEE Spectrum, 37(6):12–13, June 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zimmerman:2008:BEE

[Zim08] R. Zimmerman. Books — an exquisite ear [review of Riding the Waves: A Life in Sound, Science, and Industry (Beranek, L.; 2008)]. IEEE Spectrum, 45(7): 24, July 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2001:UNW

[ZM01] Glenn Zorpette and S. Miller. Unconventional nuclear weapons. IEEE Spectrum, 38(11):54–55, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2001:NA

[ZMR^{+}01] Glenn Zorpette, J. Mullins, W. Roush, C. Morrison, G. Weiss, S. R. Das, J. Kumagai, A. Hellemans, and F. Ferorelli. News analysis. IEEE Spectrum, 38(9): 22–39, September 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2001:MMT

[Zor01a]  Glenn Zorpette.  Megawatts

and megatons: a turning point in the nuclear age, Georges Charpak and Richard L. Garwin, Knopf, New York, $30 [book review]. IEEE Spectrum, 38 (11):68, November 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2001:TOH

[Zor01b] Glenn Zorpette. The technology of orgasm: ‘hysteria,’ the vibrator, and women’s sexual satisfaction, by Rachel P. Maines, Johns Hopkins, Baltimore, MD, $36 [book review]. IEEE Spectrum, 38(11):67–68, November 2001.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2001:CHN

[Zor01c] Glenn Zorpette. Canned heat [nuclear waste storage]. IEEE Spectrum, 38(11):44–47, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2001:NWT

[Zor01d] Glenn Zorpette. A new world of terror. IEEE Spectrum, 38 (10):24–25, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2001:QST

[Zor01e] Glenn Zorpette. The quest for the spin transistor. IEEE Spectrum, 38(12):30–35, December 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2002:MIS

[Zor02a] G. E. Zorpette. Making intelligence smarter. IEEE Spectrum, 39(1):38–43, January 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2002:BB

[Zor02b] Glenn Zorpette. Breaking into the blue. IEEE Spectrum, 39 (6):18–19, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2002:DHM

[Zor02c] Glenn Zorpette. Drilling holes in Mars. IEEE Spectrum, 39(4):22, April 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2002:LTL

[Zor02d] Glenn Zorpette. Let there be light [gallium nitride led]. IEEE Spectrum, 39(9):70–74, September 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2002:NLN

[Zor02e] Glenn Zorpette. New life for nixies [in digital clocks]. IEEE Spectrum, 39(6):44–49, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2002:SSR

[Zor02f] Glenn Zorpette. Over the solidstate rainbow. IEEE Spectrum, 39(4):62, April 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2002:SKC

[Zor02g] Glenn Zorpette. Sun kings cross the outback [solar powered vehicles marathon]. IEEE Spectrum, 39(2):40–46, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2004:LKK

[Zor04a] Glenn Zorpette. Lau kofoed kierstein: playtime in the land of lego. IEEE Spectrum, 41 (2):30–31, February 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2004:SH

[Zor04b] Glenn Zorpette. The smart hybrid. IEEE Spectrum, 41 (1):44–51, January 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2005:SCU

[Zor05] Glenn Zorpette. Super charged [ultracapacitors]. IEEE Spectrum, 42(1):32–37, January 2005. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2006:PTA

[Zor06a] Glenn Zorpette. Plane talk about cellphones [spectral lines]. IEEE Spectrum, 43(3):32–36, March 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2006:REI

[Zor06b] Glenn Zorpette. Re-engineering Iraq [power production shortage]. IEEE Spectrum, 43(2): 22–35, February 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2006:WWZ

[Zor06c] Glenn Zorpette. Working in a war zone: engineers in Iraq find high risk but high reward. IEEE Spectrum, 43 (3):64–66, March 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2008:BSD

[Zor08a] Glenn Zorpette. Bomb squad diary. IEEE Spectrum, 45(10): 40–47, October 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2008:BPS

[Zor08b] Glenn Zorpette. Bruno Putzeys — the sound of music [dream jobs 2008]. IEEE Spectrum, 45 (2):32–33, February 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2008:CI

[Zor08c] Glenn Zorpette. Countering ieds. IEEE Spectrum, 45(9): 26–35, September 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2008:WL

[Zor08d] Glenn Zorpette. Why losers? IEEE Spectrum, 45(1):30–31, January 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2009:SS

[Zor09a] Glenn Zorpette. It’s the stupidity, stupid. IEEE Spectrum, 46 (11):10, November 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2009:KBM

[Zor09b] Glenn Zorpette. Keya Banerjee: Magical realism. IEEE Spectrum, 46(2):34–35, February 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:2007:BMD

[ZR07] Glenn Zorpette and P. E. Ross.

The books that made a difference [resources]. IEEE Spectrum, 44(3):58–59, March 2007. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zubrin:2002:PPS

[Zub02] R. Zubrin. Power and purpose in space. IEEE Spectrum, 39 (11):12–14, November 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zubrin:2009:HGM

[Zub09] R. Zubrin. How to go to Mars — right now! IEEE Spectrum, 46(6):48–49, June 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).